A Complete Bibliography of The Mathematical Gazette (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

08 August 2024 Version 1.20

Title word cross-reference

$ [Cou11]. (a + b)^{p} [Jam14e]. (a + b)^{p} + (a − b)^{p} [Jam14e]. (a, b, b + 1) [HJ19]. 1/x [Geo10]. $12-95 [Rua12]. $129.00 [Sch14]. $14.95 [LP10]. $14.99 [Lev12c]. 15 [Lor12i]. $16.95 [Tol15c]. $16.99 [Tol15f]. $18.99 [Mac14d]. $19 [Lor17g]. $19.95 [Ban14b]. $19.99 [Tol19f]. $195 [Cou14]. 1: 2 [Nic14b]. $23.99 [Hal11c, Hud10a]. $24 [Lev14e]. $24.95 [Lev10f, Tol17g]. $26 [Lev14p]. $26.99 [Tol17e]. $27 [Lor18f]. $27-00 [Lev14h]. $27-95 [Gri14b]. $27.95 [Lev14n, Tol17i]. $29 [Gri10b, Hud10c]. $29.95 [Hun13d]. $29.99 [Tol17b]. 2 < e < 3 [Lor11m]. 2^{67} − 1 [Shi15d]. 3 [CP10, Had19, Rot10d, dVS18]. $30.00 [Jac19, Mac18]. $32 [Lor11g]. $32.00 [Tol15d]. $35.00 [And13b]. $36.39 [Shi16b]. $39 [Lev10e]. $39.00 [Shi10i]. $39.20 [Hun17d]. $39.95 [Lor17e]. 3 × 3 [Bea17a, Lor15f]. 4 [Rot10d]. $42.00 [Hun14a]. $45.00 [Cou10, Mac16c]. $49-00 [Gib10b]. $49.00 [Ban10a, Hun13c]. $49.95 [Tol17f]. 5 [Lor12i]. $50 [Lev11d, Lor18e]. $56.00 [Shi17a]. $59 [App11]. $59-00 [Hud11a]. $59.95 [Lev13g, Yeo16].

$60 [De 14b]. 60 [Lor12i]. $68.95 [Lev14k]. $69.95 [Lea10]. 6n ± 1 [Lor14l]. $70 [Had19]. $70.50 [Shi18b]. $75.00 [Eat12]. $79-00 [For12]. $79.95 [Lev12a]. $94.95 [App10a]. $97.50 [App10b]. [x] [Orm14]. ^{2} [LRQ10]. a [AT12]. A + B + C = π [Sub18b]. A/(x + B) + C/(x + D) [Cel17]. $A115 [Lev17e]. a^{a} [AT12]. a^{n} = b^{n} + c^{n} [Lor15b]. arctan(x) [Den10].

aX^{2} + bXY + cY^{2} = dZ^{2} [Pla14]. b [Coo19]. ( \begin{matrix}n + 1 \\ 2 \end{matrix} ) [DeM10]. · [KK10].

csc 2x = cot x − cot 2x [Cha15]. deg(0) = −∞ [Haj17e]. e [Bro12a, Mac10d]. e < 2 [Shi14a]. e < 3 [You13]. e^{1/M} [Ton16]. e^{x} ≥ 1 + x [Lor17c]. exp(−x^{2}) [Che18]. F [All18, MW18, Pla16]. f(x)/x [MW18]. Γ^{2}(1/2) [MGC13a]. I(α) = ∫_{0}^{∞} 1/(1 + x^{α})^{2} dx [Mur12]. int_{0}^{1}1/(1 + x^{2}) dx = π/4 [You11b]. ∫_{0}^{∞} 1/(1 + x^{2}) dx = π/2 [You11a]. ∫_{0}^{∞} \frac{sin x}{/}x dx [Lor16e]. ∫_{0}^{∞} sin(x)/x dx [Sik18]. ∫_{0}^{∞} | sin^{a}(x) + cos^{b}(x)|/x^{p} dx [Mur11c]. ∫_{0}^{π} log(2 sin u)du [Mer19]. ∫ csc x dx [JO14]. k [ES11, PdVF14]. lim_{m→∞} ∑_{k=0}^{m} (\frac{k}{m})^{m} [Lor16a]. ln 2 < 1\sqrt{2} [Lor14k]. log x [OJN16]. Z\sqrt{[10]} [Sim19]. m × (m + 1) [EM16]. n! [Lor17b]. N [PdVF14, Ber18, Lor17b, Ste19a]. n(x + y + z) = xyz [Dol16a]. n = 3 [Bar17]. n^{2} [DeM10]. p [Hun14b, Sch11]. p/ sin θ + q/ cos θ = 1 [Sta18]. p > 1 [Nil13]. ϕ(m) = k! [Shi13b, Sur14]. π [Hir11e, Lev12f, Lip13, Lor12a, Lor19a, Lor19e, Mac10d, MGC13a, NL16, Osl10a, Osl11b, Osl12b, SS12, Shi14a, You12, Zor11]. π ln k [Pla11a]. ∏_{n=2}^{∞} ζ(n) [Ökt15]. py^{2} = x^{4} + x [Dol13a, ZF10]. Si(x) [JLM15]. σ(n) = k! [Shi13b]. sin^{2} A + sin^{2} B + sin^{2} C = 4 sin A sin B sin C [Sub18b]. sin^{n} x/x^{m} [Tra10]. SL(2, p) [Suc11]. \sqrt{(}x), ^{3} \sqrt{3} [Hir11h]. \sqrt{2} [Lor15l, Lor17h, MGC13b, Raj18, Ste12b]. \sqrt{2π} [Lor15c]. \sqrt{abcd} [Lor12g]. \sqrt{d} [Yan14]. \sqrt{N} [Lor18a]. ∑_{1}^{n} r(r + 1) [Lor14i]. ∑_{1}^{n} r^{2} [Lor14i]. ∑_{k=1}^{∞} 1/k^{2} = π^{2}/6 [Lor16f]. ∑_{n=1}^{∞} 1/n^{2} [JL13]. ∑_{n=1}^{∞} 1/n^{p} [Nil13].

T_{1} + T_{2} + ⋯ + T_{n} = ( \begin{matrix}n + 2 \\ 3 \end{matrix} ) [Sub15a]. × [Har10a]. x^{2} + ry^{2} = z^{2} [Rea19].

x^{2} + z^{2} [Dol17a]. ζ(2) [Shi16a, Sik18]. ζ(2) = π^{2}/6 [Jam13c]. ζ(2n) [Lor14c]. ζ(3) [Hir11d, Jam14f]. ζ(n) [Jam14d]. |x| [Orm14].

-adic [Hun14b, Sch11]. -D [dVS18]. -dimensional [CP10]. -folds [Rot10d, Rot10d]. -gonal [ES11]. -gons [Lor12i]. -mean [Pla16]. -ratio [All18]. -Spinout [Coo19].

. [Jon19].

0-521-61999-8 [Lev10c]. 0-691-13576-2 [Mac10b]. 0-906588-67-7 [Hal10b]. 0-907669-46-8 [Kil10]. 0-9516701-8-2 [Rot10a].

1 [BL12]. 1-56881-234-5 [Mah10b]. 100.C [Ret17]. 100.I [Bea18a]. 1000 [Yeo17a, Yeo17a]. 101.F [Gib19a]. 117-00 [Shi11d]. 13-978-1-4020-8886-5 [Sch10]. 1439-7382 [App14b]. 15-00 [Hun10b]. 15th [Lor16d]. 17 [Gri14a].

18th [FR02]. 19-99 [Hal10b]. 1908-2008 [M^{+}08]. 1997 [Sch14]. 1997-2011 [Sch14]. 19th [GG10a, Cra07, FR02]. 1st [Fre15].

2010 [Lev12d]. 2011 [Lev14l, Sch14]. 2012 [Hun16, Lev14m]. 2013 [Lev15b]. 2013.ISBN [Lev14e]. 2014 [Ano14a]. 2015 [Ano14f, Ano15d, Lev17g]. 2016 [How16, Mac17, Lev18d, Tol15g]. 2017 [Gol17]. 2019 [Ano18f, Ano19c, Aul19, How19]. 2020 [Ano19d]. 2025 [Sto14c]. 20th [Lev15a]. 21st [Ano12d, Aul19, Rou13a]. 22-50 [De 11b]. 23-99 [Hal10b]. 24th [Kum11]. 2nd [Bay11a, Cou17b, Gib17a, Gri11d, Hal11c, Hud11a, Hun17d, Jon19, Lor14e, Shi10i, Shi10g, Sut10].

3-folds [Cor07]. 31st [Aul19]. 387 [App10b]. 3D [De 13]. 3rd [Bay14a, Gib11c, Gri17d, Mun12, Shi15c, Shi16b, Tol16c].

4 [Cri15a, Sut10]. 4-folds [Cor07]. 44-99 [Hai10]. 4th [Gri14e, Lor10e, Mac10c].

5 [Mac13b]. 5-99 [Hal10a]. 50 [Lev10a, War10a]. 55th [Ano14a]. 56-00 [Mah10b].

6 [Pea11, Hal13]. 61614-423-4 [Lev14h].

70-99 [Mun12].

8-99 [Bay10a, Wil10].

9 [Gri14b]. 9-789813-2026-10 [Lor18d]. 9-789814-651639 [Lor17f]. 90-50 [Gri14e]. 90th [Sut11, GDDR08]. 977-1-84724-008-8 [Lev10a]. 978-0-00-743808-2 [Hal13]. 978-0-00-747601-5 [Hal14]. 978-0-128-04153-6 [Hun18a]. 978-0-12804-317-2 [Tol18b]. 978-0-12814-759-7 [Hun19a]. 978-0-14-103023-4 [Lor10e]. 978-0-19-280722-9 [Wil10]. 978-0-19-531076-4 [Lev12a]. 978-0-19-536790-4 [War10b]. 978-0-19-856599-4 [Cou10]. 978-0-19-857061-5 [Rot10d]. 978-0-19-870259-7 [Rua17]. 978-0-19-875335-3 [Gib17b]. 978-0-19-875523-4 [Hop18b]. 978-0-19-879492-9 [Lor18f]. 978-0-19-880454-3 [Gib18b, McB18]. 978-0-19-911779-6 [Dav10]. 978-0-19-911780-2 [Rot10b]. 978-0-19-911781-9 [Gol10]. 978-0-19-911782-6 [Hal10c]. 978-0-19-913430-4 [Gri11d]. 978-0-19-921291-0 [Cou12]. 978-0-19-921986-5 [And10a, And10b]. 978-0-19-922730-3 [Shi10e, Shi10f]. 978-0-19-923769-2 [War11a]. 978-0-19-928201-2 [Mac10c]. 978-0-19-953987-1 [De 11b]. 978-0-19-953988-8 [De 11b]. 978-0-19-956207-7 [Haw11]. 978-0-19-956307-4 [Shi10g]. 978-0-19-956684-6 [And13c]. 978-0-19-957400-1 [Sch13].

978-0-19-958405-5 [Cri13]. 978-0-19-959968-4 [Lev14o]. 978-0-19-960139-4 [Lev13e]. 978-0-19-960542-2 [Lev14i]. 978-0-19-960587-3 [Gib13a]. 978-0-19-964008-9 [Bay14b]. 978-0-19-965444-4 [Mac16a]. 978-0-19-965659-2 [Lev14b]. 978-0-19-966132-9 [Fos14b]. 978-0-19-966437-5 [Mac15b]. 978-0-19-967085-7 [Mac14b]. 978-0-19-967086-4 [Mac14b]. 978-0-19-968197-6 [Lev14g]. 978-0-19-973259-3 [Fre15]. 978-0-19-974044-4 [Gri14c]. 978-0-19-975656-8 [Bev13, Kee18]. 978-0-19-989100-9 [And13b]. 978-0-19-993161-3 [Tol14b]. 978-0-195-36789-8 [Bay11b]. 978-0-19666-541-1 [Hun15]. 978-0-19870643-4 [Sto17b]. 978-0-19872-349-3 [Row19]. 978-0-19874-989-9 [Hun17f]. 978-0-19875-404-6 [Mac19]. 978-0-19875-547-0 [Bay19b]. 978-0-19878-828-7 [Tol18a]. 978-0-230-25208-0 [Gri13b, Mac13a]. 978-0-262-51429-3 [Haw12]. 978-0-273-72891-7 [Ten12]. 978-0-309-25473-1 [Rou13a]. 978-0-309-28457-8 [Sto14c]. 978-0-387-72475-1 [Lor10c]. 978-0-387-74640-1 [Lor11h]. 978-0-387-74646-3 [Bay11a]. 978-0-387-79714-4 [Gib10c]. 978-0-387-79851-6 [Hun12b]. 978-0-387-85528-8 [Shi10d]. 978-0-387-87858-4 [Hai12]. 978-0-387-89491-1 [Shi11d]. 978-0-393-07195-5 [Lev14f]. 978-0-444-63555-6 [Gib17a]. 978-0-465-00910-7 [Lev10f]. 978-0-465-01123-0 [Lev10d]. 978-0-465-02198-7 [Zet13]. 978-0-465-02331-8 [Lev14c]. 978-0-465-02837-5 [Mac14d]. 978-0-465-03292-1 [Lev13f]. 978-0-465-05074-1 [Lev14d]. 978-0-465-05497-8 [Tol15f]. 978-0-465-09377-9 [Lor18f]. 978-0-465-09463-9 [Tol19f]. 978-0-465-09612-1 [Dav18]. 978-0-46505-373-5 [Tol17b]. 978-0-46506-095-5 [Tol17e]. 978-0-46509-334-2 [Jac19]. 978-0-46509-758-6 [Mac18]. 978-0-470-14619-7 [Shi10c]. 978-0-470-23288-0 [App10b]. 978-0-470-74532-8 [Hal11a]. 978-0-470-87890-3 [Lor14e]. 978-0-486-44593-9 [Rua12]. 978-0-521-71978-0 [Lev10b]. 978-0-521-73524-7 [Sut11]. 978-0-521-73525-4 [Bay10a]. 978-0-521-75994-6 [War11b]. 978-0-674-04661-0 [Gib11b]. 978-0-674-05755-5 [Ban13]. 978-0-674-08891-7 [Tol17k]. 978-0-691-04955-7 [Shi10h]. 978-0-691-11713-3 [Dav17]. 978-0-691-11927-4 [Hoa17a]. 978-0-691-12067-6 [Cou11]. 978-0-691-12677-7 [Rot10c]. 978-0-691-12745-3 [Lev11d]. 978-0-691-12895-5 [For11]. 978-0-691-12956-3 [Hoa10]. 978-0-691-12973-0 [Rua11]. 978-0-691-13444-4 [Mac11b]. 978-0-691-13759-9 [Hud10b]. 978-0-691-14014-8 [Hud13a]. 978-0-691-14020-9 [And11]. 978-0-691-14175-6 [Cou18]. 978-0-691-14247-0 [Hun13d]. 978-0-691-14342-2 [Gri13f]. 978-0-691-14520-4 [And13d]. 978-0-691-14714-7 [Gri14b]. 978-0-691-14715-4 [Mac15c]. 978-0-691-14775-8 [Lev13a].

978-0-691-14841-0 [Lev12d]. 978-0-691-14854-0 [Ban14b]. 978-0-691-14892-2 [Rua14b]. 978-0-691-14905-9 [Hoa15b]. 978-0-691-15039-0 [Hun17e]. 978-0-691-15099-4 [Lev14a]. 978-0-691-15315-5 [Lev14l]. 978-0-691-15423-7 [Tol16a]. 978-0-691-15463-3 [Lor15g]. 978-0-691-15570-8 [Cri15b]. 978-0-691-15574-6 [Shi14b]. 978-0-691-15655-2 [Lev14m]. 978-0-691-15685-9 [Cri17a]. 978-0-691-15755-9 [Bay15b]. 978-0-691-16012-2 [Tol15b]. 978-0-691-16041-2 [Lev15b]. 978-0-691-16044-3 [Hoa14b]. 978-0-691-16060-3 [Tin17]. 978-0-691-16129-7 [Mac16b]. 978-0-691-16140-2 [Pas18]. 978-0-691-16173-0 [Lev17b]. 978-0-691-16244-7 [Von17]. 978-0-691-16350-5 [Tol15c]. 978-0-691-16751-0 [De 17]. 978-0-691-16767-1 [Tol18f]. 978-0-691-16965-2 [Lev17g]. 978-0-691-17019-0 [Shi17b, Tol17i]. 978-0-691-17168-5 [Lor17e]. 978-0-691-17486-0 [Tol17d]. 978-0-691-17529-4 [Lev18d]. 978-0-69115-866-2 [Hun18c]. 978-0-69116-828-9 [Tol17g]. 978-0-69116-863-0 [Haw18b]. 978-0-69117-171-5 [Yeo18b]. 978-0-69117-293-4 [Yeo17d]. 978-0-69117-416-7 [Tol18d]. 978-0-69117-690-1 [Tol19d]. 978-0-69117-879-0 [Bay19a]. 978-0-69118-285-8 [Hun19b]. 978-0-713-99757-6 [Rua10]. 978-0-8018-8714-7 [Hun10a]. 978-0-8018-9018-5 [Hun10b]. 978-0-8018-9299-8 [Gri11c]. 978-0-8176-4694-3 [Rua13]. 978-0-8218-4761-9 [Sto12c]. 978-0-8218-4807-4 [Gri11e]. 978-0-8218-4808-1 [Lor11f]. 978-0-8218-4814-2 [Gri12f]. 978-0-8218-4817-3 [Lor11g]. 978-0-8218-4873-9 [Lev11b]. 978-0-8218-4928-6 [Gri11b]. 978-0-8218-4999-6 [Hun14a, Lor14f]. 978-0-8229-4418-8 [Lev14n]. 978-0-8264-9606-5 [Zar10b]. 978-0-85709-223-6 [Gib11c]. 978-0-85709-224-3 [Gib11c]. 978-0-85729-928-4 [Gib13b]. 978-0-88385-580-5 [Haw18a]. 978-0-88585-790-8 [Lor18e]. 978-0-888385-562-1 [Eat11]. 978-0-90658-884-0 [Hal17b]. 978-0-90658-891-8 [Lev18c]. 978-0-906588-68-0 [Lor11i]. 978-0-906588-71-0 [Bay12]. 978-0-906588-78-9 [Lor14d]. 978-0-906588-79-6 [Tol15e]. 978-0-906588-80-2 [Tol15e]. 978-0-906588-82-6 [Lev18b]. 978-0-906588-97-0 [Jon19]. 978-0-907944-11-9 [Ban10c]. 978-0-955-59142-6 [Yeo18a]. 978-0-9555477-1-3 [For10a]. 978-0-9555477-3-7 [Tol13]. 978-0-9558042-3-6 [Lev14j]. 978-0-961-13540-3 [Hun12a]. 978-0-965706-3 [App14a]. 978-0-965707-0 [App14a]. 978-0-9799269-0-7 [Lea10]. 978-0-98856-225-7 [Yeo16]. 978-0-9926294-0-3 [Cri15a]. 978-0-99906-350-7 [Zet19]. 978-019-966132-9 [Sto14a]. 978-1- [Lev14h]. 978-1-10-00266-1 [Hoa15a]. 978-1-107-03948-3 [Sto17a]. 978-1-107-06392-1 [Tol15a]. 978-1-107-10381-8 [Shi18a]. 978-1-107-11850-8 [Bev17]. 978-1-107-44453-9 [Tol16c]. 978-1-107-49499-2 [Tol16e].

978-1-107-50049-5 [Cou17b]. 978-1-107-61418-5 [Bay15a]. 978-1-107-64252-2 [Tol16b]. 978-1-107-65848-6 [Tol16d]. 978-1-107-66330-5 [Bay15a]. 978-1-107-66854-6 [Hal16]. 978-1-107-67532-2 [Bay15a]. 978-1-107-67866-8 [Lee17]. 978-1-107-69091-2 [Mac15a]. 978-1-10712-844-6 [Cam18]. 978-1-107142774-7 [Lev17a]. 978-1-10762-944-8 [Hun18b]. 978-1-10840-630-7 [Tol19e]. 978-1-10841-641-2 [Cou19]. 978-1-118-11775-0 [Mac14c]. 978-1-118-11776-7 [Mac14c]. 978-1-118-13535-8 [Gri14e]. 978-1-118-22962-0 [Hun13a]. 978-1-119-12792-5 [Hai17a]. 978-1-13803-016-9 [Tol19b]. 978-1-200-6681-4 [App10a]. 978-1-316-60589-9 [Tol17j]. 978-1-316-60788-6 [Hop19]. 978-1-31651-896-0 [Tol19c]. 978-1-31661-483-9 [Tol19g]. 978-1-4092-5670-0 [Ban10b]. 978-1-421-41565-9 [Mac16c]. 978-1-421-42035-6 [Had19]. 978-1-4214-0315-1 [Hai13]. 978-1-4214-0438-7 [And13a]. 978-1-4214-0484-4 [Lev13d]. 978-1-4214-0637-4 [De 14b]. 978-1-4214-0724-1 [De 14b]. 978-1-4214-1098-2 [Tol15d]. 978-1-4214-1176-7 [Shi15b]. 978-1-4214-1380-8 [Lev15a]. 978-1-4214-1442-3 [Dav15]. 978-1-4214-2230-5 [Shi18b]. 978-1-42141-918-3 [Tol17f]. 978-1-42142-267-1 [Zet18]. 978-1-4292-1623-4 [Zar11]. 978-1-4363-0180-0 [Hal11c]. 978-1-4363-3416-7 [Hud10a]. 978-1-4363-3419-8 [War13a]. 978-1-4398-1624-0 [Mun12]. 978-1-4398-4896-8 [Shi13c]. 978-1-4398-5336-8 [Cou13]. 978-1-4398-6337-4 [Bev14, Rua14a]. 978-1-4398-9834-5 [Bay14a]. 978-1-4411-7975-3 [Sto14b]. 978-1-4419-0252-8 [Hai10]. 978-1-4419-1595-5 [Meg12]. 978-1-4419-1731-7 [Shi12c]. 978-1-4419-2313-4 [Hun17b]. 978-1-4419-6949-1 [App14b]. 978-1-4419-9499-8 [Gri13a]. 978-1-4471-2172-5 [Shi14c]. 978-1-4471-4615-5 [Pas14]. 978-1-4471-6418-0 [Shi16b]. 978-1-4538-1057-6 [Rua14c]. 978-1-4614-9592-5 [Shi15c]. 978-1-4665-0155-3 [Mac13d]. 978-1-4665-0429-5 [Gib14b]. 978-1-4665-5464-1 [Shi14e]. 978-1-4704-4828-8 [Gib19b]. 978-1-47043-584-4 [Tol18c]. 978-1-4729-3999-9 [Gib19c]. 978-1-5416-7248-2 [Hoa19]. 978-1-56881-220-5 [Eat12]. 978-1-56881-245-8 [Sut11]. 978-1-56881-316-5 [LP10]. 978-1-56881-331-8 [For10b]. 978-1-56881-340-0 [Hud11a]. 978-1-56881-341-7 [Lev10e]. 978-1-56881-348-6 [Hud10c]. 978-1-56881-410-0 [Ban10a]. 978-1-56881-412-4 [Shi10i]. 978-1-56881-430-8 [Gri10b]. 978-1-56881-435-3 [For13]. 978-1-56881-436-0 [Mac13e]. 978-1-56881-441-4 [Gib10a]. 978-1-56881-444-5 [Lev10e]. 978-1-56881-449-0 [Gib10b]. 978-1-56881-456-8 [Hun13c]. 978-1-56881-457-5 [Mac13c]. 978-1-56881-466-7 [Hoa12]. 978-1-56881-579-4 [For12]. 978-1-56881-584-8 [Nai12].

978-1-56881-710-1 [War13b, War13c]. 978-1-56881-714-9 [Mac13b]. 978-1-56881-721-7 [Gib14c]. 978-1-58488-793-5 [War10c]. 978-1-59691-522-0 [Lev13b]. 978-1-61614-421-0 [Lev13c]. 978-1-61614-546-0 [Lev17c]. 978-1-61614-587-3 [Lev14p]. 978-1-61614-745-7 [Mac14a]. 978-1-61614-747-1 [Lev14e]. 978-1-62040-593-2 [Gri17c]. 978-1-63388-148-8 [Lor17g]. 978-1-63388-167-9 [Lev17h]. 978-1-63388-297-3 [Tol18e]. 978-1-781-25030-3 [Tol17c]. 978-1-781-25288-8 [Hem19]. 978-1-78240-471-2 [And19]. 978-1-786-34184-6 [Gib18a]. 978-1-78607-264-1 [Haw18c]. 978-1-78634-068-9 [Hal17a]. 978-1-84354-525-5 [Hud11b]. 978-1-84419-261-8 [Sto12b]. 978-1-84419-262-5 [Sto12a]. 978-1-84419-269-4 [Sto12b]. 978-1-84419-286-1 [Hal12b]. 978-1-84419-287-8 [Hal12b]. 978-1-844191-91-8 [Hal10a]. 978-1-844192-39-7 [Hal11b]. 978-1-84628-790-9 [GG10a]. 978-1-84668-064-9 [Tol10]. 978-1-84668-198-1 [Gri14f]. 978-1-84668-1998 [Bay14d]. 978-1-84668-292-6 [Lev11c]. 978-1-84668-531-6 [Gri14a]. 978-1-84724-008-8 [War10a]. 978-1-84800-987-5 [Shi11e]. 978-1-84831-238-8 [Gri13c]. 978-1-84882-532-1 [App11]. 978-1-84916-240-1 [Lev11f]. 978-1-84931-066-6 [Hoa14a]. 978-1-84996-485-2 [Gri13d]. 978-1-85539-147-5 [Sto10b]. 978-1-87642-075-8 [Lev17e]. 978-1-90600-1193 [Rou14]. 978-1-90600-124-7 [Gri18d]. 978-1-90600-129-2 [Mor19]. 978-1-90600-130-8 [Tsa18]. 978-1-906001-03-2 [Zar10a]. 978-1-906001-04-9 [Bay10b]. 978-1-906001-06-3 [Sto10a]. 978-1-906001-11-7 [Mac12b]. 978-1-906001-16-2 [Hun13b]. 978-1-906001-17-9 [Fox15]. 978-1-906338-01-5 [Lev11e]. 978-1-909-98664-2 [Hor17]. 978-1-911616-03-0 [Lev19b]. 978-1-92197-231-7 [Pat17]. 978-1-92197-232-4 [Pat17]. 978-1-92197-233-1 [Pat17]. 978-1-92197-234-8 [Pat17]. 978-1-92197-235-5 [Pat17]. 978-1-92197-242-3 [Hal15a, Hal15b]. 978-1-935638-02-5 [Lev12c]. 978-1-93813-486-9 [Bay17d]. 978-14214-1728-9 [Bay17b]. 978-2-34001-108-3 [Yeo17a]. 978-2-34001-109-0 [Yeo17a]. 978-2-7298-86738 [Lor15i]. 978-3-0346-0473-4 [Gib12b]. 978-3-0346-0536-6 [Jac14]. 978-3-0348-0229-1 [De 14c]. 978-3-0348-0381-6 [Sch14]. 978-3-0348-0553-7 [Tol14a]. 978-3-0348-0693-0 [Sch16b]. 978-3-03719-104-0 [Hoa13a]. 978-3-319-02044-0 [Lor15h]. 978-3-319-04860-4 [Hai15]. 978-3-319-05653-1 [Lor16b]. 978-3-319-11945-8 [Rou16]. 978-3-319-14426-9 [Lor16c]. 978-3-319-20435-2 [Hun17c]. 978-3-319-25566-8 [Tol17h]. 978-3-319-28339-5 [Hai17b]. 978-3-319-39374-2 [Tol19a]. 978-3-319-90319-4 [Hal19]. 978-3-31927-806-3 [Bay17a]. 978-3-31929-044-7 [Bay17c]. 978-3-31978-630-8 [Shi19c].

978-3-527-40758-8 [Mac10a]. 978-3-540-70996-1 [Lev12b]. 978-3-642-01372-0 [Hoa11]. 978-3-642-03544-9 [Lor13d]. 978-3-642-04832-6 [War12a]. 978-3-642-05091-6 [Lev11a]. 978-3-642-11671-1 [Gib12a]. 978-3-642-13605-4 [GG12c]. 978-3-642-14033-4 [Eat13]. 978-3-642-21146-1 [Hun14b]. 978-3-642-22103-3 [Ban14a]. 978-3-642-22463-8 [Tol14c]. 978-3-642-22464-5 [Tol14c]. 978-3-642-22557-4 [Lev13g]. 978-3-642-23791-1 [Shi14d]. 978-3-642-23847-5 [Gri14d]. 978-3-642-38895-8 [Cou15, Cou17a]. 978-3-642-39448-5 [Hun16]. 978-3-642-54074-5 [Sch16c]. 978-3-7643-7538-6 [De 10b]. 978-3-7643-8332-9 [De 10a]. 978-4-431-55841-5 [Fox17]. 978-81-322-1598-1 [Shi15a]. 978-88-12-00015-9 [GG10b]. 978-88385-577-5 [Lev14k]. 978-9-481474-043-2 [Lev17d]. 978-9-81314-364-7 [Yeo18c]. 978-9-81314-623-5 [Shi17a]. 978-9-81314-634-1 [Hal17d]. 978-9-81314-754-6 [Tol17a]. 978-9-81314-802-4 [App18]. 978-9-81314-884-0 [Gri18c]. 978-9-81320-311-2 [Lev17i]. 978-9-814-72581-1 [Mac17]. 978-9-81429-149-1 [Hop18a]. 978-9-81466-746-3 [Lev17f]. 978-9-81469-645-6 [Yeo17b]. 978-9-81472-572-9 [Hoa18a]. 978-9-814723-375-5 [Gib17c]. 978-9-81473-003-7 [Yeo17c]. 978-93-80250-17-5 [Cou14]. 978-981-3222-92-2 [Gri18e]. 978-98128-3541-3 [Hal17c]. 9780-691-13994-4 [Hud13b]. 981-04-2759-X [Lev11d]. 987-0-19-958848-0 [Bay14c]. 987-1-84837-902-2 [Lev13h]. 987-9814-68909-0 [Hun17a].

= [CF15, Den14].

A* [Sto12b, SD10]. A-level [Lor11i, Nic08a, Tol14d]. A. [Dou13]. A4 [Lor17h]. A4-sized [Lor17h]. Aassila [Yeo17a]. Åstrom [Mac10b]. Abbott [Lev10d, War11b]. Abel [Hoa11, Hun16, HP10b]. ability [Hal17d]. above [Yeo17c]. Abstract [FGR14, Gri14e, Nic12a, Shi15b]. abused [SC13, Lev13f]. Academic [Hun18a, Hun19a, Tol18b, BEU12, Tol14c]. Academies [Rou13a, Sto14c]. Accelerating [Mur11a]. acceleration [Jew10, LN17, Lor15k, Sim11c]. Acheson [Gib18b, McB18]. Acknowledgements [Ano11a, Ano12a, Ano14b, Ano18a, Ano19a]. across [And13a, Cri10, Swe12]. activities [BCSS19, Leu18]. Aczel [Tol17h]. ad [Sco18c]. Adam [For11]. Adams [Hal11c, Hud10a, Lor11g, War13a]. add [Hop16]. addendum [Raj18]. addition [Gau19]. Additional [HHHS14, Hal15a, Hal15b]. additive [Hop13]. address [Gol17]. Adhikari [Shi15a]. adic [Hun14b, Sch11]. Adler [Mac15c]. Adrian [Lev13e]. Advanced [Duf11, Raj13, Ric08a, Bay14a, Fyf11, JM09a, JM09b, Kra12, Lev12a, Tol13, App10b, Bay17a, Mac14b, Mun12]. adventure [McB18]. Adventures [Rop18, Gla19, Lev18b, Mic09, Gri11c]. affect [Min12]. affirmative [Dol12a]. Africa [Cri17a]. after [PdVF14]. Afzal [Sto10b]. again [Sco17a]. Agency [Cou14]. ages [Swe12, And13a]. Agnes [For12]. Agnijo [Haw18c]. agony [Wil08]. Ahmed [Sto10b]. air [Lor19d]. AK [Lev10e, Shi14e]. Akemann [Sch13]. Akiyama [Fox17]. al [Hal17c]. Alain [Mac19]. Alan [Bev17, Fre10a, Yeo17c, Ano14p, AF12, How16, Shi14b]. Albers [Hud11a]. Alberti [WMW10, Gib12b]. Alberto [Lev14n]. Alcock [Fos14b, Sto14a]. Alessandro [Tol19a]. Alessio [Rot10d]. Alex [Hun14a, Lev12a]. Alexander [And13d, Bay11a, De 11b, Gib11b, Lor11h, Shi18b]. Alexanderson [Hud11a]. Alexandre [Sto12c]. Alfred [Hal17d, Lev14e, Lev14h, Lev14p, Lev17h, Tol18e]. Algebra [FGR14, HF04, AA13, ADLM18, Ano13m, Ban16, Dav12, FH05, Gri14e, Hoa15b, KP14, Kem11, Mac13d, Mac16a, Mah10b, MP11, Nic12a, Pat17, Qua12a, Ran18, Rou16, San15, Shi15a, Shi15b, Ten08, Tol15e, Rib18, Gri10b, Lor10d, Lor13d, Lor16c, Shi18a, Tol19c, Bay17b]. algebraic [Cam18, JJ13b, Pla12]. algebras [Hun17b, Kan09a, Lor10c]. algèbre [Yeo17a]. Algorithm [LP16, OW14, OJN16]. Algorithmic [LL11, Gri14c]. Algorithms [Orm14, Abe16, BC13, Gri14b, Hir11h, Mac12a, Cou15, Cou17a]. Alice [De 10a, Gla19]. aliens [Adl14, Mac15c]. alignment [Nic12b]. Allen [Rua10]. Almanac [War12b]. almost [Ben12a]. alphabets [Lew14]. also [Bay17a, Bay17c, Dav18, Gib17c, Hai17a, Hai17b, Hun16, Hun17b, Hun17c, Mac17, Sch16b, Sch16c, Shi15c, Shi15a, Shi16b, Tol17h, Tol18b, Tol18e, Tol19a, Tol19f]. Alteration [Nur13]. alternating [Coo11a, GB12, Mur11a, Nur13, Nur14a, Pla18]. alternation [HK14]. alternative [Cer16, Dou13, Kan17, Lor18b, Ste17f]. altitude [Jos16a]. altitudes [ASHH18]. AM [BS12a, Shi11a]. ambitious [Kör14, Tol15a]. America [Bay10c, Eat11, Lev14k, Tol18c]. American [Gib19b, Gri11e, Gri11b, Gri12f, Hun14a, Hun17d, Lev11b, Lor11g, Lor11f, Lor14f, Sto12c]. Amir [Gib11b, Tol17h]. Amol [Lev17d]. amongst [Gri18b]. Amrein [App10a]. AMS [Tol18c]. AMS/Mathematical [Tol18c]. Amy [Bay17b, Eat11]. analogues [Ste18a]. analyse [Yeo17a]. Analysis [Gar13, App11, App12, App14a, Bay14b, For12, KS14, Lor14e, Mac09, MD10, ML13, Rib19, Ric08a, Shi19c, Sta11, VV09, Yeo17d, App10b, Bay15a, Bay19a, Cou17b, Hun17a, Lor18d, Rou16, Shi10c, Shi10d, Tol16c]. Anamorphoscopes [SNH11]. Anany [Gri14c]. ancient [Dav17, Gib14a, Gra13, Lan12, Lev14b, Mac13e, Rei14, WW13, YM11, Rua13]. Andreas [De 10a]. Andreescu [Lea10]. Andrew [Mac15b, Mor19, Rou14, Shi14b, Tol13, Tsa18]. Andrews [Sto14b]. Angelo [GG12c, Tol19a]. angle [BCL10, Chi17, CR14, Job18b, Kos16, Leu17, Shi11c, WO11]. angle-bisector [CR14]. angled [Leu12a, Osl11a]. angles [Nic14b, O’L10, Sub15b]. Annette [Dav17]. annotated [Abb08, Lev10d]. annular [Che16a]. Answering [Ano06]. anthology [War12c]. Anthony [Shi15b, Tol15d]. Anthropologic [BC15]. antiquity [Hoa15b, KP14]. Antonio [Rou16]. anxieties [Pas11b]. any [CO11, Koo11]. anywhere [Jam13b]. AP [Sco18d]. Apéry [Mel13]. apologies [Tol16a]. Appeal [Lev18a]. appearance [ASHH18, Hir11d, Lor17a]. Appel [Shi14b]. Applebaum [Bay14b]. application [Gut11, Haj17a, Lor11b, Luk18, PG17, Rua14c]. Applications [Mol12, AA13, Dav12, Hal11c, Kan09b, Kra15, Mac14c, RJPG12, R^{+}12, Cou19, Gib17a, Mac13d, Shi11d, Shi15a, Shi18b, Shi19c, Bev14, Rua14a]. Applied [Cou12, Hun13a, Hun17e, Hun18a, Reb12, Tol19c, Mac19]. appraisal [Sta19b]. Appreciation [Lev19a]. approach [Dou13, Egg15, Gau19, Jac11, JO14, Lev12f, Lor14e, Mer19, Mil12, Ran18, Rza10, Sta11, TL16, Tra13]. approaches [Gri12e, Lor12a, Cam18]. Approaching [Leu17]. Approximate [DL14, Szy17]. approximating [Che18]. approximation [Gib14a, Jam14a, PG17, SS12, VSR13]. approximations [SB19, Mac10d, Ste12b]. April [Gol17, How19]. AQA [Gri11d, Hal12b, Ano10a, Kra09]. Arabic [Lo 13]. Arbelos [For10a, Tol13, Job16, Oku13a, Oku13b, Sci19, Shi11b]. arc [Gor18, Osl16, Shi10b]. Archimedean [Die19, Oku13b, OW14, OJN16]. Archimedes [Hoa18c, Kir13, Shi11b]. Arcturus [Lev13h]. Area [Jos19c, Spo18, Bea17b, CF15, Jos15, Jos17a, Jos19d, KN19, Lor12g, Lor13b, Lor17d, Lor18b, Oku13a, Sil17, Lev18b]. areal [Sco13a, Sco18b]. areals [Smi16]. Areas [Lor15f, Sew14, Lor12b]. argument [Gla19, Mar11]. arguments [Tre17]. Arianrhod [Tol14b]. Arild [Gib12a]. arise [Sci19]. arising [Gri15g, Luk19]. Arithmetic [Shi10a, AM11, Dol19a, Gri12g, Jam14a, Jam18b, Pur19, OC14, Shi12c]. Arlen [Bay17c]. Arnold [Raj13]. Art [Jac19, Gri11c, Haw12, Hoa19, Lan12, Mah10a, Mic09, Rua14b, Tub14, Van13, Lev15a, Mac13e]. artful [Lev17b]. Arthur [How15]. article [Bur10a]. arts [BY11]. Artstein [Lev17c]. Arturo [Shi10h]. Artzt [Sha15]. Arzarello [GG10b]. Ash [Shi17b, Tol15c, Tol17i]. Ashby [Kil10]. Ashish [Hun15]. asinorum} [Haj15a]. ask [Lev18c]. Askar [Hun15]. Asok [Gri18e]. aspects [Bor10b, Sto12c]. ASS [Haj10]. Assessing [Fre10a]. Assessment [BHJ10, Hal11a]. assistant [Fie07, Zar10b]. associated [EF10, Gri17a, Kri18, Mah10c, Oai15]. Association [Ano14m, Bay10c, Bay12, Eat11, Gol17, Hal10b, Hal17b, Jon19, Lev14k, Lev18c, Lev18b, Lev19b, Lor11i, Lor14d, Tol15e, Tol15g, Tol18c, Ano17m, Ano17n, Ano17o, Ano18n, Ano18o, Ano18m, PW13, Tol15g]. associativity [Bur17, Mac14e]. asymmetric [Sch16a]. asymptotes [Mit12, Sco15b]. Atiyah [How19]. Atlantic [Hud11b]. attached [Ada10, War13a]. attractive [CK19]. attractors [GG08c, Lev10e]. Attridge [Hal17a]. AU$41.00 [Pat17]. AU$50.00 [Hal15a, Hal15b]. AU$54.50 [Pat17]. Aubel [OS15]. AUD$86.32 [Cou17b]. Audin [Gib13b]. audited [San15]. August [Aul19]. Austen [Von17]. Australian [Lev17e]. author [Dol15b, Hal13]. Automata [CSC10, Eat13]. Available [Gib11a, Bay17a, Bay17c, Dav18, Dol15b, Gib17c, Hai17a, Hai17b, Hun16, Hun17b, Hun17c, Mac17, Sch16b, Sch16c, Shi15c, Shi15a, Shi16b, Tol17h, Tol18b, Tol18e, Tol19a, Tol19f, Yeo18a]. average [Edg16]. averages [Lor10b, Lor11a, Lor13c]. Avishek [Shi15a]. Avner [Shi17b, Tol15c, Tol17i]. avoid [Bay15c]. Axes [Far18]. axioms [ADLM18, Cul11].

B [Bay17b, Gri10b, Lor16c, Lor18e, Rua13, Rua14c]. Bach [And17]. Back [Ano10h, Ano10j, Ano10l, Ano11g, Ano11i, Ano11k, Ano12g, Ano12i, Ano12k, Ano13f, Ano13h, Ano13j, Ano14g, Ano14i, Ano14k, Ano15e, Ano15g, Ano15i, Ano16c, Ano17f, Ano17g, Ano17i, Ano17k, Ano18g, Ano18i, Ano18k, Ano19f, Ano19h, Ano19j, AW07, Sto10b]. backbone [Bay10b, Gri08]. background [Shi18b]. backward [McC15]. Badger [dV14]. Baggott [And13c]. Baik [Sch13]. bake [Hem19]. Balaji [Cou14]. Balkan [Yeo16]. Ball [Gib19c, Wap12, Hol17]. ballet [And13b, Ben12b]. Balls [Sim11a]. Baltimore [De 14b]. Banach [Lor10c, Kan09a]. Banchof [War11b]. Banchoff [Hun13c]. Banerjee [Haw18c]. Barbara [Shi10d]. Barbeau [Haw18a]. Barel [Hun10a]. Barnabas [Cri17a]. Barnaby [Lee17]. Barnard [Tol19b]. Barnes [Lev11a]. Bartocci [GG12c]. barycentre [Abe19]. based [And17, Hos17b, McG13, Rza10]. Basel [De 10a, Lor14c, Pat13b, Rib19]. Basic [AA13, Dav18, Haw18c, Hoa19, Jac19, Lev10d, Lev10f, Lev11c, Lev13f, Lev14c, Lev14d, Lor18f, Mac14d, Mac18, Tol10, Tol15f, Tol17b, Tol17e, Tol19f, Zet13, Hal11c, Shi15a, Tol17a]. basics [Shi18b]. Bataille [Lor15i]. Battista [Gib12b, WMW10]. Bauer [Tol18f]. Bayesian [Mar16]. be [CT14, De 10a, HHPM07, Hop12, Osl12a, Shi18c, Shi19d, Sim11a]. Beardon [Ano14p]. Beatty [Gri18b]. Beautiful [MJ14, Lev14a]. beauty [And19, Ban14a, Fre15, Gri13d, Her11, LG10, Pas11b, Sen12]. Becheanu [Yeo16]. Before [BT09, War10b]. beginner [Hoa18a]. beginners [RJPG12, R^{+}12]. behave [Di 13]. behaviour [LY11]. behind [Bev13, RT11]. being [Tay07]. Belcastro [For10b, For13]. Bello [Tol15d]. belong [Fra19a]. bend [Kuc12]. Benford [Zor11]. Benjamin [Hud13a, Lev13a, Lev14i, Shi15b, Pas08, Hoa10]. Benson [And13b, Lev10c]. Bent [Lor11g, Ada09b]. Bercovici [Bay17c]. Berger [Lev12b]. Berggren [Cri17a]. Bergmann [Tol14c]. Bernoulli [Jam19b, Vîj19]. Berry [Hal14]. Besicovitch [Shi10b]. best [Cri12b, Lev12d, Lev14l, Lev14m, Lev15b, Lev17g, Lev18d, Pit11, Pit12, PM13, Pit14]. better [AW07, Sto10b]. Betti [GG12c]. between [Bau13, Eul15, ES16, Jam19a, Koo11, Lor11j, Rea10b, Shi13d, Ste17a, Sub19]. Bevan [Jon19]. Beyond [FMS10, Jam18c, Lor14c, Gib19c, Lev17h]. Bezuszka [Gri17d]. bicentric [Jos15]. Bidgood [Hal11a]. big [FF19, Lev11f]. bigger [HK14]. Biggs [Gib17b]. bilateral [WO13]. Bill [Ano17b]. Bin [Yeo17b]. Binomial [Lor10b, Lor11e, Lor13c, Den10, GB12, HP10a, Kos11, Rob10, Rob10]. biographies [Ram09]. biological [Cou13, Sma12]. Biology [Cou13]. birds [JJ17]. Birgit [Tol14c]. Birkhauser [De 14c, Hai17b, Jac14, Rou16, De 10b, Gib12b, Rua13, Sch14, Sch16b, Tol14a, De 10a]. birth [Ano17b, GV17, Ric08b]. birthday [GDDR08, HB16, Sut11]. bisect [Jos17e]. bisect-diagonal [Jos17e]. Bisecting [Osl16]. bisector [CR14]. black [Sim18]. Blackjack [Wer09, Hai10]. Block [War10b]. Blocks [McL18].

Bloomsbury [Gib19c, Gri17c, Lev13b, Sto14b]. blunder [Smy10]. Blythe [Pat17]. body [Sto17a]. Bogdan [Yeo16]. Bombelli [Tra13]. Bóna [Gri18c]. Bondi [Cla18]. Book [And10a, And10b, And11, And13c, And13a, And13b, And13d, And19, App10b, App10a, App11, App14b, App14a, App18, Ban10a, Ban10c, Ban10b, Ban13, Ban14b, Ban14a, Bay10c, Bay10b, Bay10a, Bay11b, Bay11a, Bay12, Bay14a, Bay14b, Bay14c, Bay14d, Bay15b, Bay15a, Bay17b, Bay17a, Bay17c, Bay17d, Bay19b, Bay19a, Bev13, Bev14, Bev17, Cam18, Cou10, Cou11, Cou12, Cou13, Cou14, Cou15, Cou17a, Cou17b, Cou18, Cou19, Cri13, Cri15b, Cri15a, Cri17a, Dav10, Dav15, Dav17, Dav18, De 10b, De 10a, De 11b, De 14c, De 14b, De 17, Eat11, Eat12, Eat13, For10a, For10b, For11, For12, For13, Fos14b, Fox15, Fox17, Fre10a, Fre10b, Fre15, Gib10a, Gib10c, Gib10b, Gib11b, Gib11c, Gib11a, Gib12b, Gib12a, Gib13b, Gib13a, Gib14b, Gib14c, Gib17a, Gib17c, Gib17b]. Book [Gib18a, Gib18b, Gib19b, Gib19c, Gol10, GG10a, GG12c, Gri10b, Gri11d, Gri11e, Gri11b, Gri11c, Gri12f, Gri13a, Gri13b, Gri13d, Gri13c, Gri13f, Gri14f, Gri14a, Gri14e, Gri14b, Gri14c, Gri14d, Gri17c, Gri17d, Gri18d, Gri18e, Gri18c, Had19, Hai10, Hai12, Hai13, Hai15, Hai17a, Hai17b, Hal10c, Hal10a, Hal10b, Hal11b, Hal11c, Hal11a, Hal12b, Hal13, Hal14, Hal15a, Hal15b, Hal16, Hal17a, Hal17b, Hal17c, Hal17d, Hal19, Haw11, Haw12, Haw18b, Haw18a, Haw18c, Hem19, Hoa10, Hoa11, Hoa12, Hoa13a, Hoa14a, Hoa14b, Hoa15a, Hoa15b, Hoa17a, Hoa18a, Hoa19, Hop18a, Hop18b, Hop19, Hor17, Hud10c, Hud10b, Hud10a, Hud11a, Hud11b, Hud13a, Hud13b, Hun10b, Hun10a, Hun12b, Hun12a, Hun13d, Hun13c, Hun13b, Hun13a, Hun14b, Hun14a, Hun15, Hun16, Hun17a, Hun17b, Hun17c]. Book [Hun17d, Hun17e, Hun18b, Hun18a, Hun18c, Hun19b, Hun19a, Hun17f, Jac14, Jac19, Jon19, Kee18, Kil10, LP10, Lea10, Lee17, Lev10d, Lev10f, Lev10e, Lev10a, Lev10b, Lev10c, Lev11b, Lev11d, Lev11a, Lev11e, Lev11c, Lev11f, Lev12a, Lev12d, Lev12b, Lev12c, Lev13e, Lev13c, Lev13b, Lev13h, Lev13a, Lev13g, Lev13f, Lev13d, Lev14e, Lev14i, Lev14n, Lev14o, Lev14k, Lev14l, Lev14j, Lev14c, Lev14h, Lev14p, Lev14b, Lev14m, Lev14a, Lev14d, Lev14f, Lev14g, Lev15b, Lev15a, Lev17g, Lev17h, Lev17a, Lev17b, Lev17c, Lev17e, Lev17d, Lev17f, Lev17i, Lev18d, Lev18c, Lev18b, Lev19b, Lor10e, Lor10d, Lor10c, Lor11g, Lor11h, Lor11f, Lor11i, Lor13d, Lor14f, Lor14e, Lor14d, Lor15g, Lor15i, Lor15h, Lor16c, Lor16b, Lor17g, Lor17e, Lor17f, Lor18f, Lor18d, Lor18e, Mac10b]. Book [Mac10c, Mac10a, Mac11b, Mac12b, Mac13e, Mac13a, Mac13c, Mac13b, Mac13d, Mac14b, Mac14c, Mac14d, Mac14a, Mac15a, Mac15b, Mac15c, Mac19, Mac16a, Mac16b, Mac16c, Mac17, Mac18, Mah10b, McB18, Meg12, Mor19, Mun12, Nai12, Pas14, Pas18, Pat17, Rot10c, Rot10d, Rot10b, Rot10a, Rou13a, Rou14, Rou16, Row19, Rua10, Rua11, Rua12, Rua13, Rua14c, Rua14a, Rua14b, Rua17, Sch10, Sch13, Sch14, Sch16b, Sch16c, Shi10e, Shi10f, Shi10c, Shi10i, Shi10g, Shi10d, Shi10h, Shi11e, Shi11d, Shi12c, Shi13c, Shi14d, Shi14b, Shi14c, Shi14e, Shi15c, Shi15a, Shi15b, Shi16b, Shi17a, Shi17b, Shi18b, Shi18a, Shi19c, Sto10b, Sto10a, Sto12c, Sto12b, Sto12a, Sto14a, Sto14b, Sto14c, Sto17b, Sto17a, Sut10, Sut11, Ten12, Tin17, Tol10, Tol13, Tol14b, Tol14a, Tol14c, Tol15b, Tol15e, Tol15c]. Book [Tol15g, Tol15a, Tol15d, Tol15f, Tol16a, Tol16d, Tol16c, Tol16b, Tol16e, Tol17a, Tol17b, Tol17c, Tol17d, Tol17e, Tol17g, Tol17f, Tol17h, Tol17i, Tol17k, Tol17j, Tol18a, Tol18b, Tol18e, Tol18c, Tol18d, Tol18f, Tol19a, Tol19d, Tol19g, Tol19b, Tol19c, Tol19f, Tol19e, Tsa18, Von17, War10a, War10c, War10b, War11a, War11b, War12a, War13a, War13b, War13c, Wil10, Yeo16, Yeo17a, Yeo17b, Yeo17c, Yeo17d, Yeo18a, Yeo18b, Yeo18c, Zar10b, Zar10a, Zar11, Zet13, Zet18, Zet19, dV14, AD08, GG08a, Soi09a, Lea10]. Bookguild [Hor17]. Books [Ano14c, Ano16a, Ano17a, Bay14d, Dav18, Gri13c, Haw18c, Hem19, Hoa19, Hud11b, Jac19, Lev10d, Lev10f, Lev11c, Lev13c, Lev13f, Lev14e, Lev14c, Lev14h, Lev14p, Lev14d, Lev17h, Lev17c, Lor17g, Lor18f, Mac14d, Mac14a, Mac18, Tol10, Tol15f, Tol17b, Tol17c, Tol17e, Tol18e, Tol19f, Zet13]. booster [Hal13, Pea11]. bootstrapping [Shi11a]. Borel [Col19]. born [CT14]. Borovik [Sto12c]. Borwein [Ban10a]. Botelho [Shi19c]. both [Gri16b]. bound [Nic11, SS11]. boundaries [AZY12, Per12]. boundary [Job13a]. bounds [Gus15, KK10, Sch16c, Tor17]. bowls [HI14]. box [Sim16]. Boyd [Tol19c]. boys [CT14]. Bradley [Mac12b]. brain [Grä11, War13b, War13c]. brainteaser [Ros09, Bay11b]. Brams [Hud10c]. Breach [Gri13b, Mac13a]. breaking [Yeo18b]. breakthrough [Ges11, Gri13c]. Brian [Fre10b, Gol10, Hun17b, Rot10b, Tol18d]. Bride [LRQ10]. bridge [Lev08, Sto10a]. brief [Gib19c, Rua17]. Brilliance [Kan17]. bringing [Hol17]. Britain [FRW11, War12b, Lev13e]. British [GG10a, Cra07]. Brouncker [Hir11e, Osl11b, Osl12b]. Brouwer [Pas14, vDD13]. Brown [Bay17c]. Bruce [Lev17f, Zar11]. Brummelen [Rua11, Rua14b]. Bryan [Dol15b]. Bug [Tal13]. buildings [Hah12, And13d]. Bulajich [Rou16]. bundles [Cou14, Tau11, Gib13a]. Burgiel [Eat12]. Bürgisser [Cou15, Cou17a]. Burkhard [Lev13d]. Burn [Tol16c, Yeo18a, Tol17b]. Burness [Hun18b]. Burning [Vil11]. Burns [dV14]. Burton [Ten12]. Butler [Hop19]. bytes [Tin17]. byways [Eat11].

C [And13b, Bay19a, Cou14, Hoa10, Hun17b, Hun18b, Mac12b, Mac14c, Tol16b]. C. [SB^{+}12, Sil10]. C.J [dV14]. C1 [Dav10, Hin08]. C2 [Dav10, Hin08]. cabinet [Ste08, Ste09a, Ste09b, Tol10]. Cable [Pur18]. Cakes [Hem19]. calc [Ada09b, Lor11g]. calculating [Abe16, Hir11h]. Calculation [Lor19e]. calculations [vW19]. calculator [Lor14k, SLMRR10]. Calculus [HB16, Kör14, App10b, AM15, Bay15b, Bay17a, DN11, Fer14, Haw18b, Hir11b, Hun19b, Lor15m, Lor17i, Mac14c, Mac16c, MP11, McA10, McB18, Ric08a, RJPG12, R^{+}12, Rza10, Shi16b, Shi17a, Pat17, Spo18, App14b, Gib18b, Hop18a, Shi18b, Tol15a]. calculus-based [Rza10]. Calendar [Tol15g, Sta19a]. Calinger [Hoa17a]. call [Lor17g]. Cambridge [Bay10a, Bay15a, Bev17, Cam18, Cou17b, Cou19, Fre10a, GG10a, Hal16, Haw18a, Hoa15a, Hop19, Hun18b, Lee17, Lev10c, Lev17a, Mac15a, Shi18a, Sto17a, Sut11, Tol15a, Tol16d, Tol16c, Tol16b, Tol16e, Tol17j, Tol19g, Tol19c, Tol19e, War11b, Cra07, Dol15b, HHHS14, Hal15b, Hal15a]. cameos [Lor18i]. Can [And12, Shi18c, CK19, Shi19d, Sim11a, Kuc12]. cannot [Osl12a]. can’t [Gri13f, Hav12]. Cantor [Sco17c]. cap [Lor18b]. Cape [Ano14a]. card [NG15]. cards [Hal17b, Hal17b]. Carl [Bay17c]. Carmen [Hun18a]. Carmichael [Jam11a, Sur18]. Carolyn [For10b, For13]. carriers [She16]. Carroll [Rua10, Wil08]. carry [Gri15f]. carry-on-until-one-girl [Gri15f]. Cartwright [Lor19a]. case [Hun17f]. casebook [Ste14, Tol15f]. cases [McN14, OS15]. Casino [Wer09, Hai10]. Catalan [Gau12b, GL11, JL17, Kos11, KG11, Kos12b, Kos12a, Lev13i, Lev17a]. category [Hem19]. Catherine [Pat17]. Cats [Lev12e, Ban14b]. Cauchy [Lor15j, Mer17]. Ceccherini [Eat13]. Ceccherini-Silberstein [Eat13]. celebration [Hop19]. Cellular [CSC10, Eat13]. Centenary [Ano17b]. Central [BGGGR18, Kos11]. centre [Abe19, Bra11b, Job15, Sco19b, Sim10]. centred [ES11, ES16]. centres [Luk19]. centroids [Lor18h]. centuries [FFW13, FR02, Lev14g]. century [Ano12d, B^{+}11, Dev08, Ges11, GG10b, KP14, Lev15a, M^{+}08, Tub14, Cra07, GG10a, GG12c, Gri13c, Hoa15b, Rou13a]. Cerone [Shi13c]. certain [ASH19b, GG13, Gri18b, Jos19c, Kum10, Kum11, Lor15a]. certainty [Bar10]. Cesàro [Nur16]. cevians [Haj17b]. Châtelet [Tol14b, Ari12]. Chad [Lev14c]. Chain [Eat12]. chains [CLM13, Sur18]. Chair [LRQ10]. Chaitin [Hud11b]. Challenge [Lor11c, Cam16, Lor10g, Lor14k, Lor18k]. challenges [Ano13m, Hal17c, Hor14, Tol15e, Yeo17a, Yeo17c, Tol15e]. challenging [HJS11, Hun13d]. chance [Tol18d]. Chancing [Tol17c]. Chandler [Lev10e]. Chang [Shi17a]. change [BS12b, FMK18, Hor17]. changed [Gri14a, Gri14b, Mac12a, Ste12c]. changes [SA16]. chaos [Raj13]. Chapman [Cou13, War10c]. characterisation [Haj18a, Mah13b, VM19]. Characterisations [KF19, ASH19b, Jos17c, Jos18a]. Charles [Lev14f, Mac15c, Pla11c]. Chartier [Tin17]. Chebyshev [Gri16a]. Cheng [Hem19, Hoa19]. Chicone [Hun18a]. children [PdVF14]. Chinese [Coo19]. Ching [Shi17a]. choice [Mcn16]. choose [CF10b]. Chris [Bur10a, Hal15a, Hal15b, Jon19, Lev18b, Lev19b, Pat17]. Christian [Tol18e]. Christopher [Cam18, Kil10, Ste16b]. Chunhui [Bay17d]. Chwe [Von17]. Cinderella [Gib11a]. cinlar [Hai12]. cipher [Hol12a, Lew10]. ciphers [Bay10d]. circle [Bra11a, Fra19a, Job11b, Lev17h, Sco17a, Sco19a, SNH11, Spo18, Sud11, Szy17, Szy19]. Circles [Smi16, Bra11b, Bra12, Mah18a, Mat18b, Nun13, Oku13b, Sew15a, Sil18b, Sil18a, Ste17b, Wil18]. circular [AZY12, Ell17, Lor10i, Per12, Shi13a]. circum [HE19]. circum-medial [HE19]. circumcentre [Bra11b]. circumscribed [De 11a, GG12b]. circumscriptible [ASH19b]. Clarke [KA19]. class [Bra17, Kac18, Lev19c, Mez14, Sto14b, Tol17b]. classes [Dol18a]. classic [Rib19, Hai17a]. classical [Gri13a, Jac19, Leh11, Sch16c, Hun18b, Lor10d]. classification [Dol11b, Jos16c]. classroom [BCSS19, Hun17f, PF08, SD10, She10a, Sto17a]. Claudio [GG12c]. Claudius [Hol12a]. Clausen [Rza10]. Clay [Hun18c]. Clerk [FMW14, Mac15b].

Clifford [Gib13a]. Clones [She13]. Closed [BT17, Gau12b, Sco14d, Sco19c, Lev19c]. Closing [Tol18a]. closure [Hol19]. clothoid [Lor16e]. Co [Hop18a, Hun17a]. coastline [BM19]. Codes [Bay10d, Coo19]. coefficients [Den10, FK19, HP10a, Kos11, Vil14b]. cognitive [Bor10b, Sto12c]. cohomology [Cou10, Gue08]. coin [Gri12h]. coin-tossing [Gri12h]. coinage [Shi18c]. Coincidence [Abe19, Tol17e]. coincident [Gla13]. Colin [Gib11c, Lev13c, Lev18c, Lor11g, Mac14a]. collaboration [Bau13]. Collected [SB^{+}12, Cou14]. collection [GDDR08, Sut11, Bay19b]. collinear [HJ18b]. Collins [Hal13, Hal14, Hal13]. collisions [Lor15j, Sim11a]. Colmez [Lev13f]. colorful [Soi09a]. coloring [Soi09a]. colour [Ste16a]. colourful [Lor11h]. colouring [Lor11h]. colours [Gri13d, LG10]. column [Gri15b]. Combinatorial [Hal12a, Bev14, Mol12, Rua14a, Wei12, Shi14c]. Combinatorics [WW13, Gri18d, Gri18c, Lev14b, Row19]. come [Iva09, Lor11f]. Coming [Yeo18a]. Commensurable [Hir11a]. comment [Fos11]. commentary [ALB09, ALB10b, Knu14b, Knu14a, War11b]. Commission [GG10b, M^{+}08]. common [Den13, Lev19b]. communicate [Wai09]. Communicating [BRR08, Ban10a]. commutative [Kan09a, Kem11, Lor10c, Lor13d]. companion [Hou09, Hun17e, PF08, SD10, She10a, Tol19g]. Company [Zar11]. comparative [FMK18]. compare [Far13]. competitions [ZH09, De 11b]. complete [Tay07]. Complex [Dol15a, Bay17d, GS16, Jam14g, Mur12, WO11]. composite [Lor18g, MM15]. Composition [Sur13]. compound [Far13]. comprehensive [Lo 13]. comprising [Tay07]. Computational [Cou13, Gib18a]. Computations [VvBM08, Hun10a]. Computer [GD14, Dav12, LHS14, Mac13d]. Computer-generated [GD14]. computers [Mac12a, Gri14b]. Computing [Yan10]. Concepts [Tub09, BY11, Hun10b, Yeo18a, For13, Hun17c]. concerning [Bra17, Kan17]. Concise [Hop18a]. Concrete [DN11, Har14, App14b]. concurrences [Haj15b]. Condition [BC13, BS12a, Cou15, Cou17a]. conditionally [Nur16, JLL15]. conditions [Nur14b, Pur19]. cone [Lor17i]. confidence [Kil10, Orm08]. configuration [Oai15]. confusion [Lor11j]. congruence [Dol18a, Haj10, Pla11b]. congruences [Kos11]. congruent [Job18a]. conic [Gho11, Mah19b]. conical [DW19]. Conics [Bra12, DW19, McL14, Alv17, Hos17b]. conjecture [Dol17f, KA19]. conjectures [Gri17d]. Conjecturing [GR13]. Conjeevaram [SB^{+}12]. connecting [Gri16a]. connections [Gib13a, Tau11, Wel12, Hop19]. consecutive [Hir11f, JLU10, Ste19a]. consequence [Kos12a]. considerations [SA16]. Constant [Sim11c, Dol11b, FK19, Hir12, Hir13, JL17, Lor15k, Mel13, Sco10a, Sco14a, Sco17b, Sîn10]. constants [Pla11a]. constraint [Lor10a, Lor12h]. constructed [Mac11a]. Constructing [Lor16d]. construction [Lor12i, Mah18a, McL19a]. constructions [DL14, Sil12]. content [CC10, DSZ08, Lev10e, Tub14]. contentious [Bay11b, Ros09]. context [Shi18b, Bay17b]. contextual [Dav17]. continued [BT17, Der19, Hir11e, Osl11b, Osl12b, Spi10, Ton16, Ton17, Ton18, Yan14]. continuity [Ran18]. Continuous [Jam13b]. Continuum [Sto10b, Zar10b]. contradiction [Sub18a]. contributors [Ram09]. control [Wai09]. conundrum [Lor17d, Sim18]. convention [Haj17e]. convergence [McN14, Mur11a, Nur14b]. convergent [Coo11a, JLL15, Nil13, Nur16, Sco14c, Shi19a]. converging [Gri15h]. Conversations [Gib10b, Seg09]. converse [Lor18h]. Converses [Nur16]. conversions [CP10]. Convex [She16, Far18, HI14, Jos16c, Jos19c, Jos19a, Lor15h, MW18, Osb13, PU11, Sil17, IS12]. convexity [McL14, PF10, Yeo18c]. conviction [Stu10, Gib12a]. convolutions [Gau12b]. Conway [Eat12, Kar18, Gri17c]. Cooke [Lor10d]. Cooper [Hop19]. coordinate [BCL10]. coordinates [Kin10]. Coornaert [Eat13]. copycats [Geo15]. Coralie [Lev13f]. Core [Hal14, BL12, Hin08, Dav10]. Corner [L.10a, L.10b, L.11, L.15a, L.15b, L.15c, L.16, L.17a, L.17b, L.17c, L.18a, L.18b, L.18c, L.19a, L.19b, Lor10h, Lor11k, Lor11l, Lor12d, Lor12e, Lor12f, Lor13e, Lor13f, Lor13g, Lor14m, Lor14n, Lor14o, Lor19c]. corners [Lor10a]. Corporation [War13a]. correction [VSR13]. correctly [Hop16]. Correspondence [Chi12, Coo11b, CF10a, Fos10, JJ18, Lor10f, Mah11, Mah15a, McL11, MD16, PH17, Pla10a, Ste19b]. Corry [Rua17]. Corti [Rot10d]. cosecants [Sub15b]. Cosine [Leu12b, Jam15e, Kha13]. cosines [CO11, Rou13b]. Cosmic [Ste11a, Zet13]. Cotangent [Sub15b]. cotangents [Ste17a]. Could [Gra13]. Council [Rou13a, Sto14c]. Count [Rei14, Gri13f, Hav12, Jon19, Tol15b]. countable [Sub14a]. Counting [Jam15a, AG14, SS11, Sub14b]. Course [Gar13, And17, Bay15a, Hun13a, Hun17a, JM09a, JM09b, Kan09a, Kem11, Lor10c, Lor13d, Reb12, Sch16b, Shi10c, Shi18a]. coursebook [MC09, Hal11b]. courtroom [SC13]. cousins [Sof10]. Cover [Ano10g, Ano10h, Ano10i, Ano10k, Ano10j, Ano10m, Ano10l, Ano11f, Ano11h, Ano11g, Ano11j, Ano11i, Ano11l, Ano11k, Ano12f, Ano12h, Ano12g, Ano12j, Ano12i, Ano12l, Ano12k, Ano13e, Ano13g, Ano13f, Ano13i, Ano13h, Ano13k, Ano13j, Ano14h, Ano14g, Ano14j, Ano14i, Ano14l, Ano14k, Ano15f, Ano15e, Ano15h, Ano15g, Ano15j, Ano15i, Ano16d, Ano16c, Ano17h, Ano17g, Ano17j, Ano17i, Ano17l, Ano17k, Ano18h, Ano18g, Ano18j, Ano18i, Ano18l, Ano18k, Ano19g, Ano19f, Ano19i, Ano19h, Ano19k, Ano19j, KN19]. covers [MC09]. Cows [Ste10, Haw11]. crack [Kul13a]. Crafting [BY11, For13]. Craig [Tol18f, Zar11, Ste17f]. Craik [GG10a]. Crapoulette [NG15]. craps [NG15]. Crato [War12a]. Crawford [Hal10b]. crazy [Pas11b, Coo19]. CRC [App10a, Bay14a, Bev14, Cou13, Gib14b, Gib14c, Mac13e, Mac13b, Mac13d, Mun12, Rua14a, Shi13c, Shi14e, Tol19b, War13b, War13c]. Creating [Lev17b]. creation [Gri13a, Leh11]. creative [Bay12, DSZ08, Hol11, Hol13b, Lev10e, Lor14d]. creators [Soi09a, Lor11h]. Crilly [Lev10a, Lev11f, War10a]. crisis [Wel08, Ban10c]. Cristina [Cou12]. critical [Hoa15a, Ken14]. Cross [Dob11, Rot10a]. crossed [Jos17b, Jos18b]. Crossing [Lev08, Gus12, Lev13g, Sto10a]. crossnumber [Cut14].

Crosspacks [Tay07, Rot10a]. crossroads [De 14c, Tob12]. crossword [CMS18]. Cryptic [Cut14]. crystal [Wap12]. CSE [Sto12b]. cube [DL12, GG08b, Kuc15, Sut11]. cubes [Ber18, Chu18, Loo11, Ste19a, Sub14b]. cubic [Kul13a, TL16, Tra13, Vil14a]. Cubics [Lor14g, Fos12a, Joh11, Kul16]. cuboidal [ML13]. Cucker [Cou15, Cou17a]. cult [Lev14n, Mar12]. culture [BEU12, Tol14c]. cultures [Sen12]. Cunningham [Sur18]. CUP [Lev10b]. curiosities [Ste08, Ste09a, Ste09b, Wap12, Tol10]. curiou [Nur16]. curious [Gri17c, Luk18, Qua12b, War12b]. Curricular [Hun17f]. curriculum [And12]. curvature [Gib13a, Sco18b, Tau11]. curves [AG14, BL10, Ber17, Far18, Gla12, Hun13c, KL11]. Curzon [Gib18a]. custard [Hem19]. cut [Fos12c]. cutting [Leu18, Lor10a]. Cvetkovski [Shi14d]. cycle [Haj11]. cycles [Dol17b]. Cyclic [JST12, Ste18b, ASH19b, De 11a, Fra19b, Jos17b, Hun15]. Cyclotomy [Shi16c]. cylinders [Bea13].

D [Bay15a, Cou12, GG10a, Had19, Lev17e, Mac10c, Mun12, Shi10d, Sut11, Tol17g, Tol17h, Zet13, dVS18]. D1 [Rot10b, Jef08a]. Dahan [Eat11]. Dahan-Dalmedico [Eat11]. Dalen [Pas14]. Dalmedico [Eat11]. Dan [Hun18c]. Dana [Lor14f]. Dance [Coo19]. Dandelin [Ste17b]. Daniel [Hal19, Lor16b]. Darling [Haw18c]. Darren [Tol19e]. Dars [Gib10a]. data [Whe13]. Datti [Cou19]. David [Ban10c, Bay14b, De 14b, For12, Fre10b, Gib18b, Hal10b, Hal16, Haw18c, Lev10c, Lev11b, Lor10e, Lor16b, Lor18f, Mac15a, McB18, Rot10c, Shi10c, Sto17b, Tol15b, Tol17j, Yeo17c, Zar11]. Davis [Lev10e, Mac13d]. Dawn [Ale10, Gib11b, Mah18b]. day [Vil11]. dead [Lew13]. Debate [Mer15]. decade [Dol15b]. Déchiffreurs [DLP08]. decimal [Gri15b, McL18]. Decision [Jef08a, Rot10b]. deck [NG15]. decline [Kil10, Orm08]. Decoding [Ved10, War11a]. decompositions [DeM10]. define [Ste11a, Ste12c, Zet13]. Defining [GS16]. definition [FL14]. Deformation [Har10b, Meg12]. Degree [Alc12, Alc13, Fos14b, Sto14a]. degrees [Shi11c]. Delbourgo [Lev17i]. Deleting [BKS17]. Delgado [Rou16]. Delhi [Cou14]. delightful [SV15]. delivery [Tol13]. della [GG10b]. Demaine [Sut11]. democracy [Hud13b, Szp10b]. demonstrate [Leu18]. Dénes [Gib17a]. denominators [Jam19b]. dense [Mez14]. Derangement [Sof13, Lew13]. derangements [Hun18b]. Derek [Bay12, Lor14d]. derivation [Lev13i, Lor14c]. derivations [Osl15, Ste17f]. derivatives [Cel17, Luc10, Luc13, Ran18, San11]. derived [Gre19, JJ13b]. Derych [Sto12b]. Desarguesian [Bur17]. descent [Lor13i]. descente [Dol11a]. descriptive [Gao08, Gao09, War10c]. Desheng [Bay17d]. design [Lan12, Mac13e]. Desolneux [For12]. determinants [ES10]. Determining [BB12]. determinism [Raj13]. develop [Hal17a]. developments [And12]. devil [OD13]. Devlin [Lev10f, Tol17d]. Diaconis [Tol18d]. diagonal [Jos17e]. diagonals [Gri12d]. diagram [Dol19b, Mac11a]. diagrams [Rou13c]. dialogue [Rea10b]. diameters [Osl11a]. Diane [Shi14e]. Dianne [Mah10b]. Diarmuid [Lev11e]. diary [Lev14k]. Dice [May16, Jár16, Rob13]. dice-and-numbers [Jár16]. Dick [Gib17c]. dictionary [Lo 13, Lor10e, Nel08]. Did [OD13]. died [Fre15, Sen12]. difference [Dol12b, Gri18b, LY11, Mac14e, Mah13b, Mcn16, vW19]. differences [HJ19]. different [Gri12e]. differentiable [Jam13b]. Differential [BL10, BD15, Lov10, Tau11, Vol11, Cou19, De 14b, FK19, Hun19a, Jac14, JH12, Mac14c, Mac16b, RJPG12, Gib13a, Hun13c, Mac13c]. Differentiating [FN13]. differentiation [Mur19]. diffusion [App14a, KS14]. Digging [Shi12b]. digit [Gri11f, Rea10a, Shi11f]. digital [Ban10a, BRR08, Lor15n]. digits [BKS17, Lor17b, McL18]. Dimensional [Ste18a, CP10, Haj18b, Job10]. dimensions [Abb08, ALB10a, HTSE09, Kha13, Ste17c, YN10, YN12, Cou12]. Dineen [Shi16b]. dining [Kee13]. Diophantine [Baj13, Dol16a, Hva11, Pla10c]. direct [Lev13i]. direction [Koo17]. Dirichlet [Gri12g]. Dirk [Pas14]. disc [Sco14e]. discovered [Gra13]. discovering [Bay15b, Fer14, Tol19b]. discovery [Ano12d, Osl10a, Rou13a]. Discrete [HJ18a, Lor14h, Pat17, Gri11c, Hop19, Mic09, Tol17a]. display [Wai09]. dissection [Job11a]. dissections [Mat18a, Ste12b]. Distance [GOD18a, Bar11, Bar13, Cul11, Gri19, Nic16, Sco17e]. distant [Sof10]. distribution [vdV12]. distributions [Gri10a]. divergence [Ayd18]. divergent [Shi19a]. diversions [GG08b, Sut11]. Divided [Pop12, Mcn16, Gib14b]. Dividing [Leu12a]. Divisibility [SE13, Gri18a, Has19, KG11]. divisor [Shi16d]. Divisors [Jam15a]. DIYMathematics [Mah19a]. Dmitry [Hun14a]. Do [Lev17e, Haj11, Lor15d]. Docent [Rua14c]. dodecagon [Job11a]. Dodecahedral [FJ10]. Dodgson [Pla11c]. Does [McL19a, Lor17b, Pas11b, Hal17a]. dog [Orz12, Lev14c]. Domains [Vol11, Jac14, Wil17]. domes [FMS10]. Dominic [Gri18d]. Donald [And13b, Gib17a, Hoa14b, Hud11a]. done [Dev08, Kin15]. Dora [Lev14k]. Dorothy [Sen12]. Dospinescu [Lea10]. Double [Ber17, HK14, Jam13c, JL13, Jam14f, Jam16b, Pla10b]. doublets [Fos12d]. doubly [Sao17]. Doug [Ric10]. Douglas [How15]. Dover [Rua12]. down [Gla12, KL11]. download [Gib11a]. Dr [Lev15c]. Dragomir [Shi13c]. drawer [Ste09b]. dread [Whe13]. dream [Cri15a, McG13]. drive [Gri14b, Hol12b, Mac12a]. Droz [Sil12]. Dual [BD12]. duality [Jos17d]. Dudley [App14b, Bay10c]. due [Jew10]. Duel [Ale10, Gib11b]. Duffy [Mun12]. dull [Gib19b]. Dunham [Hun19b]. Dürer [Sud13]. dusk [Mah18b]. Dwight [Mac16c, Zet18]. dynamic [Gro18, SSJ19]. Dynamics [Sma12, Cou13].

e-book [Bay17a, Bay17c, Dav18, Gib17c, Hai17b, Hun16, Hun17b, Hun17c, Mac17, Sch16b, Sch16c, Shi15c, Shi15a, Shi16b, Tol14c, Tol17h, Tol18b, Tol18e, Tol19a, Tol19f]. e-ISBN [De 14b]. each [Bay15a, Hal17b]. Earl [Tol19g]. Early [Lev11e, Hoa15b, KP14, Qua12b, Rua11, Van09]. earth [Mah13a, Rua11, Van09, Col11a]. easy [Gri11b, Jam14g, LP16, Szp10a]. ebook [Hai17a, Tol14c]. Ed [Shi14b, Fre10a, Gib10b, Gri13d, Hoa14a, Hoa15a, Lev14l, Lev17g, Lev17f, Lev18d, Lor10e, Rot10d]. Edexcel [Hal10a, Sto12b, BL12, Hin08, Nic08a, PF08, She10b, SD10, She10a, Dav10, Gol10, Hal10c, Rot10b, Hal14, Sto12a]. edges [Sch16a]. edited [Cou14, Cri17a, Hal10a, Hal17c, Hop19, Hud11a, Hun16, Hun17e, Hun18c, Lev12d, Lev14b, Lev14m, Lev15b, Mac15b, Rua12, Sto14b, Sut11, Tol14c]. Edition [Gri18c, Ten12, ALB09, ALB10b, And10a, And10b, Hud10a, Sto17b, Tol16d, Tol16b, Von17, War11b, Zet18]. Editions [Yeo17a]. Editorial [Lev11g, Lev15d]. Editors [GG12c, Hal11a, Hoa11]. edn [Bay11a, Bay14a, Cou17b, Fre15, Gib11c, Gib17a, Gri11d, Gri14e, Gri17d, Hal11c, Hud11a, Hun17d, Jon19, Lor10e, Lor14e, Mac10c, Mun12, Shi10i, Shi10g, Shi15c, Shi16b, Sut10, Tol16c]. eds [For10b, Gib10a, Gib12b, GG10b, Lev10e, Lev13e, Lev14g, Mac13b, Rua13, Sch10, Sch13]. educated [Haw12, Mah10a]. Education [Sto10b, Ten12, WILY11, AR14, BHJ10, Hal11a, Sto14b, Tol14e, M^{+}08, GG10b, Hal17c]. Edward [Gib14b, Haw18a, Lev14d]. Edwin [Lev10d, War11b, Qua12c]. effect [FFGH09, Sch10]. effectively [Hol13a]. effects [Lor19d]. Egypt [Dav17]. Egyptian [Tol15b, Rei14]. Eigencircles [EF10]. eigenvector [Yan15]. Eight [Jos19d, FFW13, Lev14g, Wil08]. elastic [Sim11a]. election [Bra08, Hud10c]. electric [Hun12a]. elegance [And13b, Ben12b]. Elegant [OS15, Lor18f]. Elementary [Mac09, AM15, Kar18, Lor16f, McA10, Mur19, Pat15, San15, Sao17, Ton18, Wil17, You11b, You12, Zor10, Shi10d]. Elements [Kra12, Suc11, Bay14a, Cou17b, Lor17e]. Elena [De 10a, Hun14a]. Elephant [Coo19]. Eleusis [Sut11, GG08b]. Eli [Lev14a, Tol19d]. élite [Gri11h]. Elizabeth [Sto17a]. Ellen [Lev13b]. Ellipse [Sci15, Gus15, Jam14c, Mah10c, Mat18b]. Ellipses [Lor15i, Yeo17a, Sil17]. ellipses.fr [Lor15i]. Elliptic [AG14, Vol11, Jac14, Jac14, Tol15c]. elliptical [Gau12a, Lor14a]. else [Pas11b, Sci15]. Elsevir [Hun18a, Tol18b]. Emil [De 10b]. Émilie [Tol14b, Ari12]. Emmental [CDF18]. Emmy [Ten08, Gri10b, Zet18]. Enciclopedia [GG10b]. encounters [Cra10, War12a]. End [Rob13]. ending [Blu11]. ends [Lew13]. energy [Sim11a]. Enescupp [Yeo16]. Engineering [Cou12, Bre11, Duf11, Gri13b, JS08, Mac10c, Mac13a, Mac14c, Mun12, RJPG12, R^{+}12]. engineers [ÅM08, Vel09, Mac10b, Mac10a]. English [Dav15, Knu14b, Knu14a]. enhance [Hal17d]. Enlightening [Maz14, Lor15g]. Enlightenment [Hoa17a]. entanglements [Sto17a]. entertaining [Cra10, War12a]. Entry [MGC13b]. enumeration [Koo11]. Envelopes [Sew15b]. Epple [Tol14c]. equable [Dol16b]. Equal [Fos14a, Lor14h, Lor19b, Shi11c]. equalities [Gri13e]. equality [BS12a]. equally [Haj17b]. equally-spaced [Haj17b]. Equals [NS87, Mar11]. equation [Dol12d, Dol13a, Dol16a, Nic12b, Rea19, Sta18, SB19, YF12, Lor18f, Lor18f]. Equations [Kre14, Vol11, Baj13, Bea12a, BD15, Cou19, De 14b, Fer16, FK19, Gri14a, JH12, Kan09b, Kul13c, Kul13b, Leu17, Pla14, Shi11d, Ste12c, TL16, Tra13, Hun19a, Jac14, Mac16b, Shi15c]. Equiangular [De 11a]. equilateral [Ano14r, De 11a, DZH14, Gri15d, Kri18, Sci19, Tod15]. equilibrium [De14]. equivalent [Nic12b]. era [BRR08, Ban10a]. Erdos [Dol17f, Cam18]. Erhan [Hai12]. Erich [Gri13a]. Erik [Ste16b, Sut11]. Ernest [Mac13d]. erratum [Lor13c]. Error [MS13]. Escher [And17]. essays [Ken14, Hoa15a]. Essentials [Hal11b]. Esteban [Gri14d]. Estermann [Raj18, Siu13]. estimates [vW19]. Etingof [Hun14a]. Euclid [Lor17e, Har14, Lor14b, McL19a, Rea10b]. Euclidean [LP16, Wil17]. Eugen [Lev14a]. Eugenia [Hem19, Hoa19]. Euler [De 10a, Hoa17a, Die19, DO12, FN13, Fel07, Gle15, GOD18a, HHPM07, Hir12, Jam12, Jam14b, Jam15b, Kha10, Lor12j, Lor14c, Lor16e, Luk17, Osl11a, OD13, Pat13b, Raj13, Ric08b, Sco10a, Sco11b, Sco14a, Sco17b, Ton16, Ton17, Ton18, WO13, De 10b, Lor18f, Rot10c]. Eulerian [Tsa11]. EUR101.60 [Sch14]. EUR19 [Lor15i]. EUR42.00 [Yeo17a]. EUR49.00 [Gib11a]. EUR60 [GG10b]. Europe [Cri17a]. European [Hoa13a, LPE^{+}12, Gri14d]. Euros [Lor10c, Lor11h]. evaluated [JL17]. Evaluating [Che16a, Che18, Jam15c, Lor14i, Sik18, You11c, Den13, Lor12b, Mur12, Sco10a]. evaluation [Lev19c, Lor16a, Shi16a, JL13]. Evans [Hun17a]. Évariste [Neu11, Hoa13a]. Even [AW07, Jam10b, Sto10b]. event [Wio19]. ever [Haj15a]. Everyday [Fer14, Ban14a, Cra10, Her11, War12a, War12b, Bay15b]. everyone [CT14]. everything [Pas11b]. everywhere [App12, Bay14b]. exam [Ada09b, Lor11g]. examples [Gla12]. exc [Cou17b]. excentral [Luk18]. excentres [Haj12, Haj17d]. excursions [And13d, Hah12]. exercices [Bat14]. exercises [App18, Lev17i]. Exeter [Sco12]. existence [Haj18b, Ste11b, ZF10, Gri14f]. Exotic [Orm08, Kil10]. expansion [Gri16b]. expectations [Wap12, Gib14c]. expected [Tor17]. expeditions [And13a, Swe12]. expenditures [Kro18]. Experimental [Vil07, Shi10e, Shi10f]. explained [Cra09, Hal10b]. Explicit [Sur14]. exploration [Lev17h]. explorations [Haw11]. explorer [KL14]. exploring [And13a, Mac14a, Pas13, Swe12, Hal19]. explosion [Gib17c]. exponential [Jam15e, Lor17c]. exponents [Vîj19]. expressed [Sub15b]. Expressing [Jam17]. Expressions [SB19, BT17, Gri11a, Tra10]. exradii [Haj18b, Lor15d, Luk17]. Extended [GG10b, LP16]. Extending [Hos17a, Lor15l, Lor18g]. extension [FM10, Kum11, Vig19]. extraordinary [Hol12a]. Extremal [Nic16, Sil17, Haj17d, Haj12]. extremum [Sim13].

F [App10b, Ban10a, Gib10a, GG10b, Hor17, War11b]. Fabio [Shi19c]. Faced [Roc14, Hai15]. facing [Pas11b]. factor [Lor15c]. factorial [Cag18, VSR13]. factorials [Cag18, Gri15c, Mac15d, RG14]. factorisation [Bau14, BB12, Shi15d, Shi15e]. factorise [Lor14g]. Factorising [Shi19b]. facts [Gri17d]. fallacies [Haw18a]. fallacy [Lor14j]. families [Gri17d]. Family [Pas11a, GB12, Kos14, Lev19b, Lor13a, Oai15]. Famous [Pet09, Gri12f]. fancy [Mat18a]. fantastical [Rua10, Wil08]. fantasy [Adl14]. Farey [Rua14c, Gut11]. Farin [Mah10b]. Farny [Sil12]. Farris [Lev17b]. fascinating [Gri11e, dK09]. Fast [Coo11a, Hol12b]. Faulhaber [Chu18]. Fauvel [Lev14g]. fears [Spr09, Ban10b]. featuring [Shi12a]. Feb [How16]. Feedback [ÅM08, Ano10b, Ano10c, Ano10d, Ano11b, Ano11c, Ano11d, Ano12b, Ano12c, Ano13a, Ano13b, Ano13c, Ano14d, Ano14e, Ano15a, Ano15b, Ano15c, Ano16b, Ano17c, Ano17d, Ano17e, Ano18c, Ano18d, Ano19b, Ste18c, Mac10b]. feeding [Gar18]. Feet [Lev12e, Ban14b]. Felder [Sch10]. Felipe [Cou15, Cou17a]. Fellmann [De 10b]. Fermat [Bar17, Bau14, Bur10b, Dev08, Dol11a, Dol12a, Dol12b, Dol12d, Shi15e]. Fernandez [Haw18b, Bay15b]. Fernando [Shi10e, Shi10f]. Ferreirós [De 17]. Ferry [Col13b]. Feuerbach [GOD18b, Sco19d]. few [Bra17]. fiber [BY11]. Fibonacci [Di 13, Gau11, Gri13e, Gri17b, Jam18d, Kos14, MMYA10, Mil11, RG14, Rou11, SE13, Sco18a, Sub19, Tol17d]. Fibonacci-like [MMYA10]. Fibonacci’s [Gri15g]. Fibonacci’s-rabbits [Gri15g]. Fibonometry [CR13]. fiction [Adl14]. field [Jac19]. Fielder [Zar10b]. Fields [FGR14, BGGGR18]. Fifth [WILY11]. fifty [Gri11b, Szp10a]. fighting [Mah10a, Haw12]. figurate [Ste18a]. figure [Mac11a]. figures [Dav18, FFW13, Lev14g]. Figuring [Cra10, War12a]. fill [Job11b]. find [Bau14, Pas11b, Rou13b]. Finding [BCL10, Jam11a, Jam11b, Shi11f, Tre17, Wio19, Gar18, JO14, Ste17e, Tol17d]. finding-feeding [Gar18]. Fine [Shi15b]. Fingers [Bau14]. finite [Cou12, HTSE09, HJ19, Shi11e, Wil09, Hud10a]. Fiona [Hal11b, Hal12b]. first [Bay10a, Ber18, GG08a, GG10b, Hoa11, HP10b, Hun13a, Lev11e, M^{+}08, Reb12, Shi10c, Ste19a]. Fisber [Gri13a]. Fisher [All18, Leh11]. fit [Tol14e]. fits [Wil08]. five [Hoa11, HP10b, Leu18, Lev11e, Osl15]. five-fold [Leu18]. fixed [CK19, Gri15b]. fixed-column [Gri15b]. Flatland [ALB09, ALB10b, ALB10a, Abb08, Lev10d, War11b]. Flavia [Hal11a]. flaws [Haw18a]. flimflam [Haw18a]. Flips [Cor07, Rot10d]. Flood [Lev13e, Lev13h, Lev14g, Mac15b]. flotation [Rob12]. Fluke [Tol17e]. fold [Leu18]. Folding [Job13b, Hir11g, Job18b, Leu18]. folds [Cor07, Rot10d, Rot10d]. Folkerts [Cri17a]. Fontes [Gri14d]. fool [OD13]. footnote [Pla10b, Shi10b]. footnotes [Lor19a]. force [BGGGR18]. forcing [Hal12a, Shi14c]. forgotten [Rua14b, Van13]. form [AT12, BT17, Dol17a, Lev19c, Lor15d, Pur19, Sco14d, Sco19c, Tra10, Tub14]. Forming [Har10a]. forms [Gau12b, Shi10a, Shi12c]. formula [Cer16, Cri12a, ES11, Far13, GL11, Hir13, Jam14b, Jam15b, Jam15d, Kha10, Kha12, Lev13i, Lev12f, Lor12g, Lor14c, Lor15c, Lor18b, Mcn16, Pla11c, Ric08b, Sao17, Ste17f, You12]. formulae [Jos19d, Lor13a, Lor15k, Pla12]. formulas [Coo11a, Gau19, Jos16b, Jos19c, OC14, Osl15, WO11]. Fortuitous [Ano14q, Lit13]. forward [AW07, McC15, Sto10b]. Foster [Hal10a, Lev18c]. found [Bur10b]. foundation [Hal17b, PF08]. foundations [FFGH09, Sch10, Sto17b]. Fountains [SG12]. Four [Job18a, Jam10a, Job10, Sta19a, Ste16a, BM19, JJ13a]. four-colour [Ste16a]. four-dimensional [Job10]. four-hundred [Sta19a]. Fourier [Rib19]. Fourth [Gri18c]. fractal [BM19]. fractals [LG10, Gri13d]. fraction [Hir11e, Osl11b, Ton16, Ton17, Yan14]. Fractional [Gau11]. Fractional-power [Gau11]. fractions [BT17, Jam17, Lor15a, Osl12b, Spi10, Tol19h, Ton18]. Frameworking [Hal13, Pea11]. Francesco [Sch13]. Francis [How19]. Franck [Tol14a]. Frank [And13a, Lev17b, Sut10]. Frankel [Lev14d]. Frankland [Lev15c]. Franklin [Hoa10, Pas08]. freaks [Pas11b]. Fredholm [Jac14, Vol11]. free [Dol15b, Jam10b, Lor14b, McA10, San11]. Freeman [Zar11]. Freitas [Sto17a]. French [Ric10]. fresh [Jam14b]. Fresnel [Jam15c]. Fresnel-type [Jam15c]. frets [RR11]. friction [Kin15, Koo17]. Friedman [Jac19]. friends [Mac17]. Frieze [Bev17, Bea13]. frightened [Pas11b, Lev13c]. Fröhlich [FFGH09]. Fröhlich. [Sch10]. Front [Ano10g, Ano10i, Ano10k, Ano10m, Ano11f, Ano11h, Ano11j, Ano11l, Ano12f, Ano12h, Ano12j, Ano12l, Ano13e, Ano13g, Ano13i, Ano13k, Ano14h, Ano14j, Ano14l, Ano15f, Ano15h, Ano15j, Ano16d, Ano17h, Ano17j, Ano17l, Ano18h, Ano18j, Ano18l, Ano19g, Ano19i, Ano19k]. FRS [Ste16b]. frustration [Lor11c]. frustum [Lor15m]. Fueling [Ano12d]. Fuelling [Rou13a]. Fukagawa [Lev11d]. Fulton [Fre10b]. Fun [Zho13]. function [Bor10a, FM10, Gib14a, Gri11f, Jam14b, Jam15d, MW18, May16, SS11, Shi16d]. Functional [Kan09b, App14b, DN11, Mac09, Shi10c, Shi10d, Shi11d]. Functions [Jam13a, Ano14q, App11, Bay17d, Cae15, Col13a, Del19, Dol19b, GS16, Hop13, Jam13b, Jam16a, Lit13, Lor11j, Luc10, Luc13, Pat15, Tol16c, VV09]. Fundamental [Ano06, Bre11, MP11, Yan10, Rib18, Spo18, Gri13b, Mac13a]. Fundamentals [MNS10, Gib11c]. Furinghetti [GG10b]. Further [Jos17c, Hoa18a, Gla12, Gri12c, Luc13]. future [Gri14b, Mac12a]. fuzzy [Osl15]. Fyfe [Tol13].

G [And10a, And10b, App14a, Bay14a, Gri11b, Gri14d, Hud13b, Mac14c]. Gabriel [Hun17f, Lea10]. Gage [Tol17j]. Gaglione [Shi15b]. gallery [Gri11c, Mic09, Hun19b]. Galois [Shi10i, FGR14, Neu11, Ser08, Hoa13a]. game [Bra08, Cri17b, Dev08, Gri11g, Jár16, Min12, Rob13, Von17, Hud10c, Lev10f]. Games [Pas16, Wel12, GG08a, Hal17d, Bay10a, Mac15a]. gamma [Jam14b, Jam15d, Jam16a]. Gao [War10c]. gap [Tol18a]. gaps [Jam19a, Shi13d]. Garcia [Shi18a]. Gardner [Bay10a, Sut11, GG08a, GDDR08, GG08b]. Gareth [Tol17f]. Garland [Lev14j]. Garling [Bay15a]. Gaston [Kee11]. Gauging [Mah13a]. Gavin [De 11b]. Gazette [Ano18e, Ano10f, Ano11e, Ano12e, Ano13d, Ano18b]. GCHQ [Ano14q]. GCSE [Gri11d, Hal12b, Hal17b, Sto12a, Ano10a, Kra09, PF08, She10b, SD10, She10a, Hal10a, Sto12b]. Gellasch [Bay17b]. gem [Ric08b, Rot10c]. Gems [Bar09, Lev12c, Rea10b, Lev11a]. Genealogical [Qua12a]. general [Lor14c]. Generalisation [Jos17d, Ret17, Bra11a, De 14a, Kir13, Kos12b, Kum10, Nic15, Red15, SL17, Shi11b, Sîn10, WO11]. Generalisations [De 13, Jos16a, Abb15]. generalised [CO11, DW19, GL11, Mil11, Oku13a, Rou11]. Generalising [De 12, Job16, Ret18, WO13, MvdB12]. generate [Haj11]. generated [GD14, Mas14]. generating [Ber17, May16]. genius [Ges11, Gri13c, Hoa17a, Gri17c]. gentle [Hal12a, Shi14c]. genuine [Kul16]. geodesics [Pop12]. Geoff [Zar10a]. Geometric [Ger08, HTSE09, OC14, Abe19, CQ19, Hun18c, Jam14a, Jam18b, Gor19b, Koo11, Cou12, For10a]. geometrical [Jac11, Mah18a, Pla18, Sle12]. Geometry [Ber10, HF04, HS13, BL10, Bar09, BCL10, BC13, Cou14, Cou15, Cou17a, FH05, Fox15, Fox17, FR08, Gib11a, Gib13a, GG13, Gri11g, Hun13c, Jos18a, Lev13j, Lov10, Mac13c, Mah10b, MJ14, Rea10b, Shi14e, SSJ19, Tau11, YN10, YN12, Pat17, Lev11a, Lev12c, Lev14a, Lev17d, Shi16b, Lev12b]. George [Cou19, Gri11b, Hud13b, War13b, War13c, Zar11]. Georgian [War12b]. Gerald [Hud11a, Mah10b]. Geretschläger [For10a, Lev17h, Tol18e]. Gerhard [Shi15b]. German [Tol14c, BEU12]. German-speaking [Tol14c, BEU12]. Gernot [Sch13]. Gerovitch [Hun14a]. Gerry [Fox15, Gri18d, Sto10a]. Gessen [Gri13c]. Get [Ben12a, Lev13f, SC13]. Getting [Sim10]. Geveci [Bay17a]. Ghosh [Mac14c]. Giacardi [GG10b]. Gian [Sch10]. Giovanni [Sch10]. girl [Gri15f]. girls [CT14, PS14, PdVF14]. Giudici [Hun18b]. give [Mac17]. given [Dol15f, Haj11]. glass [Gau12a, Wil12, Lev14j]. Glaz [Lev10e]. Glen [Rua11, Rua14b]. Glenn [Hop19]. glorious [Lev14h]. GM [BS12a, Shi11a]. Goats [JJ17]. Gödel [Hoa15a, Ken14, Lor17e, And17]. gods [Rea10b]. Godsil [Cam18]. goes [Lev13d, PR12]. Going [AZY12, Per12]. Golberg [Hun14a]. Gold [Tal13, Fos18]. Gold-Bug [Tal13]. Goldbach [Mac11c]. golden [ASHH18, Gho11, Gri13e, Gri15d, Lev14h, Lor17a, Sci19, Mol17]. golden-ratio [Gri13e]. Golf [Min12, Kan17, vdV12, Hai13]. Gómez [Rou16]. gonal [ES11]. Gong [Hop18a]. gons [Lor12i]. Goodman [Eat12, Hun12b]. Goodman-Strauss [Eat12]. Goodstein [Hoa13b, PW13]. Gordon [Gri13d]. Goriely [Mac19]. Goro [Gib10c, Shi12c]. Gorroochurn [Hai17a]. Gösta [Stu10, Gib12a]. GP [Sco18d]. Graduate [App14a, Cou10, Gib13a]. Graf [Sch10]. Graham [Hop19]. grandaunts [Sof10]. Grandma [Sof10]. Graph [Kos14, Gri19]. Graph-theoretic [Kos14]. graphical [Wai09]. graphically [All18]. graphing [Lor10g]. graphs [BEH17, Mol12, Bev14, Bev17, Rua14a]. Grätzer [War13b, War13c]. gravity [Jew10]. Gray [Coo19]. great [And13d, Bay14d, Dav18, FW11, Gri12f, Hah12, Lev13h, Pet09, Ste13a, Ste13b, Tol18d]. greatest [Ron07, Wil10]. Greek [Lo 13]. Greeks [Gra13]. Gregor [Lor13d]. Gregory [Hud11b]. Greiser [Hal19]. gridpoint [Ste18b]. Griffiths [Bay10b, Hun13b]. Grinberg [Yeo17d]. Gross [Shi17b, Tol15c, Tol17i]. Group [HS13, Kil10, Hun18c, Pur19, Shi14e, Tol19b, VV09, Ano14q, Sur10, App11]. Groups [CSC10, FGR14, Bea13, Har10a, Hun17b, Hun17d, Hun18b, Sch11, Wil09, Eat13, Hun14b, Pat17, Shi11e]. Growing [Ano18e]. Growney [Lev10e]. GST [Cou17b]. Gua [Kha13]. guard [Gri11c, Mic09]. Guerraggio [GG12c]. guessing [Haw12, Mah10a]. Guest [Cou10]. Guide [Ano10a, Fie07, Hoa18a, Iva09, Job13c, Jon19, Lor11f, Rou14, Sto12b, Zar10b, Hal10a, Hal12b, Sto12b]. guitar [RR11]. Gunning [Bay19a]. Gupta [Cou17b]. Gustafson [Lev13g]. Guthery [Rua14c].

H [And10a, And10b, Bay15a, Fre10a, Gri14d, Rua12, Zar11]. Haese [Hal15a, Hal15b, Pat17]. Hahn [And13d, Shi18b]. Haigh [Bay14c]. Halbeisen [Shi14c]. half [Job11b, WO11, O’L10]. half-angle [WO11]. halfway [AZY12, Per12]. Hall [Bay11b, Cou13, Hun17b, Sch10, War10c, FFGH09, Ros09]. Hall/CRC [Cou13]. Halpern [Mac18]. Hamilton [McG13]. Hampton [De 14b]. hand [Sur10]. handbook [ABD11, Job13c, Rou14, Sch13]. hands [Rei14]. hands-on [Rei14]. Handshaking [Cri10]. Hanoi [Bay10a, FF16, GG08a]. Hansford [Mah10b]. happening [CMZ93, Lor14f, Mac11d]. happiness [Haw18b]. hard [Bay15b, Bay17b, Bay17a, Bay17c, Bay19a, Bev17, Cam18, Cou15, Cou17a, Cou18, Cri17a, Dav17, Dav18, De 17, Fox17, Fre15, Gib17a, Gib17b, Gib18b, Gib19c, Gri17c, Gri18e, Gri18c, Had19, Hai17a, Hal17a, Haw18b, Hoa15a, Hoa15b, Hoa17a, Hoa19, Hor17, Hun16, Hun17b, Hun17e, Hun18a, Hun17f, Lor15h, Lor16c, Lor16b, Lor18d, Mac15b, Mac15c, Mac16b, Mac18, McB18, Pas18, Rua17, Sch16c, Shi15b, Shi17b, Shi18b, Shi18a, Shi19c, Sto17a, Tin17, Tol14c, Tol15a, Tol16a, Tol17c, Tol17g, Tol17f, Tol17h, Tol17i, Tol18a, Tol18b, Tol18c, Tol18d, Tol18f, Tol19a, Tol19d, Tol19c, Yeo16, Yeo18b]. Hardback [And13b, And13d, Ban10a, Bev14, Cou13, De 10b, De 11b, De 14c, For11, GG12c, Gri10b, Gri13d, Gri13c, Gri13f, Gri14f, Hoa11, Hoa12, Hud13b, Hun12b, Hun12a, Hun13d, Hun13c, Lev12a, Mac11b, Rua13, Rua14a, Shi11e, Shi11d, Shi14c, Cou10, De 10a, Gib10c, Gib14c, Hud11a, Lev14k, Rua11, Shi10c, Sut10]. hardcover [Lev14b]. Hardy [And10a, And10b, Kan17, vdV12]. Hari [Bay17c]. Harmonic [VV09, Ayd18, BT17, Jam15b, Jam17, Kha10, Nur14a, Sco11a, App11]. harmonies [KK11, Lev13b]. Harris [Tol16a]. Hartshorne [Meg12]. Harvard [Ban13, Gib11b, Tol17k]. Hassler [Gib19b]. hating [HJS11, Hun13d]. Hauling [Sim12]. Havil [Cri15b, Gri13f]. having [Lor14h]. Hawkes [dV14]. hbk [And11, And13c, App10b, App10a, App11, App14a, Bay14a, Bev13, Eat11, Eat12, Eat13, For12, For13, Gib10b, Gri14e, Gri14b, Hai10, Hai12, Hai13, Hai15, Hoa13a, Hud10b, Hud13a, Hun13a, Hun14b, Hun15, Jac14, Lev11a, Lev12c, Lev13e, Lev14e, Lev14i, Lev14n, Lev14a, Mac13c, Mac14b, Mac14c, Meg12, Shi10g, Shi12c, Shi13c, Shi14b, Shi15c, Shi15a, Sto12c, Sut11, Tol14b, Tol15b]. heads [Jon19]. health [Kro18]. heart [CC10, Fre13]. heaven [Gus12, Lev13g]. Heavenly [Van13, Rua14b]. heavens [Rua11, Van09]. Heidi [Eat12]. height [Lor12h]. heights [Jos19a]. Heinz [Sch14]. Heisenberg [App11, VV09]. Helge [Hoa11, Hun16]. help [Fre10b]. hemispheroid [De14]. Henry [Gib13a, Had19]. Hensel [Hun14a]. Hepp [Sch10]. heptadecagon [Shi16c]. Herbert [Tol16e]. Hermite [Zho11]. Heroes [Ale10, Gib11b]. Heron [BS15, Jos16b, Lin11, Lor13a].

Heron-like [Jos16b]. Heron-type [Lor13a]. Heronian [Dol16b, Dol19a, Rea16]. Herrmann [Ban14a, Shi14e]. Hersh [Hun13d]. Hess [Gib17c]. Hexaflexagons [GG08a, Bay10a]. hexagon [Kri18, dVS18]. Heyne [De 10a]. Hichem [App18]. Hidden [KK11, Bay15b, Bro12a, Fer14, Fre13, Lor15g, Maz14, YN10, YN12, Lev13b]. Hidetoshi [Lev11d]. Higgins [Bay19b, Cri13]. High [Abe16, Ran18]. Higham [Hun17e]. Higher [Ano10a, Hal17b, She10a, Ber10, Haj18b, Hal12b, She10b, Sto12b, Sto12a, Tol19g, Hal17b]. higher-dimensional [Haj18b]. HighPerception [Lev11e]. highways [Eat11]. Hilbert [B^{+}11, FS10, App10a]. Hill [Tol16b, Tol18c]. Hind [Dav10]. Hindustan [Cou14]. Historical [BCSS19, Hud13a, Lor14e, Sta11, War10d]. History [Jac12, And13c, Bag11, Bea12d, Gib19c, Gut11, Hai17a, Hoa14a, Hoa15b, KP14, Lev14o, Lor15g, Maz14, Qua12b, Ram09, Rua11, Rua14c, Rua17, Ste12a, Van09, Dav17, Eat11]. Hitchcock [De 11b]. hitchhiker [CF15]. HL [Pat17]. hoard [Lev11c, Ste09b]. hoc [Sco18c]. Hodel [Hoa14b]. Holden [Cou18, Hoa11, Hun16]. Hölder’s [Raz19]. hole [FF19, Lev18b, Sim18, vdV12]. Holland [Gib17a]. Holm [Cou12]. Holmes [Den14]. Holton [Bay12, Lor14d]. homothetic [Nic17]. Honor [Sto10b, GDDR08, Sut11]. Honsberger [Haj17a]. hook [GL11]. hook-length [GL11]. Hopkin [GG10a]. Hopkins [And13a, Bay17b, Dav15, De 14b, Gri11c, Had19, Hai13, Hun10b, Hun10a, Lev13d, Lev15a, Mac16c, Shi15b, Shi18b, Tol15d, Tol17f, Zet18, Cra07]. Horn [Shi18a]. Hornby [Tol19e]. Horsman [Tol15e]. Horton [Gri17c]. hours [Hun18c]. Houston [Lev10b]. Howard [Hud10b]. Hoyle [Cla18]. Hugh [Tol19b]. Hughes [Cri17a]. humanity [Ste12c]. humans [Hal16]. Humour [Mac12c]. Humphries [Hal15a, Hal15b]. hundred [BM19, Sta19a]. Hunger [Hoa15b]. Hunt [Hal11a]. Hurlbert [Hop19]. hyperbola [GD14, Sco13b, Sco15b, Sco16, Sco18e, Sco19d]. hyperbolas [Mit12]. hyperbolic [Haj17c, Lam19]. hyperexponential [Bor10a]. hypersolids [Tay07, Rot10a]. hypotenuses [HJ19].

Ian [Bay14d, Dav18, Gri14f, Gri14a, Haw11, Hop18b, Lev10d, Lev11c, Sto17b, Tol10, Tol15f]. Icon [Gri13c]. ideal [Wil17]. ideas [Cri07, Gri14b, Lev10a, Mac12a, Tol18d, War10a]. Identifying [Sci19]. Identities [Gri16a, Cag19, Gau11, Gri13e, Kos12a, Ste17d, Ste18a, Ton19]. identity [Den10, Gle15, Gri16b, Gri17b, JJ13b, Nic15, Ton17]. If [CT14, Shi14a, Hop16]. Iggy [Cri15a]. III [Lor18e]. illogical [Hoa19]. illustrate [Lor17h]. Illustrated [Jac12, All18, Bea12d, Hoa14a]. illustrations [De 10a]. image [DL12]. images [Lor15f]. imaginary [Rea10b]. Imhausen [Dav17]. IMO [Kum11]. implication [Nie10]. implies [MP11]. imply [Lor15k]. importance [Jos18a]. improper [Den13, Lev19c]. Improved [Mur17, Shi18c, KK10]. Improving [Mur19]. inarc [Luk19]. incentre [GOD18a, Haj12, Haj17d]. inclined [Lor18j]. incomplete [Jam16a]. Indefinite [Del19]. Index [Ano10f, Ano11e, Ano12e, Ano13d].

India [Plo09, Cou11]. Indian [Rua13, Ste12b, YM11]. indifference [Wap19]. indivisibility [Cha11]. induction [Bay15c]. inductive [BS12a, Cag18]. industrial [Gri14d, LPE^{+}12]. industry [De 14c, Tob12]. Inequalities [Cve12, Jam14c, Mer10, MS13, Bea11, Bra10, CD11, Dol19b, IS12, Jam14e, Kac17, Kan09b, Luk18, PF10, Shi11d, Shi13c, Sik12, Hir12, Mac12b, Yeo17b, Shi14d]. Inequality [Luk17, Shi11a, Fin14b, JLM15, Kra15, Kum11, Lor11b, Lor13b, Lor17c, Lor18k, Luk19, Raz19, Ret18, Sco14f, SL17, Sle12, Vîj19]. inertia [LY11, Lor14a]. infinie [Dol11a]. infinite [Bra17, Cou12, Des12, HTSE09, Kac18, OJN16, Sco11c]. infinitely [Mac13f, McL19a, Mac13f]. infinitude [Har14]. infinity [Gri13d, Hoa12, Kil10, LG10, Orm08, Ste13b, Sti10, Lee17, Hop18b]. inform [And12]. Information [Ano18b, PS14, Ved10, War11a]. ingenious [Gri14b, Mac12a]. Ingleside [Kil10]. Ingleside-Ashby [Kil10]. Inglis [Hal17a]. Ingmar [Lev14e, Lev14h, Lev14p]. inner [Mac14d, YN10, YN12]. innovation [Ano12d, Dol15b, Rou13a]. Inprasitha [Hal17c]. inradius [Lor15d, O’L10]. inscribed [Fra19b, GG12b, Mat18b, Nic14a, Nun13]. Insert [Kul13a]. insight [Ste12c]. inspirations [Muk10, Nai12]. inspired [Ton17]. Inspiring [Neu16]. Instruction [GG10b, M^{+}08]. instructive [Lor11b]. Integer [Di 13, Kin10, Lor10b, Osl12a, Rea14, Rea15, ZF10]. integers [AM11, Bur10b, Cer16, GG12a, JLU10, Lor15a, MM15, MGC19, SB19]. integrable [Cou10, Gue08]. Integral [Kre14, Che16a, Che16b, Che18, Den13, Geo10, Hir11c, JL13, JJ13a, Kra15, Lor12j, Mac14c, R^{+}12, Shi15c, You11c, Bea11]. Integrals [JL17, JJ13b, Jam15c, Jam15e, Jam16b, Lev19c, Lor12b, Mer13, Pla10b, Ste17e]. Integrating [FN13, Tra10, Mur11c]. Integration [Hos17b, Ayd18, Del19, DO12, Hos17a, Jam13c, Kön12, Mur11b, Mur19, Sch14, Sch16b, KK17, Bay17c]. Interactive [Gib11a]. Intercepts [Cae15]. interest [Far13, Mah19b]. interesting [Hir11c, Kos12a, Lor10a, Smy10, Ste17e, WO11, WO13]. intermediate [Job13c, Rou14, Mor19]. International [Ano14a, GG10b, M^{+}08, WILY11, AR14, BHJ10, Hal11a, Lev17f, Sto14b]. interplay [De 17]. interpolation [HR17, Mcn16]. Interpreting [Ken14, Hoa15a]. intersecting [Mah18a]. intersection [Pat13a]. interviews [AA08, Hud11a]. intriguing [Die19, HK14, Hun12a, Red15, Sco14f]. Introduction [Ano06, Bra10, EGH^{+}11, FGR14, HS13, JH12, Mac14c, MMC09, Nic12a, RJPG12, R^{+}12, Sut09, ÅM08, And10a, And10b, App10b, Bay19a, Hal12a, HWHBS08, Hig11, JS08, Lev14o, Lor16b, Lor18d, Mac10b, Mac10c, Mac10a, Rei14, Ric08a, Shi14c, Shi14e, Ste12a, Ter09, Vel09, Yeo18a, Bay14c, Cri13, Hop18b, Mac11b, Mac19, Row19, Bev17, De 14b, Gri14e, Gri18d, Hun14a, Mac12b, Mac14c, Shi10g, Shi15b, Tol19c, Zar11]. Introductory [App18, Tol18b]. intuitionist [Pas14, vDD13]. intuitive [Fox17]. Invariant [Gao08, Gao09, Cou14, War10c]. invariants [GW09, Hun12b]. inverse [FS10, Gla19, Lor11j, Sco11d, Ste17a]. inverses [McL19b]. inverting [SNH11]. investigations [Gri17d]. invitation [Hun18a]. involving [Ara15, Den10, Gri15b, Jam14d, Lor14j, Lor18i, Mah18a, dVS18]. Ioachimescu [Jam12, Sîn10]. Ipswich [Rot10a]. Iris [Tob12, De 14c]. irrational [AT12, Gri12b, GM15, Lor15l, Lor17h, Lor18a, Lor19a, MGC13b, Shi14a, Gri15c, Jac11, Kom10]. Irrationality [Zho11, Lor18c, Raj18]. irrationals [Gri13f, Hav12]. Irwin [Hal11c]. Isaac [Mac14a, Pas13, Rua12]. ISBN [And10a, And10b, And11, And13c, And13a, And13b, And13d, And19, App10b, App10a, App11, App14b, App14a, App18, Ban10a, Ban10c, Ban10b, Ban13, Ban14b, Ban14a, Bay10c, Bay10b, Bay10a, Bay11b, Bay11a, Bay12, Bay14a, Bay14b, Bay14c, Bay14d, Bay15b, Bay15a, Bay17b, Bay17a, Bay17c, Bay17d, Bay19b, Bay19a, Bev13, Bev14, Bev17, Cam18, Cou10, Cou11, Cou12, Cou13, Cou14, Cou15, Cou17a, Cou17b, Cou18, Cou19, Cri13, Cri15b, Cri15a, Cri17a, Dav10, Dav15, Dav17, Dav18, De 10b, De 10a, De 11b, De 14c, De 14b, De 17, Eat11, Eat12, Eat13, For10a, For10b, For11, For12, For13, Fos14b, Fox15, Fox17, Fre10a, Fre10b, Fre15, Gib10a, Gib10c, Gib10b, Gib11b, Gib11c, Gib12b, Gib12a, Gib13b, Gib13a, Gib14b, Gib14c, Gib17a, Gib17c, Gib17b, Gib18a]. ISBN [Gib18b, Gib19b, Gib19c, Gol10, GG10a, GG10b, GG12c, Gri10b, Gri11d, Gri11e, Gri11b, Gri11c, Gri12f, Gri13a, Gri13b, Gri13d, Gri13c, Gri13f, Gri14f, Gri14a, Gri14e, Gri14b, Gri14c, Gri14d, Gri17c, Gri17d, Gri18d, Gri18e, Gri18c, Had19, Hai10, Hai12, Hai13, Hai15, Hai17a, Hai17b, Hal10c, Hal10a, Hal10b, Hal11b, Hal11c, Hal11a, Hal12b, Hal13, Hal14, Hal15a, Hal15b, Hal16, Hal17a, Hal17b, Hal17c, Hal17d, Hal19, Haw11, Haw12, Haw18b, Haw18a, Haw18c, Hem19, Hoa10, Hoa11, Hoa12, Hoa13a, Hoa14a, Hoa14b, Hoa15a, Hoa15b, Hoa17a, Hoa18a, Hoa19, Hop18a, Hop18b, Hop19, Hor17, Hud10c, Hud10b, Hud10a, Hud11a, Hud11b, Hud13a, Hud13b, Hun10b, Hun10a, Hun12b, Hun12a, Hun13d, Hun13c, Hun13b, Hun13a, Hun14b, Hun14a, Hun15, Hun16, Hun17a, Hun17b]. ISBN [Hun17c, Hun17e, Hun18b, Hun18a, Hun18c, Hun19b, Hun19a, Hun17f, Jac14, Jac19, Jon19, Kee18, Kil10, LP10, Lea10, Lee17, Lev10d, Lev10f, Lev10e, Lev10a, Lev10b, Lev10c, Lev11b, Lev11d, Lev11a, Lev11e, Lev11c, Lev11f, Lev12a, Lev12d, Lev12b, Lev12c, Lev13e, Lev13c, Lev13b, Lev13h, Lev13a, Lev13g, Lev13f, Lev13d, Lev14i, Lev14n, Lev14o, Lev14k, Lev14l, Lev14j, Lev14c, Lev14h, Lev14p, Lev14b, Lev14m, Lev14a, Lev14d, Lev14f, Lev14g, Lev15b, Lev15a, Lev17g, Lev17h, Lev17a, Lev17b, Lev17c, Lev17e, Lev17d, Lev17f, Lev17i, Lev18d, Lev18c, Lev18b, Lev19b, Lor10e, Lor10d, Lor10c, Lor11g, Lor11h, Lor11f, Lor11i, Lor13d, Lor14f, Lor14e, Lor14d, Lor15g, Lor15i, Lor15h, Lor16c, Lor16b, Lor17g, Lor17e, Lor17f, Lor18f, Lor18d, Lor18e, Mac10b, Mac10c]. ISBN [Mac10a, Mac11b, Mac12b, Mac13e, Mac13a, Mac13c, Mac13b, Mac13d, Mac14b, Mac14c, Mac14d, Mac14a, Mac15a, Mac15b, Mac15c, Mac19, Mac16a, Mac16b, Mac16c, Mac17, Mac18, Mah10b, McB18, Meg12, Mor19, Mun12, Nai12, Pas14, Pas18, Pat17, Rot10c, Rot10d, Rot10b, Rot10a, Rou13a, Rou14, Rou16, Row19, Rua10, Rua11, Rua12, Rua13, Rua14c, Rua14a, Rua14b, Rua17, Sch10, Sch13, Sch14, Sch16b, Sch16c, Shi10e, Shi10f, Shi10c, Shi10i, Shi10g, Shi10d, Shi10h, Shi11e, Shi11d, Shi12c, Shi13c, Shi14d, Shi14b, Shi14c, Shi14e, Shi15c, Shi15a, Shi15b, Shi16b, Shi17a, Shi17b, Shi18b, Shi18a, Shi19c, Sto10b, Sto10a, Sto12c, Sto12b, Sto12a, Sto14a, Sto14b, Sto14c, Sto17b, Sto17a, Sut10, Sut11, Ten12, Tin17, Tol10, Tol13, Tol14b, Tol14a, Tol14c, Tol15b, Tol15e, Tol15c, Tol15a]. ISBN [Tol15d, Tol15f, Tol16a, Tol16d, Tol16c, Tol16b, Tol16e, Tol17a, Tol17b, Tol17c, Tol17d, Tol17e, Tol17g, Tol17f, Tol17h, Tol17i, Tol17k, Tol17j, Tol18a, Tol18b, Tol18e, Tol18c, Tol18d, Tol18f, Tol19a, Tol19d, Tol19g, Tol19b, Tol19c, Tol19f, Tol19e, Tsa18, Von17, War10a, War10c, War10b, War11a, War11b, War12a, War13a, War13b, War13c, Wil10, Yeo16, Yeo17a, Yeo17b, Yeo17c, Yeo17d, Yeo18a, Yeo18b, Yeo18c, Zar10b, Zar10a, Zar11, Zet13, Zet18, Zet19]. isles [Qua13]. isoperimetric [Fin14b]. ISSN [App14b]. issue [Ano10g, Ano10h, Ano10i, Ano10k, Ano10j, Ano10m, Ano10l, Ano11f, Ano11h, Ano11g, Ano11j, Ano11i, Ano11l, Ano11k, Ano12f, Ano12h, Ano12g, Ano12j, Ano12i, Ano12l, Ano12k, Ano13e, Ano13g, Ano13f, Ano13i, Ano13h, Ano13k, Ano13j, Ano14h, Ano14g, Ano14j, Ano14i, Ano14l, Ano14k, Ano15f, Ano15e, Ano15h, Ano15g, Ano15j, Ano15i, Ano16d, Ano16c, Ano17h, Ano17g, Ano17j, Ano17i, Ano17l, Ano17k, Ano18h, Ano18g, Ano18j, Ano18i, Ano18l, Ano18k, Ano19g, Ano19f, Ano19i, Ano19h, Ano19k, Ano19j]. Istituto [GG10b]. Italiana [GG10b]. Iterated [Ste12b]. Iterating [HE19, Rou13b]. Ivanov [Lor11f]. Iverson [Mac13b]. Ivy [And19].

J [And13d, App10b, Ban10a, Bay15a, Bay17b, Bev14, Cri17a, Gib10a, Gri17d, Hal11c, Haw18a, Hoa15b, Hor17, Hud10c, Hud10a, Hud11a, Hun17e, Hun17f, Lev10c, Lev14b, Lor15g, Lor16c, Mac10b, Mac11b, Mac12b, Mac13b, Pas14, Rua14a, Shi14c, Shi14e, Shi18b, Sut10, Tol16b, War13a, dV14, vDD13]. J.-F [Gib10a]. Jackson [Hoa14a]. Jacob [Hun17a, Ber10]. Jacqueline [Lev14o]. Jain [Hun15, Mac14c]. James [Hal10c, Hun19a, Tol17g, Zet13, FMW14, Mac15b]. Jan [Lev10e, Lor18d]. Jänānarāja [Knu14b]. Jane [Von17]. January [How19]. Janus [Roc14, Hai15]. Janus-Faced [Roc14, Hai15]. Japanese [Lev11d, FR02, FR08, Sil18c]. Jason [Bay11b, Bev13, Bev14, Kee18, Rua14a, Tol17b]. Jayne [Gri11d]. Jean [Gri11e, Shi10i]. Jean-Marie [Gri11e]. Jean-Pierre [Shi10i]. Jeanne [Eat11]. Jedrzejewski [Tol14a]. Jefferson [Gol10, Rot10b]. Jenny [Tol17j]. Jensen [Mer10, MS13]. Jewish [Tol14c, BEU12]. Jim [And13c]. Jin [Fox17]. Jinho [Sch13]. Jñānarāja [Dav15, Knu14a]. JoAnne [Lev10e]. Jobbings [Mor19, Rou14, Tol13, Tsa18]. Joel [Gib10b]. John [Dav15, De 14b, Eat12, For11, Gri14b, Gri17c, Hal10a, Hal14, Hoa12, Hun13d, Lev11d, Lev11a, Lev14b, Lev14g, Lev15c, Lor17e, Mac14c, Ten12, Hav14, Osl10a, Cri15b]. John-Steiner [Hun13d]. Johnny [Gib19c]. Johns [And13a, Bay17b, Gri11c, Had19, Hai13, Hun10b, Hun10a, Lev13d, Lev15a, Mac16c, Shi15b, Shi18b, Tol15d, Tol17f, Zet18]. Johnson [Tol14a]. Johnston [Lev12a]. Jolliffe [Hal11a]. Jonathan [Gib11c]. Jordan [Mac10c]. Jorge [Hun13a]. José [De 17, Rou16]. Jósef [Gib17a]. Joseph [Tol17e]. Joshua [Cou18, Hop19]. Jost [Hun17c, Lev14a]. journey [FFGH09, PL12, Sch10]. joy [Lor11c, Tol18e]. Joyner [De 14b]. Jr [Shi14e]. Julian [Cri15b, Gri13f]. Juliette [Hoa15a]. July [Ano14a, How16, Ano15d, Ano19c]. Junior [Tsa18]. Jürg [FFGH09, Sch10]. Jürgen [Hun17c]. just [MGC13a]. Justifying [Lor15c].

K.Peters [Gib14b]. Kabus [Hal11c]. Kahn [Lor16c]. Kai [Mac10a]. kaleidoscope [Pla10c]. Kallianpur [App14a]. Kaniuth [Lor10c]. Kannappan [Shi11d]. Kaplan [Lev13b]. Kaprekar [Dol11b]. Karen [Cam18, Hoa15b]. Karl [Lev13g]. Karonski [Bev17]. Katz [Cri17a, Hoa15b]. Kaye [Ten12]. Keedwell [Gib17a]. Keith [Gib19b, Gri14e, Lev10f, Tol17d]. Kemper [Lor13d]. Ken [Tol17h]. Kendig [Gib19b]. Kennedy [Hoa15a]. Kenney [Gri17d]. Kepler [BGGGR18, li17]. Kerr [Qua12c]. Kevin [Lev10b]. Khajuraho [Ber15]. Kicking [Kuc15]. Kiepert [Fox16, GD14, Oai15, Sco13b, Sco15b, Sco16, Sco18e]. Kilday [Tol13]. Kilmister [Sil10]. Kim [Cou11, Gib12b]. kind [Cra09, Hal10b]. Kingdom [Bay10b, Fox15, Gri18d, Hun13b, Mac12b, Sto10a, Zar10a]. Kirkwood [Hun19a]. kissing [Ara15]. Kitty [Tol13]. Kiyoko [Fox17]. Klaus [Sch10]. knockout [Gri10a]. know [Cri07, Yeo18a, Lev10a, War10a]. knowledge [De 17]. Knudsen [Dav15]. Ko [Cam18]. Kohar [Tol17a]. König [Sch14]. Koninck [Gri11e]. Körner [Tol15a]. Kortenkamp [Gib11a]. Kovalevskaya [Aud11, Gib13b]. Kranat [Gri11d]. Krantz [Bay14a]. Kress [Shi15c]. Kristian [Hun17a]. Kristopher [Hun17d]. Krulik [Hal17d, Lor17f]. KS3 [Hal11b, MC09]. Kuldeep [Mac16a]. Kumar [Gri18e]. Kung [Shi17a].

L [GG10b, Hoa13b, Hud11a, Mac14c, Mac15c, PW13, Shi14e, Sut11]. L. [vDD13, Pas14, Tol17g]. labyrinth [Mac18]. ladder [Ber10, De13]. ladders [GS11]. Lagrange [HR17]. Lakshmibai [Cou14]. Lalanne [Sta19b]. Land [Lev12e, Ban14b, Gla19]. Lane [Rua10]. Lang [Mac13e, Mac13b]. Langham [Hal14]. Langley [Che19]. language [Gla19]. Laplacian [Bev14, Rua14a, Mol12]. Lara [Fos14b, Sto14a]. largest [Ano14r, DZH14, Tod15]. last [Bar17]. late [Gib11c]. Latin [Lo 13, McL19b, Gib17a]. Lattice [Mat18a, Zor10]. lattices [Bea12b]. Launching [KN19]. Laura [Bev13]. Law [CO11, Zor11]. lawn [Mat16]. laws [BGGGR18]. Leadership [Jon19]. leading [Gor19b]. leads [Lor11c]. leaps [Rua13, YM11]. learn [Hal16]. Learning [And12, AR14, Sto14b]. least [Sco17e]. Leaving [Wil14]. Lebesgue [PG17]. Lecture [Shi17a]. Ledwick [Hal17b]. Leffler [Gib12a, Stu10]. Legacy [Hav14, Cri15b, Sut10]. Lehmann [Gri13a, Lev14e, Lev14h, Lev14p]. Lehmer [Gra13]. Lehmus [ASH19a, Bea19, CR14, Dac16, Haj17c]. Leibniz [You12]. Leila [Lev13f]. lemma [Che19, PG17]. lemniscate [Osl16]. lends [Sur10]. length [Gor18, GL11]. lengths [Haj17b]. Lennart [Cri17a]. Leo [Rua17]. Leon [Gib12b, WMW10]. Leonard [App10b, Gib14c, Jac19]. Leone [Ten12].

Leonhard [Fel07, HHPM07, De 10b, De 10a, Hoa17a]. Lerner [Sch16b]. Lery [Gri14d]. Lesmoir [Gri13d]. Lesmoir-Gordon [Gri13d]. Lesne [Gib10a]. Less [Dol16b, Cag18, Mat18a, Shi19d]. Lesson [Hal17c]. Lester [Sco17a, Sco18f]. Letter [Pas16, Dev08]. Letts [Hal10a, Hal12b]. Level [Gol10, Hal10c, Hal13, Rot10b, Dav10, Hal14, Hin08, Lor11i, Nic08a, Tol14d, BL12, Pea11]. Leversha [Gri18d, Sto10a, Fox15]. Levi [And11, Ban14b]. Levinson [Mer10, MS13]. Levitin [Gri14c]. Levy [Mac16b]. Lewis [Wil08, Rua10]. Li [Tol18e]. Library [Sut11]. licence [Gib11a]. Lie [Hun14b, Hun17b, Sch11, Hun17b]. Lieven [Tol19c]. Life [Hav14, Cri15b, De 14c, FMW14, Gri14f, HJS11, Iva09, Lor11h, Lor11f, McG13, Shi08, Soi09a, Ste11b, Tob12, Wil08, vDD13, Gib10c, Hun13d, Rua10]. lifesaver [Yeo17d]. lifetime [GDDR08, Sut11]. like [Gri11f, Hou09, Jos16b, Lev10b, MMYA10, Osl10b, Rei14, Tol15b]. limacons [VM19]. Limit [San11, Dow15, GR13, Jam14b, McA10, Pla16]. Limit-free [San11, McA10]. limiting [Kac18, Koo17, Sco18c]. Limits [App12, Bay14b, Kum10, Tol18c, Bay14b]. Lindegaard [Dav15]. Lindgren [War11b]. line [Ban18, Gri15a, GOD18a, Lor13i, Mez14, Sco11b, Sco17e, Smi15, Qua12b, Lev17h]. Linear [Dav12, FK19, HF04, Hun10a, Kre14, Kro18, Pat17, Sco16, ADLM18, App10b, FH05, Fer16, Gri17a, Mah10b, Ric08a, Shi18a, Tol19c, Mac13d, Mac16a, Shi15c]. lines [Fox16, GG12a]. link [Sub19]. Linker [Yeo17c]. Linnebo [Pas18]. Lionel [Gib12b]. Listener [CMS18]. literature [Lev15a, Tub14]. little [Kar18]. Liu [Hun14a]. lives [B^{+}11, Dav18, GG12c, KK11, Lev13b, Ste12c]. Lobb [Kos12b]. Locally [Osb13, Lor15h]. Locating [Mah19b]. Lockhart [Ban13]. locus [Shi13a]. log [Pur11]. Logarithmic [Die19, Mer17]. Logarithms [Hav14, Bur10c, Cri15b]. logic [AF12, Ari12, Hoa13b, Hoa19, Lee17, LHS14, Shi14b, Tol14b, Hoa14b, Hoa18a]. logical [Hal17a, Nie10, Rua10, Wil08]. logsine [Lor12j]. London [App11]. loneliness [Ber15]. Long [ASHH18, Kil10, Orm08]. Lonsdale [Hal10a, Hal11b, Sto12b, Sto12a]. look [Jam14b, Rob12, Sco14e, SNH11]. Looking [Ano17f, JJ14, McC15, AW07, Gau12a, Sto10b, Wil12, Lev14j, Tol14a]. looking-glass [Gau12a, Wil12, Lev14j]. Lord [Ano14r, Ano14o, Hir11e, Osl11b]. Lorenz [Shi14c]. lose [Geo15]. lost [Sim11a]. Louise [Kee18]. Love [Fre13, GG08c, Lev10e, Lev14d]. Loveland [Hoa14b]. Lovett [Hun13c, Mac13c]. Loving [HJS11, Hun13d]. lower [Sch16c, SS11]. Ltd [App11, Ban10a, Hoa12, Tol10]. Lucas [Di 13, ES16, Gau11, Gra13, Gri13e]. Lucchetti [GG12c]. Lulu [Ban10b]. lulu.com [Yeo18a]. Lunes [Sew14]. Luoshu [Sut10]. Lyness [Dol17b].

M [And10a, And10b, App14b, Ban10a, Cou14, Gib14c, GG10b, Gri10b, Gri12f, Gri14d, Hai17b, Hoa13a, Lev12a, Lor15h, Mac10b, Shi15b, Tol17k, Tol18b]. M.J [Gri14d]. M.S [dV14]. M1 [Gol10, Jef08b]. MA [Lev18c, Lor14d].

MAA [Gib19b, Lor18e]. MacCluer [Shi10d]. MacCormick [Gri14b]. Macey [Tol19e]. MacHale [Lev11e]. Mackenzie [Lor14f]. Maclaurin [Jam15b, Sik12]. Macmillan [Gri13b, Mac13a]. Maday [Gri14d]. made [Dev08, Jam14g, LP16]. Madhav [Cou14]. MAG [Ano10g, Ano10h, Ano10i, Ano10k, Ano10j, Ano10m, Ano10l, Ano11f, Ano11h, Ano11g, Ano11j, Ano11i, Ano11l, Ano11k, Ano12f, Ano12h, Ano12g, Ano12j, Ano12i, Ano12l, Ano12k, Ano13e, Ano13g, Ano13f, Ano13i, Ano13h, Ano13k, Ano13j, Ano14h, Ano14g, Ano14j, Ano14i, Ano14l, Ano14k, Ano15f, Ano15e, Ano15h, Ano15g, Ano15j, Ano15i, Ano16d, Ano16c, Ano17h, Ano17g, Ano17j, Ano17i, Ano17l, Ano17k, Ano18h, Ano18g, Ano18j, Ano18i, Ano18l, Ano18k, Ano19g, Ano19f, Ano19i, Ano19h, Ano19k, Ano19j]. magic [Bay10c, Bea17a, BT09, Cra09, EH11, Hal10b, Sud13, Sut10, War10b]. Magnificent [Pas13, PL13, Lev14e, Mac14a]. Mahajan [Haw12]. Mahima [Shi15a]. Maitree [Hal17c]. Make [Pur11]. Making [BY08, Iva09, For10b, Lor11f, Yeo18b]. Mallik [Gri18e]. Mamikon [Ste17c]. Man [HHPM07, Rua13, De 10a, Gib12a, Stu10]. Manfrino [Rou16]. manifolds [Lov10, Mac13c]. many [Abb08, ALB10a, CK19, Hop12, Lew13, Mac13f, McL19a]. Maor [Lev14a, Tol19d]. map [Gib10c, Shi08, Tol16d]. Mapp [Hal11b, Hal12b]. mappings [Fos12b]. marbles [Aud14]. Marc [Hun10a, Mac15b]. Marcel [Lev12b]. March [Gib12b, Ano14f, Ano18f, Ano19d]. Margalit [Hun18c]. Margaret [Gri17d, Aul19]. Maria [Gri14c]. Marie [For10b, For13, Gri11e]. Marjorie [Fre15, Lev10e]. Mark [And11, Ban14b, Gri13b, Hal15a, Hal15b, Mac13a, Mac13b, Wil10]. Markov [CLM13, Hoa18b]. Marshall [De 14b]. Martin [Bay10b, Bay10a, Cou10, Hun13b, Sut11, GG08a, GDDR08, GG08b]. Martinez [Lev14n]. Marty [Lev13d]. Martyrs [Ale10, Gib11b]. Marvelous [Muk07, LP10]. marvels [Ste13a]. Mary [Hal17b, Tol13, Tol14b, Ari12]. Maryam [Hoa17b]. Masha [Gri13c]. Mason [Ten12, dV14]. masses [Hol17]. Master [Sto14b]. Masterclass [AR14]. masterpiece [Pas13, Mac14a]. Mastronardi [Hun10a]. matematiche [B^{+}11]. material [Sto17a, Hal14]. materials [BL12]. Math [Ada10, Pas11b, PR12, SC13, Adl14, Bay11b, Bay15b, Bev13, Cra10, Fer14, Fre13, Lev14n, Mar12, Min12, Ros09, RT11, Tol17b, Tol17e, Lev14d, Tol17g, Haw18c, War12a, Lev13c, Lev13d, Tin17]. Mathematical [AA08, Ano14a, Ano14m, Ano17m, Ano17n, Ano17o, Ano18n, Ano18o, Bay10c, Bay12, CD11, Cou13, Eat11, Gar13, Gib19b, Gol17, GG10b, Gri11e, Gri12f, Hah12, Hal17b, Hun10b, Hun14a, Hun17d, Jon19, JS08, Lev11b, Lev14k, Lev18c, Lev18b, Lev19b, Lor11g, Lor11f, Lor11i, Lor11c, Lor14d, M^{+}08, PR10, Swe12, Tol15e, Tol15g, Tol18c, Vel09, Yeo16, Zar10a, Ada09a, And11, And13b, Ano12d, Ano13l, App12, Bay10a, Bay14b, Bay14d, Bay15a, BCSS19, Ben07, Ben12b, BS12b, Bor10b, CMZ93, Cra09, Cri07, DLP08, FFW13, For11, Fre10a, GG08a, GG08b, Ges11, Gib10a, Gib12b, Gla19, Gri11b, Gri11c, Gri11h, Gri13c, Hai17a, Hal10b, Hal17a, Haw11, HJS11, Hoa10, Hoa13a, Hoa13b, Hoa17a, Hoa18a, Hun12a, Hun13d, Job13c, Lan12, Lev10a, Lev10c, Lev11c, Lev14g, Lev17h]. mathematical [Lev09, Lo 13, Lor11h, Lor14f, Lor15g, Mac10c, Mac13e, Mac11d, Maz14, McB18, Mic09, Neu11, Pas11b, Pas08, Pla11a, PL12, Rou13a, Rou14, Rua10, San09, Shi10h, Smi08, Soi09a, Ste08, Ste09a, Ste09b, Ste13a, Ste13b, Ste14, Sto12c, Sto14c, SA16, Sut11, Szp10a, Tol10, Tol15d, Tol15f, Wap12, War10a, WMW10, Wil08, Yeo18b, Ano18m, B^{+}11, PW13, Tub09, Gib19c, And13a, And13d, Ano10f, Ano11e, Ano12e, Ano13d, Ano18b, De 17, GG12c, Hud11a, Hun17c, Hun19a, Mac10c, Mac10a, Shi13c]. Mathematically [Lor11g, Ada09b, Lev17f, MBS10, Ten12, Hal16]. mathematician [Gus12, Hou09, Lev10b, Lev13g]. mathematicians [BEU12, FW11, Pet09, Seg09, Tol14c, Dav18, Gib10b, Gri12f, Hai17b, Lev13h]. Mathematics [Alc12, Ale10, Ano06, App11, App14b, App14a, Bay10b, Bea12d, Bor10b, Cou10, Cou12, Cri11, Den14, Dol15b, FRW11, Fox15, Gib13a, Gri18d, HHHS14, Hal15a, Hal15b, Hun13b, Jac12, Lev17e, Lev18d, Mac12b, Mac19, Mac12c, MNS10, Mor19, Orm13, Pat17, Plo09, Soi09b, Spr09, Ste11b, Sto10a, Tsa18, Tub14, WILY11, Wri09, Zar10a, dS13, Ada10, AW07, Alc13, And12, AR14, Ban10a, Ban10c, BY08, BY11, BRR08, Cou18, Cra07, Cri17a, DSZ08, Dav10, De 11b, Dev08, Duf11, Eat11, For10b, Fos14b, FR02, FR08, Fyf11, Gib11b, Gib17b, Gib19b, GG08c, Gol10, Gol17, GG10a, GG10b, GD14, Gut11, Had19, Hal10c, Hal11b, Hal12b, Hal17b, Hal17c, Hal19, Haw12, Her11, HJS11, Hin08, Hoa12, Hol11, Hol13b, Hop19, Hou09, Hud13b, Hun13d, Imr10, Iva09]. mathematics [JM09a, JM09b, Kra12, LPE^{+}12, Lev10e, Lev11d, Lev12d, Lev14l, Lev14m, Lev15b, Lev17g, Lev17b, Lev18b, Lip13, Lor11h, Lor11f, Lor11i, Lor17e, LHS14, Mah10a, M^{+}08, Mun12, Nel08, Nic08a, Orm08, Pas11b, Pit11, Pit12, PM13, Pit14, PL13, Ram09, Reb12, Rei14, Ron07, Rot10b, Rua11, Rua14c, Rua14b, She10b, SD10, She10a, Soi09a, Ste12a, Ste13a, Sti10, Sto10b, Sto12b, Sto12a, Sto14a, Sto14b, Szp10a, Szp10b, Tob12, Tol16d, Tol19g, Van09, Van13, War13a, War10d, War12b, War12c, Wel08, Wel12, YM11, ZH09, vDD13, Pat17, Lev10e, Hal15a, Hal15b, Ban14a, Bay12, Bay14a, De 14c, Gib11c, Gri11b, Gri14d, Hal17d, Hud10a, Hud13a, Hun13a, Hun17e, Hun18a, Hun17f, Jon19, Kil10, Lev11d, Lev11f, Lev12a, Lev14e, Lev14o, Lor10e, Lor14d, Lor16b, Lor17f]. mathematics [Mac15a, Pas18, Rua13, Sto17b, Tol13, Tol16b, Tol17a, Tol17f, Tol18e, Wil10, Yeo18a, dV14, Ban10b, Bay11a, Cou11, Dav17, Gri14f, Hoa14a, Lev11b, Lev13e, Lev15a, Lev17c, Pat17, Sto12c, Sto17a, Tol16a, War13a]. mathématiques [Bat14, Lor15i, Yeo17a]. Mathews [Lev17e]. Maths [Ach11, Ano10a, Hal13, BL12, Bre11, Cha05, Cha07, Gri13b, Hal10a, Hal14, Hud11b, Mac13a, MC09, Pea11, PF08, Lev13f]. MATLAB [Duf11, Mun12]. Matrices [VvBM08, Bea17a, Bev14, EH11, Har10a, Hun10a, McL19b, Mol12, Rua14a]. Matrix [VvBM08, ABD11, Bev14, BB12, FS10, Lor15f, Mol12, Rua14a, Sch13, Yan15, Hun10a, Hun17d]. Matsunaga [Fox17]. Matt [Hun18c]. matter [Ano10g, Ano10h, Ano10i, Ano10k, Ano10j, Ano10m, Ano10l, Ano11f, Ano11h, Ano11g, Ano11j, Ano11i, Ano11l, Ano11k, Ano12f, Ano12h, Ano12g, Ano12j, Ano12i, Ano12l, Ano12k, Ano13e, Ano13g, Ano13f, Ano13i, Ano13h, Ano13k, Ano13j, Ano14h, Ano14g, Ano14j, Ano14i, Ano14l, Ano14k, Ano15f, Ano15e, Ano15h, Ano15g, Ano15j, Ano15i, Ano16d, Ano16c, Ano17h, Ano17g, Ano17j, Ano17i, Ano17l, Ano17k, Ano18h, Ano18g, Ano18j, Ano18i, Ano18l, Ano18k, Ano19g, Ano19f, Ano19i, Ano19h, Ano19k, Ano19j]. matters [Sim12]. Matthew [Hal17a]. Matthews [Tol17c]. Maximal [MM15, KN19]. maximisation [Lor13h, Hir11b]. Maximisation/minimisation [Hir11b]. MaxProduct [Der12]. Maxwell [FMW14, Mac15b]. May [Aul19, Bur10b]. maze [Haw11, Ste10]. Mazur [Tol17e, Lor15g]. Mazzotti [Tol19a]. McAllister [Lev12a]. McCabe [Zar11]. McCartney [Mac15b]. McGovern [Cri15a]. McGregor [Gib11c]. McKay [Tol16e]. McOwan [Gib18a]. MD [De 14b]. me [Cla18]. Meagher [Cam18]. mean [App11, Jam14a, Lor10b, Lor14h, Mer13, Mer17, Pla16, Shi16d, VV09, OC14]. mean-square [Shi16d]. meaning [Sut10, Tub14]. means [Gor19b, Jam18b, Lor11b, Sco18d]. Measure [Kön12, Bay17c, Sch14]. Measurement [Loc12, Wil18, Gib17b, Ban13]. Measuring [Col11a]. mechanic [Lev09, And11]. mechanical [Tod15]. Mechanics [Hol17, Jef08b, Cou12, HTSE09, KL11, Luc10, Luc13, Raj13, App10a, Mac14b, Gol10]. medial [HE19]. mediant [Gut11, Rua14c]. medieval [Cri17a]. medley [Gri11b, Szp10a]. Meenakshi [LP10, Nai12]. Meeting [WILY11, CF14]. meets [HB16, Mer17]. Mehrmann [Gri14d]. Méla [Gib10a]. Melluish [Tol16d]. Members [Hal10b, Bay12, Lev18c, Lor14d, Tol15e]. memoir [Cla18]. memoirs [Gus12, Lev13g]. Memoriam [Ano10e]. men [Cra07, GG10a]. ménages [KK18]. Mendoza [Fos11, Mil12]. Menghini [GG10b]. Menso [Cri17a]. Mental [Lip13]. Mersenne [Cam16]. Meta [Cha05, Cha07, Hud11b]. metacentre [Rob12]. metamathematical [Pas16]. Metcalf [Jon19]. method [Dol11a, DO12, Hos17a, Kir13, Kul16, Shi15e, Sik18, McN14]. methods [App10a, BHJ10, Hal11a, JJ13a, Lan12, Mac13e, Sch16c, Yeo17b]. Metric [Cul11, Jos17b, Jos18b, Shi10g, Sut09]. metrics [Gib13a, Tau11]. Michael [Cou13, Gri11c, Hal15a, Hal15b, How19, Hun18b, Mac16b, Pat17, Tol16a, Von17]. Michal [Bev17]. Michel [Eat13, Lor15i, Sch16c]. Michele [Sch10, Gib13b]. microscope [Bor10b, Sto12c]. might [CT14]. Mikkel [Yeo18c]. Miklós [Gri18c]. Mikusiński [Lor18d]. Millman [Lor16c]. mind [Gri17c]. mine [Mat16]. Minimal [Dol18a, Jos15, MM12]. minimisation [Hir11b]. minimum [Sco14b, Sco18f]. Minton [Hai13]. Miquel [De 14a]. Mircea [Lev12d, Lev14l, Lev14m, Lev15b, Lev17g, Lev18d, Yeo16]. mirror [DL12]. mirrors [SNH11, Vil11]. Mirzakhani [Hoa17b]. misattribution [Qua12b]. Miscellaneous [Sco10b]. miss [McL19a]. missing [Osl12b, Shi11f]. mistakes [Lev14e, PL13]. Misunderstandings [Tol14d]. Mitchell [Hal11c]. Mittag [Gib12a, Stu10]. Mittag-Leffler [Gib12a, Stu10]. Mixing [Kos16]. model [CP10, Kan17, Tol19f]. modeling [Mac10a, Vel09]. modelling [BS12b]. models [Hud10a, Kos14, Kro18]. Modern [Ale10, AA13, Ber10, Dev08, Gib11b, Gri10b, Hai17a, Sch16c, Shi15a, Shi17b, Ten08, Tol17i, WW13, Lev14b]. modes [Lor11e]. modular [LP10, Muk07, Pur19]. modules [Hun15, JST12]. Mohammed [App18, Yeo17a]. Mohan [Rua13]. Moivre [Dou13]. Molitierno [Bev14, Rua14a]. moment [Gib19b, Lor14a]. moments [Bag11, LY11, Lor14i, And13c]. money [Gib17b]. monkey [Gro18]. Monographs [App11, App14b]. monotonocity [VSR13]. monster [Ron07, Wil10]. Monty [Bay11b, Ros09]. Moore [Zar11]. Moriarty [Den14]. Moritz [Tol14c]. Morley [Mac11a, Sco13a]. Morphing [Osl11b]. morphology [Sim19]. morsel [Haj17a]. Mortad [App18]. most [Bau14, Bay11b, Bea12c, Bev13, Lor16f, Lor18f, Ros09, RT11]. mother [Gri10b, Ten08]. motif [Gut11]. Motifs [Rua14c]. motion [Ben12b, Ell17, And13b]. Motivation [Ban10c, Wel08]. move [AW07, Sto10b]. moves [Rob13]. movies [PR12, Lev13d]. moving [DL12]. Mr [Cra07, GG10a]. Mrs [Hun12a]. MSOR [Yeo18a]. much [Shi19d]. Mukerji [LP10, Nai12]. multigrade [Pla11b]. multimagic [Kee11]. Multinomial [Gri10a]. multiple [Sub15b, Tol15g]. multiples [Jam18d, Lip13]. multiplication [Pis16]. multiplicative [Col13a, ES10, Pur19]. multiplicative-type [ES10]. Multisection [BD16]. multivariate [Kac17, Shi16b]. Mumford [For12]. Murray [Mac10b]. Murthy [Cou14]. Music [Ben07, Tol17f, Wri09, Lev11b, Lev10c, Tol19d]. Musielak [Lev14k]. my [Gib10c, Hop16, Shi08, Tol17h]. mysteries [FW11, Ste13a, Ste14, Tol15f]. mystery [San09, Shi10h]. mystic [Cri15a, McG13]. myth [Tol17e]. myths [HJS11, Hun13d, Lev14n, Mar12].

N [Hai10, Lev17e]. Nadis [Mac14d]. Nahin [Hun12a]. Naked [Whe13, Lev14f]. Nandakumaran [Cou19]. Napier [Cri15b, Bur10c, Hav14]. Napoleon [Oai15]. Nathalie [Sto17a]. National [Mer15, Rou13a, Sto14c, Kro18]. Nattygrafix [Rot10a]. natural [Jac11, Mol17, Ran18]. Nature [Gar18, Ada09a, Die19, For11, Kha10, Kha12, Rua12, Tha05, And19]. Nature-nurture [Gar18]. Nautical [Tol16b]. Neale [Tol18a]. Necessary [Nur14b]. need [Cri07, Hol19, Lev10a, Pas11b, War10a]. Needed [Lev15e]. needle [Lor19e]. needlework [BY08, For10b]. negative [Jam18a, Vîj19]. Neill [Tol19b]. Nelsen [Lor18e]. Nelson [Lor10e]. Nesbitt [SL17]. nested [Dow15, Osl10b, OJN16, Sco14a]. Network [Sto10b, Yeo18a]. Neuenschwander [Mac16c, Zet18]. Neumann [Hoa13a]. Never [Gib19b]. Neville [Hal11a]. Newton [Mac14a, Pas13, Rua12, Che16b, Hol17, Lor11d, McN14, Mcn16, Rea10b, Tha05, Rua12]. Newtonian [Tol14b, Ari12]. Neyman [Gri13a, Leh11]. Nice [Joh11]. Nicholas [Hun17e]. Nicholson [Gri14e, Hal10c]. Nick [Ano14r, Ano14o]. Nicola [Hun10a]. Nicolas [Sch16b]. Niels [Hun17a]. Nielsena [Yeo18c]. Nigel [Gri13d]. Nimmo [Gib11c]. Nina [Hal17a]. Nine [Mac12a, Fra19a, Sco19a, Sud11]. nine-point [Fra19a, Sco19a, Sud11]. Nint [Dol14a]. No [Gib13a, Hal13]. Noether [Gri10b, Zet18, Ten08]. Nolan [Hun12b]. nomogram [Sta18, Sta19b]. nomograms [Hop13]. Non [Bur17, Hal10b, Bay12, HJ18b, Kro18, Lor14k, Tol15e, ZF10]. non-calculator [Lor14k]. non-collinear [HJ18b]. Non-Desarguesian [Bur17]. non-existence [ZF10]. non-linear [Kro18]. Non-members [Hal10b, Bay12, Tol15e]. Norbert [Ban14a]. Nori [Cou14]. normal [Col19]. normals [Gla13, Sco11b]. Norman [Gib17b]. North [Cri17a, Gib17a]. North-Holland [Gib17a]. Norvaisa [App14b]. Not-on-the-line [Gri15a]. notation [Lor15g, Maz14]. Note [GOD18b, Col11b, GB12, Gib19a, Gri17a, Haj17d, Jam18a, Jam18b, Tor17]. Notebooks [MGC13b]. Notes [HR17, Har14, ALB09, ALB10b, Bor10b, Lev17i, Shi17a, Sto12c, War11b]. novel [JO14]. November [Ano19e]. Number [HS13, Ano13m, AT12, AG14, Col19, Egg15, Gri12d, GM15, Gri17d, Hun10b, Shi10e, Shi10f, Shi14e, Shi17b, Sur10, Tol15e, Tol17i, Tub09, Vil07, Zor10, Gri17d, Shi14e]. Numberland [Rua10, Wil08]. Numbers [FGR14, Hig11, Jac12, Szp10b, And10a, And19, Bay10c, Bea12d, Bre15, Cag19, Col11b, Dol15a, ES11, Eul15, ES16, Gau11, Gau12b, GB12, Gib19c, GL11, Gri13f, HWHBS08, Has19, Hav12, Hoa10, Hoa14a, Hoa18b, Hop16, Jac11, Jam10b, Jam11a, Jam14g, Jam19b, Jár16, JJ14, JLU10, KK10, Koo11, KG11, Lev13f, Lor14b, Lor14l, Lor16c, Lor16g, Lor18g, Mac17, Min12, Mur12, Pas08, Pis16, RG14, SE13, Sao17, SC13, Shi12a, Shi19b, Ste11a, Sub19, Sur18, Tsa11, War12c, WO11, Zet13, dK09, And10b, Bay17d, Gri11e, Gri18e, Hai13, Lev13a, Lev17a, Rua17, Tol14a, Tol16e, Tol19d, Cri13, Hud13b, Tol16c]. numeracy [Fie07, Fre10b, Zar10b]. numerical [BC13, Cou15, Cou17a, Cou17b, Mur19, Mur11b]. Nunnally [Gib12a]. Nuno [War12a]. nurture [Gar18]. NY [Tol10].

O [App10a, Lor11f]. O.M [How19]. object [Hol13a]. oblique [Osl11a]. oblique-angled [Osl11a]. observations [Lor18j, Gri12a]. Obtaining [Luc10, Kac18]. obtuse [GOD18a]. occurrence [Mol17]. octet [Sew15a]. odd [FM10, Jam10b, JLU10, Mil11, NL16, Ste19a]. odyssey [Pas08, Hoa10]. offering [Ben07, Lev10c]. Office [Hun18c]. OIH [Jos19d]. old [Dol12c, Rea10b]. Oleg [Hun14a]. Olympiad [Ano14a, Zar10a, Smi08, Lor11d]. olympiads [Job13c, Yeo16, Rou14]. omega [Cha05, Cha07, Hud11b]. once [Sco17a]. One [Haj17d, Mar11, Cag18, Gri15f, NG15, Ron07, Shi11c, Shi17b, Tol17i, Wil10]. online [Gib11c]. Ono [Tol17h]. Open [BEH17, Haj15b]. opening [Vil11]. operational [AM15]. opportunistic [Haw12, Mah10a]. optimisation [GG13, Kac17]. Optimising [Jew10]. optimization [Gro11]. Optimum [Mcn16]. oracles [Shi19b]. Orbits [Jam14g]. order [Abe16, Hir11h, McN14, Suc11, Sut10]. ordered [Gri12c]. Ordering [Aud14]. orderings [Egg15]. orderly [Pop12]. Ordinary [Cou19]. Orieami [Mac13b]. Origami [GG08b, Lan12, Muk10, Sut11, WILY11, Ger08, Muk07, For10a, LP10, Mac13e, Nai12]. origin [Kul13c]. originaux [Bat14]. Origins [Tub09, Lo 13, Hun10b, Tol15d]. Ormell [Kil10, Bur10a]. Ortega [Rou16]. orthic [Nic17]. orthocentroidal [Sco14e]. Orthogonal [Fos12b]. Orzel [Lev14c]. Osborne [Lor15h]. Oscar [Bay15b, Haw18b]. O’Shea [Bay10c, Lor17g]. Other [Lev12e, Ada09b, Ban14b, Gib17c, Gri11c, Haw11, Hun12a, Lor11g, Mic09, Pla11a]. our [Shi18c, Ste11a, Ste12c, Zet13]. outer [Pur18]. outscribed [Ano14r, DZH14, Tod15]. oval [Tol19a]. overcome [Pas11b]. Owen [Bay10c, Lor11i, Lor17g]. own [Pur11]. Oxford [And10a, And10b, And13c, And13b, App14a, Bay11b, Bay14b, Bay14c, Bay19b, Bev13, Cou10, Cou12, Cri13, Dav10, De 11b, Fos14b, Fre15, Gib13a, Gib17b, Gib18b, Gol10, Gri11d, Gri14c, Hal10c, Haw11, Hop18b, Hun15, Hun17f, Kee18, Lev12a, Lev13e, Lev14i, Lev14o, Lev14b, Lev14g, Lor18f, Mac10c, Mac14b, Mac15b, Mac19, Mac16a, McB18, Rot10d, Rot10b, Row19, Rua17, Sch13, Shi10e, Shi10f, Shi10g, Sto14a, Sto17b, Tol14b, Tol18a, War10b, War11a, Wil10, ABD11, FFW13, Lev14g]. Øystein [Pas18].

P [App14a, Cou19, Hal11c, Hun17a, Lev17a, Lor16c, Mac10c, Tol18f, Yeo18a, Zar11, KK10]. package [Gib11a]. Packing [EM12, EM16]. Padlock [Ano13m, Hor14, Tol15e, Tol15e]. Page [Tol19f]. pairs [Di 13, Gri18b]. Palgrave [Gri13b, Mac13a]. Palindromic [Har12]. Pamfilos [Haj18a]. Panaretos [Hai17b]. Panning [Fos18]. panorama [Lev17f]. Paolo [Hai15]. Papa [Mac17]. Paper [Leu18, And19, App18, Bay15a, Bay17d, Bay19b, Cou17b, Gib17c, Gib18a, Gib19b, Gri17d, Gri18d, Hai17b, Hal15a, Hal15b, Hal16, Hal17c, Hal17d, Hal19, Haw18c, Hem19, Hoa18a, Hop18a, Hop18b, Hun17a, Hun17c, Hun17d, Hun18c, Hun19b, Hun19a, Jac19, Jon19, Lee17, Lor17h, Mac15a, Mac19, Mac16a, Mac16c, Mac17, Mor19, Pat17, Rou16, Row19, Sch16b, Shi16b, Shi17a, Sto17b, Tol15e, Tol15c, Tol15d, Tol15f, Tol16d, Tol16c, Tol16b, Tol16e, Tol17a, Tol17b, Tol17e, Tol17k, Tol17j, Tol18e, Tol19g, Tol19b, Tol19f, Tol19e, Tsa18, Von17, Yeo17a, Yeo17b, Yeo17c, Yeo17d, Yeo18c, Zet18, Zet19]. Paper-folding [Leu18]. Paperback [De 11b, Gib11c, And10a, And10b, Hop19, Tol16d, Tol16b]. papers [BY08, Cou14, SB^{+}12]. Papillault [Gib10a]. Pappus [Ste17c]. parabola [Gla13, Hoa18c, KF19]. parabolas [HJ18b, Sha15]. paradox [Bea18b, Bel19]. Paradoxes [Lev12e, Ban14b, Bay10a, GG08a]. Paradoxical [FN13]. parallel [Osl11a]. parallelogram [Lor15e]. parameters [BS15, Mah19b]. Parametric [Sew15b]. parametrisation [Hos17b]. parenthesisation [Kos12b]. parity [Lor15l]. Parshall [Hoa15b]. part [Tay07]. Partial [Tol19h, Vol11, Hun19a, Jac14, Osl12a, Mac16b]. particular [Gri12a, Gri12e, Gri19]. Partition [Der12, Che16a]. Partitioning [Shi12d]. partitions [Gri12c, Gri18a, CQ19]. parts [Del19, DO12, Gri18a, Shi19d]. party [Shi12a, Bel19]. Pascal [Bay10b, Bro12b, Bro12a, Dev08, Fos12c, Fra19b, Gri08, MMYA10, Shi12b, VM19]. Pascal-like [MMYA10]. Pascal-points [Fra19b]. Pask [Lev13c, Mac14a]. Pasles [Hoa10]. Pasquale [Lev17e]. pastures [Mah18c]. path [Ban18, Lor13i]. paths [Lor11n]. Patrick [Rot10a]. Patsy [Mac13b]. Pattern [MD10, Haj11, For12]. patterns [Die19, Mas14, Pas11b, Lev17b]. Paul [Ban13, Gib18a, Hoa10, Hun12a, Jon19, Mac18, Shi14e, Sto14b, Ano14n]. Paulo [Mac17, Mac17]. Pavaman [Cou14]. Pavel [Hun14a]. Pbk [Cou12, App14a, Gib11c, Hud10a, Mac14b, Rou14, Shi10e, Shi10f, Shi14d]. Pearcy [Bay17c]. Pearson [Ten12]. peculiar [Gle15]. peg [Lev18b]. Peiffer [Eat11]. Pell [Dol12d, Eul15, ES16, Kos13, SB19]. pencil [Bev13, RT11]. pendulum [DW19, Mur17]. Penelope [Hal11a]. Penguin [Lor10e, Nel08, Lor10e]. pension [SA16]. pentagon [Fos18, Lor16d]. people [AA08, Hud11a]. Perfect [Ges11, Lew10, Gri13c]. Perfectly [Lev12c]. perimeter [CF15, Gus15, Jam14c]. period [Ada10, War13a]. periodic [App11, VV09]. periods [Jam18d]. Perkins [Hun12a]. permutations [Haj11, Wei12]. Persi [Tol18d]. person [SS12]. perspective [BHJ10, CD11, Cut14, Hal11a, Shi13c]. perspectives [AR14, FMW14, Sto14b]. Peter [Ban10b, Bay19b, Cou15, Cou17a, Cri13, Dav10, Hoa13a, Hun14b, Lor16b, Pat17, Sto12b, Sto12a]. Peters [Ban10a, Eat12, For10b, For12, For13, Gib10a, Gib10b, Gri10b, Hoa12, Hud10c, Hud11a, Hun13c, LP10, Lev10e, Mac13c, Mah10b, Nai12, Shi10i, Shi14e, Sut10, Sut11]. Petković [Gri12f]. Phi [Lev14h]. Philippe [Sch13]. philosopher [vDD13, Pas14]. philosophy [LHS14, Rua12, Tha05, Pas18]. Physical [Lev12e, Ban14b, JS08, Lev09, Mac10c, Tre17]. physics [Adl14, FFGH09, Hun19a, Min12, Sch10, Hun12a, Mac16c]. PI [Shi11d, OW14, Hem19]. Picking [Kul16]. Picturing [Wai09, Hud10b]. pieces [FJ10, Gri11b, Szp10a]. Piecing [Mah15b]. Piene [Hoa11, Hun16]. Pierre [Shi10i]. Pietro [Shi13c]. pillars [Tol17k]. Pini [De 10a]. pioneer [Gib19b]. pioneering [Lor18f]. Piotr [Lor18d]. pipeline [Ano14f, Ano15d, Ano18f, Ano19c, Ano19d, Ano19e]. PISA [And12]. Pitici [Lev12d, Lev14l, Lev14m, Lev15b, Lev17g, Lev18d]. Pitot [Jos19b]. Pittsburg [Lev14n]. Plain [Lev17d]. plainer [Mat18a]. planar [Ban16]. plane [GG13, Lev17d, Lor16d]. planes [Bur17, Lor18j]. planet [Jew10]. planetary [And13b, Ben12b]. planets [And13b, Ben12b]. planning [Pas11a]. Plato [Hud13b, Szp10b]. play [Cri17b, Gri17c, Mat16]. plc [War10a]. Plofker [Cou11]. plus [Hal10a, Shi17b, Sto12b, Tol13, Tol17i, dV14]. Poddar [Mac14c]. poems [GG08c, Lev10e]. point [Ban10c, Dow15, Fra19a, Sco12, Sco17e, Sco17d, Sco19a, Sud11, Wel08]. points [Abe19, Bra11b, Cae15, CK19, Dol12c, Fra19a, Fra19b, GD14, HJ18b, Mcn16]. polar [Lor12b]. polarisation [Mah10c]. policy [PdVF14]. Polster [Lev13d]. polygon [CO11, Job15]. polygons [Bar13, De 11a, Fox17, Kin10, She16, Ste17d, Ste18b, Wil18, Zor10]. Polyhedra [CC18, Tay07, Fox17]. Polyhedral [She13, FJ10]. polyhedron [CO11, Ric08b]. polynomial [Cae15, Chu18, CK19, Gri11a]. Polynomials [FGR14, Fin14a, Gau11, Gri16a, Har12, McA10, Nic11, Sur13, Vil14b]. pons [Haj15a]. poor [SS12, War12b, Lev14i]. Pope [dV14]. Popko [Gib14b]. popular [Bev13, RT11, War12c]. Popularising [Haj10]. population [BS12b, Gib17c]. Posamentier [Hal17d, Lev14e, Lev14h, Lev14p, Lev17h, Lor17f, Tol18e]. posent [Bat14, Lor15i]. positive [Jam18a, Nur14b, Vil14b, Yan15]. postage [dV14]. pouring [SV15]. power [BD15, FM10, Gau11, Gib18a, Gor19b, Gri12a, Gri13d, LG10, Lor10g, Wap19, Baj13]. powers [Bur10b, Cer16, Der15, Der19, Lor14b, Lor15g, Maz14, Sub15b, Tre17, Tsa11, Blu11, EH11, JLU10]. Pp [And10a, And10b, And11, And13c, And13b, And13d, App10a, App11, App14b, Ban10a, Ban10c, Ban13, Ban14b, Ban14a, Bay10c, Bay10b, Bay10a, Bay11b, Bay11a, Bay12, Bay14a, Bay14c, Bay14d, Bev14, Cou10, Cou11, Cou12, Cou13, Cou14, Cri13, Dav10, De 10b, De 11b, De 14c, De 14b, Eat11, Eat12, Eat13, For10a, For10b, For12, For13, Fre10a, Fre10b, Gib11b, Gib11c, Gib12b, Gib12a, Gib13b, Gib13a, Gib14b, Gib14c, Gol10, GG10a, GG10b, Gri10b, Gri11d, Gri11e, Gri11b, Gri11c, Gri12f, Gri13a, Gri13b, Gri13d, Gri13c, Gri13f, Gri14f, Gri14a, Gri14e, Gri14b, Gri14c, Hai10, Hai12, Hai13, Hal10c, Hal10a, Hal11b, Hal11c, Hal11a, Hal12b, Hal13, Hal14, Haw12, Hoa11, Hoa12, Hoa13a, Hoa14a, Hud10b, Hud10a, Hud11a, Hud13a, Hud13b, Hun10b]. Pp [Hun10a, Hun12b, Hun12a, Hun13d, Hun13c, Hun13b, Hun13a, Hun14b, Hun14a, Jac14, LP10, Lea10, Lev10d, Lev10e, Lev10a, Lev10b, Lev10c, Lev11b, Lev11d, Lev11a, Lev11e, Lev11c, Lev11f, Lev12a, Lev12d, Lev12b, Lev12c, Lev13e, Lev13c, Lev13b, Lev13h, Lev13a, Lev13g, Lev13f, Lev13d, Lev14e, Lev14i, Lev14n, Lev14o, Lev14k, Lev14l, Lev14j, Lev14c, Lev14h, Lev14p, Lev14b, Lev14f, Lor10e, Lor10d, Lor11g, Lor11h, Lor11i, Lor13d, Lor14f, Lor14e, Mac10b, Mac10c, Mac10a, Mac12b, Mac13e, Mac13a, Mac13c, Mac13b, Mac13d, Mac14b, Mac14c, Mac14d, Mac14a, Mah10b, Meg12, Mun12, Nai12, Rot10c, Rot10d, Rot10b, Rot10a, Rou13a, Rou14, Rua11, Rua12, Rua13, Rua14c, Rua14a, Sch10, Sch13, Shi10e, Shi10f, Shi10c, Shi10i, Shi10g, Shi10d, Shi10h, Shi11e]. Pp [Shi11d, Shi12c, Shi13c, Shi14d, Shi14b, Shi14c, Sto10b, Sto10a, Sto12c, Sto12b, Sto12a, Sut10, Sut11, Ten12, Tol10, Tol13, Tol14b, Tol14a, War10c, War10b, War11a, War12a, War13a, War13b, War13c, Wil10, Zar10b, Zar10a, Zet13, And13a, And19, App10b, App14a, App18, Ban10b, Bay14b, Bay15b, Bay15a, Bay17b, Bay17a, Bay17c, Bay17d, Bay19b, Bay19a, Bev13, Bev17, Cam18, Cou15, Cou17a, Cou17b, Cou18, Cri15b, Cri15a, Cri17a, Dav15, Dav17, Dav18, De 10a, De 17, For11, Fos14b, Fox15, Fox17, Fre15, Gib10c, Gib10b, Gib17a, Gib17c, Gib17b, Gib18a, Gib18b, Gib19b, Gib19c, GG12c, Gri14d, Gri17c, Gri17d, Gri18d, Gri18e, Gri18c, Had19, Hai15, Hai17a, Hai17b, Hal10b, Hal15a, Hal15b, Hal16, Hal17a, Hal17c, Hal17d, Hal19, Haw11, Haw18b, Haw18c, Hem19, Hoa10, Hoa14b, Hoa15a]. pp [Hoa15b, Hoa17a, Hoa18a, Hoa19, Hop18a, Hop18b, Hop19, Hor17, Hud10c, Hud11b, Hun15, Hun16, Hun17a, Hun17b, Hun17c, Hun17d, Hun17e, Hun18a, Hun18c, Hun19b, Hun19a, Hun17f, Jac19, Jon19, Kee18, Kil10, Lee17, Lev10f, Lev14m, Lev14a, Lev14d, Lev14g, Lev15b, Lev15a, Lev17g, Lev17h, Lev17a, Lev17b, Lev17c, Lev17e, Lev17d, Lev17f, Lev17i, Lev18d, Lev18c, Lev18b, Lev19b, Lor10c, Lor11f, Lor14d, Lor15g, Lor15i, Lor15h, Lor16c, Lor16b, Lor17g, Lor17e, Lor17f, Lor18f, Lor18d, Lor18e, Mac11b, Mac15a, Mac15b, Mac15c, Mac19, Mac16a, Mac16b, Mac16c, Mac17, Mac18, McB18, Mor19, Pas14, Pas18, Pat17, Rou16, Row19, Rua10, Rua14b, Rua17, Sch14, Sch16b, Sch16c, Shi14e, Shi15c, Shi15a, Shi15b, Shi16b, Shi17a, Shi17b, Shi18b, Shi18a, Shi19c, Sto14a, Sto14b, Sto14c, Sto17b]. pp [Sto17a, Tin17, Tol14c, Tol15b, Tol15e, Tol15c, Tol15a, Tol15d, Tol15f, Tol16a, Tol16d, Tol16c, Tol16b, Tol16e, Tol17a, Tol17b, Tol17c, Tol17d, Tol17e, Tol17g, Tol17f, Tol17h, Tol17i, Tol17k, Tol17j, Tol18a, Tol18b, Tol18e, Tol18c, Tol18d, Tol18f, Tol19a, Tol19d, Tol19g, Tol19b, Tol19c, Tol19f, Tol19e, Tsa18, Von17, War10a, War11b, Yeo16, Yeo17a, Yeo17b, Yeo17c, Yeo17d, Yeo18a, Yeo18b, Yeo18c, Zar11, Zet18, Zet19, dV14]. Practical [FH05, HF04, Mah10b]. practice [Bor10b, MMC09, Zar11, Sto12c]. practices [De 17]. Prakash [Hai17a]. pre [Lor17i]. pre-calculus [Lor17i]. precise [Ada10, War13a, War13a]. preference [Egg15]. Preiss [Hal11c]. prescient [Wap19]. prescribed [Haj18b]. present [Szp10b, Hud13b]. preserve [Gri11h]. presidential [Bra08, Hud10c, Gol17]. Press [And10a, And10b, And11, And13c, And13a, And13b, And13d, And19, App10a, App14a, Ban10c, Ban13, Ban14b, Bay10a, Bay11b, Bay14a, Bay14b, Bay14c, Bay15b, Bay15a, Bay17b, Bay19b, Bay19a, Bev13, Bev14, Bev17, Cam18, Cou10, Cou11, Cou12, Cou17b, Cou18, Cou19, Cri13, Cri15b, Cri15a, Cri17a, Dav10, Dav15, Dav17, De 11b, De 14b, De 17, For11, Fos14b, Fre10a, Fre15, Gib11b, Gib13a, Gib14b, Gib14c, Gib17c, Gib17b, Gib18b, Gib19b, Gol10, Gri11d, Gri11c, Gri13f, Gri14b, Gri14c, Had19, Hai13, Hal10c, Hal16, Haw11, Haw12, Haw18b, Haw18a, Hoa10, Hoa14a, Hoa14b, Hoa15a, Hoa15b, Hoa17a, Hop18b, Hop19, Hud10b, Hud13a, Hud13b, Hun10b, Hun10a, Hun12a, Hun13d, Hun15, Hun17e, Hun18b, Hun18a, Hun18c, Hun19b, Hun19a, Hun17f, Kee18, Lea10, Lee17, Lev10c]. Press [Lev11d, Lev12a, Lev12d, Lev12c, Lev13e, Lev13b, Lev13a, Lev13d, Lev14i, Lev14n, Lev14o, Lev14l, Lev14b, Lev14m, Lev14a, Lev14g, Lev15a, Lev17a, Lev17b, Lev18d, Lor15g, Lor17e, Lor18f, Lor18e, Mac10c, Mac11b, Mac13e, Mac13b, Mac14b, Mac15a, Mac15b, Mac15c, Mac19, Mac16a, Mac16b, Mac16c, McB18, Mun12, Pas18, Rot10c, Rot10d, Rot10b, Rou13a, Row19, Rua11, Rua14c, Rua14a, Rua14b, Rua17, Sch13, Shi10e, Shi10f, Shi10g, Shi10h, Shi13c, Shi14b, Shi14e, Shi15b, Shi17b, Shi18b, Shi18a, Sto14a, Sto14c, Sto17b, Sto17a, Sut11, Tin17, Tol14b, Tol15b, Tol15c, Tol15a, Tol15d, Tol16a, Tol16d, Tol16c, Tol16b, Tol16e, Tol17d, Tol17g, Tol17f, Tol17i, Tol17k, Tol17j, Tol18a, Tol18b, Tol18d, Tol18f, Tol19d, Tol19g, Tol19b, Tol19c, Tol19e, Von17, War10b, War11a, War11b, War13b]. Press [War13c, Wil10, Yeo16, Yeo17d, Yeo18b, Zet18, Mac13d, Dol15b]. previously [Hem19]. Price [Cou14]. prices [GS11]. primality [Cha11]. Prime [Lor14l, BB12, Col11b, Gri12g, Has19, Hun13b, Sao17, SS11, Mac17]. primer [De 11b, Smi08, ZH09, Zar10a]. Primes [Dol17a, Cha11, Gri12g, Har14, Mac13f, MM13b, McL19a, Osl15, Sco11c, SLMRR10, Hun18b, Lor17g]. Primicerio [Gri14d]. Primitive [Rea15]. primorials [Gri15e]. Princeton [And11, And13d, Ban14b, Bay15b, Bay19a, Cou11, Cou18, Cri15b, Cri17a, Dav17, De 17, For11, Gri13f, Gri14b, Haw18b, Hoa10, Hoa14b, Hoa15b, Hoa17a, Hud10b, Hud13a, Hud13b, Hun12a, Hun13d, Hun17e, Hun18c, Hun19b, Lev11d, Lev12d, Lev13a, Lev14l, Lev14m, Lev14a, Lev15b, Lev17g, Lev17b, Lev18d, Lor15g, Lor17e, Mac10b, Mac11b, Mac15c, Mac16b, Pas18, Rot10c, Rua11, Rua14b, Shi10h, Shi14b, Shi17b, Tin17, Tol15b, Tol15c, Tol16a, Tol17d, Tol17g, Tol17i, Tol18d, Tol18f, Tol19d, Von17, Yeo17d, Yeo18b, AF12]. principal [Wil17]. Principia [Mac14a, Pas13]. principles [Cou19, Hal11c]. print [Gib11c]. printing [Had19]. Prism [Bro12b]. Pritchard [Jon19, Lev18b, Lev19b]. Prize [Hoa11, Hun16, HP10b]. probabilistic [Kha12]. Probability [Çin11, NS87, Roc14, Bay10a, Bea11, Che16a, Che16b, Che18, Dav12, GG08a, Hai15, Hor14, Kra15, Mac13d, Tol15e, Wio19, You11c, ML13, Pat17, Tol17j, Bay14c, Hai12]. probable [Bau14]. Problem [Ach11, Ano14q, Bay11b, EO12, Gib19a, Hol11, Kee13, L.10a, L.10b, L.11, L.15a, L.15b, L.15c, L.16, L.17a, L.17b, L.17c, L.18a, L.18b, L.18c, L.19a, L.19b, Lor10h, Lor11k, Lor11l, Lor12d, Lor12e, Lor12f, Lor13e, Lor13f, Lor13g, Lor14m, Lor14n, Lor14o, Lor19c, Ara15, Bar11, Bar13, Bay11a, Bay19b, Che19, CQ19, CF15, De 12, Dou13, GV15, Haj19, Hal17d, Hal19, HB16, Haw12, Hir11g, Hol13b, Job13c, Kos12b, Lor11d, Lor11c, Lor14c, Lor14d, Lor15m, Lor15n, Mah10a, Pat13b, Ros09, Rou14, Rou13c, SV15, Sim13, Soi09b, Vig19, dV14, Bea18a, Ret17, Rib19, Bay12, Lev17e, Lor17f, Yeo17c]. problem-solving [Hal17d, Hal19, dV14, Lor17f, Yeo17c]. problème [Bat14, KK18, Lor15i]. Problems [AD08, Cro10, Dol10a, Dol10b, Dol11c, Dol11d, Dol11e, Dol12e, Dol12f, Dol12g, Dol13b, Dol13c, Dol13d, Dol14b, Dol14c, Dol14d, Dol15c, Dol15d, Dol15e, Dol16c, Dol17c, Dol17d, Dol17e, Dol18c, Dol18d, Dol18e, Dol19e, Dol19c, Dol19d, Vol11, And13a, BCSS19, Cve12, FR02, Gor19b, Haj15b, Jac14, Kac17, Lam19, Lev17i, Lev09, Lor13h, Mac17, May16, Shi14d, Ste13a, Ste13b, Swe12, Yeo18c, Lev11d, Bay14d, Mor19, Sch16c, Tsa18, Lea10, Lor16c]. process [Sco18c]. processes [KS14, Sch16c, App14a]. Proctor [Hal10a]. product [Ban16, Gre19, Lev12f, Lor12a, Ökt15, Osl10a, Osl11b, OW14, OJN16]. products [Col13a, Des12, GB12, Kac18, Mac15d, Osl10b, Bea17a]. Professor [Ste08, Ste09a, Ste09b, Ste14, Bay10c, Bay19b, Tol10, Tol15f, Lev11c]. proficiency [Fre10a]. Profile [Bay14d, Gri14f, Gri14a, Hem19, Tol10, Tol17c]. profiles [AA08, Hud11a]. Progress [Fyf11, Tol13]. Progressing [Imr10]. progression [AM11, Dol19a]. progressions [CQ19, Gri12g, Lor15d]. Proizvolov [Nic15]. Project [Dol15b]. projectile [KN19]. Projectiles [Ell17]. projections [Tol16d]. projective [Smi15]. projects [BY08]. Prometheus [Lev13c, Lev14e, Lev14h, Lev14p, Lev17h, Lev17c, Lor17g, Mac14a, Tol18e]. Promislow [Shi10c]. Proof [Cag18, Edg16, GV17, Lor11m, Mah13b, Pla18, SV15, SS16, Ste17a, Sub15b, Sub18b, BS12a, Dac16, Gla19, Gle15, Haj15a, Haj18a, Har14, Jam10a, Jam13c, Jam15d, Kar18, KA19, Lor15l, Lor16f, Lor18c, Lor18a, Lor19a, Nil13, Pla16, Raj18, Rib18, Rib19, Shi11a, Sti10, Sub14a, Sub18a, Tod15, Ton16, Ton18, Wei12, Wil12, Wil17, You11b, You11a, You12, You13, Zor10, Cha15, DeM10, GV15, PF10, Sub15a, Hal19, Hoa12]. Proofs [Ber18, Cag19, Job18b, ASH19b, IS12, Lor15a, Lor16c, Mac13f, MM13c, Mer10, SSJ19, Zho11, Job11a, Lor18e]. Properties [GG12b, Jos16d, Jos17e, Rou11, FS10, Fox16, Gri11f, Haj12, Haj17d, KG11, Mah10c, SE13, Sco10b, li17]. property [ES10, Gri18b, Gri19, Mil11, Mit12, Nic17]. prophecies [War12b]. prophesies [Lev14i]. proportion [Gri15f, Szy19]. proportionality [Lam19]. protagonists [B^{+}11, GG12c]. proved [PF10]. Proving [Ber15, GG13, Yeo17b]. pseudo [Haj10]. pseudo-congruence [Haj10]. pseudoprimes [Jam11b]. Ptolemy [Wil12, Lev14j]. Publications [Lev14j, Sch14, Kön12]. published [Hem19]. Publishing [Bay17d, Gib11c, Gri17c, Gri17d, Hoa18a, Hop18a, Hor17, Hun17a, Lev11d, Lev17i, War10a]. pull [Hol13a]. pupils [Lev18c]. purchases [Tol15g]. purpose [Tol14e]. Pushing [Tol18c]. Putnam [CQ19]. putt [Lor12h]. puzzle [Coo19, FF19, Gri12g, RT11, Bev13, Hun13b]. Puzzles [Lev12e, Bay10a, Cut14, GG08a, GDDR08, Grä11, Gri12f, LL11, Pet09, Sut11, Ban14b, Gib17c, Gri14c, War13b, War13c]. pyramidal [Edg16, Pis16]. Pythagora [Shi10h]. Pythagoras [Lev14n, Gri15h, Luz16, Mar12, San09, Siu13, Sub18a]. Pythagorean [Lev13b, Bea12a, BD12, HJ18a, HJ19, Jos16d, KK11, Kos16, MvdB12, O’L10].

Q [Zet19]. Quadling [How15]. Quadratic [Jam13a, Dol18b, Fer16, Leu17, Shi10a, Shi12c, Yan10, Fos12d]. quadrature [Hoa18c, Szy17]. quadrilateral [Bea12c, Bea17b, Fra19b, Jos15, Jos17a, Jos17d, Jos19a, Lor13b, Sil17]. Quadrilaterals [Lor12g, ASH19b, Fin14b, Fra19b, Jos16c, Jos16b, Jos16d, Jos17c, Jos17b, Jos17e, Jos19c, Leu12a, Lor14j, PU11]. quadrisection [Sco16]. Quadritangent [Fox14]. Quadros [Gri14d]. Quand [Bat14, Lor15i]. quantum [And13c, App10a, Bag11, Cou10, FFGH09, Gue08, Mac18, Sch10, Ved10, War11a, Zet19]. quartic [Kul13c, Nic12b, YF12]. Quaternia [Cri15a]. Quercus [Lev10a, Lev11f, War10a]. quest [Cha05, Cha07, Hud11b]. question [Lor18h, Sik12]. Questions [Ano06, BEH17, Dol12a, Lev11f, Lev18c]. quests [Ron07, Wil10]. quick [Lor16a, Lor18j, You13, Lor15a]. quickest [Lor13i]. quilt [Hun12a]. Quinn [Pat17]. quintic [Col15, Kul13b]. Quite [Gib17b]. Quotient [Vil14b, Bea13]. Quotients [Bre15, HJ19].

R [App14b, Gib12b, Lev17f, Lor16c, Mac10b, Tol16d, Yeo18a]. R. [Hoa13b, Hol17, PW13]. R.P [Tol16c, dV14]. rabbits [Gri15g]. Rachel [Mac16b, Tol15e]. Radhey [Cou17b]. radian [DL14]. radical [Sco14a]. radicals [Col15, Osl10b, OJN16]. radius [Col11a]. Radmila [Rou16]. Rado [Cam18]. Raf [Hun10a]. Raffi [Yeo17d]. Ragni [Hun16]. Ragui [Hoa11]. Rain [Ban10c]. Rainer [Shi15c]. Rajeev [Mac14b]. Raju [Cou19]. Ramanujan [Hir13, MGC13b, Pla12, VSR13, Tol17h]. Ramon [Shi18a]. Random [Gri15g, ABD11, Bev17, Lor14h, Sch13, Tor17]. Ranjan [Shi15a]. rare [Wio19]. rates [FMK18]. ratio [ASHH18, All18, Bra11a, Dob11, Gho11, Gri13e, Gri15d, Jam18c, Lev14h, Lor17a, Nic14b, Red15, Mol17]. Rational [Gor18, AT12, Bar11, Bar13, Gri11a, Gri12b, Koo11, Pla14, SB19]. rationals [Koo11, Sub14a]. ratios [Nic16, FM10]. Raymond [Hoa18a, Lev13e, Lev13h, Lev14g, Mac15b]. Rayner [Aul19]. Rays [Bro11]. Read [Tol13, Hud13a, War10d]. readable [Lor16b]. Real [Sta11, CK19, Fin14a, For12, Lev17c, Lor16b, Mez14, MD10, Nic11, Wio19, Yeo17d, Lor14e, Shi19c]. real-world [For12, MD10]. Reale [Lev12c]. reality [Fre13, Ved10, War11a]. really [And12, Cri07, Dev08, Lev10a, War10a]. rearrangement [Nur13]. Rearranging [Nur14a]. reasoning [Lev09]. Rebaza [Hun13a]. Rebecca [Hal10a]. Received [Ano14c, Ano16a, Ano17a]. reciprocal [Lin11, Lor11j]. reciprocals [Gri15c, Gri15e, Sco18a]. reciprocity [Dol14a]. Reckoned [HHPM07, De 10a]. recognize [Pas11b]. Recountings [Seg09, Gib10b]. rectangle [EM16]. rectangles [EM12, EM16, GG12b, Pat13a]. rectangular [FF19]. recurrence [Dow15, Pla16]. recurrences [Gri17a]. Recursive [OC14, Cer16, Egg15, Kha10]. Recycling [LN17]. reductions [Tol15g]. Redundant [ADLM18]. referees [Lev14q, Lev18a]. Reflecting [GG10b, M^{+}08]. Reflection [Lor11n]. Reflections [Col13b, Ada10, Sik12, Bea12b]. reform [Cra07, GG10a]. Regarding [Sîn10]. Regular [She13, Fos18, Lor12i, Ste17d, Kin10]. Reimer [Tol15b]. related [FM10, Jam14f, OC14, SE13, Sim12]. relation [Dow15]. Relations [MM13a, CF10b, Jos17b, Jos18b, Pat17]. relationship [Eul15, ES16, Ste17a]. relative [FMK18, Nic14a]. relativity [Jac19, Lev14c, Orz12]. remarkable [Bay11b, JJ13b, Osl15, Ros09]. Remarks [Bea18a, Bra17, Pla11c]. Remembering [Aud11, CR16, Gib13b]. Removing [Bea15b]. Renate [De 14c]. Renato [GG12c]. repetition [Gri18a, Mas14]. representation [Cou14, EGH^{+}11, GS11, Hun14a, Mel13]. representations [GW09, Gri15b, Hun12b, Hun17b]. representing [GG12a]. reprise [Cri12a]. Request [Lev14q]. Research [Rou13a, Sto14c]. residues [Dol18b]. resistance [Lor19d]. resources [Hun17f]. restitution [Lor11n, Lor19d]. result [MvdB12, Wei12, WO13, dVS18]. results [Loo11, Mur17]. return [KL14]. returning [KL14]. returns [GS11]. Reuben [Hun13d]. revealed [Ber10, Lev12b]. revenge [San09, Shi10h]. Review [And10a, And10b, And11, And13c, And13a, And13b, And13d, And19, App10b, App10a, App11, App14b, App14a, App18, Ban10a, Ban10c, Ban10b, Ban13, Ban14b, Ban14a, Bay10c, Bay10b, Bay10a, Bay11b, Bay11a, Bay12, Bay14a, Bay14b, Bay14c, Bay14d, Bay15b, Bay15a, Bay17b, Bay17a, Bay17c, Bay17d, Bay19b, Bay19a, Bev13, Bev14, Bev17, Cam18, Cou10, Cou11, Cou12, Cou13, Cou14, Cou15, Cou17a, Cou17b, Cou18, Cou19, Cri13, Cri15b, Cri15a, Cri17a, Dav10, Dav15, Dav17, Dav18, De 10b, De 10a, De 11b, De 14c, De 14b, De 17, Eat11, Eat12, Eat13, For10a, For10b, For11, For12, For13, Fos14b, Fox15, Fox17, Fre10a, Fre10b, Fre15, Gib10a, Gib10c, Gib10b, Gib11b, Gib11c, Gib11a, Gib12b, Gib12a, Gib13b, Gib13a, Gib14b, Gib14c, Gib17a, Gib17c, Gib17b]. Review [Gib18a, Gib18b, Gib19b, Gib19c, Gol10, GG10a, GG10b, GG12c, Gri10b, Gri11d, Gri11e, Gri11b, Gri11c, Gri12f, Gri13a, Gri13b, Gri13d, Gri13c, Gri13f, Gri14f, Gri14a, Gri14e, Gri14b, Gri14c, Gri14d, Gri17c, Gri17d, Gri18d, Gri18e, Gri18c, Had19, Hai10, Hai12, Hai13, Hai15, Hai17a, Hai17b, Hal10c, Hal10a, Hal10b, Hal11b, Hal11c, Hal11a, Hal12b, Hal13, Hal14, Hal15a, Hal15b, Hal16, Hal17a, Hal17b, Hal17c, Hal17d, Hal19, Haw11, Haw12, Haw18b, Haw18a, Haw18c, Hem19, Hoa10, Hoa11, Hoa12, Hoa13a, Hoa14a, Hoa14b, Hoa15a, Hoa15b, Hoa17a, Hoa18a, Hoa19, Hop18a, Hop18b, Hop19, Hor17, Hud10c, Hud10b, Hud10a, Hud11a, Hud11b, Hud13a, Hud13b, Hun10b, Hun10a, Hun12b, Hun12a, Hun13d, Hun13c, Hun13b, Hun13a, Hun14b, Hun14a, Hun15, Hun16, Hun17a]. Review [Hun17b, Hun17c, Hun17d, Hun17e, Hun18b, Hun18a, Hun18c, Hun19b, Hun19a, Hun17f, Jac14, Jac19, Jon19, Kee18, Kil10, LP10, Lea10, Lee17, Lev10d, Lev10f, Lev10e, Lev10a, Lev10b, Lev10c, Lev11b, Lev11d, Lev11a, Lev11e, Lev11c, Lev11f, Lev12a, Lev12d, Lev12b, Lev12c, Lev13e, Lev13c, Lev13b, Lev13h, Lev13a, Lev13g, Lev13f, Lev13d, Lev14e, Lev14i, Lev14n, Lev14o, Lev14k, Lev14l, Lev14j, Lev14c, Lev14h, Lev14p, Lev14b, Lev14m, Lev14a, Lev14d, Lev14f, Lev14g, Lev15b, Lev15a, Lev17g, Lev17h, Lev17a, Lev17b, Lev17c, Lev17e, Lev17d, Lev17f, Lev17i, Lev18d, Lev18c, Lev18b, Lev19b, Lor10e, Lor10d, Lor10c, Lor11g, Lor11h, Lor11f, Lor11i, Lor13d, Lor14f, Lor14e, Lor14d, Lor15g, Lor15i, Lor15h, Lor16c, Lor16b, Lor17g, Lor17e, Lor17f, Lor18d, Lor18e]. Review [Mac10b, Mac10c, Mac10a, Mac11b, Mac12b, Mac13e, Mac13a, Mac13c, Mac13b, Mac13d, Mac14b, Mac14c, Mac14d, Mac14a, Mac15a, Mac15b, Mac15c, Mac19, Mac16a, Mac16b, Mac16c, Mac17, Mac18, Mah10b, McB18, Meg12, Mor19, Mun12, Nai12, Pas14, Pas18, Rot10c, Rot10d, Rot10b, Rot10a, Rou13a, Rou14, Rou16, Row19, Rua10, Rua11, Rua12, Rua13, Rua14c, Rua14a, Rua14b, Rua17, Sch10, Sch13, Sch14, Sch16b, Sch16c, Shi10e, Shi10f, Shi10c, Shi10i, Shi10g, Shi10d, Shi10h, Shi11e, Shi11d, Shi12c, Shi13c, Shi14d, Shi14b, Shi14c, Shi14e, Shi15c, Shi15a, Shi15b, Shi16b, Shi17a, Shi17b, Shi18b, Shi18a, Shi19c, Sto10b, Sto10a, Sto12c, Sto12b, Sto12a, Sto14a, Sto14b, Sto14c, Sto17b, Sto17a, Sut10, Sut11, Ten12, Tin17, Tol10, Tol13, Tol14b, Tol14a, Tol14c, Tol15b, Tol15e, Tol15c]. Review [Tol15g, Tol15a, Tol15d, Tol15f, Tol16a, Tol16d, Tol16c, Tol16b, Tol16e, Tol17a, Tol17b, Tol17c, Tol17d, Tol17e, Tol17g, Tol17f, Tol17h, Tol17i, Tol17k, Tol17j, Tol18a, Tol18b, Tol18e, Tol18c, Tol18d, Tol18f, Tol19a, Tol19d, Tol19g, Tol19b, Tol19c, Tol19f, Tol19e, Tsa18, Von17, War10a, War10c, War10b, War11a, War11b, War12a, War13a, War13b, War13c, Wil10, Yeo16, Yeo17a, Yeo17b, Yeo17c, Yeo17d, Yeo18a, Yeo18b, Yeo18c, Zar10b, Zar10a, Zar11, Zet13, Zet18, Zet19, dV14].

Reviewers [Lev15e]. Reviews [Lor18f, Pat17]. revised [Hud10a, Zet18]. Revision [Ano10a, SD10, She10a, Hal10a, Hal12b, Hal17b, PF08, Sto12b]. revisited [Bar17, FF16, KK18, Kom10, Kos12b, Lor11a, Raz19, Sco11a, Sco12, Sco13b, Sco17d, Sco19d, Sco19c, Sil18b, Sim11c, Sof13]. Revisiting [De13, Jam18b]. revolution [Ari12, Fos14a, Lor19b, Vil14a, Tol14b]. Reward [Wer09, Hai10]. rhombus [Haj18a]. Ribenboim [Mac17]. Rice [Lev13e]. rich [Ben12a, Lor15m, Lor15n]. Richard [Hai10, Hoa14b, How16, Lev17a, Tol17a, Tol19g]. Richardson [App10b]. Richeson [Rot10c]. Riemann [PG17]. Rigby [Lev11d, Lev15c]. right [Bro11, Gib17b, Job18a, Leu12a, Leu17]. right-angled [Leu12a]. rigour [Ges11, Gri13c]. Riley [Hor17]. Ring [Coo19]. Rings [FGR14, JST12, Hun15]. Riot [Ada09b, Lor11g]. Rise [Ale10, Gib11b]. Risk [Wer09, Hai10]. road [Qua13]. Roads [Sti10, Hoa12]. Robert [Bay19a, For10a, Hun10b, Lev13b, Lev15a, Lev17h, Mac13e, Mac13b, Pat17, Shi11e, Shi17b, Tol15c, Tol17c, Tol17i, Tol18e]. Roberto [GG12c]. Roberts [Gri17c, Tol17f]. Robin [Jon19, Lev13e, Lev13h, Lev14i, Lev14b, Lev14g, Lor18f, Meg12, Row19, Rua10, War12b]. Robyn [Tol14b]. Rocchi [Hai15]. Rocha [Ban10a]. Rodgers [Sut11]. Rodrigues [Ban10a]. Rodriguez [Shi10e, Shi10f]. Roe [Hun12b]. Rogelio [Rou16]. Roger [Lev17i, Lor10d, Lor18e, Shi18a]. Rohde [Mac14c, Yeo18c, Mac14c]. Roi [Cri17a, Yeo18b]. Roland [Hai13]. rolls [ML13]. romance [Abb08, ALB10a]. Rome [GG10b]. Ron [Hop19]. Ronald [Hoa17a]. Ronan [Wil10]. root [Kul13a]. rooted [vDD13]. roots [Abe16, Cam16, Fin14a, GG12a, Jac11, Kom10, KR17, Lo 13, Lor15n, Lor18c, Nic11, Shi12b, SB19]. Rosenberger [Shi15b]. Rosenhouse [Bay11b, Bev13, Kee18]. Rosenthal [Lor16b]. Ross [Lev13d, Tol18b]. rotation [BCL10]. Rothman [Lev11d]. rough [HI14, Lor18j, Sim11b]. Round [NS87, Hop16, Lev18b]. Rounding [vW19]. Rounds [NS87]. Routh [Abb15]. Routh-Steiner [Abb15]. Rowan [McG13]. Rowland [Gri18d, Sto14b]. Rudi [Tol14c]. Rudolph [Zet19]. Rugby [CP10, PR10]. Rule [Leu12b, Hud13b, Jam12, Kha13, Szp10b]. rules [Hop13]. Runge [De 14c, Tob12]. Runs [Lew14, Lor18g, MM15].

S [Cou14, Cou17b, Cou19, Gib14b, Gri11c, Hal17d, Hoa17a, Hol17, Hun15, Lev14e, Lev14h, Lev14p, Lev17h, Lor14e, Lor16c, Lor17f, Rot10c, Rua12, Rua13, SB^{+}12, Shi10c, Shi13c, Tol16b, Tol18e, War10b, Zar11, dV14, Bat14, Lor15i, KK10]. S.G [Hoa14b, Mac14b]. S1 [Hal10c, Nic08a]. Sacred [FR08, Lev11d]. Sage [JH12, De 14b]. Salisbury [Gar10]. Sally [Shi14e]. sample [FL14]. Sandra [Hal15a, Hal15b]. sangaku [Ara15]. Sangalli [Shi10h]. Sangwin [Hal15a, Hal15b, Pat17, dV14]. Sanjoy [Haw12]. Santiago [War10b]. Sara [Zar10b]. Sarah [For10b, For13, Lev10e]. Sarah-Marie [For10b, For13]. Sasane [Lev17d]. Sastry [Haj19]. save [Hol12b]. scalar [Gre19]. scale [Aud14]. scalene [Sco19b]. scary [Pas11b]. scenario [Gri12h, Gri15g]. Schilders [Gri14d]. Schmah [Cou12]. Schneider [Hun14b]. Schnell [Fos11, Mil12]. Schneps [Lev13f]. Schoenfeld [Fre10a]. school [Gol17, Ran18, Yeo17c, Dol15b]. schools [Lev17f]. Schubert [Cou14]. Schuppert [Gri14d]. Schwarz [Lor15j]. Schwendner [Hun14a]. Science [WILY11, Wer09, Adl14, Bre11, DSZ08, Dav12, De 14c, Hai10, LHS14, Mac13d, Sen12, Tob12, Gri13b, Lev10e, Mac13a]. sciences [Ano12d, Ano13l, CMZ93, FFW13, JS08, Mac11d, Rou13a, Sto14c, Lev14g, Lor14f, Mac10c]. Scientific [App18, Bay17d, Gib17c, Gib18a, Gri17d, Gri18e, Gri18c, Hal17a, Hal17c, Hal17d, Hoa18a, Hop18a, Hun17a, Lev12c, Lev17d, Lev17f, Lev17i, Lor17f, Lor18d, Mac17, Shi17a, Tol17a, Yeo17b, Yeo17c, Yeo18c]. scientists [ÅM08, Mac10b, Mac10a, Vel09]. scores [vdV12]. Scotland [BM19]. Scott [Lor15h, Rua14c, Tol19f]. SCT [Lev11d]. Seán [Shi16b]. search [Der15, Der19, Sut10, Tol17h]. Sebastian [Hun14a]. Second [Ten12, Shi18a, Sto17b, Vig19]. secondary [Lev17f, Yeo17c]. secrets [Gri14f, Lan12, Lev14p, Mac13e, PL12, Ste11b, Cou18]. section [Mah19b]. sections [Gho11]. sector [Lor10i]. Seduced [Ari12, Tol14b]. see [DL12]. Segel [Gib10b]. Segerman [Had19]. Selecta [Sch10, FFGH09]. selected [Cve12, Shi14d]. selections [Rua12, Tha05]. self [Ste17e]. self-similar [Ste17e]. semicircle [Che16b]. semicircles [Job11b]. Semiseparable [VvBM08, Hun10a]. Senechal [Fre15, Lev10e]. Senior [Lor11c]. sense [Yeo18b]. Separating [Lor19d]. sequel [Dol15b]. sequence [Gri15h, Gut11, McG13, Mil11, Osl12b, Pla12, Pur19, Rua14c]. sequences [Di 13, Gri11a, Gri18b, GS16, Kum10, Lor14l, MMYA10, Pla16, Lor11a]. Serenus [Haj19]. Series [Cou13, Jam14d, Ayd18, BD15, BD16, Coo11a, Gri12a, Gri12e, Gri12b, Jam14f, Jam18a, JLL15, Kha10, LN17, Lor10g, Mel13, Mur11a, Nil13, NL16, Nur13, Nur14b, Nur14a, Nur16, Osl12a, OD13, Pla18, Sco11d, Sco11a, Sco11c, Sco14c, Sco15a, Sco17b, Sco17c, Sco18a, Shi19a, Sub15b, Tol19h, WO13, Gri15b, Pla11a]. seriously [Bev13, RT11, Kee18]. Serre [Shi10i]. Seshadri [Cou14, SB^{+}12]. set [Cou14, Gao08, Gao09, Hal12a, Mac14e, Shi14c, Szy17, War10c]. Sets [MM13a, Pat17, Dol18a]. setter [Cut14]. seven [Sil18b, Tol17k]. Seventeen [Ste12c]. seventeenth [Dev08]. seventeenth-century [Dev08]. seventh [Tay07]. Sever [Shi13c]. several [Tay07]. Seymour [War10b]. Shakespeare [dS13]. Shanghai [Mer15]. Shape [HS13, DSZ08, Lev10e, Mac14d, Rop18, Shi14e, YN10, YN12]. shaped [MW18]. Shapes [Job13a]. shaping [GG10b, M^{+}08]. Sharp [Fre10b]. sharpening [Ret17]. Shearer [Mac16b]. Sheets [Ano18b]. Sheldon [Tol18b]. Shell [Bay17b]. Shell-Gellasch [Bay17b]. Shen [Bay17d]. Sheng [Hop18a]. Sheppard [Lee17]. Sherran [Sto12b, Sto12a]. Shifting [Kul13c]. Shimura [Gib10c, Shi12c]. Shing [Mac14d]. Shing-Tung [Mac14d]. Shiue [Lor16c]. SHM [Sim11b]. Short [Osl15, Bay14c, Cri13, Fos12c, Haj18a, Hig11, Hop18b, Kar18, Lev14o, Lor15g, Mac19, Maz14, Ram09, Row19, Ste12a, Mer10]. shortest [Ban18]. Shot [Lor12h]. Should [Mat16, Cri12b, Gra13]. show [Lor14k]. Showing [Ayd18]. Siava [Hun14a]. Siddhāntasundara [Knu14b]. Siddhantāsundara [Dav15, Knu14a]. side [Bay12, Hol11, Hol13b, Lor14d]. sides [Tol19a]. Sigma [Gib19c]. Sign [Zho13, HP10a]. signals [MD10, For12]. Significant [Bra11b, Dav18]. signs [Nur14a]. Silberstein [Eat13]. Silva [Shi19c]. Similar [Leu12b, Leu12a, Lor14j, Ste17e]. simple [Far13, JJ13b, Jam15d, Mel13, Mur17, Nur13, Shi11e, Sle12, TL16, Ton16, Wil09]. simplest [Haj15a, Jam10a]. simplices [Haj18b]. Simpson [Bea18b]. Simson [Smi15, Qua12b]. simulation [Mac10a, Vel09]. Simultaneous [Lor12i]. Sinclair [Sto17a]. Sine [Jam15e, Gib14a]. sines [Rou13b]. Singapore [Lev11d, Mer15]. Singh [Mac16a]. single [Bay17a, Gib11a, Jam16b]. singular [Lor15f]. Siobhan [Gri17c]. Sir [How19, Ste16b]. Six [Rea10b, Lev12c]. sixth [And10a, And10b]. size [Nic14a]. sized [Lor17h]. sizes [KK10]. skew [Job13a]. Skinner [Hal10a]. Skyrms [Tol18d]. sledge [Sim12]. slices [FJ10]. Slide [Hop13]. Sliding [HI14, De13, Gla12, KL11, Luz16]. Slot [Yeo18c]. slowly [Ben12a, Coo11a, Gri15h, Shi19a]. Small [Cou13]. Smith [Mac10c, Zar10a]. Smullyan [Hoa18a]. snakes [GS11]. snapshots [DLP08, Gib10a]. sociable [Kee13]. Society [Gib19b, Gri11e, Gri11b, Gri12f, Hoa13a, Hun14a, Hun17d, Lev11b, Lor11g, Lor11f, Lor14f, Sto12c]. Sofya [Gib13b, Aud11]. Soifer [Bay11a, Lor11h]. solid [Mac11a]. solution [Kul13b, Sta18, YF12]. solutions [Baj13, May16, Pla10c, Pla14, Sur14, ZF10, Shi13b, App18, Yeo18c]. solve [Che19, Fer16, Jam16b, Kul13c, Lev09]. solver [Job13c, Rou14]. Solving [Ano14q, Kac17, Bay11a, Bay12, Hal17d, Hal19, Hol11, Hol13b, Lev17e, Lor14d, Mah10a, Soi09b, SLMRR10, TL16, Tra13, dV14, Col15, Haw12, Lor17f, Yeo17c]. Soma [Sut11, GG08b]. Some [And17, Fox16, Haj15b, JJ13b, Jam14e, NL16, Pla11c, Ste17f, li17, Abb15, Baj13, Gri12a, IS12, Mur11a, Pat15, Pla11a, Pla14, Ram09, Ste17e, SA16, Gri12e, Jam14f, KG11, Kos12a, Siu13]. Somerville [Tol14b, Ari12]. sonnet [McG13]. Sons [App10b, Hai17a, Mac14c]. Sophie [Lev14k]. Sourcebook [Cri17a]. space [App10a, Pur18, Rop18, YN10, YN12, Mac14d]. spaced [Haj17b]. spaces [App11, Osb13, Shi10g, Sut09, VV09, Lor15h]. spade [Shi12b]. spatial [VM19]. speaking [BEU12, Tol14c]. special [Bra12, Lor15e, McN14, OS15, Jac19, Lev17f]. specified [Hal13]. specious [Mar11]. Speigelhalter [Tol17j]. sphere [Pop12]. spheres [Ara15, Fox14, Nun13, Pop12, Gib14b]. spherical [CR16, Lor18b, Rua14b, Van13]. Spieker [Sco19b]. Spindown [CC18]. Spinout [Coo19]. spirals [Die19]. spirographic [Ber17]. Splitting [Gor13]. spreadsheet [Mur11b]. Spreitzer [Tol18e]. Sprent [Ban10b]. Springer [App11, App14b, Ban14a, Bay11a, Bay17a, Bay17c, Cou15, Cou17a, Eat13, Fox17, Gib10c, Gib12a, Gib13b, GG10a, GG12c, Gri13a, Gri13d, Gri14d, Hai10, Hai12, Hai15, Hal19, Hoa11, Hun12b, Hun14b, Hun16, Hun17b, Hun17c, Lev11a, Lev12b, Lev13g, Lor10c, Lor11h, Lor13d, Lor15h, Lor16c, Lor16b, Meg12, Pas14, Sch10, Sch16c, Shi10d, Shi11e, Shi11d, Shi12c, Shi14d, Shi14c, Shi15c, Shi15a, Shi16b, Shi19c, Tol14c, Tol17h, Tol19a, War12a]. Springer-Verlag [App11, Gib13b, Pas14]. Spurt [Gro11]. Spurtacular [SG12]. Square [Pis16, Ber15, Ber18, Cam16, Cri12a, Edg16, Eul15, ES16, GG12a, GV17, Hir11f, Jac11, Jam10b, Job13a, KR17, Lev18b, Lor18c, McL19b, ML13, Sch16a, Sco15a, Shi16d, Ste19a, Sub14b, SB19, Sut10, Szy17]. square-free [Jam10b]. squared [Haj17b]. squares [AM11, BT09, CDF18, Dol12b, Dol18b, Dol19a, EM12, Fos12e, Gib17a, GV15, Hir11f, Jam10a, Jam19a, Kee11, Lor10i, Mah13b, Nic14a, Sco18f, Shi13d, Fos12e, Lor12c, War10b]. Squaring [Szy19]. Srirangachari [SB^{+}12]. Srivastava [Hun15]. Stability [Ter09, Rob12, Mac11b]. stabilization [Mac11b, Ter09]. Stable [Bea15a, Gri11g]. Stacey [Ten12]. stacking [Chu18]. Stahl [Lor14e]. Stanley [Gri17d, Lev17a]. starships [Adl14, Mac15c]. state [Rou13c]. static [De14]. statistical [BHJ10, FFGH09, Hal11a, Sch10, Tol17k]. Statistics [Kra09, Nic08b, Pat17, Tol14e, Gri13a, Hor14, Leh11, Mar16, MMC09, Nic08a, Tol14d, Tol15e, Whe13, Hai17a, Lev14f, Lev19b, Lor11i, Tol18b, Tol19e, Zar11, Gri11d, Hai17b, Hal10c, Hal11c]. stats [Min12]. Staudt [Rza10]. steadiness [Kan17]. Stedall [Lev14o]. Stein [Tol17g, Zet13]. Steiner [Hun13d, Abb15, ASH19a, Bea19, CR14, Dac16, Haj17c, Hva11, Sco17d]. Steinhaus [Ste16a]. step [Cag18, Mac16a, Lor18h, Sik12, Mac16a]. Stephan [Shi18a]. Stephen [Gib12b, Hal17d, Hun13c, Mac13c, Tol17k, Tol19c]. Stephenson [Ano14n]. steps [Tol16c]. Sterrett [Hoa14b]. Steve [Hop19, Mac14d]. Steven [Bay14a, Hud10c]. Stewart [And19, Bay14d, Dav18, Gri14f, Gri14a, Haw11, Hop18b, Lev10d, Lev11c, Sto17b, Tol10, Tol15f, Ste08, Ste09a, Ste09b, Ste14]. Stigler [Tol17k]. Stillwell [Hoa12, Lor17e]. stimulants [Bat14]. Stipp [Lor18f]. Stirling [Gri12d, Jam15b, Jam15d, LN17, Lor15c]. Stochastic [KS14, For12, MD10, Sch16c, App14a]. stochastics [Çin11, Hai12]. stock [GS11]. Stoica [Cou12]. Stop [HP10a, PdVF14]. Stop-sign [HP10a]. stor [Bro12a]. stor- [Bro12a]. stories [Ada09b, Gri14d, Hun12a, LPE^{+}12, Lor11g]. story [And13c, Bag11, Bay11b, Gib17b, Gib18b, Gri13f, Gri18e, Hav12, McB18, Ros09]. Stothers [Gib11c]. Stråhle [RR11]. straight [Ban18, GG12a, Lor13i]. Strange [GG08c, Lev10e]. strategic [Smy10]. Strategies [Joh12, And17, Lor17f]. Strategy [Hal17d]. Strauss [Eat12]. streams [Fos18]. Street [Mah10a, Haw12]. Street-fighting [Mah10a, Haw12]. strictly [MM15]. string [YN10, YN12]. strings [Ada10, Lew14, War13a]. stripping [Whe13]. structure [Hud10a, Hun15, JST12]. Stubhaug [Gib12a]. Student [BL12, Cro10, Dol10a, Dol10b, Dol11c, Dol11d, Dol11e, Dol12e, Dol12f, Dol12g, Dol13b, Dol13c, Dol13d, Dol14b, Dol14c, Dol14d, Dol15c, Dol15d, Dol15e, Dol16c, Dol17c, Dol17d, Dol17e, Dol18c, Dol18d, Dol18e, Dol19e, Dol19c, Dol19d, Hal14]. students [Bau13, Iva09, Lor17i, Yeo17c, Lor11f]. studies [Mac14c, RJPG12, R^{+}12]. Study [Alc12, Alc13, FMK18, Fos14b, Hal17a, Hal17c, Sto14a]. Studying [Cel17]. Stylianides [Hun17f]. subdivision [Pop12]. subsets [Mez14, Pla14]. substitution [Hos17a, Ste17e]. subtle [Wel12]. subtraction [Gau19]. subtriangles [Lor13b]. Success [Hal12b, Gri14d, LPE^{+}12, Hal12b]. Sudoku [Bev13, BT09, Kee18, FF19, RT11, War10b, Job10, MM12]. sudokus [SLMRR10]. Sue [Hal14]. suggestion [Kan17]. suitable [BCSS19]. Suk [Von17]. Suk-Young [Von17]. Sulbasutras [Ste12b]. Sultan [Yeo17c]. Sum [NS87, Ber18, Gri12b, Gri15e, Haj17b, Hir11f, Job18b, Lor16g, Nur13, Sco14c, Sco18f, Shi19d, Ste19a]. Summing [Sub14b, Gri12e, Shi17b, Tol17i]. Sums [AM11, Bur10b, KR17, MGC19, Bra17, Cag18, Cer16, Cri12a, Der15, Der19, GB12, Gri15c, Jam15b, Jam17, Jam19a, JLU10, Loo11, Lor15a, Osl12a, Shi13d, Tre17, Tsa11]. Sundar [App14a]. Superbrain [Lev11e]. support [BL12, Hal14]. surd [Vil15]. surely [Ben12a]. surface [Jew10, Sim11b, Vil14a]. surfaces [BL10, EF10, Hun13c]. surprising [ES16, Hir11d, Lor13h, dVS18]. survey [JM09a, JM09b]. surveying [Lam19]. Susskind [Jac19]. Sutherland [Shi10g]. Swetz [And13a, Sut10]. Sylvester [De 12]. symbolic [Lev12f]. symbols [Lor15g, Maz14, Pas11b]. symmetric [App11, Bea12c, Dol19b, Far18, Mac14e, VV09, Gri11a]. Symmetrical [Fos12a]. symmetries [CBGS08, Eat12]. Symmetry [GW09, HS13, Ron07, Yan14, Cou12, HTSE09, KK17, Leu18, Lev17b, Shi14e, Hun12b, Wil10]. system [Pla10c, Shi14b, Shi18c]. Systematic [Mac14c, RJPG12, R^{+}12]. Systems [Hun10a, ÅM08, AF12, Gue08, Mac10b, Sma12, Cou10, Cou13]. Szpiro [Gri11b, Hud13b].

T [Gri11c, Gri14d, Tol15a]. T.O [dV14]. Taalman [Bev13, Kee18]. table [Bur10c, Cri10, Nie10, Pis16, Pur11]. tables [Bea15a, Osl10a]. tackles [Lor11d]. tactics [Lev17e]. Taking [RT11, Fos12c, Bev13, Kee18]. Talagrand [Sch16c]. tale [Bau13, Dev08, Gro18]. talented [Lev17f]. tales [AG14, Tol15c]. Tall [Hal16, Sto17b]. Taming [KP14, Hoa15b]. Tammadge [How16]. Tan [Shi17a, Gla19]. tangent [Sco11d]. tangential [GOD18a, Jos17a, Jos17c]. Tangents [Gla13]. Tantrix [BEH17]. Tanya [Cou12]. Tapp [Hun17d]. Tarry [Kee11]. taster [Pas16]. tasty [CDF18]. Taubes [Gib13a]. Tavares [War10b]. Taylor [Rot10a, Far13]. tbe [For12, Gri13a]. teach [Lev14c, Orz12]. teacher [Ada10, Bau13]. Teachers [Neu16, Iva09, Lor11f]. Teaching [Jam13a, Ran14, SSJ19, dS13, Ada10, ADLM18, AR14, And17, Fie07, Sto14b, Zar10b, Tol17j, Tol19e, dV14]. technique [LN17]. techniques [Cve12, JS08, Mac10c, Shi14d, Yeo17b]. Ted [Tol18c]. telescoping [Tol19h]. template [Sch16a]. temple [FR08, Sil18c]. ten [BY08, Tol18d]. Tensor [Mac16c]. Tent [Gri10b]. Tenth [Hir11h]. tents [FMS10]. Teresa [Tol13]. term [VSR13]. terms [JL17, Jam18a, MS13, Mil11, Nur14b, Pur19, Sco18d]. Terrell [Mac11b]. Terrence [Zet19]. Terry [Zet19]. tesseract [Sud13]. test [Cha11, Gra13, Jam18c]. Tests [Has19]. Tethering [Mah18c]. tetrahedra [Haj18b]. tetrahedral [Edg16]. tetrahedron [Bea15b, Kuc12, Lin11, Nic12b]. Tewkesbury [Gri14d]. Textbooks [Mer15]. Texts [App14a, Cou10, Cou12, Gib13a]. Thayer [Rua12]. Thébault’s [Vig19]. Their [Lev12e, Ban14b, Gib17a, Hun10b, McL19b, Mit12, Tub09]. them [Bur10b, Hop16]. theme [ASH19a, Gor19a, Lor10i, Lor12j]. Theorem [Rib18, Shi11b, Spo18, Abb15, Bar17, Col19, CR14, Dac16, De 13, De 14a, Gri12g, Haj10, Haj15a, Haj17c, Haj19, HJ18a, Jos16a, Jos17d, Jos19b, KK11, Kar18, Kha13, Lam19, Luz16, Mac11a, MP11, Mer13, Mer17, Nur16, OS15, PG17, Rza10, Sco18d, Sil18b, Sil18c, Sub18a, Wil12, Lev13b, Zet18]. theorems [Bea19, Cam18, Cve12, GG13, HP10a, Shi14d]. theoretic [Kos14]. theorist [Hun18c, Von17]. Theory [HS13, Vol11, ABD11, And10a, And10b, AG14, Bea11, Bev14, Cou14, EGH^{+}11, FGR14, For12, Gao08, Gao09, Hal12a, HWHBS08, Har10b, Jac14, Kra15, Mol12, MD10, Pla10b, Rua14a, Ser08, Shi10i, Shi14c, Shi14e, Sur10, Vil07, YN10, YN12, Hem19, Cou14, Hun14a, Jac19, Meg12, Sch13, Sch16b, Shi10e, Shi10f, Shi14e, Shi17b, Tol17i, Tol19b, War10c]. ThePackage [Mah19a]. there [Hop12, Mac13f]. thesis [AF12, Shi14b]. things [CBGS08, Gib19c, Eat12]. think [Hal16, Hou09, Lev10b]. thinker [Tol19f]. Thinking [MBS10, Hal17a, Gib18a, Ten12]. third [Orm13]. Thirty [Fre10b]. Thomae [Gri11f]. Thomae-like [Gri11f]. Thomas [Hun13c, Lev12c, War11b]. Thoo [Bay17b]. Those [dK09, Pas11b, Gri11e]. thought [Ram09]. Thoughts [Dol17f, Gri12c]. Three [Lor18i, LHS14, Aud14, Bea19, Fre10b, GV15, HJ18b, Kha13, MC09, Shi11a, Ste17c, Sut10, Des12, Lor19a, Hoa14b]. three-way [Aud14]. Thwaites [Dol15b]. Tiankai [Hun14a]. tide [Mah14]. tier [PF08]. Tiina [Gib12a]. tiles [Lor16d]. tiling [Gri16b, Ste16a]. Tilings [Sch16a, Gri17a]. Tim [Sto14b, Tin17]. time [Hol12b, Rop18]. times [KK11, Lev13b]. Timothy [Hun18b]. TIMSS [And12]. Tina [Hal10a]. Titu [Lea10]. Tobies [De 14c]. today [Gri14b, Mac12a]. together [Mah15b]. Toke [Dav15]. told [Ste12c]. Toller [Lor11i]. Tom [Hoa14a, Sut11, Tol14a]. Tony [Lev10a, Lev11d, Lev11f, Tol19b, War10a]. too [Hol12b]. Toolbox [HF04, FH05, Mah10b]. tools [App12, Bay14b]. Topics [Ser08, Hai17a, Rou16, Shi10i]. topological [Shi10g, Sut09]. Topologist [Pas14, vDD13]. topology [App18, Ric08b]. Toss [Joh12]. tossing [Gri12h]. touching [Sil18a]. tournaments [Gri10a]. tower [Bay10a, GG08a]. town [Ano14a]. tradition [BEU12, Tol14c]. Traditional [FR02, Lev11d]. Train [Grä11, War13b, War13c]. trajectories [Gro18]. trans [Gib10a]. Transcending [BEU12, Tol14c]. transfers [Rea10a]. transformation [Gri15a]. transformations [Fer16, Lor15f, Wil12]. transition [JM09a, JM09b, Lev12a]. Translated [Gib12a, Bay17d, Shi17a]. translates [Lor14g]. translation [Del19, Knu14b, Knu14a, Dav15]. trapezia [Jos18b]. trapezium [Jam12, Leu12a]. treasures [Ste09b, Lev11c]. treasury [Gri17d]. Treks [Fox17]. trial [Lev13f, SC13]. trialogue [Mac17]. triangle [Ano14r, BS15, Bro12a, Bro11, CF15, Cri17b, Dol15f, DZH14, Fos12c, Fox15, Fox16, Gor13, Gri08, Gri12d, GOD18a, GOD18b, Haj12, Haj17d, Hir11g, Job13b, Job18b, Jos19d, Kri18, Lev13j, Lor13a, Luk17, Luk18, Luk19, Luz16, Nic14a, O’L10, Ret18, Sco10b, Sco13a, Sco14b, Sco19b, Sha15, Tod15, li17, Gau12a, Bay10b]. triangle-folding [Hir11g]. Triangles [CF14, Dol15f, Nic14b, Shi11c, Bau13, Dol15a, Dol16b, Dol19a, Dow15, Gri15d, Haj18b, Hir11a, HE19, Job18a, Leu12b, Lor12c, MMYA10, Nic17, Oai15, PL12, Rea14, Rea15, Rea16, Sci19, Shi12d, Lor15b, Lev14p]. triangular [Bre15, Cag19, Edg16, Eul15, Lor16g, Shi12a, Sub19]. Tribulations [Ran14]. trick [Shi12a]. tricks [Cra09, Hal10b]. tricorn [Mac11a]. Trigonometric [Ste17d, GS16, JJ13a, Luc10, Luc13, Gau19, Bay17d]. trigonometry [CR16, Gri11h, Van09, Van13, You11a, Tal13, Rua11, Rua14b, Lev17i]. trinomial [Gri16b]. triple [Ban16]. triples [BD12, HJ19, Hva11, Kos16, MvdB12]. triplicate [Bra11a]. Tripos [Vil15]. trisecting [Osl16]. trisection [Kos16]. Triumphs [Ran14]. trivariate [Gri17b]. truncated [Kuc12]. truncation [vW19]. Trust [Bay10b, Fox15, Gri18d, Hun13b, Lev17e, Mac12b, Mor19, Sto10a, Tsa18, Zar10a]. truth [Hoa12, Nie10, Sti10]. Tschirnhaus [Kul16]. Tubbs [Hun10b, Lev15a]. Tuganbaev [Hun15]. Tullio [Eat13]. Tunc [Bay17a]. Tung [Mac14d]. Turing [Shi14b, AF12]. turning [Cae15]. Tutte [Ano17b]. twentieth [B^{+}11, GG12c, Hoa15b, KP14, Tub14]. twentieth-century [Tub14]. twenty [Lev11e]. twenty-five [Lev11e]. twin [MM13b, Oku13b, Cha11]. twinned [Bra12]. Two [Gro18, Lor13h, Lor18j, PS14, PF10, Bau13, Ber18, Bra12, BT17, Dol12b, FS10, GV17, Jam19a, Koo11, Leu12a, Lor16g, Mah18a, Mah13b, Nic14b, Shi13d, Ste17c, Ste18a, Sub15b, Jam10a, Job11b, Loo11, MM13c, Pat13a]. type [ES10, Jam15c, Lor13a].

U [Mac14c]. U$ [Cou12]. ubiquitous [Die19]. UK [Mor19, Tol10, Tsa18, SA16]. UKMT [Rou14, Lor11c]. Ulrich [Gib11a, Yeo18c]. umbrellas [FMS10]. Unbounded [Vol11, Jac14]. uncertain [Hud10b, Wai09]. uncertainty [Wai09]. unchanged [Wil14]. undergraduate [Hou09, Yeo18a, dV14, Yeo18c]. undergraduates [Hun17d]. understand [Wai09]. Underwood [Bay10c]. Unexpected [Wap12, ASHH18, Lor17a, Sco14d, Sco19c, SLMRR10, Gib14c]. unfinished [Dev08, Lev10f]. Ungar [Tol14c]. unified [Gau19]. unique [Ste12c]. Uniqueness [Mas14]. unit [Gri19]. unit-distance [Gri19]. United [Bay10b, Fox15, Gri18d, Hun13b, Mac12b, Sto10a, Zar10a]. units [Yan10]. universe [FW11, Ste11a, Ved10, War11a, YN10, YN12, Zet13]. University [And10a, And10b, And11, And13c, And13a, And13b, And13d, App14a, Ban13, Ban14b, Bay10a, Bay11b, Bay14b, Bay14c, Bay15b, Bay15a, Bay19b, Bay19a, Bev13, Bev17, Cam18, Cou10, Cou11, Cou12, Cou17b, Cou18, Cou19, Cri13, Cri15b, Cri17a, Dav10, Dav15, Dav17, De 11b, De 14b, De 17, Dol15b, For11, Fos14b, Fre10a, Fre15, Gib11b, Gib13a, Gib17b, Gib18b, Gol10, Gri11d, Gri11c, Gri13f, Gri14b, Gri14c, Had19, Hai13, Hal10c, Hal16, Haw11, Haw18b, Haw18a, Hoa10, Hoa14b, Hoa15a, Hoa15b, Hoa17a, Hop18b, Hop19, Hud10b, Hud13a, Hud13b, Hun10b, Hun10a, Hun12a, Hun13d, Hun15, Hun17e, Hun18b, Hun18c, Hun19b, Hun17f, Kee18, Lee17, Lev10c, Lev11d, Lev12a, Lev12d, Lev13e, Lev13a, Lev13d, Lev14i, Lev14n, Lev14o, Lev14l, Lev14b, Lev14m, Lev14a].

University [Lev14g, Lev15a, Lev17a, Lev17b, Lev18d, Lor15g, Lor17e, Lor18f, Mac10c, Mac11b, Mac14b, Mac15a, Mac15b, Mac15c, Mac19, Mac16a, Mac16b, Mac16c, McB18, MNS10, Pas18, Rot10c, Rot10d, Rot10b, Row19, Rua11, Rua14b, Rua17, Sch13, Shi10e, Shi10f, Shi10g, Shi10h, Shi14b, Shi15b, Shi17b, Shi18b, Shi18a, Sto14a, Sto17b, Sto17a, Sut11, Tin17, Tol14b, Tol15b, Tol15c, Tol15a, Tol15d, Tol16a, Tol16d, Tol16c, Tol16b, Tol16e, Tol17d, Tol17g, Tol17f, Tol17i, Tol17k, Tol17j, Tol18a, Tol18d, Tol18f, Tol19d, Tol19g, Tol19c, Tol19e, Von17, War10b, War11a, War11b, Wil10, Yeo17d, Yeo18b, Zet18, Gib11c]. unknown [Hoa15b, KP14]. unlikely [Bar10]. unlocking [Gri14f, Ste11b]. unravelers [DLP08, Gib10a]. Unravelling [Bea18b, FW11]. Unsolved [Tol18f]. unsung [Hoa10, Pas08]. until [Gri15f]. unusual [Lor18k]. updated [Von17, Zet18]. Upper [Sch16c]. US$27.95 [Shi17b]. USA [Kro18]. USD [De 14b]. use [Dob11, Hun17f, SLMRR10]. used [Lev13f, SC13]. user [Gib11a]. Using [Fer16, Jam16b, Lor17h, Ran18, Ano13m, Ayd18, De 14b, JH12, Kac17, Lev09, Lor12b, Lor14i, Luc10, Luc13, Mur12, Mur19, Shi12b, Ste17e, SSJ19, SB19, Szy17, Szy19, Tol15e, Tre17, Wil12, Far13].

V [App11, Cou14, Gib10a, Gri14d, Sto12c, BD15, Mah19a, MW18]. v-shaped [MW18]. Vaintrob [Hun14a]. Valdez [Rou16]. Valery [App11]. value [Mer13, Mer17, PS14, Shi16d, Tor17]. values [Kac18]. Vandebril [Hun10a]. Vandenberghe [Tol19c]. variable [Tor17, Bay17a]. variables [Lor14h, Shi11a]. variance [FL14, LY11, Lor14h, Wil14]. Variants [Sur18]. variation [De 14a]. Variations [Ara15, Lor10i, Lor12j, Sil18c, ASH19a, Mac11c, Gor19a, MM13b, Shi17a]. VCH [Mac10a]. Vector [Cou14, Ban16]. Vectors [Sud11]. Vedral [War11a]. Velten [Mac10a]. Vera [Hun13d]. Verlag [App11, Bay17a, Bay17c, Cou15, Cou17a, De 10b, De 10a, Gib13b, Hai15, Hal19, Hun16, Hun17b, Hun17c, Lor15h, Pas14, Sch16c, Shi16b, Shi19c, Tol14c, Tol17h, Tol19a]. versatile [Lor17c]. Version [Gib11a, Haj17c, HJ18a]. versus [vW19]. vertex [Bea15b]. vertical [Lor14g]. vertices [Sco18e]. very [Bay14c, Cri13, Haj18a, Hig11, Hop18b, Kar18, Lev14o, Mac19, Row19, Shi19d, Ste12a]. vexing [Hud13b, Szp10b]. via [Den10, Gri15h, GS16, Jam13c, Lor15j, Spo18, Yeo18a]. Vicky [Tol18a]. Victor [Cri17a, Hai17b, Hoa15b, Yeo18c]. Victorian [Lev13e, FRW11]. Vieta [Osl10b, OW14]. Vieta-like [Osl10b]. view [Har14, Sco13a, Sco18b]. views [Hoa14b, LHS14]. Vikraman [Cou14]. Villegas [Shi10e, Shi10f]. vintage [CMS18]. Visions [Ste13b]. Visual [Pla16, Gau19, Lor18a, Mil12]. Visualising [Spi10]. visualize [Pas11b]. Visualizing [Had19]. Vitali [Jac14]. Vitaly [App11]. Vite [B^{+}11]. Viviani [De 13]. Vlatko [War11a]. Vogeli [Lev17f]. Vol [Hun10a]. Volchkov [App11]. Volpert [Jac14]. Volume [Lor14f, Sut11, Vol11, Ano10g, Ano10h, Ano10i, Ano10k, Ano10j, Ano10m, Ano10l, Ano11f, Ano11h, Ano11g, Ano11j, Ano11i, Ano11l, Ano11k, Ano12f, Ano12h, Ano12g, Ano12j, Ano12i, Ano12l, Ano12k, Ano13e, Ano13g, Ano13f, Ano13i, Ano13h, Ano13k, Ano13j, Ano14h, Ano14g, Ano14j, Ano14i, Ano14l, Ano14k, Ano15f, Ano15e, Ano15h, Ano15g, Ano15j, Ano15i, Ano16d, Ano16c, Ano17h, Ano17g, Ano17j, Ano17i, Ano17l, Ano17k, Ano18h, Ano18g, Ano18j, Ano18i, Ano18l, Ano18k, Ano19g, Ano19f, Ano19i, Ano19h, Ano19k, Ano19j, Cou14, Jac14, Lin11, Lor17i]. volumes [Bay15a, FM10, Fos14a, Lor19b]. vote [Cri12b]. vs [Ada10, War13a].

W [Gib18a, Gri10b, Gri14e, Gri14d, Hoa14b, Mac10c, Shi14b, Sil10, Tol15a, Zar11]. W.W.Norton [Lev14f]. Wagner [Cri17a]. Wainer [Hud10b]. walk [Ada09a, Gri18c, For11]. Walks [Kos13, Gri15g]. Wallach [Hun12b]. Wallis [Kha12, Lev12f, Lor12a, Osl10a, Osl11b]. wallpaper [Lev17b]. walls [Sim11a]. Wang [Mac13b]. Wang-Iverson [Mac13b]. Wanger [Yeo18b]. wanted [CT14]. Wardhaugh [Hud13a, Lev13a, Lev14i]. Warner [Gib14c]. Wassell [Gib12b]. water [SV15]. Watkins [Lev14b]. Watling [Tol16b]. wave [Sco15a]. way [Aud14, Orm13]. ways [Fre10b]. weak [Bur17]. weakness [Lev19b]. wealth [Lev13a, War12c]. Weierstrass [PG17]. weighted [Abe19]. Weird [Haw18c]. Wells [Ban10c, Mac15a]. Werner [App10a]. Werthamer [Hai10]. Wheelan [Lev14f]. wheels [BD10, BD10]. Where [Job15, Sub18b]. Which [HK14, Lor15b]. Whitaker [Mac15b]. Whitney [Gib19b]. who [Mac17]. Whoever [Ram09]. whole [Shi19d]. whose [Lor14g]. Wikler [Lev10e]. Wiles [B^{+}11]. Wiley [App10b, Gri14e, Hai17a, Hal11a, Hun13a, Lor10d, Lor14e, Mac10a, Mac14c, Shi10c]. Wiley-VCH [Mac10a]. Wilkes [Tol17b]. Will [Hop16, Tol19e, Hop12]. William [Hal11c, Hud10a, Hun19b, Lev12a, Lev14j, Mac11b, McG13, War11b, War13a, Gar10]. Williams [Gib12b, Sto10b]. Willson [Lev14j, Gar10]. Wilson [Gib11c, Lev13e, Lev13h, Lev14b, Lev14g, Lor18f, Row19, Rua10, Shi10g, Shi11e]. Wind [Mah14]. winners [Hop12]. wire [Lor14a]. wisdom [Tol17k]. within [BD10, Fos12e, Rob10]. Without [GV17, SS16, Ban10b, Ber18, Cag18, Cag19, Cel17, Cha15, DeM10, Edg16, GV15, Hir11b, Lor11m, Lor18e, MM13c, Mah13b, PF10, Pla18, SV15, Spr09, Ste17a, SSJ19, Sub15a, Sub15b, Sub18b, Tol16a, You11a]. Wizards [Adl14, Mac15c]. wonderful [Zet18]. Wonders [Gib19c]. Woodhead [Gib11c]. Woolwich [Col13b]. word [And13a, Swe12]. Words [Ber18, Cag18, GV17, Pla18, SS16, Sub18b, Cag19, Cha15, DeM10, Edg16, GV15, Lo 13, Lor11m, Lor18e, MM13c, Mah13b, PF10, SV15, Ste17a, SSJ19, Sub15a, Sub15b, Tol15d]. work [Dav18, FMW14, Hop19, Kin15]. workbook [Hal12b, Hal13, Pea11, She10b, Sto12a]. works [Gib12b, WMW10]. World [App18, Bay17d, Gib17c, Gib18a, Gri17d, Gri18e, Gri18c, Hal17a, Hal17c, Hal17d, Hoa18a, Hop18a, Hun17a, Lev17d, Lev17f, Lev17i, Lor17f, Lor18d, Mac17, Shi17a, Tol17a, Yeo17b, Yeo17c, Yeo18c, And13d, Bev13, BT09, Dev08, For12, Fox17, GG10b, Hah12, M^{+}08, MD10, Pas11b, RT11, Ste12c, Tol16e, Wai09, War10b, Wio19, Gri14a, Hoa19, Hud10b, Lev17c]. Worth [Hoa14a, Grä11, War13b, War13c]. Wright [And10a, And10b, Lev11b]. Wrinch [Sen12]. writes [Ano14p, Ano14q, Ano14r, Ano14o, Ano14n]. writing [DSZ08, Lev10e, Lev12d, Lev14l, Lev14m, Lev15b, Lev17g, Lev18d, Pit11, Pit12, PM13, Pit14, War12c]. writings [Hoa13a, Neu11, Tha05, Rua12]. www.cinderella.de [Gib11a]. www.editions [Lor15i]. www.editions-ellipses.fr [Lor15i]. Wynne [Gar10, Lev14j].

X [Lev11d]. Xiong [Yeo17b]. Xlibis [Hud10a]. Xlibris [Hal11c, War13a]. XYZ [Lea10, Yeo16].

Yackel [For10b, For13]. Yadav [Rua13]. Yang [Bay17d]. Yau [Mac14d]. year [Grä11, Sta19a, Sut10, War13b, War13c, Ton19]. years [BM19, HP10b, MC09, War12c, Hoa11, Lev11e]. yields [Rob10]. Yim [Mac13b]. Yong [Yeo17b]. Young [Von17]. Yudinova [Hun14a].

Zawaira [De 11b]. Zdravko [Shi14d]. Zeeman [Ste16b]. zero [Mar11, Sco14c]. zeros [Kul16]. Zhang [Shi17a]. zigzag [Wei12]. Zvi [Lev17c]. Zwicky [Lev10e].

References

Albers:2008:MPP

[AA08] Donald J. Albers and Gerald L. Alexanderson, editors. Mathematical people: profiles and interviews. A. K. Peters, Ltd., Wellesley, MA, USA, second edition, 2008. ISBN 1-56881-340-6. xxvi + 386 pp. LCCN QA28 .M37 2008.

Adhikari:2013:BMA

[AA13] Mahima Ranjan Adhikari and Avishek Adhikari. Basic modern algebra with applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 81-322-1598-2 (hard cover). xix + 637 pp. LCCN ????

Abbott:2008:AFR

[Abb08] Edwin Abbott Abbott. The annotated flatland: a romance of many dimensions. Basic Books, New York, NY, USA, 2008. ISBN 0-7382-0541-9, 0-465-01123-3. xxvii + 3 + 239 pp. LCCN QA699

           .A115 2008. URL http://catdir.loc.gov/catdir/toc/fy036/

2001098249.html. Preface by Ian Stewart.

Abboud:2015:RST

[Abb15] Elias Abboud. On the Routh-Steiner theorem and some generalisations. The Mathematical Gazette, 99(544):45–53, March 2015. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Akemann:2011:OHR

[ABD11] Gernot Akemann, Jinho Baik, and Philippe Di Francesco, editors. The Oxford handbook of random matrix theory. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-957400-6. xxxi + 919 pp. LCCN QA188 .O94 2011. URL http:

//www.loc.gov/catdir/enhancements/fy1402/2011029624-b.
html;   http://www.loc.gov/catdir/enhancements/fy1402/
2011029624-d.html;   http://www.loc.gov/catdir/enhancements/
fy1402/2011029624-t.html.

Abel:2016:HOA

[Abe16] Ulrich Abel. High order algorithms for calculating roots. The Mathematical Gazette, 100(549):420–428, November 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
2FACD442DB78364B04DA2E64BA06F269.

Abel:2019:CBG

[Abe19] Ulrich Abel. Coincidence of the barycentre and the geometric centre of weighted points. The Mathematical Gazette, 103 (558):409–415, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/coincidence-of-the-barycentre-and-the-geometric-
centre-of-weighted-points/E811F40221EE8C3B19B6B1E01C3C1A4C.

Acheson:2011:WPM

[Ach11] David Acheson. What’s the problem with maths? The Mathematical Gazette, 95(534):386–394, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Andreescu:2008:PB

[AD08] Titu Andreescu and Gabriel Dospinescu. Problems from the book. XYZ Press, Allen, TX, 2008. ISBN 0-9799269-0-4. xvi + 554 pp. LCCN QA43 .A534 2008.

Adam:2009:MNW

[Ada09a] John A. Adam. A mathematical nature walk. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-12895-2 (hardcover). xx + 248 + 8 pp. LCCN QA99 .A23 2009.

Adams:2009:RCE

[Ada09b] Colin Conrad Adams. Riot at the calc exam and other mathematically bent stories. American Mathematical Society, Providence, RI, USA, 2009. ISBN 0-8218-4817-8. xvi + 271 pp.

LCCN QA99 .A235 2009. URL http://www.loc.gov/catdir/
toc/fy1001/2009009835.html.

Adams:2010:MPP

[Ada10] William J. Adams. Math is precise, period, vs. math is precise, strings attached: reflections of a math teacher on teaching mathematics. Xlibris Corporation, [S.l.], 2010. ISBN 1-4363-3419-5 (hardcover), 1-4363-3418-7 (paperback). 492 pp. LCCN QA11.2 .A32 2010. With illustrations by Ramunçe B. Adams.

Adler:2014:WAS

[Adl14] Charles L. Adler. Wizards, aliens, and starships: physics and math in fantasy and science fiction. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-14715-9 (hardcover). xi + 378 pp. LCCN PN3433.8 .A35 2014.

Amram:2018:RAT

[ADLM18] Meirav Amram, Miriam Dagan, Sagi Levi, and Artour Mouftakhov. Redundant axioms in teaching linear algebra. The Mathematical Gazette, 102(555):401–412, November 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/redundant-axioms-in-teaching-
linear-algebra/BFB203D6C27835B4E89C05B5C7B6BE75.

Appel:2012:ATS

[AF12] Andrew W. Appel and Solomon Feferman, editors. Alan Turing’s systems of logic: the Princeton thesis. Princeton University Press, Princeton, NJ, USA, facsimile edition, 2012. ISBN 0-691-15574-7 (hardcover). xv + 142 pp. LCCN QA9.2 .T86 2012.

Ash:2014:ETC

[AG14] Avner Ash and Robert Gross. Elliptic tales: curves, counting, and number theory. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-16350-2 (paperback). xvii + 253 pp. LCCN QA343 .A97 2012.

Abbott:2009:FEN

[ALB09] Edwin Abbott Abbott, William F. Lindgren, and Thomas Banchoff. Flatland: an edition with notes and commentary. Cambridge University Press, Cambridge, UK, 2009. ISBN 1-4114-2828-5, 1-4043-0321-9, 0-521-76988-4 (hardcover), 0-521-75994-3 (paperback). 124 pp. LCCN QA699 .A13 2009.

Abbott:2010:FRM

[ALB10a] Edwin Abbott Abbott, William F. Lindgren, and Thomas F. Banchoff. Flatland: a romance of many dimensions. MAA spectrum. Cambridge University Press, Cambridge, UK, 2010. ISBN 1-139-19492-5 (e-book), 0-521-75994-3, 0-521-76988-4 (hardcover), 1-139-85529-8 (e-book). ix + 294 pp. LCCN QA699 .A13 2010.

Abbott:2010:FEN

[ALB10b] Edwin Abbott Abbott, William F. Lindgren, and Thomas F. Banchoff. Flatland: an edition with notes and commentary. Cambridge University Press, Cambridge, UK, 2010. ISBN 1-139-85529-8 (e-book), 0-521-76988-4 (hardcover), 0-521-75994-3 (paperback). ix + 294 pp. LCCN QA699 .A13 2010. URL http://lib.

myilibrary.com?id=538493.

Alcock:2012:HSM

[Alc12] Lara Alcock. How to Study for a Mathematics Degree. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-966132-4 (paperback), 0-19-163736-X (e-book). xvi + 272 pp. LCCN QA11.2 .A43 2012.

Alcock:2013:HSM

[Alc13] Lara Alcock. How to study for a mathematics degree. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-966132-4 (paperback). xvi + 272 pp. LCCN QA11.2 .A452 2013.

Alexander:2010:DDH

[Ale10] Amir Alexander. Duel at Dawn: Heroes, Martyrs, and the Rise of Modern Mathematics. Harvard University Press, Cambridge, MA, USA, 2010. ISBN 0-674-04661-7. 307 pp. LCCN QA10.7 .A44 2010.

Allen:2018:FRI

[All18] Rory Allen. Fisher’s F -ratio illustrated graphically. The Mathematical Gazette, 102(553):50–62, March 2018. CODEN MAGAAS.

ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/fishers-fratio-illustrated-graphically/CB219DBA8F54455FE35EA253FAB18BF8.

Alvarez:2017:C

[Alv17] Marcos Alvarez. 101.41 on conics. The Mathematical Gazette, 101(552):511–515, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10141-on-conics/7F030BB3BBA21A5B1954FF54BA969BEB.

AAstrom:2008:FSI

[ÅM08] Karl J. (Karl Johan) Åström and Richard M. Murray. Feedback systems: an introduction for scientists and engineers. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-13576-2. xii + 396 pp. LCCN TJ216 .A78 2008. URL http:

//www.loc.gov/catdir/enhancements/fy0806/2007061033-b.
html;   http://www.loc.gov/catdir/enhancements/fy0806/
2007061033-d.html;   http://www.loc.gov/catdir/enhancements/
fy0834/2007061033-t.html.

Apostol:2011:SSI

[AM11] Tom M. Apostol and Mamikon A. Mnatsakanian. Sums of squares of integers in arithmetic progression. The Mathematical Gazette, 95(533):186–196, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Atanasiu:2015:EOC

[AM15] Dragu Atanasiu and Piotr Mikusiński. An elementary operational calculus. The Mathematical Gazette, 99(544):54–59, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anderson:2010:BRBa

[And10a] Ian Anderson. Book review: An introduction to the theory of numbers (sixth edition), by G. H. Hardy and E. M. Wright. Pp. 620. 2008. £30 (paperback). ISBN: 978-0-19-921986-5 (Oxford University Press). The Mathematical Gazette, 94(529):184, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anderson:2010:BRBb

[And10b] Ian Anderson. Book review: An introduction to the theory of numbers (sixth edition), by G. H. Hardy and E. M. Wright. Pp. 620. 2008. £30 (paperback). ISBN: 978-0-19-921986-5 (Oxford University Press). The Mathematical Gazette, 94(529):184, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anderson:2011:BRB

[And11] Charlotte Anderson. Book review: The mathematical mechanic, by Mark Levi. Pp. 186. £13.95 (hbk). 2009. ISBN 978-0-691-14020-9 (Princeton University Press). The Mathematical Gazette, 95 (532):165–166, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Andrews:2012:LCP

[And12] Paul Andrews. Learning from others can PISA and TIMSS really inform curriculum developments in mathematics? The Mathematical Gazette, 96(537):386–407, November 2012. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anderson:2013:BRBb

[And13a] Charlotte Anderson. Book review: Mathematical expeditions: exploring word problems across the ages, by Frank Swetz. pp. 192. £15.50. 2012. ISBN: 978-1-4214-0438-7 (Johns Hopkins University Press). The Mathematical Gazette, 97(540):554–555, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anderson:2013:BRBc

[And13b] Charlotte Anderson. Book review: The ballet of the planets: on the mathematical elegance of planetary motion, by Donald C. Benson. Pp. 162. $35.00 (Hardback). 2012. ISBN: 978-0-19-989100-9 (Oxford University Press). The Mathematical Gazette, 97(540): 563–565, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anderson:2013:BRBa

[And13c] Charlotte Anderson. Book review: The quantum story a history in 40 moments by Jim Baggott Pp. 469 £16.99 (hbk). 2011. ISBN: 978-0-19-956684-6 (Oxford University Press). The Mathematical Gazette, 97(539):357–358, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anderson:2013:BRBd

[And13d] Ian Anderson. Book review: Mathematical excursions to the world’s great buildings, by Alexander J. Hahn. Pp. 318. £34.95.(Hardback). 2012. ISBN: 978-0-691-14520-4 (Princeton University Press). The Mathematical Gazette, 97(540):566–567, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Andrilli:2017:SST

[And17] Stephen Andrilli. Some strategies for teaching a course based on Gödel, Escher, Bach. The Mathematical Gazette, 101(550): 50–59, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/CB0EEB5A4FBB5CD1865DDD985852D237.

Anderson:2019:BNN

[And19] Ian Anderson. Book review: The beauty of numbers in nature by Ian Stewart, pp. 224, £14.99 (paper), ISBN 978-1-78240-471-2, Ivy Press (2017). The Mathematical Gazette, 103 (557):368, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/beauty-
of-numbers-in-nature-by-ian-stewart-pp-224-1499-paper-
isbn-9781782404712-ivy-press-2017/7850C1224A5F123605DDDCFE2E90E74E.

Anonymous:2006:AFQ

[Ano06] Anonymous. Answering Fundamental Questions: An Introduction to Mathematics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-531076-4. xx + 745 pp. LCCN ????

Anonymous:2010:AGM

[Ano10a] Anonymous. AQA GCSE Maths Higher Revision Guide. Letts Educational, ????, 2010. ISBN 1-84419-286-5. ???? pp. LCCN ????

Anonymous:2010:Fa

[Ano10b] Anonymous. Feedback. The Mathematical Gazette, 94(529):151– 158, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2010:Fb

[Ano10c] Anonymous. Feedback. The Mathematical Gazette, 94(530):336– 341, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2010:Fc

[Ano10d] Anonymous. Feedback. The Mathematical Gazette, 94(531):536– 543, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2010:M

[Ano10e] Anonymous. In memoriam. The Mathematical Gazette, 94(530): 353, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2010:IBG

[Ano10f] Anonymous. Index to the Mathematical Gazette. The Mathematical Gazette, 94(I1):i–ix, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2010:MVIa

[Ano10g] Anonymous. MAG volume 94 issue 529–531 cover and front matter. The Mathematical Gazette, 94(I1):f1, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2010:MVIb

[Ano10h] Anonymous. MAG volume 94 issue 529 cover and back matter. The Mathematical Gazette, 94(529):b1–b3, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2010:MVIc

[Ano10i] Anonymous. MAG volume 94 issue 529 cover and front matter. The Mathematical Gazette, 94(529):f1–f2, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2010:MVIe

[Ano10j] Anonymous. MAG volume 94 issue 530 cover and back matter. The Mathematical Gazette, 94(530):b1–b3, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2010:MVId

[Ano10k] Anonymous. MAG volume 94 issue 530 cover and front matter. The Mathematical Gazette, 94(530):f1–f2, July 2010. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2010:MVIg

[Ano10l] Anonymous. MAG volume 94 issue 531 cover and back matter. The Mathematical Gazette, 94(531):b1–b3, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2010:MVIf

[Ano10m] Anonymous. MAG volume 94 issue 531 cover and front matter. The Mathematical Gazette, 94(531):f1–f3, November 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2011:A

[Ano11a] Anonymous. Acknowledgements. The Mathematical Gazette, 95 (534):575, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2011:Fa

[Ano11b] Anonymous. Feedback. The Mathematical Gazette, 95(532):128– 132, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2011:Fb

[Ano11c] Anonymous. Feedback. The Mathematical Gazette, 95(533):359– 361, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2011:Fc

[Ano11d] Anonymous. Feedback. The Mathematical Gazette, 95(534):557– 559, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2011:IBG

[Ano11e] Anonymous. Index to the Mathematical Gazette. The Mathematical Gazette, 95(I1):i–ix, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2011:MVIa

[Ano11f] Anonymous. MAG volume 95 issue 532–534 cover and front matter. The Mathematical Gazette, 95(I1):f1, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2011:MVIc

[Ano11g] Anonymous. MAG volume 95 issue 532 cover and back matter. The Mathematical Gazette, 95(532):b1–b3, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2011:MVIb

[Ano11h] Anonymous. MAG volume 95 issue 532 cover and front matter. The Mathematical Gazette, 95(532):f1–f2, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2011:MVIe

[Ano11i] Anonymous. MAG volume 95 issue 533 cover and back matter. The Mathematical Gazette, 95(533):b1–b3, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2011:MVId

[Ano11j] Anonymous. MAG volume 95 issue 533 cover and front matter. The Mathematical Gazette, 95(533):f1–f2, July 2011. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2011:MVIg

[Ano11k] Anonymous. MAG volume 95 issue 534 cover and back matter. The Mathematical Gazette, 95(534):b1–b3, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2011:MVIf

[Ano11l] Anonymous. MAG volume 95 issue 534 cover and front matter. The Mathematical Gazette, 95(534):f1–f3, November 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2012:A

[Ano12a] Anonymous. Acknowledgements. The Mathematical Gazette, 96 (537):541, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2012:Fa

[Ano12b] Anonymous. Feedback. The Mathematical Gazette, 96(535):174– 176, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2012:Fb

[Ano12c] Anonymous. Feedback. The Mathematical Gazette, 96(537):553– 559, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2012:FID

[Ano12d] Anonymous. Fueling innovation and discovery: the mathematical sciences in the 21st century. National Academies Press, Washington, DC, USA, 2012. ISBN 0-309-25473-6 (paperback), 0-309-25474-4 (pdf). v + 57 pp. LCCN T173.4 .F84 2012.

Anonymous:2012:IBG

[Ano12e] Anonymous. Index to the Mathematical Gazette. The Mathematical Gazette, 96(I1):i–ix, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2012:MVIa

[Ano12f] Anonymous. MAG volume 96 issue 535–537 cover and front matter. The Mathematical Gazette, 96(I1):f1, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2012:MVIc

[Ano12g] Anonymous. MAG volume 96 issue 535 cover and back matter. The Mathematical Gazette, 96(535):b1–b4, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2012:MVIb

[Ano12h] Anonymous. MAG volume 96 issue 535 cover and front matter. The Mathematical Gazette, 96(535):f1–f2, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2012:MVIe

[Ano12i] Anonymous. MAG volume 96 issue 536 cover and back matter. The Mathematical Gazette, 96(536):b1–b3, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2012:MVId

[Ano12j] Anonymous. MAG volume 96 issue 536 cover and front matter. The Mathematical Gazette, 96(536):f1–f2, July 2012. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2012:MVIg

[Ano12k] Anonymous. MAG volume 96 issue 537 cover and back matter. The Mathematical Gazette, 96(537):b1–b4, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2012:MVIf

[Ano12l] Anonymous. MAG volume 96 issue 537 cover and front matter. The Mathematical Gazette, 96(537):f1–f3, November 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2013:Fa

[Ano13a] Anonymous. Feedback. The Mathematical Gazette, 97(538):172– 177, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2013:Fb

[Ano13b] Anonymous. Feedback. The Mathematical Gazette, 97(539):337– 340, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2013:Fc

[Ano13c] Anonymous. Feedback. The Mathematical Gazette, 97(540):538– 543, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2013:IBG

[Ano13d] Anonymous. Index to the Mathematical Gazette. The Mathematical Gazette, 97(I1):iii–xi, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2013:MVIa

[Ano13e] Anonymous. MAG volume 97 issue 538–540 cover and front matter. The Mathematical Gazette, 97(I1):f1, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2013:MVIc

[Ano13f] Anonymous. MAG volume 97 issue 538 cover and back matter. The Mathematical Gazette, 97(538):b1–b3, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2013:MVIb

[Ano13g] Anonymous. MAG volume 97 issue 538 cover and front matter. The Mathematical Gazette, 97(538):f1–f2, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2013:MVIe

[Ano13h] Anonymous. MAG volume 97 issue 539 cover and back matter. The Mathematical Gazette, 97(539):b1–b3, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2013:MVId

[Ano13i] Anonymous. MAG volume 97 issue 539 cover and front matter. The Mathematical Gazette, 97(539):f1–f2, July 2013. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2013:MVIg

[Ano13j] Anonymous. MAG volume 97 issue 540 cover and back matter. The Mathematical Gazette, 97(540):b1–b3, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2013:MVIf

[Ano13k] Anonymous. MAG volume 97 issue 540 cover and front matter. The Mathematical Gazette, 97(540):f1–f3, November 2013. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2013:MS

[Ano13l] Anonymous. The mathematical sciences in 2025. National Academies Press, Washington, DC, USA, 2013. ISBN 0-309-28457-0 (paperback). xv + 205 pp. LCCN QA13 .M345 2013.

Anonymous:2013:PCU

[Ano13m] Anonymous. Padlock challenges: using number and algebra. The Mathematical Association, Leicester, UK, 2013. ISBN 0-906588-79-0 (paperback). 80 pp. LCCN ????

Anonymous:2014:IMO

[Ano14a] Anonymous. 55th International Mathematical Olympiad, Cape town, 3–13 July 2014. The Mathematical Gazette, 98(543):546, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:A

[Ano14b] Anonymous. Acknowledgements. The Mathematical Gazette, 98 (543):575, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:BR

[Ano14c] Anonymous. Books received. The Mathematical Gazette, 98(543): 574, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:Fa

[Ano14d] Anonymous. Feedback. The Mathematical Gazette, 98(541):133– 135, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:Fb

[Ano14e] Anonymous. Feedback. The Mathematical Gazette, 98(542):356– 357, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:PM

[Ano14f] Anonymous. In the pipeline for March 2015. The Mathematical Gazette, 98(543):546, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:MVIb

[Ano14g] Anonymous. MAG volume 98 issue 541 cover and back matter. The Mathematical Gazette, 98(541):b1–b3, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:MVIa

[Ano14h] Anonymous. MAG volume 98 issue 541 cover and front matter. The Mathematical Gazette, 98(541):f1–f2, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:MVId

[Ano14i] Anonymous. MAG volume 98 issue 542 cover and back matter. The Mathematical Gazette, 98(542):b1–b4, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:MVIc

[Ano14j] Anonymous. MAG volume 98 issue 542 cover and front matter. The Mathematical Gazette, 98(542):f1–f2, July 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:MVIf

[Ano14k] Anonymous. MAG volume 98 issue 543 cover and back matter. The Mathematical Gazette, 98(543):b1–b2, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:MVIe

[Ano14l] Anonymous. MAG volume 98 issue 543 cover and front matter. The Mathematical Gazette, 98(543):f1–f3, November 2014. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:MA

[Ano14m] Anonymous. The mathematical association. The Mathematical Gazette, 98(543):576, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:PSW

[Ano14n] Anonymous. On 87.24: Paul Stephenson writes. The Mathematical Gazette, 98(543):528–530, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:NLW

[Ano14o] Anonymous. On 98.05: Nick Lord writes. The Mathematical Gazette, 98(543):534, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:ABW

[Ano14p] Anonymous. On 98.15: Alan Beardon writes. The Mathematical Gazette, 98(543):534–535, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:FFG

[Ano14q] Anonymous. On fortuitous functions: The GCHQ Problem Solving Group writes. The Mathematical Gazette, 98(543):531–532, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2014:LOE

[Ano14r] Anonymous. On the largest outscribed equilateral triangle: Nick Lord writes. The Mathematical Gazette, 98(543):533–534, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2015:Fa

[Ano15a] Anonymous. Feedback. The Mathematical Gazette, 99(544):157– 164, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2015:Fb

[Ano15b] Anonymous. Feedback. The Mathematical Gazette, 99(545):363– 367, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2015:Fc

[Ano15c] Anonymous. Feedback. The Mathematical Gazette, 99(546):545– 546, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2015:PJ

[Ano15d] Anonymous. In the pipeline for July 2015. The Mathematical Gazette, 99(544):74, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2015:MVIb

[Ano15e] Anonymous. MAG volume 99 issue 544 cover and back matter. The Mathematical Gazette, 99(544):b1–b2, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2015:MVIa

[Ano15f] Anonymous. MAG volume 99 issue 544 cover and front matter. The Mathematical Gazette, 99(544):f1–f2, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2015:MVId

[Ano15g] Anonymous. MAG volume 99 issue 545 cover and back matter. The Mathematical Gazette, 99(545):b1–b3, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2015:MVIc

[Ano15h] Anonymous. MAG volume 99 issue 545 cover and front matter. The Mathematical Gazette, 99(545):f1–f2, July 2015. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2015:MVIf

[Ano15i] Anonymous. MAG volume 99 issue 546 cover and back matter. The Mathematical Gazette, 99(546):b1–b4, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2015:MVIe

[Ano15j] Anonymous. MAG volume 99 issue 546 cover and front matter. The Mathematical Gazette, 99(546):f1–f3, November 2015. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2016:BR

[Ano16a] Anonymous. Books received. The Mathematical Gazette, 100 (547):189–190, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2016:F

[Ano16b] Anonymous. Feedback. The Mathematical Gazette, 100(547):161– 163, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2016:MVIb

[Ano16c] Anonymous. MAG volume 100 issue 547 cover and back matter. The Mathematical Gazette, 100(547):b1–b4, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2016:MVIa

[Ano16d] Anonymous. MAG volume 100 issue 547 cover and front matter. The Mathematical Gazette, 100(547):f1–f2, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Anonymous:2017:BR

[Ano17a] Anonymous. Books received. The Mathematical Gazette, 101 (550):190, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/85FBF3ECB90022050230E11C939F722B.

Anonymous:2017:CBB

[Ano17b] Anonymous. Centenary of the birth of Bill Tutte. The Mathematical Gazette, 101(551):279, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.

cambridge.org/core/product/EF7EDD7EFFB42BD3B051373489EFB47A.

Anonymous:2017:Fa

[Ano17c] Anonymous. Feedback. The Mathematical Gazette, 101(550): 146–149, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/15CF48512269C9089EF469BEA5D8F69B.

Anonymous:2017:Fb

[Ano17d] Anonymous. Feedback. The Mathematical Gazette, 101(551): 337–338, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/D5020E511FA9383C0DD9EDD279F91029.

Anonymous:2017:Fc

[Ano17e] Anonymous. Feedback. The Mathematical Gazette, 101 (552):542–545, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/feedback/FCDBC41281B223B4282AC049D3614E8A.

Anonymous:2017:LB

[Ano17f] Anonymous. Looking back. The Mathematical Gazette, 101 (552):400, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/BD98F9659DEF53029F0345E582C59913.

Anonymous:2017:MVIb

[Ano17g] Anonymous. MAG volume 101 issue 550 cover and back matter. The Mathematical Gazette, 101(550):b1–b2, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
2536E3C1B451B9D56A9893E8C71BD949.

Anonymous:2017:MVIa

[Ano17h] Anonymous. MAG volume 101 issue 550 cover and front matter. The Mathematical Gazette, 101(550):f1–f3, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
C3655A8625D708B4F4D0B9D5DB287461.

Anonymous:2017:MVId

[Ano17i] Anonymous. MAG volume 101 issue 551 cover and back matter. The Mathematical Gazette, 101(551):b1, July 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
2BDFF6A4DCF68F2D0E82B7B97586C4CE.

Anonymous:2017:MVIc

[Ano17j] Anonymous. MAG volume 101 issue 551 cover and front matter. The Mathematical Gazette, 101(551):f1–f3, July 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
46DF1874D0EB713A4CF8BD2D440BAF93.

Anonymous:2017:MVIf

[Ano17k] Anonymous. MAG volume 101 issue 552 cover and back matter. The Mathematical Gazette, 101(552):b1–b3, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-101-issue-552-
cover-and-back-matter/80374430CE2EA819B4BD512073D8D37E.

Anonymous:2017:MVIe

[Ano17l] Anonymous. MAG volume 101 issue 552 cover and front matter. The Mathematical Gazette, 101(552):f1–f2, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-101-issue-552-
cover-and-front-matter/28E6C16CFF4E64125BB19D6CD12A2CAD.

Anonymous:2017:MAa

[Ano17m] Anonymous. The Mathematical Association. The Mathematical Gazette, 101(550):191, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.

cambridge.org/core/product/9B03CBACB71B9D1BECD2FCAE54A523D4.

Anonymous:2017:MAb

[Ano17n] Anonymous. The Mathematical Association. The Mathematical Gazette, 101(551):384, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.

cambridge.org/core/product/230C0F8531AE1CB8A2E1F85D73D73418.

Anonymous:2017:MAc

[Ano17o] Anonymous. The Mathematical Association. The Mathematical Gazette, 101(552):575, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/mathematical-association/8529A73FF663BFD470EE9A1B8BAFAB34.

Anonymous:2018:A

[Ano18a] Anonymous. Acknowledgements. The Mathematical Gazette, 102(555):573, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/acknowledgements/F57B65DF078F7089C5F28FC4535937F6.

Anonymous:2018:MGI

[Ano18b] Anonymous. The Mathematical Gazette information sheets. The Mathematical Gazette, 102(555):493, November 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mathematical-gazette-information-
sheets/143E106960926959577506BE1703D3FC.

Anonymous:2018:Fa

[Ano18c] Anonymous. Feedback. The Mathematical Gazette, 102(554):354, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/feedback/9643B61B3F73904321A829522213F714.

Anonymous:2018:Fb

[Ano18d] Anonymous. Feedback. The Mathematical Gazette, 102 (555):540–542, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/feedback/AA3B86534765A0414AC96D51F7544AF9.

Anonymous:2018:GG

[Ano18e] Anonymous. Growing the gazette. The Mathematical Gazette, 102 (554):301, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/growing-the-
gazette/CACF7F5595CA8C4CD9D83B28ED3CDA60.

Anonymous:2018:PM

[Ano18f] Anonymous. In the pipeline for March 2019. The Mathematical Gazette, 102(555):539, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/in-the-pipeline-for-march-2019/554662DF9A281EE71D343B81CDB26DD2.

Anonymous:2018:MVIb

[Ano18g] Anonymous. MAG volume 102 issue 553 cover and back matter. The Mathematical Gazette, 102(553):b1, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-102-issue-553-
cover-and-back-matter/531C482F3AD948C5A9A161DDCA9797ED.

Anonymous:2018:MVIa

[Ano18h] Anonymous. MAG volume 102 issue 553 cover and front matter. The Mathematical Gazette, 102(553):f1–f3, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-102-issue-553-
cover-and-front-matter/7EF293FD8315816872EDEEF9406D8838.

Anonymous:2018:MVId

[Ano18i] Anonymous. MAG volume 102 issue 554 cover and back matter. The Mathematical Gazette, 102(554):b1, July 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-102-issue-554-
cover-and-back-matter/E57613DE37E1D167743C71879E949546.

Anonymous:2018:MVIc

[Ano18j] Anonymous. MAG volume 102 issue 554 cover and front matter. The Mathematical Gazette, 102(554):f1–f3, July 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-102-issue-554-
cover-and-front-matter/C0082905B894AD7F8668B39377CB4D80.

Anonymous:2018:MVIf

[Ano18k] Anonymous. MAG volume 102 issue 555 cover and back matter. The Mathematical Gazette, 102(555):b1–b2, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-102-issue-555-
cover-and-back-matter/168F7EFA9DF3CE5BFCDAAD4FE5F11260.

Anonymous:2018:MVIe

[Ano18l] Anonymous. MAG volume 102 issue 555 cover and front matter. The Mathematical Gazette, 102(555):f1–f3, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-102-issue-555-
cover-and-front-matter/9F01EEEAF6ACCA0D2B06DCBC358CDC9D.

Anonymous:2018:MAc

[Ano18m] Anonymous. The Mathematical Association. The Mathematical Gazette, 102(555):575, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/mathematical-association/1649D1D4BAEA1B35C31F62CA44B4AF72.

Anonymous:2018:MAa

[Ano18n] Anonymous. The Mathematical Association. The Mathematical Gazette, 102(553):192, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/mathematical-association/CFF099025D94C6143002DB016917ECA2.

Anonymous:2018:MAb

[Ano18o] Anonymous. The Mathematical Association. The Mathematical Gazette, 102(554):384, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/mathematical-association/6D86E45AE59FDEBBB3A113F2E2B32C41.

Anonymous:2019:A

[Ano19a] Anonymous. Acknowledgements. The Mathematical Gazette, 103(558):574, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/acknowledgements/FB818E38E36089B046DB53A8B9DA4FDD.

Anonymous:2019:F

[Ano19b] Anonymous. Feedback. The Mathematical Gazette, 103(556):171, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/feedback/E3890DD4F144723F5898E2FB25903816.

Anonymous:2019:PJ

[Ano19c] Anonymous. In the pipeline for July 2019. The Mathematical Gazette, 103(556):171, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/in-the-pipeline-for-july-2019/5B6A272F4042363BA17F873DDF4CDDFC.

Anonymous:2019:PM

[Ano19d] Anonymous. In the pipeline for March 2020. The Mathematical Gazette, 103(558):460, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/in-the-pipeline-for-march-2020/84F4AD6202350996A68F58964727EDB7.

Anonymous:2019:PN

[Ano19e] Anonymous. In the pipeline for November 2019. The Mathematical Gazette, 103(557):264, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/in-the-pipeline-for-november-2019/BB1366243C344C396F649A52F2EEDDDC.

Anonymous:2019:MVIb

[Ano19f] Anonymous. MAG volume 103 issue 556 cover and back matter. The Mathematical Gazette, 103(556):b1–b2, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-103-issue-556-
cover-and-back-matter/F8400181DAC6FF51C755BBED45033672.

Anonymous:2019:MVIa

[Ano19g] Anonymous. MAG volume 103 issue 556 cover and front matter. The Mathematical Gazette, 103(556):f1–f3, March 2019.

CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-103-issue-556-
cover-and-front-matter/C793789062CD3545D3B841ED0E58445B.

Anonymous:2019:MVId

[Ano19h] Anonymous. MAG volume 103 issue 557 cover and back matter. The Mathematical Gazette, 103(557):b1–b3, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-103-issue-557-
cover-and-back-matter/D6C1D8C5799C3E8A00510C4AFE49C5B3.

Anonymous:2019:MVIc

[Ano19i] Anonymous. MAG volume 103 issue 557 cover and front matter. The Mathematical Gazette, 103(557):f1–f2, July 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-103-issue-557-
cover-and-front-matter/0E3A31CCF6668E647242B68132482549.

Anonymous:2019:MVIf

[Ano19j] Anonymous. MAG volume 103 issue 558 cover and back matter. The Mathematical Gazette, 103(558):b1–b4, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-103-issue-558-
cover-and-back-matter/C62509ED6AA707F651DF56DEF087BE02.

Anonymous:2019:MVIe

[Ano19k] Anonymous. MAG volume 103 issue 558 cover and front matter. The Mathematical Gazette, 103(558):f1–f2, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/mag-volume-103-issue-558-
cover-and-front-matter/2CF9816287185D1F3977B95A34D67EF5.

Applebaum:2010:BRBb

[App10a] Dave Applebaum. Book review: Hilbert space methods and quantum mechanics, by Werner O. Amrein. Pp. 395. $94.95 (hbk). 2009. ISBN: 978-1-200-6681-4 (CRC Press). The Mathematical Gazette, 94(531):571, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Applebaum:2010:BRBa

[App10b] David Applebaum. Book review: Advanced calculus: an introduction to linear analysis, by Leonard F. Richardson, J. Wiley and Sons, Inc. (2008), ISBN: 978-0-470-23288-0 £78.50, $97.50 (hbk) pp. 387. The Mathematical Gazette, 94(530):372–373, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Applebaum:2011:BRB

[App11] Dave Applebaum. Book review: Harmonic analysis of mean periodic functions on symmetric spaces and the Heisenberg group, by Valery V. Volchkov and Vitaly V. Volchkov. Pp. 671 $59. 99 (hbk). 2009. ISBN 978-1-84882-532-1 (Springer Monographs in Mathematics, Springer-Verlag London Ltd.). The Mathematical Gazette, 95(534):571–572, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Applebaum:2012:LLE

[App12] David Applebaum. Limits, limits everywhere: the tools of mathematical analysis. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-964008-4. xv + 200 pp. LCCN QA300 .A578 2012. URL http:

//www.loc.gov/catdir/enhancements/fy1409/2011945216-b.
html;   http://www.loc.gov/catdir/enhancements/fy1409/
2011945216-d.html;   http://www.loc.gov/catdir/enhancements/
fy1409/2011945216-t.html.

Applebaum:2014:BRS

[App14a] Dave Applebaum. Book review: Stochastic analysis and diffusion processes by G. Kallianpur and P. Sundar, pp 352, £65 (hbk), £29.99 (pbk) ISBN 978-0-965706-3, ISBN 978-0-965707-0, Oxford Graduate Texts in Mathematics 24, Oxford University Press (2014). The Mathematical Gazette, 98(543):568–570, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Applebaum:2014:BRB

[App14b] David Applebaum. Book review: Concrete functional calculus, by R. M. Dudley and R. Norvaisa. Pp. 671, £90.00, 2011, ISSN 1439-7382, ISBN 978-1-4419-6949-1 (Springer Monographs in Mathematics). The Mathematical Gazette, 98(541):166–168, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Applebaum:2018:BRI

[App18] Dave Applebaum. Book review: Introductory topology, exercises and solutions by Mohammed Hichem Mortad, pp. 356, £40, (paper), ISBN 978-9-81314-802-4, World Scientific (2017). The Mathematical Gazette, 102(553):185–186, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/introductory-topology-exercises-
and-solutions-by-mohammed-hichem-mortad-pp-356-40-paper-
isbn-9789813148024-world-scientific-2017/46A1C95B00E0501473E6CC144338FCAB.

Andrews:2014:MME

[AR14] Paul (Paul R.) Andrews and Tim Rowland, editors. Masterclass in mathematics education: international perspectives on teaching and learning. MasterClass. Bloomsbury Academic, an imprint of Bloomsbury Publishing Plc, London, UK, 2014. ISBN 1-4411-7235-1 (hardcover), 1-4411-7975-5 (paperback). xv + 223 pp. LCCN QA11.2 .M2679 2014.

Aravind:2015:VSP

[Ara15] P. K. Aravind. Variations on a sangaku problem involving kissing spheres. The Mathematical Gazette, 99(546):504–515, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Arianrhod:2012:SLE

[Ari12] Robyn Arianrhod. Seduced by logic: Émilie du Châtelet, Mary Somerville, and the Newtonian revolution. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-993161-5. 338 pp. LCCN Q141 .A725 2012.

Abu-Saymeh:2019:MVS

[ASH19a] Sadi Abu-Saymeh and Mowaffaq Hajja. More variations on the Steiner–Lehmus theme. The Mathematical Gazette, 103(556):1–11, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/more-variations-
on-the-steinerlehmus-theme/ECB104D2233FB62B7EE182554CD78A22.

Abu-Saymeh:2019:NPC

[ASH19b] Sadi Abu-Saymeh and Mowaffaq Hajja. New proofs of certain characterisations of cyclic circumscriptible quadrilaterals. The Mathematical Gazette, 103(558):401–408, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/new-proofs-
of-certain-characterisations-of-cyclic-circumscriptible-
quadrilaterals/C787FB4BC5539210774640FB6C0775E1.

Abu-Saymeh:2018:LAU

[ASHH18] Saadi Abu-Saymeh, Mowaffaq Hajja, and Mostafa Hayajneh. 102.50 long altitudes and an unexpected appearance of the golden ratio. The Mathematical Gazette, 102(555):523–527, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10250-long-
altitudes-and-an-unexpected-appearance-of-the-golden-ratio/
5B391F2A6C89DCCA601A4BEC815D54C6.

Ash:2012:RNF

[AT12] J. Marshall Ash and Yiren Tan. 96.06 A rational number of the form a^{a} with a irrational. The Mathematical Gazette, 96(535):106– 108, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Audin:2011:RSK

[Aud11] Michèle Audin. Remembering Sofya Kovalevskaya. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 0-85729-928-X, 0-85729-929-8 (e-book). ix + 284 pp. LCCN QA29.K67 A8313 2011.

Audet:2014:OMT

[Aud14] Charles Audet. Ordering 15 marbles with a three-way scale. The Mathematical Gazette, 98(542):304–316, July 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ault:2019:MRA

[Aul19] Irene Ault. Margaret E. Rayner 21st August 1929–31st May 2019. The Mathematical Gazette, 103(558):385–387, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/margaret-e-
rayner-21st-august-1929-31st-may-2019/70EDBE69205233ECCD33CBB301114729.

Ahmed:2007:EBM

[AW07] Afzal Ahmed and Honor Williams. Even better mathematics: looking back to move forward. Network Continuum, London, UK, 2007. ISBN 1-85539-147-3. viii + 99 pp. LCCN QA11.2

.A36 2007. URL http://www.loc.gov/catdir/enhancements/
fy1213/2007029517-b.html;   http://www.loc.gov/catdir/
enhancements/fy1214/2007029517-d.html;   http://www.loc.
gov/catdir/toc/ecip0723/2007029517.html.

Aydin:2018:SDH

[Ayd18] Hazar Aydin. 102.46 showing the divergence of the harmonic series using integration. The Mathematical Gazette, 102 (555):514–515, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10246-showing-the-divergence-of-the-harmonic-series-
using-integration/B2C3B13F5FB49A545F29C967BC8B0F5A.

Allaire:2012:GHC

[AZY12] Patricia R. Allaire, Junmin Zhou, and Haishen Yao. 96.25 Going halfway with circular boundaries. The Mathematical Gazette, 96 (535):161–165, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bartocci:2011:VMM

[B^{+}11] C. (Claudio) Bartocci et al., editors. Vite matematiche. Mathematical lives: protagonists of the twentieth century from Hilbert to Wiles. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-13605-2 (hardcover). xiii + 238 pp. LCCN QA28 .M36513 2011; QA28 .M38 2011. Translated by Kim Williams.

Baggott:2011:QSH

[Bag11] Jim Baggott. The quantum story: a history in 40 moments. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-956684-4 (hardcover), 0-19-965597-9 (paperback). xix + 469 + 16 pp. LCCN QC173.98 .B34 2011.

Bajc:2013:PSS

[Baj13] Drago Bajc. 97.01 Power solutions of some Diophantine equations. The Mathematical Gazette, 97(538):107–110, March 2013. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bankov:2010:BRBa

[Ban10a] Kiril Bankov. Book review: Communicating mathematics in the digital era, by J. M. Borwein, E. M. Rocha, and J. F. Rodrigues. Pp. 325. $49.00 (Hardback), 2008. ISBN 978-1-56881-410-0 (A. K. Peters, Ltd.). The Mathematical Gazette, 94(531):557–559, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bankov:2010:BRBc

[Ban10b] Kiril Bankov. Book review: Mathematics without fears, by Peter Sprent. pp. 453. £14.59. 2009. ISBN 978-1-4092-5670-0 (Lulu). The Mathematical Gazette, 94(531):562–563, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bankov:2010:BRBb

[Ban10c] Kiril Bankov. Book review: What is the point? Motivation and the mathematics crisis, by David Wells. Pp. 309. £10.00. 2008. ISBN 978-0-907944-11-9 (Rain Press). The Mathematical Gazette, 94(531):561–562, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bankov:2013:BRB

[Ban13] Kiril Bankov. Book review: Measurement, by Paul Lockhart. Pp. 398. £20.00. 2012. ISBN: 978-0-674-05755-5 (Harvard University Press). The Mathematical Gazette, 97(540):565–566, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bankov:2014:BRBb

[Ban14a] Kiril Bankov. Book review: The beauty of everyday mathematics, by Norbert Herrmann. Pp. 138. £17.99. 2012. ISBN 978-3-642-22103-3 (Springer). The Mathematical Gazette, 98(541):160–161, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bankov:2014:BRBa

[Ban14b] Kiril Bankov. Book review: Why cats land on their feet? And 76 other physical paradoxes and puzzles, by Mark Levi. Pp. 190. US $19.95, 2012. ISBN 978-0-691-14854-0 (Princeton University Press). The Mathematical Gazette, 98(541):157–158, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bantchev:2016:VTP

[Ban16] Boyko B. Bantchev. 100.17 the vector triple product and planar vector algebra. The Mathematical Gazette, 100(548):314–317, July 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
D5F171080A6E2C9AC3BBBD80D7A41460.

Banks:2018:SLS

[Ban18] Jessica E. Banks. Is a straight line the shortest path? The Mathematical Gazette, 102(553):1–12, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/is-a-straight-line-the-shortest-path/C6F34EBFEAB29D370ECBA193BDB0D550.

Barnes:2009:GG

[Bar09] J. G. P. (John Gilbert Presslie) Barnes. Gems of geometry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-642-05091-3 (hardcover), 3-642-05092-1 (e-book). xiii + 311 pp. LCCN QA445 .B36 2009.

Barnard:2010:UC

[Bar10] Tony Barnard. 94.39 An unlikely certainty. The Mathematical Gazette, 94(531):514–517, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Barbara:2011:RDP

[Bar11] Roy Barbara. 95.01 The rational distance problem. The Mathematical Gazette, 95(532):59–61, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Barbara:2013:RDP

[Bar13] Roy Barbara. 97.11 The rational distance problem for polygons. The Mathematical Gazette, 97(538):144–149, March 2013. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Barbara:2017:FLT

[Bar17] Roy Barbara. 101.01 Fermat’s last theorem for n = 3 revisited. The Mathematical Gazette, 101(550):90–93, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
EDDB6CD93CC9D1432C27B3F4B4BEA872.

Bataille:2014:QMP

[Bat14] Michel Bataille. Quand les mathématiques posent problème(s) : 120 exercices originaux et stimulants. Références sciences. Ellipses, Paris, France, 2014. ISBN 2-7298-8673-7 (paperback). ISSN 2260-

8044. 216 pp. LCCN ???? URL http://catalogue.bnf.fr/ark:
/12148/cb438350630.

Baumback:2013:TTT

[Bau13] Randall Baumback. A tale of two triangles: a collaboration between teacher and students. The Mathematical Gazette, 97(540): 524–532, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baumback:2014:FFF

[Bau14] Randall Baumback. Fingers find Fermat’s factorisation most probable. The Mathematical Gazette, 98(543):452–458, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2010:BRBc

[Bay10a] John Baylis. Book review: Hexaflexagons, probability paradoxes and the tower of Hanoi: Martin Gardner’s first book of mathematical puzzles and games, by Martin Gardner. Pp. 193. £8-99. 2008. ISBN 978-0-521-73525-4 (Cambridge University Press). The Mathematical Gazette, 94(530):375, July 2010. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2010:BRBb

[Bay10b] John Baylis. Book review: The backbone of Pascal’s triangle, by Martin Griffiths. Pp. 163. £12.00. 2007. ISBN 978-1-906001-04-9 (The United Kingdom Mathematics Trust). The Mathematical Gazette, 94(529):184–185, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2010:BRBa

[Bay10c] John Baylis. Book review: The magic numbers of the professor, by Owen O’Shea and Underwood Dudley. Pp. 168. £21.99. 2007. ISBN 0 88385 557 7 (The Mathematical Association of America). The Mathematical Gazette, 94(529):180–181, March 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2010:CC

[Bay10d] John Baylis. Codes, not ciphers. The Mathematical Gazette, 94 (531):412–425, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2011:BRBb

[Bay11a] John Baylis. Book review: Mathematics as problem solving (2nd edn), by Alexander Soifer. Pp. 106. £35.99. 2009. ISBN 978-0-387-74646-3 (Springer). The Mathematical Gazette, 95(532):171, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2011:BRBa

[Bay11b] John Baylis. Book review: The Monty Hall Problem — the remarkable story of math’s most contentious brainteaser, by Jason Rosenhouse. Pp. 194. £15.99. 2009. ISBN 978-0-195-36789-8 (Oxford University Press). The Mathematical Gazette, 95(532):167– 169, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2012:BRB

[Bay12] John Baylis. Book review: Problem solving — the creative side of mathematics, by Derek Holton. Pp. 211. £11.99 (members), £14.99 (non-members). 2010. ISBN: 978-0-906588-71-0 (The Mathematical Association). The Mathematical Gazette, 96(537):573–574, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2014:BRBa

[Bay14a] John Baylis. Book review: Elements of advanced mathematics (3rd edn.) by Steven G. Krantz. Pp. 351. £44.99 (hbk). 2012. ISBN: 978-1-4398-9834-5 (CRC Press). The Mathematical Gazette, 98 (541):146–147, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2014:BRBb

[Bay14b] John Baylis. Book review: Limits, limits everywhere: the tools of mathematical analysis, by David Applebaum. pp. 200. £19.99. 2012. ISBN: 978-0-19-964008-9 (Oxford University Press). The Mathematical Gazette, 98(541):170–171, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2014:BRBc

[Bay14c] John Baylis. Book review: Probability — A very short introduction, by John Haigh. Pp. 128. £7.99. 2012. ISBN 987-0-19-958848-0 (Oxford University Press). The Mathematical Gazette, 98(541):183–185, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2014:BRBd

[Bay14d] John Baylis. Book review: The great mathematical problems, by Ian Stewart. Pp. 340. £15.99. 2013. ISBN: 978-1-84668-1998 (Profile Books). The Mathematical Gazette, 98(542):379–381, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2015:BRC

[Bay15a] John Baylis. Book review: A course in mathematical analysis by D. J. H. Garling, pp 939 (paper, 3 volumes, £29.99 each), ISBN 978-1-107-61418-5, 978-1-107-67532-2, 978-1-107-66330-5, Cambridge University Press (2013, 2014). The Mathematical Gazette, 99(546):569–570, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2015:BRB

[Bay15b] John Baylis. Book review: Everyday calculus — discovering the hidden math all around us by Oscar E. Fernandez, pp. 150, £16.95 (hard), ISBN 978-0-691-15755-9, Princeton University Press (2014). The Mathematical Gazette, 99(544):181–182, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2015:WAI

[Bay15c] John Baylis. Why avoid induction? The Mathematical Gazette, 99(545):290–297, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Baylis:2017:BRAb

[Bay17a] John Baylis. Book review: Advanced calculus of a single variable, by Geveci, Tunc, pp. 382, £37.99 (hard), ISBN 978-3-31927-806-3, also available as e-book, Springer Verlag (2016). The Mathematical Gazette, 101(551):367–368, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
795976666721F6DCB2D8608C9EACE516.

Baylis:2017:BRAa

[Bay17b] John Baylis. Book review: Algebra in context, by Shell-Gellasch, Amy and Thoo, J. B., pp. 536, £80.50 (hard), ISBN 978-14214-1728-9, Johns Hopkins Press (2015). The Mathematical Gazette, 101(550):177–179, March 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
8992E4A7E0954EAD46B342A0F2BAC6E4.

Baylis:2017:BRM

[Bay17c] John Baylis. Book review: Measure and integration, by Bercovici, Hari; Brown, Arlen and Pearcy Carl, pp. 300, £46.99 (hard), ISBN 978-3-31929-044-7, also available as e-book, Springer Verlag (2016). The Mathematical Gazette, 101(551):358–359, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
C7C29EF64863062D8CF6CBB62A6B609A.

Baylis:2017:BRT

[Bay17d] John Baylis. Book review: Trigonometric functions and complex numbers by Desheng Yang, translated by Chunhui Shen, pp. 411, £32 (paper), ISBN 978-1-93813-486-9, World Scientific Publishing (2017). The Mathematical Gazette, 101 (552):562–563, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/trigonometric-functions-and-complex-numbers-by-
desheng-yang-translated-by-chunhui-shen-pp-411-32-paper-
isbn-9781938134869-world-scientific-publishing-2017/11B9F1C1107C18D61D0CCCBA4242CF4F.

Baylis:2019:IAR

[Bay19a] John Baylis. Book review: An introduction to analysis by Robert C. Gunning, pp. 370, £58 (hard), ISBN 978-0-69117-879-0, Princeton University Press (2018). The Mathematical Gazette, 103(557):369–370, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/an-introduction-to-analysis-by-robert-c-gunning-
pp-370-58-hard-isbn-9780691178790-princeton-university-
press-2018/9E64E11C7919B75F59767DFEDEDB0D64.

Baylis:2019:BRP

[Bay19b] John Baylis. Book review: Professor Higgins’s problem collection by Peter Higgins, pp. 126, £14.99 (paper), ISBN 978-0-19875-547-0, Oxford University Press (2017). The Mathematical Gazette, 103(556):190, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/professor-higginss-problem-collection-by-peter-
higgins-pp-126-1499-paper-isbn-9780198755470-oxford-university-
press-2017/5BED2892A6DFD5F33E527DA37A5CDC18.

Binder-Brantley:2012:DPF

[BB12] Joshua A. Binder-Brantley. 96.34 Determining the prime factorisation of a matrix. The Mathematical Gazette, 96(536):275–280, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Burgisser:2013:CGN

[BC13] Peter Bürgisser and Felipe Cucker. Condition: the geometry of numerical algorithms, volume 349 of Die Grundlehren der mathematischen Wissenschaften = A series of comprehensive studies in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-642-38895-7 (hardcover), 3-642-38896-5 (e-book). ISSN 0072-7830. xxxi + 554 pp.

LCCN QA297 .B845 2013. URL http://link.springer.com/
book/10.1007/978-3-642-38896-5/.

Buchanan:2015:A

[BC15] Andrew Buchanan and John Conway. 99.37 anthropologic. The Mathematical Gazette, 99(546):537–538, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Blight:2010:FAR

[BCL10] Barry Blight, Mike Cornelius, and Nick Lord. Finding an angle of rotation in coordinate geometry. The Mathematical Gazette, 94 (529):149–150, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ben-Chaim:2019:HMP

[BCSS19] David Ben-Chaim, Yechiel Shalitin, and Moshe Stupel. Historical mathematical problems suitable for classroom activities. The Mathematical Gazette, 103(556):12–19, March 2019.

CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/historical-mathematical-problems-
suitable-for-classroom-activities/BFA123243C49FEF418509D8633F891CB.

Bryant:2010:WWW

[BD10] Victor Bryant and John Duncan. 94.36 Wheels within wheels. The Mathematical Gazette, 94(531):502–505, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bhanu:2012:DPT

[BD12] K. S. Bhanu and M. N. Deshpande. 96.02 Dual Pythagorean triples. The Mathematical Gazette, 96(535):96–99, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Beauregard:2015:DEV

[BD15] Raymond A. Beauregard and Vladimir A. Dobrushkin. Differential equations v. power series. The Mathematical Gazette, 99(546):499– 503, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Beauregard:2016:MS

[BD16] Raymond A. Beauregard and Vladimir A. Dobrushkin. Multisection of series. The Mathematical Gazette, 100(549):460– 470, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/AD84414F418A881EBBA1BFB76C9602AE.

Beardon:2011:IIP

[Bea11] Alan F. Beardon. 95.20 Integral inequalities in probability theory. The Mathematical Gazette, 95(532):119–123, March 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Beardon:2012:PE

[Bea12a] A. F. Beardon. 96.04 Pythagorean equations. The Mathematical Gazette, 96(535):102–103, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Beardon:2012:RL

[Bea12b] A. F. Beardon. 96.35 Reflections on lattices. The Mathematical Gazette, 96(536):280–283, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Beardon:2012:WMS

[Bea12c] A. F. Beardon. What is the most symmetric quadrilateral? The Mathematical Gazette, 96(536):207–212, July 2012. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Beatty:2012:MIH

[Bea12d] Richard Beatty. Mathematics: an illustrated history of numbers. Worth Press, Cambridge, UK, 2012. ISBN 1-84931-066-1 (hardcover). LCCN QA21 .M36118 2012.

Beardon:2013:FGC

[Bea13] A. F. Beardon. Frieze groups, cylinders, and quotient groups. The Mathematical Gazette, 97(538):95–100, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Beardon:2015:ST

[Bea15a] A. F. Beardon. 99.36 stable tables. The Mathematical Gazette, 99 (546):534–536, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Beardon:2015:RVT

[Bea15b] A. F. Beardon. Removing a vertex from a tetrahedron. The Mathematical Gazette, 99(544):90–96, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Beardon:2017:PMM

[Bea17a] A. F. Beardon. 101.03 Products of 3 × 3 magic matrices. The Mathematical Gazette, 101(550):96–98, March 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
5CF601D45C426258D11FFDFF6E9A4A05.

Beardon:2017:AQ

[Bea17b] A. F. Beardon. 101.34 the area of a quadrilateral. The Mathematical Gazette, 101(552):492–494, November 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
7C1B2438BAEC1368341BBE537D7E7C36.

Beardon:2018:RP

[Bea18a] A. F. Beardon. 102.05 remarks on Problem 100.I. The Mathematical Gazette, 102(553):102–104, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10205-remarks-on-problem-
100i/FC48D7E90509472510C331A61AB622CF.

Beardon:2018:USP

[Bea18b] A. F. Beardon. 102.54 unravelling Simpson’s paradox. The Mathematical Gazette, 102(555):534–535, November 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10254-unravelling-simpsons-
paradox/E094BD0EBF3A427372ADFBF869B3C0C3.

Beardon:2019:TSL

[Bea19] A. F. Beardon. The three Steiner–Lehmus theorems. The Mathematical Gazette, 103(557):293–304, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/three-steinerlehmus-theorems/732D9DC073DFD0A3138A5F95834D1DC0.

Burgiel:2017:OQT

[BEH17] Heidi Burgiel and Mahmoud El-Hashash. Open questions on Tantrix graphs. The Mathematical Gazette, 101(550):83–89, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
9E815C4E6F6D653A8013862D1EE4E268.

Belcher:2019:PP

[Bel19] Paul Belcher. 103.14 the Party Hat Paradox. The Mathematical Gazette, 103(556):159–162, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10314-the-party-hat-paradox/16F6365039999563134C7E96065B6242.

Benson:2007:MMO

[Ben07] D. J. (David J.) Benson. Music: a mathematical offering. Cambridge University Press, Cambridge, UK, 2007. ISBN 0-521-85387-7 (hardcover), 0-521-61999-8 (paperback). xiii + 411 pp. LCCN ML3805 .B35 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0708/2007295519-d.
html;   http://www.loc.gov/catdir/enhancements/fy0708/
2007295519-t.html;   http://www.loc.gov/catdir/enhancements/
fy0733/2007295519-b.html.

Benko:2012:GRS

[Ben12a] David Benko. Get rich slowly, almost surely. The Mathematical Gazette, 96(536):226–235, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Benson:2012:BPM

[Ben12b] Donald C. Benson. The ballet of the planets: on the mathematical elegance of planetary motion. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-989100-1 (hardcover). xiv + 178 pp. LCCN QB361 .B46 2012.

Berger:2010:GRJ

[Ber10] Marcel Berger. Geometry revealed: a Jacob’s ladder to modern higher geometry. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-540-70996-7 (hardcover), 3-540-70997-5 (e-book). xvi + 831 + 5 pp. LCCN QA641 .B47 2010; QA641 .B48 2010. URL http://d-nb.info/

98389647X/04;   http://site.ebrary.com/id/10404636.

Berendonk:2015:PLK

[Ber15] Stephan Berendonk. Proving the loneliness of the Khajuraho square. The Mathematical Gazette, 99(544):33–39, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Berendonk:2017:DGS

[Ber17] Stephan Berendonk. Double generating spirographic curves. The Mathematical Gazette, 101(550):27–37, March 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
184E505742356F663149B290CA12A689.

Berendonk:2018:SFC

[Ber18] Stephan Berendonk. 102.20 the sum of the first n cubes is a square — two proofs without words. The Mathematical Gazette, 102(554):308, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10220-the-sum-of-the-first-n-cubes-is-a-square-
two-proofs-without-words/52012A9A5C32E5C52B63EFFD757484FE.

Bergmann:2012:TTJ

[BEU12] Birgit Bergmann, Moritz Epple, and Ruti Ungar, editors. Transcending tradition: Jewish mathematicians in German-speaking academic culture. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-22463-6 (paperback), 3-642-22464-4 (e-book). 289 pp. LCCN QA28 .J8313 2012;

QA28 .T73 2012. URL http://d-nb.info/1012892530/04.

Bevan:2013:BRB

[Bev13] David Bevan. Book review: Taking Sudoku seriously: The math behind the world’s most popular pencil puzzle, by Jason Rosenhouse and Laura Taalman. pp. 214. £13.99.(hbk) 2011. ISBN: 978-0-19-975656-8 (Oxford University Press). The Mathematical Gazette, 97 (540):574–575, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bevan:2014:BRB

[Bev14] David Bevan. Book review: Applications of combinatorial matrix theory to Laplacian matrices of graphs, by Jason J. Molitierno. Pp. 405. £57.99 (Hardback). 2012. ISBN: 978-1-4398-6337-4 (CRC Press). The Mathematical Gazette, 98(541):180–181, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bevan:2017:BRI

[Bev17] David Bevan. Book review: Introduction to random graphs, by Frieze, Alan and Karonski, Michal, pp. 464, £49.99 (hard), ISBN 978-1-107-11850-8, Cambridge University Press (2016). The Mathematical Gazette, 101(551):376–377, July 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
BD5B55148B1924DA0A81CEB5E95FBC2D.

Blanco:2018:CFF

[BGGGR18] Luis Blanco, María García-García, Joaquín Gutíerrez, and Andrea Rios. Central force fields and Kepler’s laws. The Mathematical Gazette, 102(554):270–279, July 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/central-force-fields-and-
keplers-laws/3698D18A0B004CFB58171265E0F36B93.

Bidgood:2010:AMS

[BHJ10] Penelope Bidgood, Neville Hunt, and F. R. (Flavia R.) Jolliffe, editors. Assessment methods in statistical education: an international perspective. John Wiley, New York, NY, USA, 2010. ISBN 0-470-74532-0 (paperback). xvi + 286 pp. LCCN QA276.18 .A785 2010.

Baoulina:2017:DD

[BKS17] Ioulia N. Baoulina, Martin Kreh, and Jörn Steuding. Deleting digits. The Mathematical Gazette, 101(550):60–68, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
25E86E6727FB2CBA28CE7ABB875B6EAA.

Banchoff:2010:DGC

[BL10] Thomas Banchoff and Stephen (Stephen T.) Lovett. Differential geometry of curves and surfaces. A. K. Peters, Ltd., Wellesley, MA, USA, 2010. ISBN 1-56881-456-9. xvi + 331 pp. LCCN

QA641 .B32 2010. URL http://www.zentralblatt-math.org/
zmath/en/search/?an=1200.53001.

Berry:2012:SSM

[BL12] J. S. (John Stephen) Berry and Sue Langham. Student support materials for Edexcel A Level maths. Core 1. Collins student support materials. Collins, London, UK, 2012. ISBN 0-00-747601-9 (paperback). 101 pp. LCCN ????

Bluskov:2011:PE

[Blu11] Iliya Bluskov. 95.46 Powers ending in 00 … 01. The Mathematical Gazette, 95(533):342–347, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bradley:2019:FHY

[BM19] Eoghan Bradley and Mark McCartney. 103.38 Four hundred years of the fractal coastline of Scotland. The Mathematical Gazette, 103(558):518–521, November 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10338-four-hundred-years-
of-the-fractal-coastline-of-scotland/9FFCC96F466036680A76E11206B40692.

Boreland:2010:HF

[Bor10a] Gareth Boreland. 94.28 The hyperexponential function. The Mathematical Gazette, 94(531):473–476, November 2010. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Borovik:2010:MUM

[Bor10b] Alexandre V. Borovik. Mathematics under the microscope: notes on cognitive aspects of mathematical practice. American Mathematical Society, Providence, RI, USA, 2010. ISBN 0-8218-4761-9. xxvi + 317 pp. LCCN BF456.N7 B67 2010.

Brams:2008:PEG

[Bra08] Steven J. Brams. The presidential election game. A. K. Peters, Ltd., Wellesley, MA, USA, 2008. ISBN 1-56881-348-1. xxii + 194

pp. LCCN JK528 .B73 2008. URL http://www.loc.gov/catdir/
toc/ecip0724/2007031961.html.

Bradley:2010:II

[Bra10] C. J. (Christopher John) Bradley. Introduction to inequalities, volume 2 of Handbooks. The United Kingdom Mathematics Trust, Leeds, UK, second edition, 2010. ISBN 1-906001-11-1. LCCN QA295 .B73 2010.

Bradley:2011:GTR

[Bra11a] C. J. Bradley. 95.16 The generalisation of the triplicate ratio circle. The Mathematical Gazette, 95(532):110–113, March 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bradley:2011:SPC

[Bra11b] Christopher J. Bradley. Significant points on circles centre the circumcentre. The Mathematical Gazette, 95(534):404–406, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Bradley:2012:CTT

[Bra12] Christopher Bradley. Conics twinned with two special circles. The Mathematical Gazette, 96(536):236–242, July 2012. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Brading:2017:FRC

[Bra17] Bill Brading. 101.30 a few remarks concerning a class of infinite sums. The Mathematical Gazette, 101(552):480–485, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
82EDE3677B0C0947D22E76835DC82891.

Breach:2011:FME

[Bre11] Mark Breach. Fundamental maths: [for engineering and science]. Palgrave Macmillan, Basingstoke, UK; New York, NY, USA, 2011. ISBN 0-230-25208-7 (paperback). xi + 380 pp. LCCN QA39.3 .B74 2011.

Breiteig:2015:QTN

[Bre15] Trygve Breiteig. Quotients of triangular numbers. The Mathematical Gazette, 99(545):243–255, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Brown:2011:RRT

[Bro11] Peter G. Brown. Rays in a right triangle. The Mathematical Gazette, 95(532):126–127, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Brothers:2012:PTH

[Bro12a] Harlan J. Brothers. 96.20 Pascal’s triangle: The hidden stor-e. The Mathematical Gazette, 96(535):145–148, March 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Brothers:2012:PP

[Bro12b] Harlan J. Brothers. Pascal’s prism. The Mathematical Gazette, 96(536):213–220, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Borwein:2008:CMD

[BRR08] Jonathan M. Borwein, E. M. (Eugenio M.) Rocha, and José-Francisco Rodrigues, editors. Communicating mathematics in the digital era. A. K. Peters, Ltd., Wellesley, MA, USA, 2008. ISBN 1-56881-410-0. xii + 325 pp. LCCN QA76.95 .C59 2008. URL

http://www.loc.gov/catdir/toc/fy0903/2008022183.html.

Barratt:2012:IPC

[BS12a] Carl Barratt and Ramesh Sharma. 96.16 An inductive proof of the condition for the AM–GM equality. The Mathematical Gazette, 96 (535):131–133, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Berry:2012:MMP

[BS12b] Stuart Berry and John Stubbs. The mathematical modelling of population change. The Mathematical Gazette, 96(536):243–250, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Beardon:2015:HPT

[BS15] Alan F. Beardon and Paul Stephenson. The Heron parameters of a triangle. The Mathematical Gazette, 99(545):205–212, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Block:2009:BSW

[BT09] Seymour S. Block and Santiago Alves Tavares. Before Sudoku: the world of magic squares. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-536790-1 (paperback). x + 239 pp. LCCN QA165 .B56 2009. URL http://www.loc.

gov/catdir/toc/ecip0817/2008018537.html;   http://www.
zentralblatt-math.org/zmath/en/search/?an=1163.00005.

Bunder:2017:CFE

[BT17] Martin Bunder and Joseph Tonien. Closed form expressions for two harmonic continued fractions. The Mathematical Gazette, 101(552):439–448, November 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
2B8C596F537765955BF3161162B5B182.

Burn:2010:COA

[Bur10a] Bob Burn. Chris Ormell’s article. The Mathematical Gazette, 94 (529):112–113, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Burn:2010:SPI

[Bur10b] Bob Burn. Sums of powers of integers — how Fermat may have found them. The Mathematical Gazette, 94(529):18–26, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Burn:2010:NTL

[Bur10c] R. P. Burn. 94.15 Napier’s table of logarithms. The Mathematical Gazette, 94(530):299–303, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Burn:2017:NDP

[Bur17] R. P. Burn. Non-Desarguesian planes and weak associativity. The Mathematical Gazette, 101(552):458–464, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
959E6ECAC23256835D0EB354473EDD6C.

Belcastro:2008:MMN

[BY08] Sarah-Marie Belcastro and Carolyn Yackel, editors. Making mathematics with needlework: ten papers and ten projects. A. K. Peters, Ltd., Wellesley, MA, USA, 2008. ISBN 1-56881-331-7. xii + 184 pp. LCCN TT751 .M27 2008. URL http://www.loc.gov/

catdir/toc/ecip0719/2007022382.html.

Belcastro:2011:CCF

[BY11] Sarah-Marie Belcastro and Carolyn Yackel, editors. Crafting by concepts: fiber arts and mathematics. A. K. Peters, Ltd., Wellesley, MA, USA, 2011. ISBN 1-56881-435-6 (hardcover). xii + 248 pp. LCCN TT653 .C73 2011.

Caesar:2015:ITP

[Cae15] Samantha Caesar. 99.01 Intercepts and turning points of polynomial functions. The Mathematical Gazette, 99(544):121–125, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Caglayan:2018:PWS

[Cag18] Günhan Caglayan. 102.22 proof without words: sums of sums of factorials is one less than a factorial: inductive step. The Mathematical Gazette, 102(554):309–310, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10222-proof-without-words-sums-of-sums-of-factorials-
is-one-less-than-a-factorial-inductive-step/2630F64B40C7B75F8B0C4E08D431E5FA.

Caglayan:2019:PWI

[Cag19] Günhan Caglayan. 103.02 proofs without words: identities in triangular numbers. The Mathematical Gazette, 103 (556):131–132, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10302-proofs-without-words-identities-in-triangular-
numbers/6A6F736564E9FEF58766A8144C0CFEA0.

Cameron:2016:MCS

[Cam16] John B. Cameron. Mersenne’s challenge and square roots. The Mathematical Gazette, 100(547):150–152, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Cameron:2018:BRE

[Cam18] Peter Cameron. Book review: Erdős–Ko–Rado theorems: algebraic approaches by Christopher Godsil and Karen Meagher, pp. 335 (hard), £49.99, ISBN 978-1-10712-844-6, Cambridge University Press (2015). The Mathematical Gazette, 102 (553):189–190, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/erdoskorado-theorems-algebraic-approaches-by-christopher-
godsil-and-karen-meagher-pp-335-hard-4999-isbn-9781107128446-
cambridge-university-press-2015/6B1F8D42023FA4AB9F789739602EE5BE.

Conway:2008:ST

[CBGS08] John Horton Conway, Heidi Burgiel, and Chaim Goodman-Strauss. The symmetries of things. A. K. Peters, Ltd., Wellesley, MA, USA, 2008. ISBN 1-56881-220-5. xviii + 426 pp. LCCN

QA174.7.S96 C66 2008. URL http://www.loc.gov/catdir/toc/
ecip084/2007046446.html.

Chen:2010:YHC

[CC10] Hang Chen and Curtis Cooper. 94.27 To your hearts’ content. The Mathematical Gazette, 94(531):467–473, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Constantinides:2018:SP

[CC18] Anthony F. Constantinides and George A. Constantinides. Spindown polyhedra. The Mathematical Gazette, 102(555):447–453, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/spindown-polyhedra/
72A9EC61607A31A6A3727AD33C19C0B0.

Cerone:2011:MIP

[CD11] Pietro Cerone and Sever Silvestru Dragomir. Mathematical inequalities: [a perspective]. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4398-4896-3 (hardcover). x + 391 pp. LCCN QA295 .C37 2011. URL http:/

/www.crcnetbase.com/isbn/9781439848975.

Crilly:2018:EST

[CDF18] Tony Crilly, Stan Dolan, and Colin R. Fletcher. Emmental squares are tasty. The Mathematical Gazette, 102(553):13–22, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/emmental-squares-are-tasty/
15ECED39B37D9167F425C9A983D72DA1.

Celli:2017:SD

[Cel17] Martin Celli. Studying A/(x + B) + C/(x + D) without derivatives. The Mathematical Gazette, 101(551):319–321, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
05C2886C5D3668732D6C365536F29AD9.

Cereceda:2016:ARF

[Cer16] José Luis Cereceda. An alternative recursive formula for the sums of powers of integers. The Mathematical Gazette, 100 (548):233–238, July 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/A08483E18360303EB94BD25DE1D8F287.

Crilly:2010:C

[CF10a] Tony Crilly and Colin R. Fletcher. Correspondence. The Mathematical Gazette, 94(531):535, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Crilly:2010:HCY

[CF10b] Tony Crilly and Colin R. Fletcher. How to choose your relations. The Mathematical Gazette, 94(530):193–202, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Crilly:2014:TMT

[CF14] Tony Crilly and Colin R. Fletcher. Triangles meeting triangles. The Mathematical Gazette, 98(543):432–451, November 2014. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Crilly:2015:HTP

[CF15] Tony Crilly and Colin R. Fletcher. The ‘hitchhiker triangle’ and the problem of perimeter = area. The Mathematical Gazette, 99 (546):402–415, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Chaitin:2005:MMQ

[Cha05] Gregory J. Chaitin. Meta maths: the quest for omega. Atlantic Books, London, UK, 2005. ISBN 1-84354-525-X. 220 pp. LCCN QA276.

Chaitin:2007:MMQ

[Cha07] Gregory J. Chaitin. Meta maths: the quest for omega. Atlantic Books, London, UK, paperback (with revisions and corrections) edition, 2007. ISBN 1-84354-525-X (paperback). xiv + 220 + 4 pp. LCCN ????

Chaves:2011:TPP

[Cha11] M. Chaves. 95.23 Twin primes and a primality test by indivisibility. The Mathematical Gazette, 95(533):266–269, July 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Chandrashekhar:2015:PWC

[Cha15] Shikha Chandrashekhar. 99.26 Proof without words: csc 2x = cot x − cot 2x. The Mathematical Gazette, 99(545):357, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Chen:2016:EPI

[Che16a] Hsuan-Chi Chen. 100.03 evaluating the probability integral by annular partition. The Mathematical Gazette, 100(547):111–113, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Chen:2016:NSP

[Che16b] Hsuan-Chi Chen. 100.18 Newton’s semicircle and the probability integral. The Mathematical Gazette, 100(548):317–321, July 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
90D096FFFC17736661988FFF63B010B9.

Chen:2018:EPI

[Che18] Hsuan-Chi Chen. 102.07 evaluating the probability integral by approximating exp(−x^{2}). The Mathematical Gazette, 102(553):111–113, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10207-evaluating-the-probability-integral-by-approximating-
expx2/840F9C7C9FD8C108ECA3CE6A27CAFAD5.

Chen:2019:LSL

[Che19] Yongcheng Chen. 103.39 A lemma to solve Langley’s problem. The Mathematical Gazette, 103(558):521–524, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10339-a-lemma-to-solve-langleys-
problem/4E33119838493DD4974CD9C6C9732218.

Chick:2012:C

[Chi12] John Chick. Correspondence. The Mathematical Gazette, 96(535): 174, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Chick:2017:AAA

[Chi17] J. M. Chick. 101.39 another angle on angle? The Mathematical Gazette, 101(552):502–508, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10139-another-angle-on-angle/5F713E10C9223AE7F168C7FA96E118B9.

Chu:2018:FPS

[Chu18] Adrian Chu. 102.15 a Faulhaber polynomial by stacking cubes. The Mathematical Gazette, 102(553):138–140, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10215-a-faulhaber-polynomial-
by-stacking-cubes/2D880D97DE4FCD5B225265A6CB7F4EB5.

Cinlar:2011:PS

[Çin11] Erhan Çinlar. Probability and stochastics, volume 261 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 0-387-87858-0, 0-387-87859-9 (e). xiii + 557 pp. LCCN QA274 .C563 2011.

Coelho:2019:HMR

[CK19] Terence Coelho and Bahman Kalantari. How many real attractive fixed points can a polynomial have? The Mathematical Gazette, 103(556):65–76, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/how-many-real-attractive-fixed-points-can-a-polynomial-
have/E8E45AD41919E6EB0EDF1AF46EAC3F58.

Clarke:2018:HBM

[Cla18] Robin T. Clarke. 102.01 Hoyle, Bondi and me: a memoir. The Mathematical Gazette, 102(553):89–93, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10201-hoyle-bondi-and-me-
a-memoir/2F1BD349FF34FC605BF09B8D48C8E71A.

Chan:2013:MC

[CLM13] K. C. Chan, C. T. Lenard, and T. M. Mills. 97.49 On Markov chains. The Mathematical Gazette, 97(540):515–520, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Cuthbertson:2018:VLC

[CMS18] Alastair Cuthbertson, Adam McBride, and Doug Stanford. A vintage Listener crossword. The Mathematical Gazette, 102(553):31–37, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/vintage-listener-crossword/0D7B65D8E7403CA1FD1078EF10D33933.

Cipra:1993:WHM

[CMZ93] Barry Cipra, Dana Mackenzie, and Paul Zorn, editors. What’s happening in the mathematical sciences. American Mathematical Society, Providence, RI, USA, 1993. ISBN 0-8218-8999-0 (vol. 1), 0-8218-8998-2 (vol. 2), 0-8218-0355-7 (vol. 3), 0-8218-0766-8 (vol. 4), 0-8218-2904-1 (vol. 5), 0-8218-3585-8 (vol. 6), 0-8218-4478-4 (vol. 7), 0-8218-4999-9 (vol. 8), 0-8218-8739-4 (vol. 9). ???? pp. LCCN QA3 .C57 1993.

Collins:2011:LCG

[CO11] Lee Collins and Thomas J. Osler. Law of cosines generalised for any polygon and any polyhedron. The Mathematical Gazette, 95 (533):240–243, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Colman:2011:MRE

[Col11a] W. J. A. Colman. 95.05 Measuring the radius of the Earth. The Mathematical Gazette, 95(532):72–76, March 2011. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Colman:2011:NPN

[Col11b] W. J. A. Colman. 95.52 A note on prime numbers. The Mathematical Gazette, 95(534):497–501, November 2011. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Collis:2013:PMF

[Col13a] J. R. Collis. 97.17 On products of multiplicative functions. The Mathematical Gazette, 97(539):263–264, July 2013. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Colman:2013:RWF

[Col13b] W. J. A. Colman. Reflections on the Woolwich ferry. The Mathematical Gazette, 97(539):198–204, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Colman:2015:SQR

[Col15] W. J. A. Colman. 99.13 Solving the quintic in radicals. The Mathematical Gazette, 99(545):319–323, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Colman:2019:BNN

[Col19] W. J. A. Colman. Borel’s normal number theorem. The Mathematical Gazette, 103(556):77–86, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/borels-normal-number-theorem/C87823022D469A3BF89A825950A219F5.

Cooker:2011:FFS

[Coo11a] M. J. Cooker. Fast formulas for slowly convergent alternating series. The Mathematical Gazette, 95(533):218–226, July 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Cooker:2011:C

[Coo11b] Mark J. Cooker. Correspondence. The Mathematical Gazette, 95 (534):560–561, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Cooper:2019:CRP

[Coo19] Curtis Cooper. The Chinese Ring puzzle, the Crazy Elephant Dance puzzle, the b-Spinout puzzle, and Gray codes. The Mathematical Gazette, 103(558):431–441, November 2019.

CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/chinese-ring-puzzle-the-crazy-
elephant-dance-puzzle-the-bspinout-puzzle-and-gray-codes/
37D48DF58EBEAD490A2D1E7F4E824C5B.

Corti:2007:FFF

[Cor07] Alessio Corti, editor. Flips for 3-folds and 4-folds, volume 35 of Oxford lecture series in mathematics and its applications. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-857061-9 (hardcover). vii + 189 pp. LCCN

QA564 .F55 2007. URL http://www.loc.gov/catdir/toc/
fy0801/2007280056.html.

Coutinho:2010:BRB

[Cou10] S. C. Coutinho. Book review: From quantum cohomology to integrable systems, by Martin A. Guest. Oxford Graduate Texts in Mathematics. Pp. 328. $45.00 (hardback). 2008. ISBN 13: 978-0-19-856599-4 (Oxford University Press). The Mathematical Gazette, 94(530):376–378, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Coutinho:2011:BRB

[Cou11] S. C. Coutinho. Book review: Mathematics in India, by Kim Plofker. Pp. 384. $ 39.50. 2009. ISBN: 978-0-691-12067-6 (Princeton University Press). The Mathematical Gazette, 95(532):152– 154, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Coutinho:2012:BRB

[Cou12] S. C. Coutinho. Book review: Geometric mechanics and symmetry: from finite to infinite dimensions, by D. D. Holm, Tanya Schmah and Cristina Stoica. Pp. 384. U$ 29.95 (Pbk). 2009. ISBN: 978-0-19-921291-0. (Oxford Texts in Applied and Engineering Mathematics, Oxford University Press). The Mathematical Gazette, 96(536):374–376, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Coutinho:2013:BRB

[Cou13] S. C. Coutinho. Book review: Dynamics of biological systems, by Michael Small. Pp. 262. (2012). £49.99. ISBN 978-1-4398-5336-8 (Hardback), (Mathematical and Computational Biology Series, Chapman & Hall/CRC). The Mathematical Gazette, 97(540):572– 574, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Coutinho:2014:BRB

[Cou14] S. C. Coutinho. Book review: Collected papers of C. S. Seshadri volume 1: Vector bundles and invariant theory, volume 2: Schubert geometry and representation theory, edited by Vikraman Balaji, V. Lakshmibai, M. Pavaman Murthy and Madhav V. Nori. Pp. 1008, 633 Price $195 2012. ISBN: 978-93-80250-17-5; (2 volume set), (Hindustan Book Agency, New Delhi). The Mathematical Gazette, 98(542):376–377, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Coutinho:2015:BRC

[Cou15] S. C. Coutinho. Book review: Condition: The geometry of numerical algorithms by Peter Bürgisser & Felipe Cucker, pp 554, £79.00 (hard), ISBN 978-3-642-38895-8, Springer Verlag (2013). The Mathematical Gazette, 99(546):572–574, November 2015. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Coutinho:2017:BRC

[Cou17a] S. C. Coutinho. Book review: Condition: The geometry of numerical algorithms, by Bürgisser, Peter & Cucker, Felipe, pp 554, £79.00 (hard), ISBN 978-3-642-38895-8, Springer Verlag (2013). The Mathematical Gazette, 101(550):179–181, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
1BB26DE1EA354FBEEC25BF8D14DC2BBF.

Coutinho:2017:BRE

[Cou17b] S. C. Coutinho. Book review: Elements of numerical analysis, by Gupta, Radhey S. (2nd edn.), pp. 759, AUD$86.32 exc GST (paper), ISBN 978-1-107-50049-5, Cambridge University Press (2015). The Mathematical Gazette, 101(550):189, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
E0515CE35E557B9E9793B658B2FB074A.

Coutinho:2018:BRM

[Cou18] S. C. Coutinho. Book review: The mathematics of secrets by Joshua Holden, pp. 374, £22.95 (hard), ISBN 978-0-691-14175-6, Princeton University Press (2017). The Mathematical Gazette, 102(555):559, November 2018. CODEN MAGAAS.

ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/mathematics-of-secrets-by-joshua-holden-pp-374-
2295-hard-isbn-9780691141756-princeton-university-press-
2017/DDFF61DAADB2FFAFFD6B34E5B0F7856E.

Coutinho:2019:ODE

[Cou19] S. Collier Coutinho. Book review: Ordinary differential equations, principles and applications by A. K. Nandakumaran, P. S. Datti and Raju K. George, £59.99, ISBN 978-1-10841-641-2, Cambridge University Press (2017). The Mathematical Gazette, 103(557):370–371, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/ordinary-differential-equations-principles-and-
applications-by-a-k-nandakumaran-p-s-datti-and-raju-k-
george-5999-isbn-9781108416412-cambridge-university-press-
2017/72525C6C0B5FB9ECB30DA93B81EBF798.

Crilly:2010:RCD

[CP10] Tony Crilly and Alan Pryor. Rugby conversions: a 3-dimensional model. The Mathematical Gazette, 94(530):233–238, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Chern:2019:PGP

[CQ19] Shane Chern and Shiqiu Qiu. 103.24 Partitions, geometric progressions and a Putnam problem. The Mathematical Gazette, 103(557):337–343, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10324-partitions-geometric-progressions-and-a-
putnam-problem/BCFA0EF0EA3D438389ED3C9F1B626BBA.

Conway:2013:F

[CR13] John Conway and Alex Ryba. 97.39 Fibonometry. The Mathematical Gazette, 97(540):494–495, November 2013. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Conway:2014:SLA

[CR14] John Conway and Alex Ryba. The Steiner–Lehmus angle-bisector theorem. The Mathematical Gazette, 98(542):193–203, July 2014.

CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Conway:2016:RST

[CR16] John Conway and Alex Ryba. Remembering spherical trigonometry. The Mathematical Gazette, 100(547):1–8, March 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Craik:2007:MHM

[Cra07] Alex D. D. Craik. Mr Hopkins’ men: Cambridge reform and British mathematics in the 19th Century. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. ISBN 1-84628-790-1 (hardcover), 1-84628-791-X (e-book). xiv + 405 + 24 pp. LCCN QA14.G7 C73 2007. URL http:

//www.loc.gov/catdir/enhancements/fy0825/2006940325-b.
html;   http://www.loc.gov/catdir/enhancements/fy0825/
2006940325-d.html;   http://www.loc.gov/catdir/enhancements/
fy0825/2006940325-t.html.

Crawford:2009:KMM

[Cra09] David Crawford. It’s a kind of magic: mathematical magic tricks explained. The Mathematical Association, Leicester, UK, 2009. ISBN 0-906588-67-7 (paperback). vi + 258 pp. LCCN ????

Crato:2010:FIE

[Cra10] Nuno Paulo Crato. Figuring it out: entertaining encounters with everyday math. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-04832-3, 3-642-04833-1 (e-book). ix + 1 + 227 pp. LCCN QA93 .C7313 2010.

Crilly:2007:MIY

[Cri07] A. J. (Tony) Crilly. 50 mathematical ideas you really need to know. Quercus, London, UK, thirteenth edition, 2007. ISBN 1-84724-008-9 (UK), 1-84724-147-6 (USA). 208 pp. LCCN QA39.3 .C75 2007.

Crilly:2010:HAT

[Cri10] Tony Crilly. 94.23 Handshaking across a table. The Mathematical Gazette, 94(530):317–321, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Crilly:2011:M

[Cri11] A. J. Crilly. Mathematics. The big questions. Quercus, London, UK, 2011. ISBN 1-84916-240-9 (hardcover). 208 pp. LCCN ????

Crilly:2012:SSF

[Cri12a] Tony Crilly. 96.32 The ‘sums of squares’ formula — a reprise. The Mathematical Gazette, 96(536):269–271, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Crilly:2012:HBS

[Cri12b] Tony Crilly. How best should we vote? The Mathematical Gazette, 96(537):459–465, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Crilly:2013:BRB

[Cri13] Tony Crilly. Book review: Numbers: a very short introduction, by Peter Higgins. Pp. 132 £7.99. 2011 ISBN: 978-0-19-958405-5 (Oxford University Press). The Mathematical Gazette, 97(539): 367–368, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Crilly:2015:BRM

[Cri15a] Tony Crilly. Book review: A mystic dream of 4 by Iggy McGovern, pp 92, £11.50, ISBN 978-0-9926294-0-3, Quaternia Press (2013). The Mathematical Gazette, 99(546):558–561, November 2015. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Crilly:2015:BRJ

[Cri15b] Tony Crilly. Book review: John Napier: life, logarithms, and legacy by Julian Havil, pp 279, £33.48, ISBN 978-0-691-15570-8, Princeton University Press (2014). The Mathematical Gazette, 99 (546):562–564, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Crilly:2017:BRS

[Cri17a] Tony Crilly. Book review: Sourcebook in the mathematics of medieval Europe and North Africa, edited by Katz Victor J., Folkerts Menso, Hughes Barnabas, Wagner Roi, Berggren J. Lennart, pp. 574, £70.85 (hard), ISBN 978-0-691-15685-9, Princeton University Press (2016). The Mathematical Gazette, 101(551):359–361, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
D14E8D63B5FF4A4692FC307F2B69CAD8.

Crilly:2017:HPT

[Cri17b] Tony Crilly. How to play the triangle game. The Mathematical Gazette, 101(550):1–10, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.

cambridge.org/core/product/CE2E29275D424207A95AEFD4CBFFEF57.

Cross:2010:SP

[Cro10] Tim Cross. Student problems. The Mathematical Gazette, 94 (529):165–167, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ceccherini-Silberstein:2010:CAG

[CSC10] Tullio Ceccherini-Silberstein and M. (Michel) Coornaert. Cellular Automata and Groups. Springer monographs in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-14033-5. xix + 439 pp. LCCN QA267.5.C45 C38 2010. URL http:/

/d-nb.info/1002961769/04;   http://www.loc.gov/catdir/
enhancements/fy1316/2010934641-d.html;   http://www.loc.
gov/catdir/enhancements/fy1316/2010934641-t.html.

Christian:2014:IEW

[CT14] Helen Christian and Brian Trustrum. If everyone wanted girls more boys might be born. The Mathematical Gazette, 98(543): 396–401, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Cullinane:2011:MAD

[Cul11] Michael J. Cullinane. Metric axioms and distance. The Mathematical Gazette, 95(534):414–419, November 2011. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Cuthbertson:2014:CCP

[Cut14] Alastair Cuthbertson. Cryptic crossnumber puzzles: a setter’s perspective. The Mathematical Gazette, 98(542):291–303, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Cvetkovski:2012:ITT

[Cve12] Zdravko Cvetkovski. Inequalities: theorems, techniques and selected problems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-23791-6 (paperback). x + 444 pp. LCCN QA295 .C84 2012.

Dacar:2016:YAP

[Dac16] France Dacar. 100.07 yet another proof of the Steiner–Lehmus theorem. The Mathematical Gazette, 100(547):124–126, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Davis:2010:BRB

[Dav10] Kevin Davis. Book review: Core C1 and C2: A level mathematics for Edexcel, by Peter Hind. Pp. 375. £19.50. 2008. ISBN 978-0-19-911779-6. (Oxford University Press). The Mathematical Gazette, 94(530):355, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Davis:2012:LAP

[Dav12] Ernest Davis. Linear algebra and probability for computer science applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4665-0155-3 (hardcover). xviii + 413 pp. LCCN QA76.9.M35 D38 2012.

Davis:2015:BRB

[Dav15] Kevin J. Davis. Book review: The Siddhantāsundara of Jñānarāja, an English translation by Toke Lindegaard Knudsen, pp. 351, £58.00, ISBN 013 978-1-4214-1442-3, John Hopkins University Press (2014). The Mathematical Gazette, 99(544):178–180, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Davis:2017:BRM

[Dav17] Kevin Davis. Book review: Mathematics in ancient Egypt, a contextual history by Imhausen, Annette, pp. 234, £21.05 (hard), ISBN 978-0-691-11713-3, Princeton University Press (2016). The Mathematical Gazette, 101(550):163–165, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
664442881E0E0C2B70BF5110BA0415C7.

Davis:2018:BRS

[Dav18] Kevin Davis. Book review: Significant figures: the lives and work of great mathematicians by Ian Stewart, pp. 294, £21.35 (hard), ISBN 978-0-465-09612-1, also available as e-book, Basic Books (2017). The Mathematical Gazette, 102(555):554, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/significant-figures-the-lives-
and-work-of-great-mathematicians-by-ian-stewart-pp-294-
2135-hard-isbn-9780465096121-also-available-as-ebook-basic-
books-2017/2863DA502E4C9057C6B1B96C71B492FB.

DeVilliers:2010:BRBb

[De 10a] Michael De Villiers. Book review: Leonhard Euler: a man to be reckoned with, by Andreas Heyne and Alice Heyne (illustrations by Elena Pini). pp. 45. £14.50 (hardback). 2007. ISBN: 978-3-7643-8332-9. (Basel: Birkhäuser Verlag). The Mathematical Gazette, 94 (531):555, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

DeVilliers:2010:BRBa

[De 10b] Michael De Villiers. Book review: Leonhard Euler, by Emil Fellmann. Pp. 179. £23.00 (Hardback). 2007. ISBN 13: 978-3-7643-7538-6 (Birkhäuser Verlag). The Mathematical Gazette, 94 (529):174–176, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

DeVilliers:2011:ECE

[De 11a] Michael De Villiers. 95.14 Equiangular cyclic and equilateral circumscribed polygons. The Mathematical Gazette, 95(532):102– 107, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

DeVilliers:2011:BRB

[De 11b] Michael De Villiers. Book review: A primer for mathematics competitions, Alexander Zawaira and Gavin Hitchcock. Pp.

338. £50.00 (Hardback), £22-50 (Paperback). 2009. ISBN 978-0-19-953987-1; 978-0-19-953988-8 (Paperback). (Oxford University Press). The Mathematical Gazette, 95(532):169–170, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

DeVilliers:2012:GPS

[De 12] Michael De Villiers. Generalising a problem of Sylvester. The Mathematical Gazette, 96(535):78–81, March 2012. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

DeVilliers:2013:GVT

[De 13] Michael De Villiers. 3d generalisations of Viviani’s theorem. The Mathematical Gazette, 97(540):441–445, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

DeVilliers:2014:VMT

[De 14a] Michael De Villiers. 98.16 A variation of Miquel’s theorem and its generalisation. The Mathematical Gazette, 98(542):334–339, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

DeVilliers:2014:BRBb

[De 14b] Michael De Villiers. Book review: Introduction to differential equations using Sage, by David Joyner and Marshall Hampton. Pp. 264. $60 (USD). 2012. ISBN 978-1-4214-0637-4, e-ISBN 978-1-4214-0724-1 (Baltimore, MD: The John Hopkins University Press). The Mathematical Gazette, 98(542):377, July 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

DeVilliers:2014:BRBa

[De 14c] Michael De Villiers. Book review: Iris Runge: A life at the crossroads of mathematics, science and industry, by Renate Tobies. Pp. 442. £99.00 (Hardback). 2012. ISBN: 978-3-0348-0229-1 (Birkhauser). The Mathematical Gazette, 98(541):152–155, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

DeVilliers:2017:BRM

[De 17] Michael De Villiers. Book review: Mathematical knowledge and the interplay of practices, , by Ferreirós, José, pp. 337, £31.00 (hard), ISBN 978-0-691-16751-0, Princeton University Press (2016). The Mathematical Gazette, 101(551):361–362, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
BD966DDAB3B70F0D24BB926EBA2AA5A8.

De:2013:RSL

[De13] Subhranil De. Revisiting the sliding ladder. The Mathematical Gazette, 97(539):218–223, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

De:2014:SEH

[De14] Subhranil De. On static equilibrium of a hemispheroid. The Mathematical Gazette, 98(541):73–78, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

DelNoce:2019:IIP

[Del19] Carlo Del Noce. 103.34 Indefinite integration by parts as a translation of functions. The Mathematical Gazette, 103 (558):508–512, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10334-indefinite-integration-by-parts-as-a-translation-
of-functions/74E40A080EE5D44198C30B5E2AEAD2FA.

DeMaio:2010:PWD

[DeM10] Joe DeMaio. 94.22 Proof without words: decompositions of

( \begin{matrix}n + 1 \\ 2 \end{matrix} ) and n^{2}. The Mathematical Gazette, 94(530):317, July

2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dence:2010:IIB

[Den10] Thomas P. Dence. 94.14 On an identity involving binomial coefficients via arctan(x). The Mathematical Gazette, 94(530):296–298, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dence:2013:ECI

[Den13] Thomas Dence. 97.20 Evaluating a common improper integral. The Mathematical Gazette, 97(539):270–272, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dence:2014:HMM

[Den14] Thomas Dence. Holmes + Moriarty = mathematics. The Mathematical Gazette, 98(541):51–57, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Derby:2012:MP

[Der12] Nigel Derby. 96.21 The MaxProduct partition. The Mathematical Gazette, 96(535):148–151, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Derby:2015:SSP

[Der15] Nigel Derby. A search for sums of powers. The Mathematical Gazette, 99(546):416–421, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Derby:2019:CSS

[Der19] Nigel Derby. The continued search for sums of powers. The Mathematical Gazette, 103(556):94–100, March 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/continued-search-for-sums-
of-powers/75BB1F5FCAF379692CCB80A66C256854.

Desbrow:2012:TIP

[Des12] Darrell Desbrow. 96.15 Three infinite products. The Mathematical Gazette, 96(535):130–131, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Devlin:2008:UGP

[Dev08] Keith J. Devlin. The unfinished game: Pascal, Fermat, and the seventeenth-century letter that made the world modern: [a tale of how mathematics is really done]. Basic ideas. Basic Books, New York, NY, USA, 2008. ISBN 0-465-00910-7, 0-7867-2632-6 (e-book). x + 191 + 1 pp. LCCN QA273 .D455 2008. URL http:

//www.loc.gov/catdir/enhancements/fy0913/2008012222-b.
html;   http://www.loc.gov/catdir/enhancements/fy0913/
2008012222-d.html;   http://www.loc.gov/catdir/toc/ecip0813/
2008012222.html.

DiDomenico:2013:ISB

[Di 13] A. S. Di Domenico. Integer sequences that behave as Fibonacci– Lucas pairs. The Mathematical Gazette, 97(538):1–7, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Diedrichs:2019:ALE

[Die19] Danilo R. Diedrichs. Archimedean, logarithmic and Euler spirals — intriguing and ubiquitous patterns in nature. The Mathematical Gazette, 103(556):52–64, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/archimedean-logarithmic-and-euler-spirals-intriguing-
and-ubiquitous-patterns-in-nature/25B233B64852A0DE93DFF03AF28E4B6C.

deKoninck:2009:FN

[dK09] J. M. de Koninck. Those fascinating numbers. American Mathematical Society, Providence, RI, USA, 2009. ISBN 0-8218-4807-0. xviii + 426 pp. LCCN QA241 .K685913 2009. Translated by Jean-Marie De Koninck.

Dalvit:2012:HSC

[DL12] Ester Dalvit and Domenico Luminati. How to see a cube moving into its mirror image. The Mathematical Gazette, 96(537):471– 479, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dickson:2014:ACR

[DL14] Jim Dickson and Nick Lord. Approximate constructions of 1 radian. The Mathematical Gazette, 98(542):350–353, July 2014. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dars:2008:DUM

[DLP08] Jean François Dars, Annick Lesne, and Anne Papillault, editors. Déchiffreurs. The unravelers: mathematical snapshots. A. K. Peters, Ltd., Wellesley, MA, USA, 2008. ISBN 1-56881-441-0. 182 pp.

LCCN QA99 .D4813 2008. URL http://www.loc.gov/catdir/
toc/ecip0813/2008011415.html. Translated by Vivienne Méla.

Dudley:2011:CFC

[DN11] R. M. (Richard M.) Dudley and Rimas Norvaiša. Concrete functional calculus. Springer monographs in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-6949-7. xii + 671 pp. LCCN QA300 .D828 2011. URL http:

//www.loc.gov/catdir/enhancements/fy1201/2011282333-b.
html;   http://www.loc.gov/catdir/enhancements/fy1201/
2011282333-d.html;   http://www.loc.gov/catdir/enhancements/
fy1201/2011282333-t.html.

Donahue:2012:EMI

[DO12] Steven Donahue and Thomas J. Osler. 96.09 Euler’s method of integration by parts. The Mathematical Gazette, 96(535):115–116, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dobos:2011:CRU

[Dob11] Sándor Dobos. Cross ratio in use. The Mathematical Gazette, 95 (534):444–453, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2010:SPa

[Dol10a] Stan Dolan. Student problems. The Mathematical Gazette, 94 (530):350–353, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2010:SPb

[Dol10b] Stan Dolan. Student problems. The Mathematical Gazette, 94 (531):551–552, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2011:FMD

[Dol11a] Stan Dolan. 95.24 Fermat’s method of “descente infinie”. The Mathematical Gazette, 95(533):269–271, July 2011. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2011:CKC

[Dol11b] Stan Dolan. A classification of Kaprekar constant. The Mathematical Gazette, 95(534):437–443, November 2011. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2011:SPa

[Dol11c] Stan Dolan. Student problems. The Mathematical Gazette, 95 (532):142–143, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2011:SPb

[Dol11d] Stan Dolan. Student problems. The Mathematical Gazette, 95 (533):369–371, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2011:SPc

[Dol11e] Stan Dolan. Student problems. The Mathematical Gazette, 95 (534):568–569, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2012:FAQ

[Dol12a] Stan Dolan. 96.03 Fermat’s affirmative questions. The Mathematical Gazette, 96(535):99–102, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2012:FDT

[Dol12b] Stan Dolan. Fermat and the difference of two squares. The Mathematical Gazette, 96(537):480–491, November 2012. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2012:NPO

[Dol12c] Stan Dolan. New points from old. The Mathematical Gazette, 96 (537):492–498, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2012:PEF

[Dol12d] Stan Dolan. Pell’s equation and Fermat. The Mathematical Gazette, 96(535):66–70, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2012:SPa

[Dol12e] Stan Dolan. Student problems. The Mathematical Gazette, 96 (535):183–185, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2012:SPb

[Dol12f] Stan Dolan. Student problems. The Mathematical Gazette, 96 (536):357–360, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2012:SPc

[Dol12g] Stan Dolan. Student problems. The Mathematical Gazette, 96 (537):567–569, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2013:E

[Dol13a] Stan Dolan. 97.41 The equation py^{2} = x^{4} + x. The Mathematical Gazette, 97(540):498–501, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2013:SPa

[Dol13b] Stan Dolan. Student problems. The Mathematical Gazette, 97 (538):184–186, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2013:SPb

[Dol13c] Stan Dolan. Student problems. The Mathematical Gazette, 97 (539):350–352, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2013:SPc

[Dol13d] Stan Dolan. Student problems. The Mathematical Gazette, 97 (540):551–553, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2014:NR

[Dol14a] Stan Dolan. 98.09 Nint reciprocity. The Mathematical Gazette, 98(542):317–319, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2014:SPa

[Dol14b] Stan Dolan. Student problems. The Mathematical Gazette, 98 (541):143–145, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2014:SPb

[Dol14c] Stan Dolan. Student problems. The Mathematical Gazette, 98 (542):367–369, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2014:SPc

[Dol14d] Stan Dolan. Student problems. The Mathematical Gazette, 98 (543):543–545, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2015:CNT

[Dol15a] Stan Dolan. Complex numbers and triangles. The Mathematical Gazette, 99(545):357–358, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2015:SMP

[Dol15b] Stan Dolan. The School Mathematics Project 1961–1970, a decade of innovation and its sequel by Bryan Thwaites, Cambridge University Press, available free from the author. The Mathematical Gazette, 99(545):379–380, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2015:SPa

[Dol15c] Stan Dolan. Student problems. The Mathematical Gazette, 99 (544):171–174, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2015:SPb

[Dol15d] Stan Dolan. Student problems. The Mathematical Gazette, 99 (545):376–378, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2015:SPc

[Dol15e] Stan Dolan. Student problems. The Mathematical Gazette, 99 (546):554–556, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2015:TAG

[Dol15f] Stan Dolan. Triangles around a given triangle. The Mathematical Gazette, 99(546):432–443, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2016:DEX

[Dol16a] Stan Dolan. 100.16 the Diophantine equation n(x + y + z) = xyz. The Mathematical Gazette, 100(548):314, July 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
C3271CB102B73C876B8E1E7D875C70B5.

Dolan:2016:LTE

[Dol16b] Stan Dolan. Less than equable Heronian triangles. The Mathematical Gazette, 100(549):482–489, November 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
32353B35B306159854435E2AA82B2385.

Dolan:2016:SP

[Dol16c] Stan Dolan. Student problems. The Mathematical Gazette, 100 (547):172–174, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Dolan:2017:PFX

[Dol17a] Stan Dolan. 101.26 primes of the form x^{2} + z^{2}. The Mathematical Gazette, 101(552):465, November 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
C5602DCA6C43C5A749ACABC2CE564D50.

Dolan:2017:LC

[Dol17b] Stan Dolan. Lyness cycles. The Mathematical Gazette, 101 (551):193–207, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/A847AFA4E97E1FF131E9A9F655036636.

Dolan:2017:SPa

[Dol17c] Stan Dolan. Student problems. The Mathematical Gazette, 101 (550):157–159, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/3D794FCF385265F9604E58E773DAED24.

Dolan:2017:SPb

[Dol17d] Stan Dolan. Student problems. The Mathematical Gazette, 101 (551):344–347, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/9B6ED33F0C691F6E17DA474ADA0119F0.

Dolan:2017:SPc

[Dol17e] Stan Dolan. Student problems. The Mathematical Gazette, 101(552):552–554, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/2C540D35B8D69E6F668ADF60B761AEAA.

Dolan:2017:TCE

[Dol17f] Stan Dolan. Thoughts on a conjecture of Erdős. The Mathematical Gazette, 101(552):449–457, November 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
E365E6327B643043FD8E4015511911A1.

Dolan:2018:MSC

[Dol18a] Stan Dolan. 102.02 minimal sets for congruence classes. The Mathematical Gazette, 102(553):93–97, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10202-minimal-sets-for-congruence-
classes/F2056D75515A1B064FD0E9C552FCA8F1.

Dolan:2018:WQR

[Dol18b] Stan Dolan. 102.03 when quadratic residues are squares. The Mathematical Gazette, 102(553):97–100, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10203-when-quadratic-residues-
are-squares/B674108B2B177DEAC198A3C0D3154968.

Dolan:2018:SPa

[Dol18c] Stan Dolan. Student problems. The Mathematical Gazette, 102(553):182–184, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/C5040637F24CC1CE9DE27109887A21C9.

Dolan:2018:SPb

[Dol18d] Stan Dolan. Student problems. The Mathematical Gazette, 102(554):361–363, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/648E3363C6817C802E1D95899C557505.

Dolan:2018:SPc

[Dol18e] Stan Dolan. Student problems. The Mathematical Gazette, 102(555):551–553, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/EA2A4C87F077728FB2BF306C7727F80A.

Dolan:2019:HTS

[Dol19a] Stan Dolan. 103.29 Heronian triangles and squares in arithmetic progression. The Mathematical Gazette, 103(558):490–493, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10329-heronian-
triangles-and-squares-in-arithmetic-progression/690BD0814A6428598F2437DF4E5B6060.

Dolan:2019:DIS

[Dol19b] Stan Dolan. A diagram for inequalities of symmetric functions. The Mathematical Gazette, 103(558):461–470, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/diagram-for-inequalities-
of-symmetric-functions/3EC95C473F274B5808D69E6DE63FD734.

Dolan:2019:SPa

[Dol19c] Stan Dolan. Student problems. The Mathematical Gazette, 103(556):179–181, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/BD34D4E08E38FDB9958E21F112B337B3.

Dolan:2019:SPb

[Dol19d] Stan Dolan. Student problems. The Mathematical Gazette, 103(557):365–367, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/480C65633069418A451186EBE8CE305A.

Dolan:2019:SP

[Dol19e] Stan Dolan. Student problems. The Mathematical Gazette, 103(558):567–569, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/student-problems/97EADD37A1877C667C25704433C3E68C.

Doumas:2013:AAP

[Dou13] Aristides V. Doumas. An alternative approach to a problem by A. de Moivre. The Mathematical Gazette, 97(540):446–454, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Downes:2015:RRL

[Dow15] Rob Downes. 99.14 A recurrence relation for the limit point of nested triangles. The Mathematical Gazette, 99(545):323–325, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

duSautoy:2013:TSM

[dS13] Marcus du Sautoy. Teaching the Shakespeare of mathematics. The Mathematical Gazette, 97(540):386–397, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Davis:2008:SCC

[DSZ08] Chandler Davis, Marjorie Senechal, and Jan Zwicky, editors. The shape of content: creative writing in mathematics and science. A. K. Peters, Ltd., Wellesley, MA, USA, 2008. ISBN 1-56881-444-5. xvii + 194 pp. LCCN PN6071.M3 S53 2008. URL http://www.

loc.gov/catdir/toc/ecip0819/2008022399.html.

Duffy:2011:AEM

[Duf11] Dean G. Duffy. Advanced engineering mathematics with MAT-LAB. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, third edition, 2011. ISBN 1-4398-1624-7 (hardcover). 1079 pp. LCCN TA345 .D84 2011. URL http://www.

zentralblatt-math.org/zmath/en/search/?an=1018.00003.

deVilliers:2014:BRT

[dV14] Michael de Villiers. Book review: Teaching problem-solving in undergraduate mathematics by M.S. Badger, C.J. Sangwin and T.O. Hawkes with R.P. Burns, J. Mason and S. Pope, pp 142, £5.25 plus postage. The Mathematical Gazette, 98(543):547– 549, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). Available as print on demand at:

http://www.lulu.com/shop/product-20405493.htmlor as free
PDF download at: http://mellbreak.lboro.ac.uk/problemsolving/sites/default/files/guide/Guide.pdf(2012).

deVilliers:2018:SDR

[dVS18] Michael de Villiers and Heinz Schumann. 102.30 a surprising 3-D result involving a hexagon. The Mathematical Gazette, 102(554):328–331, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10230-a-surprising-3d-result-involving-a-hexagon/
CDCF70C5950814DF230B85A6732665AA.

Daners:2019:CGC

[DW19] Daniel Daners and Theresa Wigmore. Conics and a generalised conical pendulum. The Mathematical Gazette, 103(556): 28–40, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/conics-and-
a-generalised-conical-pendulum/DE3B6FF0F7D555D1B6EE9829C7E31E50.

Dong:2014:LOE

[DZH14] Fengming Dong, Dongsheng Zhao, and Weng Kin Ho. On the largest outscribed equilateral triangle. The Mathematical Gazette, 98(541):79–84, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Eaton:2011:BRB

[Eat11] Charles Eaton. Book review: History of mathematics: highways and byways, by Amy Dahan-Dalmedico and Jeanne Peiffer. Pp.

330. £40.00 (hbk). 2010. ISBN: 978-0-888385-562-1 (Mathematical Association of America). The Mathematical Gazette, 95(532):149– 150, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Eaton:2012:BRB

[Eat12] Charles Eaton. Book review: The symmetries of things, by John Conway, Heidi Burgiel and Chain Goodman-Strauss. Pp. 426. $75.00 (hbk). 2008. ISBN: 978-1-56881-220-5 (A. K. Peters). The Mathematical Gazette, 96(535):187–190, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Eaton:2013:BRB

[Eat13] Charles Eaton. Book review: Cellular automata and groups, by Tullio Ceccherini-Silberstein and Michel Coornaert. Pp. 439. £81.00 (hbk). 2010 ISBN: 978-3-642-14033-4 (Springer). The Mathematical Gazette, 97(539):371–372, July 2013. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Edgar:2016:PWA

[Edg16] Tom Edgar. 100.35 proof without words: the average of square pyramidal and triangular is tetrahedral. The Mathematical Gazette, 100(549):516–517, November 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
17B23BE8D521D960CBE90ADE6D74655A.

Englefield:2010:EAS

[EF10] M. J. Englefield and G. E. Farr. Eigencircles and associated surfaces. The Mathematical Gazette, 94(531):438–449, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Eggleston:2015:NPO

[Egg15] Ben Eggleston. The number of preference orderings: a recursive approach. The Mathematical Gazette, 99(544):21–32, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Etingof:2011:IRT

[EGH^{+}11] P. I. (Pavel I.) Etingof, Oleg Golberg, Sebastian Hensel, Tiankai Liu, Alex Schwendner, Dmitry Vaintrob, Elena Yudovina, and Siava Gerovitch. Introduction to representation theory, volume 59 of Student mathematical library. American Mathematical Society, Providence, RI, USA, 2011. ISBN 0-8218-5351-1 (paperback). vii + 228 pp. LCCN QA155 .I586 2011. URL http://www.ams.org/ stml/059. With historical interludes by Slava Gerovitch.

Edwards:2011:PMM

[EH11] Bianca Edwards and Jim Hartman. 95.28 Powers of magic matrices. The Mathematical Gazette, 95(533):284–292, July 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Elliott:2017:PCM

[Ell17] M. G. Elliott. Projectiles from circular motion. The Mathematical Gazette, 101(550):136–141, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.

cambridge.org/core/product/879EE9EC9FA8D759376BED6A402DAE08.

Ellard:2012:PSR

[EM12] Richard Ellard and Des MacHale. Packing squares into rectangles. The Mathematical Gazette, 96(535):1–18, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ellard:2016:PRR

[EM16] Richard Ellard and Des MacHale. Packing a rectangle with m × (m + 1) rectangles. The Mathematical Gazette, 100(547):34–47, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Eggleton:2012:P

[EO12] Roger Eggleton and Krzysztof Ostaszewski. The 6666 problem. The Mathematical Gazette, 96(535):39–48, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Euler:2010:MTP

[ES10] Russell Euler and Jawad Sadek. 94.08 A multiplicative-type property for determinants. The Mathematical Gazette, 94(529):135– 136, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Euler:2011:FAG

[ES11] Russell Euler and Jawad Sadek. 95.53 A formula for all k-gonal numbers that are centred k-gonal. The Mathematical Gazette, 95 (534):501–505, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Euler:2016:SRB

[ES16] Russell Euler and Jawad Sadek. 100.33 a surprising relationship between Pell–Lucas, Pell, square and centred square numbers. The Mathematical Gazette, 100(549):510–511, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
DFD092BA5EFDF5BBEB9C9474947E9984.

Euler:2015:RBP

[Eul15] Russell Euler. 99.27 a relationship between Pell numbers and triangular square numbers. The Mathematical Gazette, 99(546):516, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Farmer:2013:UTF

[Far13] Jim Farmer. 97.12 Using Taylor’s formula to compare compound and simple interest. The Mathematical Gazette, 97(538):150–153, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Farnsworth:2018:ASC

[Far18] David L. Farnsworth. Axes for symmetric convex curves. The Mathematical Gazette, 102(553):23–30, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/axes-for-symmetric-convex-
curves/AC0E94F4F73241234AB97EA124C910F1.

Fellmann:2007:LE

[Fel07] Emil Alfred Fellmann. Leonhard Euler. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2007. ISBN 3-7643-7538-8, 3-7643-7539-6 (e-book). xv + 179 pp. LCCN

QA29.E8 F452 2007. URL http://www.loc.gov/catdir/toc/

fy0709/2006937470.html. Translated by E. Gautschi and W. Gautschi from the 1995 German original of the same title.

Fernandez:2014:ECD

[Fer14] Oscar Edward (Oscar Edward) Fernandez. Everyday calculus: discovering the hidden math all around us. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-15755-3 (hardcover), 1-4008-5008-8 (e-book), 1-306-41429-6. x + 150 pp. LCCN

QA303.2 .F47 2014. URL http://www.jstor.org/stable/10.
2307/j.ctt5hhrp2.

Ferlini:2016:ULT

[Fer16] Vincent Ferlini. 100.06 using linear transformations to solve quadratic equations. The Mathematical Gazette, 100(547):119– 124, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Forbes:2016:HR

[FF16] Tamsin Forbes and Tony Forbes. Hanoi revisited. The Mathematical Gazette, 100(549):435–441, November 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
56675C2E78B82226C229DEBF3C7BAE4D.

Forbes:2019:SPB

[FF19] Tamsin Forbes and Tony Forbes. 103.37 A sudoku puzzle with a big rectangular hole. The Mathematical Gazette, 103(558):517, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10337-a-sudoku-puzzle-with-a-big-rectangular-hole/
5052EC8A8EEAF3D6770E9EB43E7CA619.

Frohlich:2009:JTS

[FFGH09] Jürg Fröhlich, G. (Giovanni) Felder, G. M. (Gian Michele) Graf, and K. (Klaus) Hepp, editors. A journey through statistical physics: from foundations to the quantum Hall effect: selecta of Jürg Fröhlich. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4020-8886-8 (hardcover). lxxvii + 857 pp. LCCN

QC174.8 .F76 2009. URL http://www.gbv.de/dms/bowker/toc/
9781402088865.pdf;   http://www.zentralblatt-math.org/
zmath/en/search/?an=1169.01013.

Fauvel:2013:OFE

[FFW13] John Fauvel, Raymond Flood, and Robin J. Wilson, editors. Oxford figures: eight centuries of the mathematical sciences. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2013. ISBN 0-19-968197-X (hardcover). x + 406 pp. LCCN QA14.G73 O947 2013.

Fine:2014:IAA

[FGR14] Benjamin Fine, Anthony M. Gaglione, and Gerhard Rosenberger. Introduction to Abstract Algebra: From Rings, Numbers, Groups, and Fields to Polynomials and Galois theory. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1176-8 (hardcover). xiv + 2 + 566 pp. LCCN QA162 .F56 2014. URL

http://catalogue.bnf.fr/ark:/12148/cb439057775.

Farin:2005:PLA

[FH05] Gerald E. Farin and Dianne Hansford. Practical linear algebra: a geometry toolbox. A. K. Peters, Ltd., Wellesley, MA, USA, 2005. ISBN 1-56881-234-5. xvi + 384 pp. LCCN QA184.2 F37 2005.

Fielder:2007:TAG

[Fie07] Sara Fielder. The teaching assistant’s guide to numeracy. Continuum, London, UK, 2007. ISBN 0-8264-9606-7 (paperback). vii + 117 pp. LCCN QA141 .F54 2007. URL http://www.loc.gov/

catdir/toc/ecip0718/2007020999.html.

Fink:2014:PAR

[Fin14a] A. M. Fink. 98.22 polynomials with all real roots. The Mathematical Gazette, 98(543):483, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Fink:2014:IIQ

[Fin14b] A. M. Fink. 98.30 the isoperimetric inequality for quadrilaterals. The Mathematical Gazette, 98(543):504, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

French:2010:DSP

[FJ10] Doug French and David Jordan. Dodecahedral slices and polyhedral pieces. The Mathematical Gazette, 94(529):5–17, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Fralick:2019:LDE

[FK19] Bethany Fralick and Reginald Koo. Linear differential equations with constant coefficients. The Mathematical Gazette, 103 (557):257–264, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/linear-
differential-equations-with-constant-coefficients/3D9AF109A55D380F52FD47B1D2279481.

Farnsworth:2014:DSV

[FL14] David L. Farnsworth and Michael E. Long. 98.34 the definition of the sample variance. The Mathematical Gazette, 98(543):511– 517, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Feeman:2010:RVR

[FM10] Timothy G. Feeman and Osvaldo Marrero. 94.30 Ratios of volumes related to the odd extension of a power function. The Mathematical Gazette, 94(531):479–486, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Fang:2018:CSR

[FMK18] Ming Fang, Shahrooz Moosavizadeh, and Mushtaq A. Khan. 102.43 a comparative study of relative rates of change. The Mathematical Gazette, 102(555):505–511, November 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10243-a-comparative-study-
of-relative-rates-of-change/F182018C4D097BD9A35A6736C084D508.

Figueiredo:2010:BDU

[FMS10] Vera L. X. Figueiredo, Margarida P. Mello, and Sandra A. Santos. Beyond domes, umbrellas and tents. The Mathematical Gazette, 94(529):51–61, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Flood:2014:JCM

[FMW14] Raymond Flood, Mark McCartney, and Andrew Whitaker, editors. James Clerk Maxwell: perspectives on his life and work. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-966437-4. x + 364 + 4 pp. LCCN QC16.M4

J36 2014. URL http://www.loc.gov/catdir/enhancements/
fy1410/2013942195-b.html;   http://www.loc.gov/catdir/
enhancements/fy1410/2013942195-d.html;   http://www.loc.
gov/catdir/enhancements/fy1410/2013942195-t.html.

Fabian:2013:PEI

[FN13] Andrew Fabian and Hieu D. Nguyen. Paradoxical Euler: Integrating by differentiating. The Mathematical Gazette, 97(538):61–74, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Fortune:2010:BRBa

[For10a] Mary Fortune. Book review: Geometric Origami, by Robert Geretschläger. Pp. 198, £14.00. 2008. ISBN: 978-0-9555477-1-3 (Arbelos). The Mathematical Gazette, 94(529):189–190, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Fortune:2010:BRBb

[For10b] Mary Fortune. Book review: Making mathematics with needlework, Sarah-Marie Belcastro and Carolyn Yackel (eds). Pp. 184. £18.50. 2008. ISBN 978-1-56881-331-8 (A. K. Peters). The Mathematical Gazette, 94(530):378–379, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Fortune:2011:BRB

[For11] Mary Fortune. Book review: A mathematical nature walk, by John Adam. pp. 249. £16.95 (Hardback). 2009. ISBN: 978-0-691-12895-5 (Princeton University Press). The Mathematical Gazette, 95(533): 379, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Fortune:2012:BRB

[For12] Mary Fortune. Book review: Pattern theory: tbe stochastic analysis of real-world signals, by David Mumford and Agnes Desolneux. Pp. 407. $79-00 (hbk). 2010. ISBN: 978-1-56881-579-4 (A. K. Peters). The Mathematical Gazette, 96(535):190, March 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Fortune:2013:BRB

[For13] Mary Fortune. Book review: Crafting by concepts, by Sarah-Marie Belcastro and Carolyn Yackel. Pp. 248. £25.99 (hbk) 2011 ISBN: 978-1-56881-435-3 (A. K. Peters). The Mathematical Gazette, 97 (539):382–383, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Foster:2010:C

[Fos10] Colin Foster. Correspondence. The Mathematical Gazette, 94 (530):335, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Foster:2011:CSM

[Fos11] Colin Foster. A comment on Schnell and Mendoza. The Mathematical Gazette, 95(533):357, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Foster:2012:SC

[Fos12a] Colin Foster. 96.07 Symmetrical cubics. The Mathematical Gazette, 96(535):109–112, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Foster:2012:OM

[Fos12b] Colin Foster. 96.08 Orthogonal mappings. The Mathematical Gazette, 96(535):112–115, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Foster:2012:TSC

[Fos12c] Colin Foster. 96.19 Taking a short cut through Pascal’s triangle. The Mathematical Gazette, 96(535):142–145, March 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Foster:2012:QD

[Fos12d] Colin Foster. 96.30 Quadratic doublets. The Mathematical Gazette, 96(536):264–266, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Foster:2012:SWS

[Fos12e] Colin Foster. 96.47 Squares within squares. The Mathematical Gazette, 96(536):328–331, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Foster:2014:EVR

[Fos14a] Colin Foster. 98.31 equal volumes of revolution. The Mathematical Gazette, 98(543):504–507, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Foster:2014:BRB

[Fos14b] Colin Foster. Book review: How to study for a mathematics degree, by Lara Alcock, pp. 288, £12.99, ISBN 978-0-19-966132-9, Oxford University Press (2013). The Mathematical Gazette, 98(542):377– 378, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Foster:2018:PGS

[Fos18] Tony Foster. 102.32 panning for gold in the streams of a regular pentagon. The Mathematical Gazette, 102(554):332–335, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10232-panning-for-
gold-in-the-streams-of-a-regular-pentagon/731F4AA8D81E1B797F5C9794D16E7C2B.

Fox:2014:QS

[Fox14] Michael Fox. Quadritangent spheres. The Mathematical Gazette, 98(541):24–39, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Fox:2015:BRB

[Fox15] Michael Fox. Book review: The geometry of the triangle by Gerry Leversha, pp. 573, £22, ISBN 978-1-906001-17-9, The United Kingdom Mathematics Trust (2013). The Mathematical Gazette, 99(544):185–186, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Fox:2016:SPK

[Fox16] Michael Fox. Some properties of Kiepert lines of a triangle. The Mathematical Gazette, 100(547):54–67, March 2016. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Fox:2017:BRT

[Fox17] Michael Fox. Book review: Treks into intuitive geometry: the world of polygons and polyhedra by Jin Akiyama and Kiyoko Matsunaga, pp. 425, £44.99 (hard), ISBN 978-4-431-55841-5, Springer (2015). The Mathematical Gazette, 101(552):565–568, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/treks-into-
intuitive-geometry-the-world-of-polygons-and-polyhedra-
by-jin-akiyama-and-kiyoko-matsunaga-pp-425-4499-hard-isbn-
9784431558415-springer-2015/44095C844F8702BF649CA16EF9F72437.

Fukagawa:2002:TJM

[FR02] Hidetoshi Fukagawa and John F. Rigby. Traditional Japanese mathematics problems of the 18th and 19th centuries. SCT Publishers, Singapore, 2002. ISBN 981-04-2759-X. xvi + 191 pp. LCCN QA27.J3 F848 2002.

Fukagawa:2008:SMJ

[FR08] Hidetoshi Fukagawa and Tony Rothman. Sacred mathematics: Japanese temple geometry. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-12745-X (hardcover). xxiii + 348 pp. LCCN QA27.J3 F849 2008.

Fraivert:2019:NPB

[Fra19a] David Fraivert. New points that belong to the nine-point circle. The Mathematical Gazette, 103(557):222–232, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/new-points-that-
belong-to-the-ninepoint-circle/52930388DA499D8AEB7A2AA184850BF8.

Fraivert:2019:PPQ

[Fra19b] David Fraivert. Pascal-points quadrilaterals inscribed in a cyclic quadrilateral. The Mathematical Gazette, 103(557):233–239, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/

journals/mathematical-gazette/article/pascalpoints-quadrilaterals-
inscribed-in-a-cyclic-quadrilateral/6544F227FF10BE41779512702F30F48C.

French:2010:BRBa

[Fre10a] Doug French. Book review: Assessing mathematical proficiency, Alan H. Schoenfeld (ed.). Pp. 391. £17.99. 2007. ISBN 0521 69766 8 (Cambridge University Press). The Mathematical Gazette, 94 (530):368–369, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

French:2010:BRBb

[Fre10b] Doug French. Book review: Thirty three ways to help with numeracy, by Brian Sharp. Pp. 124. £19.99. 2009. ISBN 9 78 0415 46896 1 (David Fulton). The Mathematical Gazette, 94(530):371, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Frenkel:2013:LMH

[Fre13] Edward Frenkel. Love and math: the heart of hidden reality. Basic Books, a member of the Perseus Books Group, New York, NY, USA, 2013. ISBN 0-465-05074-3 (hardcover), 0-465-06995-9 (ebook). 292 pp. LCCN QA29.F725 F74 2013.

French:2015:BRD

[Fre15] Jo French. Book review: I died for beauty (1st edn), by Marjorie Senechal, pp 300, £22.50 (hard), ISBN 978-0-19-973259-3, Oxford University Press (2013). The Mathematical Gazette, 99(546):557– 558, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Flood:2011:MVB

[FRW11] Raymond Flood, Adrian C. (Adrian Clifford) Rice, and Robin James Wilson. Mathematics in Victorian Britain. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-960139-9 (hardcover). ix + 1 + 466 pp. LCCN QA27.G7 M38 2011; QA26 .M37 2011. URL http://catdir.loc.gov/

catdir/enhancements/fy1211/2011938072-b.html;   http:/
/catdir.loc.gov/catdir/enhancements/fy1211/2011938072-
d.html;   http://catdir.loc.gov/catdir/enhancements/fy1211/
2011938072-t.html.

Fabrykowski:2010:TPI

[FS10] Jacek Fabrykowski and Thomas Smotzer. 94.29 On two properties of the inverse Hilbert matrix. The Mathematical Gazette, 94(531): 476–479, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Flood:2011:GMU

[FW11] Raymond Flood and Robin Wilson. The great mathematicians: unravelling the mysteries of the universe. Arcturus, London, UK, 2011. ISBN 1-84837-902-1 (paperback). 208 pp. LCCN ????

Fyfe:2011:PAM

[Fyf11] Mary Teresa Fyfe. Progress to advanced mathematics. Arbelos, Shipley, UK, 2011. ISBN 0-9555477-3-3 (paperback). LCCN ????

Gao:2008:IDS

[Gao08] Su Gao. Invariant descriptive set theory. Pure and applied mathematics. Chapman and Hall/CRC, Boca Raton, Fla., 2008. ISBN 1-58488-793-1. xiv + 383 pp. LCCN ????

Gao:2009:IDS

[Gao09] Su Gao. Invariant descriptive set theory, volume 293 of Monographs and Textbooks in Pure and Applied Mathematics. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2009. ISBN 1-58488-793-1 (Hardcover), 1-58488-794-X (ebook). xiv + 383 pp. LCCN QA248 .G36 2009. URL http://www.

gbv.de/dms/bowker/toc/9781584887935.pdf;   http://www.
zentralblatt-math.org/zmath/en/search/?an=1154.03025.

Gardiner:2010:WSW

[Gar10] Tony Gardiner. William Salisbury Wynne Willson 1932–2010. The Mathematical Gazette, 94(531):532–534, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Garling:2013:CMA

[Gar13] D. J. H. Garling. A Course in Mathematical Analysis. Cambridge University Press Textbooks, Cambridge, UK, 2013. ISBN 1-299-54610-2 (e-book), 1-107-34580-4 (e-book), 1-139-42451-3 (ebook), 1-107-03202-4 (hardcover), 1-107-61418-X (paperback), 1-107-03204-0 (hardback: v. 3), 1-107-66330-X (paperback: v. 3), 1-107-03203-2 (vol. 2: hardcover), 1-107-67532-4 (vol. 2: paperback), 1-107-35792-6, 1-107-30185-8, 1-107-70283-6. LCCN QA300 .G276

2013. URL http://lib.myilibrary.com?id=485860;   http:/
/lib.myilibrary.com?id=573640.

Gardiner:2018:NNF

[Gar18] Tony Gardiner. Nature-nurture: finding-feeding. The Mathematical Gazette, 102(555):485–493, November 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/naturenurture-findingfeeding/
BE76699B450972025343BB4DDC8E1ADF.

Gauthier:2011:FPI

[Gau11] N. Gauthier. 95.50 Fractional-power identities for Fibonacci and Lucas polynomials and numbers. The Mathematical Gazette, 95 (534):486–494, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Gau:2012:TTE

[Gau12a] David Gau. 96.41 Triangle through an elliptical looking-glass. The Mathematical Gazette, 96(536):302–304, July 2012. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Gauthier:2012:CFC

[Gau12b] N. Gauthier. Closed forms for convolutions of Catalan numbers. The Mathematical Gazette, 96(535):82–90, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Gau:2019:TAS

[Gau19] David Gau. 103.18 Trigonometric addition and subtraction formulas, a unified visual approach. The Mathematical Gazette, 103(557):314–315, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10318-trigonometric-addition-and-subtraction-formulas-
a-unified-visual-approach/1268973E72874E4A01A6EA4EDAF944F6.

Gauthier:2012:NFA

[GB12] N. Gauthier and Paul S. Bruckman. 96.50 A note on a family of alternating sums of products of binomial numbers. The Mathematical Gazette, 96(536):337–342, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Grozdev:2014:CGM

[GD14] Sava Grozdev and Deko Dekov. 98.33 computer-generated mathematics: points on the Kiepert hyperbola. The Mathematical Gazette, 98(543):509–511, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Gardner:2008:LPC

[GDDR08] Martin Gardner, Erik D. Demaine, Martin L. Demaine, and Tom Rodgers, editors. A lifetime of puzzles: a collection of puzzles in honor of Martin Gardner’s 90th birthday. A. K. Peters, Ltd., Wellesley, MA, USA, 2008. ISBN 1-56881-245-0. x + 349 pp.

LCCN QA95 .L73 2008. URL http://www.loc.gov/catdir/toc/
ecip081/2007041016.html.

George:2010:I

[Geo10] Glyn George. The integral of 1/x. The Mathematical Gazette, 94 (530):332–334, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

George:2015:WCL

[Geo15] Glyn George. When copycats lose. The Mathematical Gazette, 99 (545):230–236, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Geretschlager:2008:GO

[Ger08] Robert Geretschläger. Geometric origami. Arbelos, Shipley, UK, 2008. ISBN 0-9555477-1-7 (paperback). xiii + 198 pp. LCCN QA482 .G47 2008.

Gessen:2011:PRG

[Ges11] Masha Gessen. Perfect rigour: a genius and the mathematical breakthrough of the century. Icon Books, London, UK, 2011. ISBN 1-84831-238-5 (paperback), 1-84831-310-1 (e-book). xi + 1 + 242 pp. LCCN QA29.P47 G47 2011; QA29.P6727 G477 2011.

Gardner:2008:HPP

[GG08a] Martin Gardner and Martin Gardner. Hexaflexagons, probability paradoxes, and the Tower of Hanoi: Martin Gardner’s first book of mathematical puzzles and games, volume 1 of The new Martin Gardner mathematical library. Cambridge University Press, Cambridge, UK, second edition, 2008. ISBN 0-521-75615-4 (hardcover), 0-521-73525-4 (paperback). xiii + 193 pp. LCCN QA95 .G247 2008. URL http:

//www.loc.gov/catdir/enhancements/fy0834/2008012533-b.
html;   http://www.loc.gov/catdir/enhancements/fy0834/
2008012533-d.html;   http://www.loc.gov/catdir/enhancements/
fy0834/2008012533-t.html.

Gardner:2008:OES

[GG08b] Martin Gardner and Martin Gardner. Origami, Eleusis, and the Soma cube: Martin Gardner’s mathematical diversions, volume 2 of The new Martin Gardner mathematical library. Cambridge University Press, Cambridge, UK, 2008. ISBN 0-521-75610-3 (hardcover), 0-521-73524-6 (paperback). xi + 234 pp. LCCN QA95 .G2975 2008. URL http:

//www.loc.gov/catdir/enhancements/fy0834/2008012534-b.
html;   http://www.loc.gov/catdir/enhancements/fy0834/
2008012534-d.html;   http://www.loc.gov/catdir/enhancements/
fy0834/2008012534-t.html.

Glaz:2008:SAP

[GG08c] Sarah Glaz and JoAnne Simpson Growney, editors. Strange attractors: poems of love and mathematics. A. K. Peters, Ltd., Wellesley, MA, USA, 2008. ISBN 1-56881-341-4, 1-4398-6518-3 (e-book). xix

+ 255 pp. LCCN ???? URL http://marc.crcnetbase.com/
isbn/9781439865187.

Grattan-Guinness:2010:BRB

[GG10a] I. Grattan-Guinness. Book review: Mr Hopkins’ men — Cambridge reform and British mathematics in the 19th century, by A. D. D. Craik. Pp. 405. 2007. £65.00. ISBN 978-1-84628-790-9 (Springer). The Mathematical Gazette, 94(530):358–359, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Grattan-Guinness:2010:ERB

[GG10b] I. Grattan-Guinness. Extended review: The first century of the International Commission on Mathematical Instruction (1908– 2008). Reflecting and shaping the world of mathematics education. M. Menghini, F. Furinghetti, L. Giacardi and F. Arzarello (eds.). Pp. 328. EUR60. 2008. ISBN: 978-88-12-00015-9 (Rome: Istituto della Enciclopedia Italiana). The Mathematical Gazette, 94(529):168–172, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Grattan-Guinness:2012:RSR

[GG12a] I. Grattan-Guinness. 96.05 On representing the square roots of integers by straight lines. The Mathematical Gazette, 96(535):104– 106, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Grattan-Guinness:2012:PIC

[GG12b] I. Grattan-Guinness. Properties of inscribed and circumscribed rectangles. The Mathematical Gazette, 96(535):71–77, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Grattan-Guinness:2012:BRB

[GG12c] Ivor Grattan-Guinness. Book review: Mathematical lives: protagonists of the twentieth century, by Claudio Bartocci, Renato Betti, Angelo Guerraggio and Roberto Lucchetti (Editors). pp. 238. £53.99 (Hardback). 2011. ISBN: 978-3-642-13605-4 (Springer). The Mathematical Gazette, 96(536):367, July 2012. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Grattan-Guinness:2013:PCO

[GG13] I. Grattan-Guinness. On proving certain optimisation theorems in plane geometry. The Mathematical Gazette, 97(538):75–80, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ghosh:2011:GRC

[Gho11] Chinmoy Kumar Ghosh. 95.17 The golden ratio in conic sections. The Mathematical Gazette, 95(532):113–115, March 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2010:BRBa

[Gib10a] Peter Giblin. Book review: The unravelers, mathematical snapshots, J.-F. Dars, A. Lesne and A. Papillault, (eds) (trans. V. Méla) £26.50. 2008. ISBN 978-1-56881-441-4 (A. K. Peters). The Mathematical Gazette, 94(529):178–180, March 2010. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2010:BRBc

[Gib10b] Peter J. Giblin. Book review: Recountings: Conversations with MIT mathematicians, ed. Joel Segel, pp. 441. $49-00 (hbk). 2009. ISBN: 978-1-56881-449-0 (A. K. Peters). The Mathematical Gazette, 94(531):555–557, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2010:BRBb

[Gib10c] Peter J. Giblin. Book review: The map of my life, by Goro Shimura. pp. 211, £19.99 (hardback). 2008. ISBN 978-0-387-79714-4 (Springer). The Mathematical Gazette, 94(531):553–554, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2011:BRBc

[Gib11a] Peter Giblin. Book review: Cinderella (Version 2): Interactive geometry package, by Ulrich Kortenkamp, EUR49.00 for single user licence, 2007. Available for download at www.cinderella.de. The Mathematical Gazette, 95(534):573–575, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2011:BRBa

[Gib11b] Peter Giblin. Book review: Duel at dawn: heroes, martyrs and the rise of modern mathematics, by Amir Alexander. Pp. 307. £21.95 2010. ISBN 978-0-674-04661-0 (Harvard University Press). The Mathematical Gazette, 95(532):146–147, March 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2011:BRBb

[Gib11c] Peter Giblin. Book review: Fundamentals of university mathematics, (3rd edn.), Colin McGregor, Jonathan Nimmo and the late Wilson Stothers. Pp. 551. £40 (pbk). 2010. ISBN 978-0-85709-223-6 (print), 978-0-85709-224-3 (online) (Woodhead Publishing Paperback). The Mathematical Gazette, 95(534):570, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2012:BRBb

[Gib12a] Peter Giblin. Book review: Gösta Mittag-Leffler: a man of conviction, by Arild Stubhaug. Translated by Tiina Nunnally. Pp. 733. £53.99. 2010, ISBN: 978-3-642-11671-1 (Springer). The Mathematical Gazette, 96(536):368–370, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2012:BRBa

[Gib12b] Peter Giblin. Book review: The mathematical works of Leon Battista Alberti, Kim Williams, Lionel March, Stephen R. Wassell (eds.). Pp. 223. £69. 2010. ISBN 978-3-0346-0473-4 (Birkhäuser). The Mathematical Gazette, 96(535):187–188, March 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2013:BRBb

[Gib13a] Peter Giblin. Book review: Differential geometry: bundles, connections, metrics and curvature, by Clifford Henry Taubes. Pp. 298. £27.50. 2011. ISBN 978-0-19-960587-3 (Oxford Graduate Texts in Mathematics No. 23, Oxford University Press). The Mathematical Gazette, 97(539):373, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2013:BRBa

[Gib13b] Peter Giblin. Book review: Remembering Sofya Kovalevskaya, by Michéle Audin. Pp. 284. £35.99. 2011 ISBN: 978-0-85729-928-4 (Springer-Verlag). The Mathematical Gazette, 97(539):363–364, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2014:AAS

[Gib14a] Peter Giblin. 98.25 an ancient approximation to the sine function. The Mathematical Gazette, 98(543):488–490, November 2014. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2014:BRBa

[Gib14b] Peter Giblin. Book review: Divided spheres, Edward S. Popko, Pp. 509. £44.99. 2012, ISBN 978-1-4665-0429-5 (CRC Press (an A. K.Peters Book)). The Mathematical Gazette, 98(541):178–179, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2014:BRBb

[Gib14c] Peter Giblin. Book review: Unexpected expectations, Leonard M. Warner. Pp. 204. £24.69 (hardback). 2012. ISBN 978-1-56881-721-7 (CRC Press). The Mathematical Gazette, 98(541):182–183, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Giblin:2017:BRL

[Gib17a] Peter Giblin. Book review: Latin squares and their applications (2nd edn.), by Donald Keedwell and Jósef Dénes, pp. 424, £45.89 (hard), ISBN 978-0-444-63555-6, North-Holland (2015).

The Mathematical Gazette, 101(552):571–573, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/latin-squares-and-their-applications-
2nd-edn-by-donald-keedwell-and-josef-denes-pp-424-4589-
hard-isbn-9780444635556-northholland-2015/61C4E240044A25D3BDC591BF354B3CCC.

Giblin:2017:BRQ

[Gib17b] Peter Giblin. Book review: Quite right: the story of mathematics, measurement and money by Norman Biggs, pp. 176, £16.85 (hard), ISBN 978-0-19-875335-3, Oxford University Press (2016). The Mathematical Gazette, 101(552):557–558, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/quite-right-the-story-of-
mathematics-measurement-and-money-by-norman-biggs-pp-176-
1685-hard-isbn9780198753353-oxford-university-press-2016/
3F999097135DD4BFA887C8AE98BB66A5.

Giblin:2017:BRP

[Gib17c] Peter Giblin. Book review: The population explosion and other puzzles, by Hess, Dick, pp. 152, £23 (paper), ISBN 978-9-814723-375-5, also available as e-book, World Scientific Press (2016). The Mathematical Gazette, 101(551):380–381, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
2218D986A8CE49744B9E21F2D2941472.

Giblin:2018:BRP

[Gib18a] Peter Giblin. Book review: The power of computational thinking by Paul Curzon and Peter W. McOwan, pp. 220, £20 (paper), ISBN 978-1-786-34184-6, World Scientific (2017). The Mathematical Gazette, 102(554):382–383, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/power-of-computational-thinking-by-paul-curzon-
and-peter-w-mcowan-pp-220-20-paper-isbn-9781786341846-
world-scientific-2017/19EAFC86B93E1AD32E4F8575EB47E392.

Giblin:2018:BRS

[Gib18b] Peter Giblin. Book review: The story of calculus by David Acheson, pp. 192, £11.99 (hard), ISBN 978-0-19-880454-3, Oxford University Press (2017). The Mathematical Gazette, 102(554):364, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/story-of-calculus-
by-david-acheson-pp-192-1199-hard-isbn-9780198804543-oxford-
university-press-2017/40B215A395D3E702337905D3ECE9EC29.

Giblin:2019:NPF

[Gib19a] Peter Giblin. 103.12 a note on problem 101.f. The Mathematical Gazette, 103(556):154–155, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10312-a-note-on-problem-101f/5427CBF250D2FFAF1487180B0689A768.

Giblin:2019:BRN

[Gib19b] Peter Giblin. Book review: Never a dull moment: Hassler Whitney, mathematics pioneer by Keith Kendig, pp. 385, £58.50 (paper), ISBN 978-1-4704-4828-8, MAA Press, American Mathematical Society (2018). The Mathematical Gazette, 103(558):571, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/never-a-dull-moment-hassler-whitney-mathematics-
pioneer-by-keith-kendig-pp-385-5850-paper-isbn-9781470448288-
maa-press-american-mathematical-society-2018/20A8F66C056851B77E7DF488CD645FB9.

Giblin:2019:BRW

[Gib19c] Peter Giblin. Book review: Wonders beyond numbers: a brief history of all things mathematical by Johnny Ball, pp. 480, £16.99 (hard), ISBN 978-1-4729-3999-9, Bloomsbury Sigma (2017). The Mathematical Gazette, 103(556):184–185, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/wonders-beyond-numbers-a-
brief-history-of-all-things-mathematical-by-johnny-ball-
pp-480-1699-hard-isbn-9781472939999-bloomsbury-sigma-2017/
ACDD382120B0C5E601AB6491DD33DFB5.

Griffiths:2011:HLF

[GL11] Martin Griffiths and Nick Lord. The hook-length formula and generalised Catalan numbers. The Mathematical Gazette, 95(532): 23–30, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Glaister:2012:FES

[Gla12] Paul Glaister. 96.54 Further examples on sliding down curves. The Mathematical Gazette, 96(537):519–523, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Glaister:2013:TCN

[Gla13] Paul Glaister. 97.31 Tangents and coincident normals to a parabola. The Mathematical Gazette, 97(539):301–303, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Glaister:2019:AAI

[Gla19] Paul Glaister. Alice’s adventures in inverse tan land — mathematical argument, language and proof. The Mathematical Gazette, 103(558):388–400, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/alices-adventures-in-inverse-tan-land-mathematical-
argument-language-and-proof/741BEC35D7B2155712060787824A1B3F.

Glebov:2015:PPI

[Gle15] Gleb Glebov. 99.06 A peculiar proof of an identity of Euler. The Mathematical Gazette, 99(544):139–143, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2015:AIN

[GM15] Martin Griffiths and Des MacHale. 99.04 Another irrational number. The Mathematical Gazette, 99(544):130–133, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Grozdev:2018:DIT

[GOD18a] Sava Grozdev, Hiroshi Okumura, and Deko Dekov. 102.13 distance from the incentre of the tangential triangle of an obtuse triangle to the Euler line. The Mathematical Gazette, 102(553):133–135, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10213-distance-from-the-incentre-of-the-tangential-
triangle-of-an-obtuse-triangle-to-the-euler-line/F525030DE002FD4D33140C406DA98C4C.

Grozdev:2018:NFT

[GOD18b] Sava Grozdev, Hiroshi Okumura, and Deko Dekov. 102.14 a note on the Feuerbach triangle. The Mathematical Gazette, 102(553):135–138, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10214-a-note-on-the-feuerbach-triangle/680316E6434529E74627F9AB91B66C35.

Golding:2010:BRB

[Gol10] Jennie Golding. Book review: Mechanics M1: A Level mathematics for Edexcel, by Brian Jefferson. Pp. 185. £10.00. 2008. ISBN 978-0-19-911781-9 (Oxford University Press). The Mathematical Gazette, 94(530):358, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Golding:2017:IMI

[Gol17] Jennie Golding. Is it mathematics or is it school mathematics?: Presidential address to The Mathematical Association, April 2017. The Mathematical Gazette, 101(552):385–400, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
6454A4F7FFE7C73BDFB1482698EE31B2.

Gordon:2013:ST

[Gor13] Russell A. Gordon. Splitting a triangle. The Mathematical Gazette, 97(540):474–493, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Gordon:2018:RAL

[Gor18] Russell A. Gordon. Rational arc length. The Mathematical Gazette, 102(554):210–225, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/rational-arc-length/DDF29F0CD3EEC70EEA78603297107381.

Gordon:2019:VT

[Gor19a] Russell A. Gordon. 103.16 Variations on a theme. The Mathematical Gazette, 103(557):306–312, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10316-variations-on-a-theme/612A7A9754F55F671E7018E46997757B.

Gordon:2019:GPL

[Gor19b] Russell A. Gordon. 103.20 Geometric problems leading to power means. The Mathematical Gazette, 103(557):318–323, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10320-geometric-problems-
leading-to-power-means/03882C1C34C179B605F504CAC4F4A871.

Griffiths:2013:CL

[GR13] Martin Griffiths and Surajit Rajagopal. Conjecturing a limit. The Mathematical Gazette, 97(540):532–535, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Gratzer:2011:TYB

[Grä11] George A. Grätzer. Train your brain: a year’s worth of puzzles. A.K. Peters/CRC Press, Boca Raton, FL, 2011. ISBN 1-56881-710-X. xviii + 235 pp. LCCN GV1507.W8 G73 2011. Translated by Tom Artin.

Granger:2013:CSA

[Gra13] Robert Granger. Could, or should, the ancient Greeks have discovered the Lucas–Lehmer test? The Mathematical Gazette, 97(539): 242–255, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Greco:2019:HSP

[Gre19] Frank A. Greco. How is the scalar product derived? The Mathematical Gazette, 103(557):357–358, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/how-is-the-scalar-product-derived/B93A2F348D47DEB40BE36BCF3782823E.

Griffiths:2008:BPT

[Gri08] Martin Griffiths. The backbone of Pascal’s triangle, volume number one of Excursions in mathematics. The United Kingdom Mathematics Trust, Leeds, UK, 2008. ISBN 1-906001-04-9 (paperback). xiii + 163 pp. LCCN QA241 .G75 2008.

Griffiths:2010:MDK

[Gri10a] Martin Griffiths. 94.37 Multinomial distributions and knockout tournaments. The Mathematical Gazette, 94(531):505–509, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2010:BRB

[Gri10b] Martin Griffiths. Book review: Emmy Noether: the mother of modern algebra, by M. B. W. Tent. Pp. 177. $29. (Hardback). 2008. ISBN 978-1-56881-430-8 (A. K. Peters). The Mathematical Gazette, 94(529):177–178, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2011:SRE

[Gri11a] Martin Griffiths. 95.27 Symmetric rational expressions in polynomial sequences. The Mathematical Gazette, 95(533):279–284, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2011:BRBc

[Gri11b] Martin Griffiths. Book review: A mathematical medley: fifty easy pieces on mathematics, by George G. Szpiro. Pp. 236. £26.95. 2010. ISBN: 978-0-8218-4928-6 (American Mathematical Society). The Mathematical Gazette, 95(532):173–174, March 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2011:BRBd

[Gri11c] Martin Griffiths. Book review: How to guard an art gallery: and other discrete mathematical adventures, by T. S. Michael. Pp. 257. £13.00. 2009. ISBN: 978-0-8018-9299-8 (Johns Hopkins University Press). The Mathematical Gazette, 95(534):572, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2011:BRBa

[Gri11d] Martin Griffiths. Book review: Statistics GCSE for AQA: (2nd edn.), by Jayne Kranat. Pp. 330. £19.50. 2009. ISBN: 978-0-19-913430-4 (Oxford University Press). The Mathematical Gazette, 95(532):144–145, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2011:BRBb

[Gri11e] Martin Griffiths. Book review: Those fascinating numbers, by Jean-Marie De Koninck. Pp. 426. £34.95. 2009. ISBN: 978-0-8218-4807-4 (American Mathematical Society). The Mathematical Gazette, 95(532):156–157, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2011:DFT

[Gri11f] Martin Griffiths. A digit function with Thomae-like properties. The Mathematical Gazette, 95(534):397–403, November 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2011:GSG

[Gri11g] Martin Griffiths. The geometry of a stable game. The Mathematical Gazette, 95(532):52–58, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2011:TPM

[Gri11h] Martin Griffiths. Is trigonometry the preserve of the mathematical élite? The Mathematical Gazette, 95(533):256–261, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2012:OSP

[Gri12a] Martin Griffiths. 96.11 Observations on some particular power series. The Mathematical Gazette, 96(535):117–121, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2012:SSR

[Gri12b] Martin Griffiths. 96.12 The sum of a series: rational or irrational? The Mathematical Gazette, 96(535):121–124, March 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2012:FTO

[Gri12c] Martin Griffiths. 96.22 Further thoughts on ordered partitions. The Mathematical Gazette, 96(535):152–153, March 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2012:DSN

[Gri12d] Martin Griffiths. 96.49 On the diagonals of a Stirling number triangle. The Mathematical Gazette, 96(536):333–337, July 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2012:SDA

[Gri12e] Martin Griffiths. 96.52 Some different approaches to summing a particular series. The Mathematical Gazette, 96(537):512–515, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2012:BRB

[Gri12f] Martin Griffiths. Book review: Famous puzzles of great mathematicians, by M. Petković. Pp. 325. £25.95. 2009. ISBN: 978-0-8218-4814-2 (American Mathematical Society). The Mathematical Gazette, 96(536):381–382, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2012:PPD

[Gri12g] Martin Griffiths. A prime puzzle: Dirichlet’s theorem on primes in arithmetic progressions, volume 2 of Excursions in mathematics. The United Kingdom Mathematics Trust, Leeds, UK, 2012. ISBN 1-906001-16-2 (paperback). xiv + 217 pp. LCCN ????

Griffiths:2012:YAC

[Gri12h] Martin Griffiths. Yet another coin-tossing scenario. The Mathematical Gazette, 96(537):543–546, November 2012. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2013:BRBa

[Gri13a] Martin Griffiths. Book review: Fisber, Neyman, and tbe creation of classical statistics, by Erich Lehmann. Pp. 115. £24.99. 2011. ISBN: 978-1-4419-9499-8 (Springer). The Mathematical Gazette, 97(538):190–191, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2013:BRBb

[Gri13b] Martin Griffiths. Book review: Fundamental maths for engineering and science, by Mark Breach. Pp. 380. £25.99. 2011. ISBN: 978-0-230-25208-0 (Palgrave Macmillan). The Mathematical Gazette, 97(539):355–356, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2013:BRBd

[Gri13c] Martin Griffiths. Book review: Perfect rigour: a genius and the mathematical breakthrough of the century, by Masha Gessen Pp. 242 £14.99 (Hardback). 2011. ISBN: 978-1-84831-238-8 (Icon Books). The Mathematical Gazette, 97(539):378, July 2013. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2013:BRBc

[Gri13d] Martin Griffiths. Book review: The colours of infinity: the beauty and power of fractals, by Nigel Lesmoir-Gordon (ed.). Pp. 207 £29.95 (Hardback). 2010. ISBN: 978-1-84996-485-2 (Springer). The Mathematical Gazette, 97(539):373–374, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2013:GRE

[Gri13e] Martin Griffiths. From golden-ratio equalities to Fibonacci and Lucas identities. The Mathematical Gazette, 97(539):234–241, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2013:BRBe

[Gri13f] Matin Griffiths. Book review: The irrationals: A story of the numbers you can’t count on, by Julian Havil. Pp. 298. £19.95 (Hardback). 2012. ISBN: 978-0-691-14342-2 (Princeton University Press). The Mathematical Gazette, 97(540):569–570, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2014:BRBb

[Gri14a] Martin Griffiths. Book review: 17 equations that changed the world, by Ian Stewart. Pp. 342. £15.99. 2012. ISBN: 978-1-84668-531-6 (Profile). The Mathematical Gazette, 98(541):159– 160, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2014:BRBd

[Gri14b] Martin Griffiths. Book review: 9 algorithms that changed the future: the ingenious ideas that drive today’s computers, by John MacCormick. Pp. 219. $27-95 (hbk). 2012. ISBN: 978-0-691-14714-7 (Princeton University Press). The Mathematical Gazette, 98 (541):181–182, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2014:BRBe

[Gri14c] Martin Griffiths. Book review: Algorithmic puzzles, by Anany Levitin and Maria Levitin. Pp. 257. £13.99. 2011. ISBN: 978-0-19-974044-4 (Oxford University Press). The Mathematical Gazette, 98(541):188, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2014:BRE

[Gri14d] Martin Griffiths. Book review: European success stories in industrial mathematics by T. Lery, M. Primicerio, M.J. Esteban, M. Fontes, Y. Maday, V. Mehrmann, G. Quadros, W. Schilders, A. Schuppert and H. Tewkesbury, pp 135, £44.99, ISBN 978-3-642-23847-5, Springer (2012). The Mathematical Gazette, 98(543): 558, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2014:BRBc

[Gri14e] Martin Griffiths. Book review: Introduction to abstract algebra (4th edn.), by W. Keith Nicholson. Pp. 535. £90-50 (hbk). 2012. ISBN: 978-1-118-13535-8 (Wiley). The Mathematical Gazette, 98 (541):176–177, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2014:BRBa

[Gri14f] Martin Griffiths. Book review: Mathematics of life: unlocking the secrets of existence, by Ian Stewart. Pp. 358. £20.00 (Hardback). 2011. ISBN: 978-1-84668-198-1 (Profile). The Mathematical Gazette, 98(541):158–159, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2015:LT

[Gri15a] Jonny Griffiths. 99.24 The not-on-the-line transformation. The Mathematical Gazette, 99(545):347–352, July 2015. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2015:SIF

[Gri15b] Martin Griffiths. 99.02 Series involving fixed-column decimal representations. The Mathematical Gazette, 99(544):125–128, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2015:ISR

[Gri15c] Martin Griffiths. 99.17 Irrational sums from reciprocals of factorials. The Mathematical Gazette, 99(545):331–335, July 2015. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2015:GRE

[Gri15d] Martin Griffiths. 99.22 The golden ratio and equilateral triangles. The Mathematical Gazette, 99(545):342–346, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2015:SRP

[Gri15e] Martin Griffiths. 99.29 on the sum of the reciprocals of the primorials. The Mathematical Gazette, 99(546):522–523, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2015:CUO

[Gri15f] Martin Griffiths. The carry-on-until-one-girl proportion. The Mathematical Gazette, 99(546):464–467, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2015:RWA

[Gri15g] Martin Griffiths. Random walks arising from a Fibonacci’s-rabbits scenario. The Mathematical Gazette, 99(544):60–67, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2015:SCS

[Gri15h] Martin Griffiths. A slowly converging sequence via Pythagoras. The Mathematical Gazette, 99(546):427–431, November 2015. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2016:ICC

[Gri16a] Jonny Griffiths. Identities connecting the Chebyshev polynomials. The Mathematical Gazette, 100(549):450–459, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
39D2879007F5EBEE4F4597EEBF266C71.

Griffiths:2016:TIB

[Gri16b] Martin Griffiths. 100.11 a trinomial identity from both a tiling and an expansion. The Mathematical Gazette, 100(547):137–139, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Griffiths:2017:NTA

[Gri17a] Martin Griffiths. 101.25 a note on tilings and associated linear recurrences. The Mathematical Gazette, 101(551):314–318, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
EFFB4A84D358186281636050BD0B4C4C.

Griffiths:2017:TFI

[Gri17b] Martin Griffiths. 101.44 on a trivariate Fibonacci identity. The Mathematical Gazette, 101(552):519–522, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10144-on-a-
trivariate-fibonacci-identity/3B7FB8358D69D30E49364C7EF70D87A5.

Griffiths:2017:BRG

[Gri17c] Martin Griffiths. Book review: Genius at play: the curious mind of John Horton Conway by Roberts, Siobhan, pp. 472, £20.00 (hard), ISBN 978-1-62040-593-2, Bloomsbury Publishing (2015). The Mathematical Gazette, 101(550):167–168, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
74643ABFC52F183D743E5094098E216E.

Griffiths:2017:BRN

[Gri17d] Martin Griffiths. Book review: Number treasury 3: investigations, facts and conjectures about more than 100 number families (3rd edn.), by Kenney, Margaret J. and Bezuszka, Stanley J., pp. 312, £25.00 (paper), ISBN 9789814603690, World Scientific Publishing (2015). The Mathematical Gazette, 101(551):348–349, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
9E262D8B65A4CA0B3FE2482B6CB74CE4.

Griffiths:2018:RDP

[Gri18a] Martin Griffiths. 102.35 on repetition and divisibility of parts in partitions. The Mathematical Gazette, 102(554):344–347, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10235-on-repetition-
and-divisibility-of-parts-in-partitions/72B3056C60D0C11F06630A426296B985.

Griffiths:2018:DPA

[Gri18b] Martin Griffiths. 102.36 a difference property amongst certain pairs of Beatty sequences. The Mathematical Gazette, 102(554):348–350, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10236-a-difference-property-amongst-certain-pairs-
of-beatty-sequences/9E15D3D269E92A943865B9C8625699CD.

Griffiths:2018:BRW

[Gri18c] Martin Griffiths. Book review: A walk through combinatorics (Fourth Edition) by Miklós Bóna, pp. 593, £91 (hard), ISBN 978-9-81314-884-0, World Scientific (2017). The Mathematical Gazette, 102(554):376–377, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/walk-through-combinatorics-fourth-edition-by-miklos-
bona-pp-593-91-hard-isbn-9789813148840-world-scientific-
2017/A88F8230B6AF65EE6328A7A42A62E288.

Griffiths:2018:BRI

[Gri18d] Martin Griffiths. Book review: Introduction to combinatorics by Gerry Leversha and Dominic Rowland, pp. 377, £18 (paper), ISBN 978-1-90600-124-7, United Kingdom Mathematics Trust (2016). The Mathematical Gazette, 102(553):187–189, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/introduction-to-
combinatorics-by-gerry-leversha-and-dominic-rowland-pp-
377-18-paper-isbn-9781906001247-united-kingdom-mathematics-
trust-2016/755CA4A28A9FF76EAD632AACBFBCCB58.

Griffiths:2018:BRS

[Gri18e] Martin Griffiths. Book review: The story of numbers by Asok Kumar Mallik, pp. 196, £42.00 (hard), ISBN 978-981-3222-92-2, World Scientific (2017). The Mathematical Gazette, 102(554):367–368, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/story-of-numbers-by-asok-kumar-mallik-pp-196-4200-
hard-isbn-9789813222922-world-scientific-2017/6E20AE9844116B33B0C81174DC42C5A3.

Griffiths:2019:PPU

[Gri19] Martin Griffiths. 103.27 A property of a particular unit-distance graph. The Mathematical Gazette, 103(557):353–356, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/

mathematical-gazette/article/10327-a-property-of-a-particular-
unitdistance-graph/C7CEAEBAD4648EC991D8AD3144E22B9B.

Groetsch:2011:SO

[Gro11] C. W. Groetsch. 95.19 Spurt optimization. The Mathematical Gazette, 95(532):118–119, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Groetsch:2018:TTD

[Gro18] C. W. Groetsch. Two trajectories: a dynamic monkey tale. The Mathematical Gazette, 102(553):164–170, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/two-trajectories-a-dynamic-
monkey-tale/268176B11070D08810DD2CCB3D679F9E.

Gager:2011:SLR

[GS11] Philip C. Gager and Mark B. Shackleton. A snakes and ladders representation of stock prices and returns. The Mathematical Gazette, 95(533):177–185, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Gustavsson:2016:DTF

[GS16] Jan Gustavsson and Mikael P. Sundqvist. Defining trigonometric functions via complex sequences. The Mathematical Gazette, 100 (547):9–23, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Guest:2008:QCI

[Gue08] Martin A. Guest. From quantum cohomology to integrable systems. Oxford graduate texts in mathematics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-856599-2. xxix + 301 pp. LCCN QA612.3 .G84 2008. URL http://www.

loc.gov/catdir/toc/fy0805/2007035101.html.

Gustafson:2012:CHM

[Gus12] Karl E. Gustafson. The crossing of heaven: memoirs of a mathematician. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-22557-8 (hardcover). xvi + 176 pp. LCCN QA29.G873 A3 2012.

Gusic:2015:BPE

[Gus15] Ivica Gusić. On the bounds for the perimeter of an ellipse. The Mathematical Gazette, 99(546):540–541, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Guthery:2011:MMH

[Gut11] Scott B. Guthery. A motif of mathematics: [history and application of the mediant and the Farey sequence]. Docent Press, Boston, MA, USA, 2011. ISBN 1-4538-1057-9 (paperback). xx + 243 pp. LCCN QA241 .G864 2011.

Grivoyannis:2015:PWP

[GV15] B. Grivoyannis and R. Viglione. 99.11 Proof without words: A problem on three squares. The Mathematical Gazette, 99(544): 154, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Grivoyannis:2017:PWB

[GV17] B. Grivoyannis and R. Viglione. 101.37 proof without words: The birth of a square (or two). The Mathematical Gazette, 101(552): 497–498, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/36B426F276068C9840C9FDE19001F455.

Goodman:2009:SRI

[GW09] Roe Goodman and Nolan R. Wallach. Symmetry, representations, and invariants, volume 255 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-79851-X, 0-387-79852-8 (ebook). xx + 716 pp. LCCN QA176 .G663 2009. URL http://

springerlink.metapress.com/content/978-0-387-79851-6/.

Haddley:2019:VMP

[Had19] Joel Haddley. Book review: Visualizing mathematics with 3D printing by Henry Segerman, pp. 186, $70 (hard), ISBN 978-1-421-42035-6, Johns Hopkins University Press (2016). The Mathematical Gazette, 103(557):382–383, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/visualizing-mathematics-with-3d-printing-by-henry-
segerman-pp-186-70-hard-isbn-9781421420356-johns-hopkins-
university-press-2016/A2D2A84D30EAE9C31127247B55FBF2A4.

Hahn:2012:MEW

[Hah12] Alexander J. Hahn. Mathematical excursions to the world’s great buildings. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 1-4008-4199-2 (e-book), 0-691-14520-2 (hardcover) 1-280-49439-5. ix + 317 + 16 pp. LCCN NA2750 .H325 2012. URL

http://www.jstor.org/stable/10.2307/j.ctt7shns.

Haigh:2010:BRB

[Hai10] John Haigh. Book review: Risk and reward: The science of casino blackjack, by N. Richard Werthamer. Pp. 131. £44-99 (hbk). 2009. ISBN: 978-1-4419-0252-8 (Springer). The Mathematical Gazette, 94(531):569–570, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Haigh:2012:BRB

[Hai12] John Haigh. Book review: Probability and stochastics, by Erhan Çinlar. Pp. 557. £53.00 (hbk). 2011. ISBN: 978-0-387-87858-4 (Springer). The Mathematical Gazette, 96(536):377–378, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Haigh:2013:BRB

[Hai13] John Haigh. Book review: Golf by the numbers, by Roland Minton. Pp. 291 £18.00 (hbk). 2012. ISBN: 978-1-4214-0315-1 (Johns Hopkins University Press). The Mathematical Gazette, 97(540):557– 558, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Haigh:2015:BRB

[Hai15] John Haigh. Book review: Janus-faced probability by Paolo Rocchi, pp 147, £60.99 (hbk), ISBN 978-3-319-04860-4, Springer Verlag (2014). The Mathematical Gazette, 99(544):187, March 2015. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Haigh:2017:BRC

[Hai17a] John Haigh. Book review: Classic topics on the history of modern mathematical statistics, by Gorroochurn, Prakash, pp. 754, £90.50 (hard), ISBN 978-1-119-12792-5, also available as ebook, John Wiley & Sons (2016). The Mathematical Gazette, 101(551):378–379, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
60AC24BD15AE82142D8FCF4B1373EF5A.

Haigh:2017:BRS

[Hai17b] John Haigh. Book review: Statistics for mathematicians by Victor M. Panaretos, pp. 196 (paper), £35.99, ISBN 978-3-319-28339-5, also available as e-book, Birkhauser (2016). The Mathematical Gazette, 101(552):573–574, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/statistics-for-mathematicians-by-victor-m-panaretos-
pp-196-paper-3599-isbn-9783319283395-also-available-as-
ebook-birkhauser-2016/20DB9E562863FD5C0AAC1D88D06B5DD1.

Hajja:2010:PAP

[Haj10] Mowaffaq Hajja. 94.35 Popularising the ASS pseudo-congruence theorem. The Mathematical Gazette, 94(531):499–502, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hajja:2011:WDP

[Haj11] Mowaffaq Hajja. 95.45 What do permutations of a given cycle pattern generate? The Mathematical Gazette, 95(533):341–342, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hajja:2012:EPI

[Haj12] Mowaffaq Hajja. 96.44 Extremal properties of the incentre and the excentres of a triangle. The Mathematical Gazette, 96(536): 315–317, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hajja:2015:SPE

[Haj15a] Mowaffaq Hajja. 99.21 The simplest proof ever of the pons asinorum theorem. The Mathematical Gazette, 99(545):341–342, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hajja:2015:SOP

[Haj15b] Mowaffaq Hajja. Some open problems in concurrences. The Mathematical Gazette, 99(544):2–10, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hajja:2017:AMH

[Haj17a] Mowaffaq Hajja. 101.11 An application of a morsel of Honsberger. The Mathematical Gazette, 101(550):123–125, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
B1C179650EA1C1C1F568B137DC00EF47.

Hajja:2017:SSL

[Haj17b] Mowaffaq Hajja. 101.12 The sum of the squared lengths of equally-spaced cevians. The Mathematical Gazette, 101(550): 126–128, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/8415BB5090CD10F444D302D5D5E3415A.

Hajja:2017:HVS

[Haj17c] Mowaffaq Hajja. 101.22 the hyperbolic version of the Steiner– Lehmus theorem. The Mathematical Gazette, 101(551):306–307, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
3136F8F72622CE83AFDF9B1B23C41B9B.

Hajja:2017:OMN

[Haj17d] Mowaffaq Hajja. 101.23 one more note on the extremal properties of the incentre and the excentres of a triangle. The Mathematical Gazette, 101(551):308–310, July 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
30B9D8EA3355FF684E9F1BA319EFE078.

Hajja:2017:CD

[Haj17e] Mowaffaq Hajja. 101.27 on the convention that deg(0) = −∞. The Mathematical Gazette, 101(552):465–470, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
43FEAB9FC221E32B8668649011D4D1FB.

Hajja:2018:VSP

[Haj18a] Mowaffaq Hajja. 102.49 a very short proof of Pamfilos’s characterisation of the rhombus. The Mathematical Gazette, 102 (555):521–523, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10249-a-very-short-proof-of-pamfiloss-characterisation-
of-the-rhombus/E8FA23D37E9F96C70209F1C6561D8D13.

Hajja:2018:ETT

[Haj18b] Mowaffaq Hajja. The existence of triangles, tetrahedra, and higher-dimensional simplices with prescribed exradii. The Mathematical Gazette, 102(554):257–263, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/existence-of-triangles-tetrahedra-and-higherdimensional-
simplices-with-prescribed-exradii/2A2B395BB265C2120C320A8C0BD56CF1.

Hajja:2019:PST

[Haj19] Mowaffaq Hajja. 103.22 On a problem of Sastry and a theorem of Serenus. The Mathematical Gazette, 103(557):328– 332, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10322-on-a-
problem-of-sastry-and-a-theorem-of-serenus/B5DB742316EB8547BFE886609D191258.

Hall:2010:BRBb

[Hal10a] Peter Hall. Book review: GCSE revision plus Edexcel maths revision guide, edited by John Proctor, Tina Foster and Rebecca Skinner. Pp. 120. £5-99. 2008. ISBN: 978-1-844191-91-8 (Letts and Lonsdale). The Mathematical Gazette, 94(531):553, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hall:2010:BRBc

[Hal10b] Peter Hall. Book review: It’s a kind of magic: mathematical magic tricks explained, by David Crawford. pp. 258. Members £19-99, Non-members £23-99. 2009. ISBN: 0-906588-67-7 (The Mathematical Association). The Mathematical Gazette, 94(531):574–575, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hall:2010:BRBa

[Hal10c] Peter Hall. Book review: Statistics S1: A Level mathematics for Edexcel, by James Nicholson. Pp. 223. £10.00. 2008. ISBN: 978-0-19-911782-6 (Oxford University Press). The Mathematical Gazette, 94(530):356–357, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hall:2011:BRBc

[Hal11a] Peter Hall. Book review: Assessment methods in statistical education: an international perspective, by Penelope Bidgood, Neville Hunt and Flavia Jolliffe (Editors). Pp. 286. £55.00. 2010. ISBN: 978-0-470-74532-8 (Wiley). The Mathematical Gazette, 95(533): 373–375, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hall:2011:BRBa

[Hal11b] Peter Hall. Book review: Essentials: KS3 mathematics coursebook, by Fiona Mapp. Pp. 288. £13.00. 2009. ISBN: 978-1-844192-39-7 (Lonsdale). The Mathematical Gazette, 95(532):144, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hall:2011:BRBb

[Hal11c] Peter Hall. Book review: Statistics: basic principles and applications (2nd edn), by William J. Adams, Irwin Kabus and Mitchell P. Preiss. Pp. 731. $23.99. 2009. ISBN: 978-1-4363-0180-0 (Xlibris). The Mathematical Gazette, 95(532):145, March 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Halbeisen:2012:CST

[Hal12a] Lorenz J. Halbeisen. Combinatorial set theory: with a gentle introduction to forcing. Springer monographs in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4471-2172-4, 1-4471-2173-2 (e-book). ISSN 1439-7382. xvi + 453 pp. LCCN QA248 .H238 2012.

Hall:2012:BRB

[Hal12b] Peter Hall. Book review: Success: AQA GCSE mathematics higher revision guide, by Fiona Mapp. Pp. 112. £4.99. 2010. ISBN: 978-1-84419-286-1 (Letts). Success: AQA GCSE mathematics higher workbook, by Fiona Mapp. Pp. 96. £4.99, 2010. ISBN: 978-1-84419-287-8 (Letts). The Mathematical Gazette, 96(536):365–366, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hall:2013:BRB

[Hal13] Peter Hall. Book review: Collins New Maths Frameworking Level 6, booster workbook (no author specified). Pp. 116. £5.99. 2011. ISBN: 978-0-00-743808-2 (Collins). The Mathematical Gazette, 97 (539):354–355, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hall:2014:BRB

[Hal14] Peter Hall. Book review: Edexcel A level maths Core 1: Students’ support material, by John Berry and Sue Langham. Pp.

102. £5.99. 2012. ISBN 978-0-00-747601-5 (Collins). The Mathematical Gazette, 98(542):370–371, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hall:2015:BRBa

[Hal15a] Peter Hall. Book review: Cambridge Additional Mathematics by Michael Haese, Sandra Haese, Mark Humphries, Chris Sangwin, pp. 504 (paper), AU$50.00, ISBN 978-1-92197-242-3, Haese Mathematics (2014). The Mathematical Gazette, 99(544):176, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hall:2015:BRBb

[Hal15b] Peter Hall. Book review: Cambridge Additional Mathematics by Michael Haese, Sandra Haese, Mark Humphries, Chris Sangwin, pp. 504 (paper), AU$50.00, ISBN 978-1-92197-242-3, Haese Mathematics (2014). The Mathematical Gazette, 99(544):176, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hall:2016:BRH

[Hal16] Peter Hall. Book review: How humans learn to think mathematically by David Tall, pp. 457, £27.99 (paper), ISBN 978-1-107-66854-6, Cambridge University Press (2014). The Mathematical Gazette, 100(547):175, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hall:2017:BRD

[Hal17a] Peter Hall. Book review: Does mathematical study develop logical thinking?, by Inglis, Matthew and Attridge, Nina, pp. 185, £68 (hard), ISBN 978-1-78634-068-9, World Scientific (2017). The Mathematical Gazette, 101(551):349, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
E5472206E9CBE2F4B601E57633890805.

Hall:2017:BRH

[Hal17b] Peter Hall. Book review: Higher and foundation GCSE mathematics revision cards, by Ledwick, Mary, each 90 cards, Higher £7.99, Foundation £7.49, ISBN 978-0-90658-884-0, The Mathematical Association (2016). The Mathematical Gazette, 101(551):350–351, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
FA8A0929E82A5EF8596F7C539D09FDEF.

Hall:2017:BRL

[Hal17c] Peter Hall. Book review: Lesson study, challenges in mathematics education, edited by Inprasitha, Maitree et al., pp. 377, £32 (paper), ISBN 978-98128-3541-3, World Scientific (2015). The Mathematical Gazette, 101(550):160–161, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
76F15AABDEAB8A4CCEC67E1CD6A02419.

Hall:2017:BRS

[Hal17d] Peter Hall. Book review: Strategy games to enhance problemsolving ability in mathematics, by Posamentier, Alfred S. and Krulik, Stephen, pp. 121, £20 (paper), ISBN 978-9-81314-634-1, World Scientific (2017). The Mathematical Gazette, 101(551):383, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
32E2D68E5960F85ADF83BCC6716392F4.

Hall:2019:EMP

[Hal19] Peter Hall. Book review: Exploring mathematics: problemsolving and proof by Daniel Greiser, pp. 320 (paper), £29.99, ISBN 978-3-319-90319-4, Springer Verlag (2018). The Mathematical Gazette, 103(557):381, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/exploring-mathematics-problemsolving-and-proof-
by-daniel-greiser-pp-320-paper-2999-isbn-9783319903194-
springer-verlag-2018/EBBFE0729DA971875BB80C0409ED7C80.

Harris:2010:FGM

[Har10a] J. R. Harris. Forming groups with 4 × 4 matrices. The Mathematical Gazette, 94(531):426–429, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hartshorne:2010:DT

[Har10b] Robin Hartshorne. Deformation theory, volume 257 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-4419-1595-8 (hardcover), 1-4419-1596-6 (e-book). ISSN 0072-5285. vi + 234 pp. LCCN QA614.58 .H37 2010.

Harris:2012:PP

[Har12] J. R. Harris. 96.31 Palindromic polynomials. The Mathematical Gazette, 96(536):266–269, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hardy:2014:NCV

[Har14] Michael Hardy. Notes 98.20 a concrete view of Euclid’s proof of the infinitude of primes. The Mathematical Gazette, 98(543):479– 481, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hassani:2019:TDP

[Has19] Mehdi Hassani. 103.30 Tests for divisibility by prime numbers. The Mathematical Gazette, 103(558):494–495, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10330-tests-for-
divisibility-by-prime-numbers/FE99D687B0412BCD021F6C9183AE311E.

Havil:2012:ISN

[Hav12] Julian Havil. The irrationals: a story of the numbers you can’t count on. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-14342-0. ix + 298 pp. LCCN QA247.5 .H38 2012.

Havil:2014:JNL

[Hav14] Julian Havil. John Napier: Life, Logarithms, and Legacy. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-15570-4, 0-691-14342-0. xv + 279 pp. LCCN QA247.5 .H38 2012.

Haworth:2011:BRB

[Haw11] Anne Haworth. Book review: Cows in the maze and other mathematical explorations, by Ian Stewart. pp. 287. £8.99. 2010. ISBN 978-0-19-956207-7 (Oxford University Press). The Mathematical Gazette, 95(533):382–383, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Haworth:2012:BRB

[Haw12] Anne Haworth. Book review: Street-fighting mathematics: the art of educated guessing and opportunistic problem solving, by Sanjoy Mahajan. Pp. 134. £18.95. 2010. ISBN: 978-0-262-51429-3. (MIT Press). The Mathematical Gazette, 96(537):572–573, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Haworth:2018:BRM

[Haw18a] Anne Haworth. Book review: More fallacies, flaws, and flimflam by Edward J. Barbeau, £23.99, ISBN 978-0-88385-580-5, Cambridge University Press (2017). The Mathematical Gazette, 102(555):569, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/more-fallacies-flaws-and-flimflam-by-edward-j-
barbeau-2399-isbn-9780883855805-cambridge-university-press-
2017/F88E42F23169C058E9685E3B9B6D3875.

Haworth:2018:BRC

[Haw18b] Anne Haworth. Book review: The calculus of happiness by Oscar E. Fernandez, pp. 159, £19.95 (hard), ISBN 978-0-69116-863-0, Princeton University Press (2017). The Mathematical Gazette, 102(554):367, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/calculus-of-happiness-by-oscar-e-fernandez-pp-
159-1995-hard-isbn-9780691168630-princeton-university-
press-2017/A031B36A48B964C9F8A0F3C9CDD00656.

Haworth:2018:BRW

[Haw18c] Anne Haworth. Book review: Weird math by David Darling and Agnijo Banerjee, pp. 320, £12.99 (paper), ISBN 978-1-78607-264-1, Basic Books (2018). The Mathematical Gazette, 102(555):560, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/weird-math-
by-david-darling-and-agnijo-banerjee-pp-320-1299-paper-
isbn-9781786072641-basic-books-2018/24E8F41119CA08C65471B79A3CC02D55.

Hathaway:2016:CMB

[HB16] Dale K. Hathaway and Joshua Barks. Calculus meets the birthday problem. The Mathematical Gazette, 100(547):86–92, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Humenberger:2019:ICM

[HE19] Hans Humenberger and Franz Embacher. Iterating circum-medial triangles. The Mathematical Gazette, 103(558):480–487, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/iterating-circummedial-
triangles/86B095C1CE562BD0EC41D27DA8DBA63E.

Hemery:2019:BRH

[Hem19] Adrian Hemery. Book review: How to bake Pi (previously published as Cakes, custard and category theory) by Eugenia Cheng, pp. 291, £9.99 (paper), ISBN 978-1-781-25288-8, Profile Books (2016). The Mathematical Gazette, 103(556): 183–184, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/how-to-
bake-pi-previously-published-as-cakes-custard-and-category-
theory-by-eugenia-cheng-pp-291-999-paper-isbn-9781781252888-
profile-books-2016/E7A400EF541DD21439D3DB2D955B1762.

Herrmann:2011:BEM

[Her11] Norbert Herrmann. The beauty of everyday mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-22103-3 (paperback). xiii + 138

pp. LCCN ???? URL http://site.ebrary.com/id/10505362.

Hansford:2004:PLA

[HF04] Dianne Hansford and Gerald E. Farin. Practical Linear Algebra: A Geometry Toolbox. A. K. Peters, Ltd., Wellesley, MA, USA, 2004. ISBN 1-56881-234-5. 384 pp. LCCN QA184.2.F37 2004.

Haese:2014:CAM

[HHHS14] Michael Haese, Sandra Haese, Mark Humphries, and Chris Sangwin. Cambridge Additional Mathematics. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-921972-42-4. 504 (est.) pp. LCCN ????

Heyne:2007:LEM

[HHPM07] Andreas K. Heyne, Alice K. Heyne, Elena S. Pini, and Tahu Matheson. Leonhard Euler: a Man to be Reckoned with. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2007. ISBN 3-7643-8332-1. 45 pp. LCCN QA29.E8 H49 2007.

Hennings:2014:SRC

[HI14] Mark Hennings and Jon Ingram. Sliding in rough convex bowls. The Mathematical Gazette, 98(541):8–23, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Higgins:2011:NVS

[Hig11] Peter M. Higgins. Numbers: a very short introduction, volume 260 of Very short introductions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-958405-2 (paperback). xiii + 1 + 132 pp. LCCN QA241 .H54 2011.

Hind:2008:CCC

[Hin08] Peter Hind. Core C1, C2: A level mathematics for Edexcel. A level mathematics for Edexcel. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-911779-9 (paperback). 375 pp. LCCN ????

Hirschhorn:2011:CT

[Hir11a] Michael D. Hirschhorn. 95.02 Commensurable triangles. The Mathematical Gazette, 95(532):61–63, March 2011. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hirschhorn:2011:MMC

[Hir11b] Michael D. Hirschhorn. 95.08 Maximisation/minimisation without calculus. The Mathematical Gazette, 95(532):83–85, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hirschhorn:2011:II

[Hir11c] Michael D. Hirschhorn. 95.11 An interesting integral. The Mathematical Gazette, 95(532):90–91, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hirschhorn:2011:SA

[Hir11d] Michael D. Hirschhorn. 95.35 A surprising appearance of ζ(3). The Mathematical Gazette, 95(533):313, July 2011. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hirschhorn:2011:LBC

[Hir11e] Michael D. Hirschhorn. 95.40 Lord Brouncker’s continued fraction for π. The Mathematical Gazette, 95(533):322–326, July 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hirschhorn:2011:WSC

[Hir11f] Michael D. Hirschhorn. 95.55 When is the sum of consecutive squares a square? The Mathematical Gazette, 95(534):511–512, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hirschhorn:2011:TFP

[Hir11g] Michael D. Hirschhorn. 95.57 A triangle-folding problem. The Mathematical Gazette, 95(534):514–517, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hirschhorn:2011:TOA

[Hir11h] Michael D. Hirschhorn. 95.61 Tenth order algorithms for calculating \sqrt{(}x), ^{3} \sqrt{3}. The Mathematical Gazette, 95(534):526–528, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hirschhorn:2012:IEC

[Hir12] Michael D. Hirschhorn. 96.17 Inequalities for Euler’s constant. The Mathematical Gazette, 96(535):133–136, March 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hirschhorn:2013:CFR

[Hir13] Michael D. Hirschhorn. 97.24 A constant in a formula of Ramanujan. The Mathematical Gazette, 97(539):276–277, July 2013. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Harminc:2018:DVP

[HJ18a] Matús Harminc and Lucia Janicková. Discrete version of the Pythagorean theorem. The Mathematical Gazette, 102(553):77–88, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/discrete-version-
of-the-pythagorean-theorem/2042EAE1B1BBB1155A9FBB84958AA759.

Huddy:2018:APT

[HJ18b]  Stanley R. Huddy and Michael A. Jones.  All parabo-

las through three non-collinear points. The Mathematical Gazette, 102(554):203–209, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/

article/all-parabolas-through-three-noncollinear-points/
3EFF0F24FCC030A3D7D559DCF93E752C.

Huddy:2019:QHP

[HJ19] Stanley R. Huddy and Michael A. Jones. 103.03 quotients of hypotenuses of Pythagorean triples (a, b, b + 1) and finite differences. The Mathematical Gazette, 103(556):132–135, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10303-quotients-
of-hypotenuses-of-pythagorean-triples-a-b-b-1-and-finite-
differences/21042D97FBB4A3A0556A1986B71CCEE5.

Hersh:2011:LHM

[HJS11] Reuben Hersh and Vera John-Steiner. Loving + hating mathematics: challenging the myths of mathematical life. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14247-5 (hardcover). x + 416 pp. LCCN QA28 .H48 2011.

Hokuto:2014:WBI

[HK14] Takeshi Hokuto and Mitsuhiro Kumano. Which is bigger? an intriguing ‘double alternation’. The Mathematical Gazette, 98(541): 67–72, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hoare:2010:BRB

[Hoa10] Graham Hoare. Book review: Benjamin Franklin’s numbers. An unsung mathematical odyssey, by Paul C. Pasles. pp. 254. £15.95. 2008. ISBN 13: 978-0-691-12956-3 (Princeton University Press). The Mathematical Gazette, 94(530):360–361, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hoare:2011:BRB

[Hoa11] Graham Hoare. Book review: The Abel Prize 2003–2007: The first five years, by Helge Holden and Ragui Piene (Editors). Pp. 327. £58.99 (Hardback). 2010. ISBN: 978-3-642-01372-0 (Springer). The Mathematical Gazette, 95(533):375–376, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hoare:2012:BRB

[Hoa12] Graham Hoare. Book review: Roads to infinity: the mathematics of truth and proof, by John Stillwell. Pp. 203. £25.99 (Hardback).

2010. ISBN 978-1-56881-466-7 (A. K. Peters Ltd). The Mathematical Gazette, 96(536):380–381, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hoare:2013:BRB

[Hoa13a] Graham Hoare. Book review: The mathematical writings of Évariste Galois, by Peter M. Neumann. Pp. 421. £74.50 (hbk.) 2011. ISBN: 978-3-03719-104-0 (European Mathematical Society). The Mathematical Gazette, 97(538):187–188, March 2013. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hoare:2013:RGM

[Hoa13b] Graham Hoare. R. L. Goodstein and mathematical logic. The Mathematical Gazette, 97(540):409–412, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hoare:2014:BRB

[Hoa14a] Graham Hoare. Book review: Mathematics, an illustrated history of numbers, Tom Jackson (ed). Pp. 144. £20. 2012. ISBN 978-1-84931-066-6 (Worth Press). The Mathematical Gazette, 98(542): 373–374, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hoare:2014:BRT

[Hoa14b] Graham Hoare. Book review: Three views of logic by Donald W. Loveland, Richard E. Hodel and S.G. Sterrett, pp 344, £34.95, ISBN 978-0-691-16044-3, Princeton University Press (2014). The Mathematical Gazette, 98(543):567–568, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hoare:2015:BRI

[Hoa15a] Graham Hoare. Book review: Interpreting Gödel: critical essays by Juliette Kennedy (ed), pp 279, £60 (hard), ISBN 978-1-10-00266-1, Cambridge University Press (2014). The Mathematical Gazette, 99(546):570–572, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hoare:2015:BRT

[Hoa15b] Graham Hoare. Book review: Taming the unknown: a history of algebra from antiquity to the early twentieth century by Victor J. Katz and Karen Hunger Parshall, pp. 485, £34.95 (hard), ISBN 978-0-691-14905-9, Princeton University Press (2014). The Mathematical Gazette, 99(546):561–562, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hoare:2017:BRL

[Hoa17a] Graham Hoare. Book review: Leonhard Euler, mathematical genius in the Enlightenment by Calinger, Ronald S., pp. 669, £37.95 (hard), ISBN 978-0-691-11927-4, Princeton University Press (2016). The Mathematical Gazette, 101(550):165–166, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
0B5A895AA617CD0BA0FC812A9C76C753.

Hoare:2017:MM

[Hoa17b] Graham T. Q. Hoare. Maryam Mirzakhani (1977–2017). The Mathematical Gazette, 101(552):545–546, November 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/maryam-mirzakhani-1977-2017/
1D79D40C88B61E750FC2DBB47EAF1943.

Hoare:2018:BRB

[Hoa18a] Graham Hoare. Book review: A beginner’s further guide to mathematical logic by Raymond Smullyan, pp. 277, £20 (paper), ISBN 978-9-81472-572-9, World Scientific Publishing (2017). The Mathematical Gazette, 102(553):190–191, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/a-beginners-further-guide-
to-mathematical-logic-by-raymond-smullyan-pp-277-20-paper-
isbn-9789814725729-world-scientific-publishing-2017/A247FCCF9F51352A1805E000315F8F3B.

Hoare:2018:MN

[Hoa18b] Graham T. Q. Hoare. 102.37 Markov numbers. The Mathematical Gazette, 102(555):494–496, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10237-markov-numbers/4C5084E88682EC0840FCCA091C270807.

Hoare:2018:AQP

[Hoa18c] Graham T. Q. Hoare. 102.51 Archimedes’ quadrature of the parabola. The Mathematical Gazette, 102(555):527–529, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/

journals/mathematical-gazette/article/10251-archimedes-
quadrature-of-the-parabola/D37894CD2CCC4FBF32E3A9AA5DA72681.

Hoare:2019:BRA

[Hoa19] Graham Hoare. Book review: The art of logic in an illogical world by Eugenia Cheng, pp. 305, £14.99 (hard), ISBN 978-1-5416-7248-2, Basic Books (2018). The Mathematical Gazette, 103(556):189–190, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/art-of-logic-in-an-illogical-world-by-eugenia-
cheng-pp-305-1499-hard-isbn-9781541672482-basic-books-
2018/AEA2330E1B37231570A06CC349FAC7AA.

Holton:2011:PSC

[Hol11] Derek Allan Holton. Problem solving: the creative side of mathematics. Mathematical Association, London, UK, 2011. ISBN 0-906588-71-5, 0-906588-70-7. x + 211 pp. LCCN QA43 .H625 2010.

Hollings:2012:CCE

[Hol12a] Christopher Hollings. I, Claudius and the cipher extraordinary. The Mathematical Gazette, 96(537):466–470, November 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Holman:2012:HDT

[Hol12b] Andrew Holman. 96.56 How to drive too fast and not save time. The Mathematical Gazette, 96(537):526–528, November 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Holman:2013:HPO

[Hol13a] Andrew Holman. How to pull an object effectively. The Mathematical Gazette, 97(539):330–333, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Holton:2013:MPS

[Hol13b] Derek Allan Holton. More problem solving: the creative side of mathematics. The Mathematical Association, Leicester, UK, 2013. ISBN 0-906588-78-2 (paperback). viii + 256 pp. LCCN QA43 .H65 2013.

Hollings:2017:RBM

[Hol17] Christopher D. Hollings. R. S. Ball’s mechanics: bringing Newton to the masses? The Mathematical Gazette, 101(551):280–288, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
C2243F6768B0577BB1932C1A711B4848.

Hollings:2019:NC

[Hol19] Christopher D. Hollings. The need for closure. The Mathematical Gazette, 103(557):248–256, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/need-for-closure/CFBF03BD79FC9F97D8D5334DCFC64784.

Hopkins:2012:HMW

[Hop12] David Hopkins. How many winners will there be? The Mathematical Gazette, 96(537):439–450, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hopkins:2013:SRN

[Hop13] David Hopkins. Slide rules, nomograms and additive functions. The Mathematical Gazette, 97(538):13–26, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hopkins:2016:WMN

[Hop16] David Hopkins. Will my numbers add up correctly if I round them? The Mathematical Gazette, 100(549):396–409, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
product/88F5753DFE9F0DDDEAD1F2552B0F8B22. The proba-

bility that rounding after fixed-point summation of n terms gives the same result as summation of rounded terms is given by p(n) = (2/π) ∫_{0}^{∞}(sin(x)/x)^{n+1} dx, and that function is always a rational number. Its values are p(n) = 1, 3/4, 2/3, 115/192, 11/20, 5887/11520, 151/315, 259723/573440, … for n = 1 to 8.

Hopkins:2018:BRC

[Hop18a] David Hopkins. Book review: Concise calculus by Sheng Gong, pp. 674, £48 (paper), ISBN 978-9-81429-149-1, World Scientific Publishing Co (2017). The Mathematical Gazette, 102(554):371–372, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/concise-calculus-by-sheng-gong-pp-674-48-paper-
isbn-9789814291491-world-scientific-publishing-co-2017/
69684EAEBFDBE8C12FE50AE6A3F13946.

Hopkins:2018:BRI

[Hop18b] David Hopkins. Book review: Infinity: a very short introduction by Ian Stewart, pp. 144, £7.99 (paper), ISBN 978-0-19-875523-4, Oxford University Press (2017). The Mathematical Gazette, 102(553): 191, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/infinity-a-
very-short-introduction-by-ian-stewart-pp-144-799-paper-
isbn-9780198755234-oxford-university-press-2017/CE1F4288E6E0D00206F5C9FA4559995B.

Hopkins:2019:CDM

[Hop19] David Hopkins. Book review: Connections in discrete mathematics: a celebration of the work of Ron Graham edited by Steve Butler, Joshua Cooper and Glenn Hurlbert, pp. 352, £44.99 (paperback), ISBN 978-1-316-60788-6, Cambridge University Press (2018). The Mathematical Gazette, 103(557):374–375, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/connections-in-discrete-mathematics-
a-celebration-of-the-work-of-ron-graham-edited-by-steve-
butler-joshua-cooper-and-glenn-hurlbert-pp-352-4499-paperback-
isbn-9781316607886-cambridge-university-press-2018/48D706E31AF7579308F4317F17ABF2CB.

Horsman:2014:PCS

[Hor14] Rachel Horsman. Padlock challenges: statistics and probability. The Mathematical Association, Leicester, UK, 2014. ISBN 0-906588-80-4 (spiral). 72 pp. LCCN ????

Hornby:2017:BRA

[Hor17] Will Hornby. Book review: All is change by J. F. Riley, pp. 454, £18 (hard), ISBN 978-1-909-98664-2, Bookguild Publishing (2014). The Mathematical Gazette, 101(552): 556, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/all-is-
change-by-j-f-riley-pp-454-18-hard-isbn-9781909986642-
bookguild-publishing-2014/053E8246623D07CE5B645A61B9483B19.

Hoseana:2017:ESM

[Hos17a] Jonathan Hoseana. Extending the substitution method for integration. The Mathematical Gazette, 101(552):538–541, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/extending-the-substitution-
method-for-integration/B6A2BA929EB33549765B12D602EA36CC.

Hoseana:2017:IBP

[Hos17b] Jonathan Hoseana. Integration based on the parametrisation of conics. The Mathematical Gazette, 101(551):261–267, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
5B99213B3E7F3902D13439288C5ED1D0.

Houston:2009:HTL

[Hou09] Kevin Houston. How to think like a mathematician: a companion to undergraduate mathematics. Cambridge University Press, Cambridge, UK, 2009. ISBN 0-521-89546-4 (hardcover), 0-521-71978-X (paperback), 0-511-50431-4 (e-book), 0-511-80825-9 (ebook), 1-139-63778-9, 1-139-12971-6 (e-book). xi + 265 pp. LCCN

QA13 .H68 2009. URL http://catdir.loc.gov/catdir/toc/
ecip0825/2008034663.html.

Howson:2015:DAQ

[How15] Geoffrey Howson. Douglas Arthur Quadling 1926–2015. The Mathematical Gazette, 99(545):193–195, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Howson:2016:ART

[How16] Geoffrey Howson. Alan Richard Tammadge 9 July 1921–25 Feb 2016. The Mathematical Gazette, 100(549):490–491, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
688D9E76DB759F0D7FC71E26ECFBBB02.

Howson:2019:SMF

[How19] Geoffrey Howson. Sir Michael Francis Atiyah, O.M., 22 April 1929–11 January 2019. The Mathematical Gazette, 103(557):193–195, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/sir-michael-francis-atiyah-om-22-april-1929-11-
january-2019/A3384965D738C8363A8496B19FEFDD32.

Hilton:2010:SST

[HP10a] Peter Hilton and Jean Pedersen. Stop-sign theorems and binomial coefficients. The Mathematical Gazette, 94(530):247–261, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Holden:2010:APF

[HP10b] H. (Helge) Holden and Ragni Piene, editors. The Abel Prize: 2003– 2007: the first five years. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-01372-4 (hardcover). xi + 327 pp. LCCN QA20.3 .A24 2010.

Hajja:2017:NLI

[HR17] Mowaffaq Hajja and Edris Rawashdeh. 101.47 notes on Lagrange’s interpolation. The Mathematical Gazette, 101 (552):528–533, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10147-notes-on-lagranges-interpolation/06B5E42F7FE48C607F895629A9AA1398.

Herrmann:2013:NSS

[HS13] Diane Herrmann and Paul J. Sally, Jr. Number, Shape, and Symmetry: an Introduction to Number Theory, Geometry, and Group Theory. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2013. ISBN 1-4665-5464-9 (hardcover). xvii + 426 pp. LCCN QA241 .H44 2013.

Holm:2009:GMS

[HTSE09] Darryl D. Holm, Schmah Tanya, Cristina Stoica, and David C. P. Ellis. Geometric mechanics and symmetry: from finite to infinite dimensions, volume 12 of Oxford Texts in Applied and Engineering Mathematics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-921291-0 (paperback), 0-19-921290-2 (hardcover), 0-19-154986-X (e-book). xvi + 515 + 4 pp. LCCN QA807.5 .H65 2009.

Hudson:2010:BRBc

[Hud10a] Paul Hudson. Book review: Finite mathematics, models and structure (revised edition) by William J. Adams. Pp. 619. $23.99 (pbk). 2009. ISBN: 978-1-4363-3416-7 (Xlibis). The Mathematical Gazette, 94(531):568–569, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hudson:2010:BRBb

[Hud10b] Paul Hudson. Book review: Picturing the uncertain world, by Howard Wainer. Pp. 244. £17.95 (hbk). 2009. ISBN: 978-0-691-13759-9 (Princeton University Press). The Mathematical Gazette, 94(531):559–560, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hudson:2010:BRBa

[Hud10c] Paul Hudson. Book review: The presidential election game, by Steven J. Brams. pp. 194. $29. 2008. ISBN 978-1-56881-348-6 (A. K. Peters). The Mathematical Gazette, 94(530):362–363, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hudson:2011:BRBa

[Hud11a] Paul Hudson. Book review: Mathematical people: profiles and interviews (2nd edn.), edited by Donald J. Albers and Gerald L. Alexanderson. Pp. 385. $59-00 (hardback). 2008. ISBN: 978-1-56881-340-0 (A. K. Peters). The Mathematical Gazette, 95 (532):147–149, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hudson:2011:BRBb

[Hud11b] Paul Hudson. Book review: Meta maths: the quest for omega, by Gregory Chaitin. pp. 220. £9.99. 2007. ISBN: 978-1-84354-525-5 (Atlantic Books). The Mathematical Gazette, 95(533):377–379, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hudson:2013:BRBa

[Hud13a] Paul Hudson. Book review: How to read historical mathematics, by Benjamin Wardhaugh, Pp. 116. £15.95.(hbk) 2010. ISBN: 978-0-691-14014-8 (Princeton University Press). The Mathematical Gazette, 97(539):358–359, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hudson:2013:BRBb

[Hud13b] Paul Hudson. Book review: Numbers rule: the vexing mathematics of democracy from Plato to the present, by George G. Szpiro. Pp. 226 £18.95 (Hardback). 2010. ISBN: 9780-691-13994-4 (Princeton University Press). The Mathematical Gazette, 97(540):559–560, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hunacek:2010:BRBb

[Hun10a] Mark Hunacek. Book review: Matrix computations and semiseparable matrices (Vol I: Linear Systems) Raf Vandebril, Marc Van Barel and Nicola Mastronardi. Pp. 575. £50. 2008. ISBN: 978-0-8018-8714-7 (Johns Hopkins University Press). The Mathematical Gazette, 94(531):565–566, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hunacek:2010:BRBa

[Hun10b] Mark Hunacek. Book review: What is a number? Mathematical concepts and their origins, by Robert Tubbs. Pp. 305. £15-00. 2009. ISBN: 978-0-8018-9018-5 (Johns Hopkins University Press). The Mathematical Gazette, 94(531):560–561, November 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hunacek:2012:BRBb

[Hun12a] Mark Hunacek. Book review: Mrs Perkins’s electric quilt: and other intriguing stories of mathematical physics, by Paul Nahin. Pp. 391. £20.95 (Hardback). 2009. ISBN: 978-0-961-13540-3 (Princeton University Press). The Mathematical Gazette, 96 (536):378–380, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hunacek:2012:BRBa

[Hun12b] Mark Hunacek. Book review: Symmetry, representations and invariants, by Roe Goodman and Nolan Wallach. Pp. 716. £53.99 (Hardback). 2009. ISBN: 978-0-387-79851-6 (Springer). The Mathematical Gazette, 96(536):372–373, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hunacek:2013:BRBd

[Hun13a] Mark Hunacek. Book review: A first course in applied mathematics, by Jorge Rebaza. Pp. 439. £76.95. (hbk) 2012. ISBN: 978-1-118-22962-0 (Wiley). The Mathematical Gazette, 97(540):570– 571, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hunacek:2013:BRBc

[Hun13b] Mark Hunacek. Book review: A prime puzzle, by Martin Griffiths. Pp. 217. £12.00. 2012. ISBN: 978-1-906001-16-2 (The United Kingdom Mathematics Trust). The Mathematical Gazette, 97(540):

568–569, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hunacek:2013:BRBb

[Hun13c] Mark Hunacek. Book review: Differential geometry of curves and surfaces, by Thomas Banchoff and Stephen Lovett. Pp. 331. $49.00 (Hardback). 2010. ISBN: 978-1-56881-456-8 (A. K. Peters). The Mathematical Gazette, 97(539):370–371, July 2013. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hunacek:2013:BRBa

[Hun13d] Mark Hunacek. Book review: Loving + hating mathematics: challenging the myths of mathematical life, by Reuben Hersh and Vera John-Steiner. Pp. 416. $29.95 (Hardback). 2011 ISBN: 978-0-691-14247-0 (Princeton University Press). The Mathematical Gazette, 97(539):364–367, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hunacek:2014:BRBb

[Hun14a] Mark Hunacek. Book review: Introduction to representation theory, by Pavel Etingof, Oleg Golberg, Sebastian Hensel, Tiankai Liu, Alex Schwendner, Dmitry Vaintrob, Elena Yudinova and Siava Gerovitch. Pp. 228, $42.00, 2011. ISBN: 978-0-8218-4999-6 (American Mathematical Society). The Mathematical Gazette, 98 (541):175–176, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hunacek:2014:BRBa

[Hun14b] Mark Hunacek. Book review: p-adic Lie groups, by Peter Schneider. Pp. 254. £72.00 (hbk). 2011. ISBN: 978-3-642-21146-1 (Springer). The Mathematical Gazette, 98(541):165–166, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hunacek:2015:BRB

[Hun15] Mark Hunacek. Book review: Cyclic modules and the structure of rings by S. K. Jain, Ashish K. Srivastava and Askar A. Tuganbaev, pp 232, £68 (hbk), ISBN 978-0-19666-541-1, Oxford University Press (2012). The Mathematical Gazette, 99(544):182–183, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hunacek:2016:BRA

[Hun16] Mark Hunacek. Book review: The Abel Prize 2008–2012, edited by Helge Holden & Ragni Piene, pp. 572, £59.99 (hard), ISBN 978-3-642-39448-5, also available as e-book, Springer Verlag (2014). The Mathematical Gazette, 100(547):176–178, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hunacek:2017:BRC

[Hun17a] Mark Hunacek. Book review: A course in analysis, by Jacob, Niels and Evans, Kristian P., pp. 744, £43 (paper), ISBN 987-9814-68909-0, World Scientific Publishing Co (2015). The Mathematical Gazette, 101(550):175–177, March 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
E61B20511AB328B5DD1395E7B49E0E3B.

Hunacek:2017:BRL

[Hun17b] Mark Hunacek. Book review: Lie groups, Lie algebras and representations, by Hall, Brian C., pp. 354, £35.99 (hard), ISBN: 978-1-4419-2313-4, also available as e-book, Springer Verlag (2015). The Mathematical Gazette, 101(551):369–371, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
72F49A80A84AD5C020A418C73330DDFF.

Hunacek:2017:BRMa

[Hun17c] Mark Hunacek. Book review: Mathematical concepts, by Jost, Jürgen, pp. 312, £53.99 (paper), ISBN 978-3-319-20435-2, also available as e-book, Springer Verlag (2016). The Mathematical Gazette, 101(551):365–366, July 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
FD3EB69F2876DCF17A4D7664B70214A0.

Hunacek:2017:BRMb

[Hun17d] Mark Hunacek. Book review: Matrix groups for undergraduates (2nd edn.), by Tapp, Kristopher, pp. 239, $39.20 (paper), American Mathematical Society (2016). The Mathematical Gazette, 101 (551):368–369, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/0C3D41E63498DA7A51F414BAEAB041B4.

Hunacek:2017:BRP

[Hun17e] Mark Hunacek. Book review: The Princeton companion to applied mathematics, edited by Higham, Nicholas J., pp. 1016, £69.95 (hard), ISBN 978-0-691-15039-0, Princeton University Press (2015). The Mathematical Gazette, 101(550):169–170, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
9E2536735F089D0B64F8573DB8632167.

Huntley:2017:BRC

[Hun17f] Ray Huntley. Book review: Curricular resources for classroom use: the case of mathematics, by Stylianides, Gabriel J., pp. 176, £55.00 (hard), ISBN 978-0-19874-989-9, Oxford University Press (2016). The Mathematical Gazette, 101(551):350, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
E9A2281F09520DF8A3C84C4766CA8703.

Hunacek:2018:BRI

[Hun18a] Mark Hunacek. Book review: An invitation to applied mathematics by Carmen Chicone, pp. 856, £67.15 (hard), ISBN 978-0-128-04153-6, Academic Press (Elsevir) (2016). The Mathematical Gazette, 102(555):571–572, November 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/an-invitation-to-applied-
mathematics-by-carmen-chicone-pp-856-6715-hard-isbn-9780128041536-
academic-press-elsevir-2016/11E4FDBEE65C5A6E893D7ECB651F24A6.

Hunacek:2018:BRC

[Hun18b] Mark Hunacek. Book review: Classical groups, derangements and primes by Timothy C. Burness and Michael Giudici, £50.00, ISBN 978-1-10762-944-8, Cambridge University Press (2016). The Mathematical Gazette, 102(554):368–369, July 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/classical-groups-derangements-
and-primes-by-timothy-c-burness-and-michael-giudici-5000-
isbn-9781107629448-cambridge-university-press-2016/632406C929FE293971F49DE9247ED286.

Hunacek:2018:BRO

[Hun18c] Mark Hunacek. Book review: Office hours with a geometric group theorist edited by Matt Clay and Dan Margalit, pp.

441, £45.95 (paper), ISBN 978-0-69115-866-2, Princeton University Press (2017). The Mathematical Gazette, 102(554): 369–371, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/office-
hours-with-a-geometric-group-theorist-edited-by-matt-clay-
and-dan-margalit-pp-441-4595-paper-isbn-9780691158662-
princeton-university-press-2017/96DD455A27F589B4FF35183E1254A65F.

Hunacek:2019:MPP

[Hun19a] Mark Hunacek. Book review: Mathematical physics with partial differential equations by James Kirkwood, pp. 478, £61.99 (paper), ISBN 978-0-12814-759-7, Academic Press (2018). The Mathematical Gazette, 103(557):379–381, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/mathematical-physics-with-partial-differential-
equations-by-james-kirkwood-pp-478-6199-paper-isbn-9780128147597-
academic-press-2018/A9810D17C4C4F27299EF2665554B8ECF.

Hunacek:2019:CGW

[Hun19b] Mark Hunacek. Book review: The calculus gallery by William Dunham, pp. 236, £14.99 (paper), ISBN 978-0-69118-285-8, Princeton University Press (2018). The Mathematical Gazette, 103(557):372–374, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/calculus-gallery-by-william-dunham-pp-236-1499-
paper-isbn-9780691182858-princeton-university-press-2018/
354973924B2BFFAAA9A7E8A59D0B2513.

Hvala:2011:DST

[Hva11] Bojan Hvala. Diophantine Steiner triples. The Mathematical Gazette, 95(532):31–39, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Hardy:2008:ITN

[HWHBS08] Godfrey Harold Hardy, Edward Maitland Wright, D. R. Heath-Brown, and Joseph H. Silverman. An introduction to the theory of numbers. Oxford mathematics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, sixth edition, 2008. ISBN 0-19-921985-0 (hardcover), 0-19-921986-9 (paperback). xxi + 1 + 621 pp. LCCN QA241 .H28 2008.

Imrie:2010:PTM

[Imr10] Jane Imrie. Progressing through mathematics. The Mathematical Gazette, 94(531):386–400, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ibadula:2012:CPS

[IS12] Denis Ibadula and Dirk Segers. 96.55 Convex proofs of some inequalities. The Mathematical Gazette, 96(537):523–526, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ivanov:2009:MMC

[Iva09] O. A. (Oleg A.) Ivanov. Making mathematics come to life: a guide for teachers and students. American Mathematical Society, Providence, RI, USA, 2009. ISBN 0-8218-4808-9. x + 337 pp. LCCN QA39.3 .I9313 2009. Translated by Robert Burns.

Jackson:2011:ISR

[Jac11] Terence Jackson. 95.42 Irrational square roots of natural numbers — a geometrical approach. The Mathematical Gazette, 95(533): 327–330, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jackson:2012:MIH

[Jac12] Tom Jackson, editor. Mathematics: An Illustrated History of Numbers. Ponderables. Worth Press Ltd, Cambridge, UK, 2012. ISBN 1-84931-066-1 (hardcover). 144 pp. LCCN QA21 .M36118 2012. Contributors: Richard Beatty, James Bow, Mike Goldsmith, Dan Green, Tom Jackson, Robert Sneddon, and Susan Watt.

Jacob:2014:BRB

[Jac14] Niels Jacob. Book review: Elliptic partial differential equations, volume 1: Fredholm theory of elliptic problems in unbounded domains, by Vitali Volpert. Pp. 639. £90.00 (hbk). 2011. ISBN: 978-3-0346-0536-6 (Birkhauser). The Mathematical Gazette, 98 (541):172–174, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jacoby:2019:SRC

[Jac19] Michael Jacoby. Book review: Special relativity and classical field theory by Leonard Susskind and Art Friedman, pp. 448, $30.00 (paper), ISBN 978-0-46509-334-2, Basic Books (2017). The Mathematical Gazette, 103(557):378–379, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/special-relativity-and-classical-field-theory-
by-leonard-susskind-and-art-friedman-pp-448-3000-paper-
isbn-9780465093342-basic-books-2017/286A35129626D754FEBEC71DA8300BCA.

Jameson:2010:TSF

[Jam10a] G. J. O. Jameson. 94.03 Two squares and four squares: the simplest proof of all? The Mathematical Gazette, 94(529):119–123, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2010:EOS

[Jam10b] G. J. O. Jameson. 94.04 Even and odd square-free numbers. The Mathematical Gazette, 94(529):123–127, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2011:FCN

[Jam11a] G. J. O. Jameson. Finding Carmichael numbers. The Mathematical Gazette, 95(533):244–255, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2011:FP

[Jam11b] G. J. O. Jameson. Finding pseudoprimes. The Mathematical Gazette, 95(534):420–432, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2012:EIT

[Jam12] G. J. O. Jameson. 96.18 Euler, Ioachimescu and the trapezium rule. The Mathematical Gazette, 96(535):136–142, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

James:2013:TQF

[Jam13a] Peter James. Teaching quadratic functions. The Mathematical Gazette, 97(539):335, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2013:CFD

[Jam13b] G. J. O. Jameson. 97.10 Continuous functions that are not differentiable anywhere. The Mathematical Gazette, 97(538):137–143, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2013:APD

[Jam13c] Tim Jameson. 97.44 Another proof that ζ(2) = π^{2}/6 via double integration. The Mathematical Gazette, 97(540):506–507, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2014:AAG

[Jam14a] G. J. O. Jameson. An approximation to the arithmetic–geometric mean. The Mathematical Gazette, 98(541):85–95, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2014:FLE

[Jam14b] G. J. O. Jameson. A fresh look at Euler’s limit formula for the gamma function. The Mathematical Gazette, 98(542):235–242, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2014:IPE

[Jam14c] G. J. O. Jameson. Inequalities for the perimeter of an ellipse. The Mathematical Gazette, 98(542):227–234, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2014:SI

[Jam14d] G. J. O. Jameson. Series involving ζ(n). The Mathematical Gazette, 98(541):58–66, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2014:SIB

[Jam14e] G. J. O. Jameson. Some inequalities for (a+b)^{p} and (a+b)^{p} +(a− b)^{p}. The Mathematical Gazette, 98(541):96–103, March 2014. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2014:SDS

[Jam14f] Tim Jameson. 98.13 Some double series related to ζ(3). The Mathematical Gazette, 98(542):327–331, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2014:OME

[Jam14g] Tim Jameson. Orbits made easy by complex numbers. The Mathematical Gazette, 98(542):353–355, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2015:CD

[Jam15a] G. J. O. Jameson. Counting divisors. The Mathematical Gazette, 99(544):11–20, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2015:EMH

[Jam15b] G. J. O. Jameson. Euler–Maclaurin, harmonic sums and Stirling’s formula. The Mathematical Gazette, 99(544):75–89, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2015:EFT

[Jam15c] G. J. O. Jameson. Evaluating Fresnel-type integrals. The Mathematical Gazette, 99(546):491–498, November 2015. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2015:SPS

[Jam15d] G. J. O. Jameson. A simple proof of Stirling’s formula for the gamma function. The Mathematical Gazette, 99(544):68–74, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2015:SCE

[Jam15e] G. J. O. Jameson. Sine, cosine and exponential integrals. The Mathematical Gazette, 99(545):276–289, July 2015. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2016:IGF

[Jam16a] G. J. O. Jameson. The incomplete gamma functions. The Mathematical Gazette, 100(548):298–306, July 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
9373A31AD28D793AB5431E35EA5C2AF6.

Jameson:2016:UDI

[Jam16b] G. J. O. Jameson. Using double integrals to solve single integrals. The Mathematical Gazette, 100(548):257–265, July 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
FCE02B0C6C9E982124D83C968F7C17C8.

Jameson:2017:EHS

[Jam17] G. J. O. Jameson. Expressing harmonic sums as fractions. The Mathematical Gazette, 101(551):236–249, July 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
D77D362425837BB88538EE2C7FEF45D7.

Jameson:2018:NSP

[Jam18a] G. J. O. Jameson. 102.28 a note on series of positive and negative terms. The Mathematical Gazette, 102(554):320–325, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10228-a-note-on-
series-of-positive-and-negative-terms/FD291E7E568250092F7410CAD023ACC2.

Jameson:2018:RNA

[Jam18b] G. J. O. Jameson. 102.45 revisiting a note on arithmetic and geometric means. The Mathematical Gazette, 102(555):513, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10245-revisiting-
a-note-on-arithmetic-and-geometric-means/7D7E5F6E956573FB16FCB8998A9E3193.

Jameson:2018:BRT

[Jam18c] G. J. O. Jameson. Beyond the ratio test. The Mathematical Gazette, 102(555):471–484, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/beyond-the-ratio-test/D10BF58DD4926AB18878FA6E088A3A65.

Jameson:2018:FPM

[Jam18d] G. J. O. Jameson. Fibonacci periods and multiples. The Mathematical Gazette, 102(553):63–76, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/fibonacci-periods-and-multiples/1CD89C358C7AD0CE86005E0BF340F766.

Jameson:2019:MGB

[Jam19a] G. J. O. Jameson. 103.32 More on the gaps between sums of two squares. The Mathematical Gazette, 103(558):499–503, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10332-more-
on-the-gaps-between-sums-of-two-squares/B95F219B9B1C22CB480D89672395AA1D.

Jameson:2019:DBN

[Jam19b] G. J. O. Jameson. 103.44 The denominators of the Bernoulli numbers. The Mathematical Gazette, 103(558):541–548, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10344-the-denominators-
of-the-bernoulli-numbers/CE9B6C5E662F7BEB6E4534A3C7D94589.

Jarai:2016:DNG

[Jár16] Antal A. Járai. The dice-and-numbers game. The Mathematical Gazette, 100(549):410–419, November 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
11CEEC49400B97CF8D2CE037808CF95A.

Jefferson:2008:DD

[Jef08a] Brian Jefferson. Decision: D1. A Level mathematics for Edexcel. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-911780-2 (paperback). ???? pp. LCCN ????

Jefferson:2008:MM

[Jef08b] Brian Jefferson. Mechanics: M1. A Level mathematics for Edexcel. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-911781-0 (paperback). ???? pp. LCCN ????

Jewess:2010:OAD

[Jew10] Michael Jewess. Optimising the acceleration due to gravity on a planet’s surface. The Mathematical Gazette, 94(530):203–215, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Joyner:2012:IDE

[JH12] David Joyner and Marshall Hampton. Introduction to differential equations using Sage. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0637-3 (hardcover), 1-4214-0724-8 (e-book). xiii + 3 + 264 pp. LCCN QA371

.J75 2012. URL http://www.loc.gov/catdir/enhancements/
fy1305/2012934439-b.html;   http://www.loc.gov/catdir/
enhancements/fy1305/2012934439-d.html.

Jameson:2013:FMT

[JJ13a] Graham J. O. Jameson and Timothy P. Jameson. 97.07 Four methods for a trigonometric integral. The Mathematical Gazette, 97(538):127–131, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2013:SRI

[JJ13b] Graham J. O. Jameson and Timothy P. Jameson. Some remarkable integrals derived from a simple algebraic identity. The Mathematical Gazette, 97(539):205–209, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Johnson:2014:LN

[JJ14] Tom Johnson and Franck Jedrzejewski. Looking at numbers. Birkhäuser/Springer, Basel, Switzerland, 2014. ISBN 3-0348-0553-5 (paperback), 3-0348-0554-3 (e-book). xvii + 120 pp. LCCN QA141 .J59 2014.

Jameson:2017:GB

[JJ17] Graham Jameson and Nicholas Jameson. 101.17 goats and birds. The Mathematical Gazette, 101(551):296–300, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
8679C58D2644CFC8BDA53870E846619A.

Jameson:2018:C

[JJ18] Graham Jameson and Pauline Jameson. Correspondence. The Mathematical Gazette, 102(553):171, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/correspondence/7BC5879A74240C83C8D8340AF586DD31.

Jameson:2013:EDI

[JL13] G. J. O. Jameson and Nick Lord. 97.43 Evaluation of ∑_{n=1}^{∞} 1/n^{2} by a double integral. The Mathematical Gazette, 97(540):504–505, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2017:IET

[JL17] Graham Jameson and Nick Lord. Integrals evaluated in terms of Catalan’s constant. The Mathematical Gazette, 101(550): 38–49, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/6AAE897C262F3FB12A081BCEB9A3D0B1.

Jee:2015:CCS

[JLL15] Dong Jun Jee, Tuo Yeong Lee, and Nathaniel Zhi-Wei Leon. 99.16 Conditionally convergent series. The Mathematical Gazette, 99 (545):328–331, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jameson:2015:I

[JLM15] Graham Jameson, Nick Lord, and James McKee. 99.05 An inequality for Si(x). The Mathematical Gazette, 99(544):133–139, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Junaidu:2010:PIS

[JLU10] S. B. Junaidu, A. Laradji, and A. Umar. 94.02 Powers of integers as sums of consecutive odd numbers. The Mathematical Gazette, 94(529):117–119, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Johnston:2009:SAM

[JM09a] William Johnston and Alex M. McAllister. A survey of advanced mathematics: a transition course. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-531076-4 (hardcover). xx + 745 pp. LCCN QA37.3 .J65 2009.

Johnston:2009:TAM

[JM09b] William Johnston and Alex M. McAllister. A transition to advanced mathematics: a survey course. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-531076-4 (hardcover). xx + 745 pp. LCCN QA37.3 .J65 2009.

Jacob:2014:NAF

[JO14] Walter Jacob and Thomas J. Osler. 98.35 a novel approach to finding ∫ csc x dx. The Mathematical Gazette, 98(543):518–519, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jobbings:2010:SFD

[Job10] Andrew Jobbings. 94.24 Sudoku is four-dimensional. The Mathematical Gazette, 94(530):321–326, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jobbings:2011:PDD

[Job11a] Andrew K. Jobbings. 95.15 Proofs by dissection of a dodecagon. The Mathematical Gazette, 95(532):107–110, March 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jobbings:2011:TSF

[Job11b] Andrew K. Jobbings. 95.64 Two semicircles fill half a circle. The Mathematical Gazette, 95(534):538–540, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jobbings:2013:SSS

[Job13a] Andrew Jobbings. 97.30 Shapes with a skew square as boundary. The Mathematical Gazette, 97(539):297–301, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jobbings:2013:FT

[Job13b] Andrew Jobbings. Folding a triangle. The Mathematical Gazette, 97(539):210–217, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jobbings:2013:PSH

[Job13c] Andrew Jobbings. A problem solver’s handbook: a guide to intermediate mathematical olympiads. The United Kingdom Mathematics Trust, Leeds, UK, 2013. ISBN 1-906001-19-7. LCCN QA43 .J63 2013.

Jobbings:2015:WCP

[Job15] Andrew Jobbings. Where is the centre of a polygon? The Mathematical Gazette, 99(544):109–120, March 2015. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jobbings:2016:GA

[Job16] Andrew Jobbings. 100.21 generalising the arbelos. The Mathematical Gazette, 100(548):329–335, July 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
3A9CD589247E5F24C0E2DCBD9B03F067.

Jobbings:2018:FCR

[Job18a] Andrew Jobbings. 102.11 four congruent right triangles. The Mathematical Gazette, 102(553):129–131, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10211-four-congruent-right-
triangles/6ECB589323302EC5CA096E0EC993DCED.

Jobbings:2018:PFA

[Job18b] Andrew Jobbings. Proofs by folding of the angle sum of a triangle. The Mathematical Gazette, 102(553):153–154, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/proofs-by-folding-of-the-
angle-sum-of-a-triangle/CDC4546D740D40632419D59FBCA01626.

Johnson:2011:NC

[Joh11] C. Johnson. 95.26 Nice cubics. The Mathematical Gazette, 95 (533):273–279, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Johnson:2012:TS

[Joh12] Roger W. Johnson. ‘toss up’ strategies. The Mathematical Gazette, 96(537):415–419, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Jones:2019:BRL

[Jon19] Andy Ashworth Jones. Book review: Leadership to count on, a guide for heads of mathematics (2nd edn.) by Robin Bevan, Paul Metcalf and Chris Pritchard, pp. 144, £14.00 (paper), ISBN 978-0-906588-97-0, The Mathematical Association (2018). The Mathematical Gazette, 103(558):570, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/leadership-to-count-on-a-
guide-for-heads-of-mathematics-2nd-edn-by-robin-bevan-
paul-metcalf-and-chris-pritchard-pp-144-1400-paper-isbn-
9780906588970-the-mathematical-association-2018/A6D227EAD90008A97D849E9BDB21F5B0.

Josefsson:2015:MAB

[Jos15] Martin Josefsson. Minimal area of a bicentric quadrilateral. The Mathematical Gazette, 99(545):237–242, July 2015. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Josefsson:2016:GAT

[Jos16a] Martin Josefsson. 100.30 generalisations of the altitude theorem. The Mathematical Gazette, 100(549):501–505, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
6718C3C63174E599396E35612AC002CA.

Josefsson:2016:HLF

[Jos16b] Martin Josefsson. 100.31 Heron-like formulas for quadrilaterals. The Mathematical Gazette, 100(549):505–508, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
9512A31CCF6D7DE95933CD573C1CE143.

Josefsson:2016:CCQ

[Jos16c] Martin Josefsson. On the classification of convex quadrilaterals. The Mathematical Gazette, 100(547):68–85, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Josefsson:2016:PPQ

[Jos16d] Martin Josefsson. Properties of Pythagorean quadrilaterals. The Mathematical Gazette, 100(548):213–224, July 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
49F9DC20F0A8CDF796B1C235C700AC7F.

Josefsson:2017:ATQ

[Jos17a] Martin Josefsson. 101.36 on the area of a tangential quadrilateral. The Mathematical Gazette, 101(552):495–497, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
B99433E9DE565D5AE7DD9FAEA88EDF15.

Josefsson:2017:MRC

[Jos17b] Martin Josefsson. 101.38 metric relations in crossed cyclic quadrilaterals. The Mathematical Gazette, 101(552):499–502, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10138-metric-relations-
in-crossed-cyclic-quadrilaterals/E8CAFBB7D7BB5E4196E474C925D1847F.

Josefsson:2017:FCT

[Jos17c] Martin Josefsson. Further characterisations of tangential quadrilaterals. The Mathematical Gazette, 101(552):401–411, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
22EF975086DC754A4017A9FB8E1BEA8B.

Josefsson:2017:GQD

[Jos17d] Martin Josefsson. Generalisation of a quadrilateral duality theorem. The Mathematical Gazette, 101(551):208–213, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
1AF05F68FF2BE038A9811B2BE5A8EFA3.

Josefsson:2017:PBD

[Jos17e] Martin Josefsson. Properties of bisect-diagonal quadrilaterals. The Mathematical Gazette, 101(551):214–226, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
F2B01E594DDB0AB7A53A4D7BCFD492CF.

Josefsson:2018:ICG

[Jos18a] Martin Josefsson. The importance of characterisations in geometry. The Mathematical Gazette, 102(554):302–307, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/importance-of-characterisations-
in-geometry/F9EB4744755CAA1BDD191C431F091426.

Josefsson:2018:MRC

[Jos18b] Martin Josefsson. Metric relations in crossed trapezia. The Mathematical Gazette, 102(554):264–269, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/metric-relations-in-crossed-trapezia/3B4A7199D4F8EA4887A6355DDF1699F7.

Josefsson:2019:HCQ

[Jos19a] Martin Josefsson. 103.09 on the heights of a convex quadrilateral. The Mathematical Gazette, 103(556):145–150, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10309-on-the-heights-of-a-
convex-quadrilateral/49997FB052A78E35564EE2CA0DAFC3A5.

Josefsson:2019:PT

[Jos19b] Martin Josefsson. 103.23 On Pitot’s theorem. The Mathematical Gazette, 103(557):333–337, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10323-on-pitots-theorem/1249344B51FA1C9C6532E13A02A35E05.

Josefsson:2019:AFC

[Jos19c] Martin Josefsson. Area formulas for certain convex quadrilaterals. The Mathematical Gazette, 103(556):20–27, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/area-formulas-for-certain-
convex-quadrilaterals/2E8D3793052F913AFAE2C95B12477AD6.

Josefsson:2019:EFA

[Jos19d] Martin Josefsson. Eight formulae for the area of triangle OIH. The Mathematical Gazette, 103(558):471–479, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/eight-formulae-for-the-area-
of-triangle-oih/B357ED7FC376E64543DD406E2A1F570D.

Jordan:2008:MTI

[JS08] D. W. (Dominic William) Jordan and Peter Smith. Mathematical techniques: an introduction for the engineering, physical, and mathematical sciences. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, fourth edition, 2008. ISBN 0-19-928201-3. xx + 976 pp. LCCN QA300 .J82 2008. URL http:

//www.loc.gov/catdir/enhancements/fy0907/2008004046-b.
html;   http://www.loc.gov/catdir/enhancements/fy0907/
2008004046-d.html;   http://www.loc.gov/catdir/enhancements/
fy0907/2008004046-t.html.

Jain:2012:CMS

[JST12] S. K. Jain, Ashish Srivastava, and Askar A. Tuganbaev. Cyclic modules and the structure of rings. Oxford mathematical monographs. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-966451-X (hardcover). x + 220 pp. LCCN QA251.4 .J35 2012.

Kilic:2019:PCC

[KA19] Emrah Kiliç and Talha Arikan. 103.26 A proof of Clarke’s conjecture. The Mathematical Gazette, 103(557):346–352, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10326-a-proof-of-clarkes-
conjecture/1760231C7670B550FF3322A059FB12FE.

Kaczkowski:2017:SMO

[Kac17] Stephen Kaczkowski. Solving multivariate optimisation problems using inequalities. The Mathematical Gazette, 101(552): 412–423, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/B49C5ABCA0932F2A06A0FA350F463812.

Kaczkowski:2018:OLV

[Kac18] Stephen Kaczkowski. On obtaining limiting values of a class of infinite products. The Mathematical Gazette, 102 (555):428–434, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/on-obtaining-limiting-values-of-a-class-of-infinite-
products/A56E0921B5BBF52890B9C663796AB5F0.

Kaniuth:2009:CCB

[Kan09a] Eberhard Kaniuth. A course in commutative Banach algebras, volume 246 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-72475-3 (hardcover), 0-387-72476-1 (e). xii + 353 pp. LCCN QA326 .K29 2009. URL http://www.loc.gov/

catdir/toc/fy0905/2008937945.html.

Kannappan:2009:FEI

[Kan09b] Palaniappan Kannappan. Functional equations and inequalities with applications. Springer monographs in Mathematics. Springer Science, Dordrecht, Germany, 2009. ISBN 0-387-89491-8 (hardcover), 0-387-89492-6 (e-book). xxiii + 810 pp. LCCN QA431 .K36

2009. URL http://www.gbv.de/dms/ilmenau/toc/592047296.
PDF;   http://www.springerlink.com/openurl.asp?genre=book%
26isbn=978-0-387-89491-1.

Kang:2017:BSS

[Kan17] Jeehoon Kang. Brilliance or steadiness? A suggestion of an alternative model to Hardy’s model concerning golf (1945). The Mathematical Gazette, 101(551):250–260, July 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
F9C01991F888131AED9FDB6FAFD75460.

Karamzadeh:2018:VES

[Kar18] O. A. S. Karamzadeh. 102.38 a very elementary short proof of Conway’s little theorem. The Mathematical Gazette, 102 (555):496–497, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10238-a-very-elementary-short-proof-of-conways-
little-theorem/DFD3D529F046D1EC34C6DDFDE170346D.

Keedwell:2011:GTM

[Kee11] A. D. Keedwell. Gaston Tarry and multimagic squares. The Mathematical Gazette, 95(534):454–468, November 2011. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Keedwell:2013:PSD

[Kee13] A. Donald Keedwell. 97.35 A problem of sociable dining. The Mathematical Gazette, 97(539):311–315, July 2013. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Keedwell:2018:BRT

[Kee18] Donald Keedwell. Book review: Taking Sudoku seriously by Jason Rosenhouse and Louise Taalman, pp. 214, £18.49, ISBN 978-0-19-975656-8, Oxford University Press (2011). The Mathematical Gazette, 102(553):186–187, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/taking-sudoku-seriously-by-jason-rosenhouse-and-
louise-taalman-pp-214-1849-isbn-9780199756568-oxford-university-
press-2011/0550D5EA59F70CA6E01A67617566FE1A.

Kemper:2011:CCA

[Kem11] Gregor Kemper. A course in commutative algebra, volume 256 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-03544-2 (hardcover). ISSN 0072-5285. xi + 246 pp. LCCN

QA251.3 .K456 2011. URL http://d-nb.info/996813195/04.

Kennedy:2014:IGC

[Ken14] Juliette Kennedy, editor. Interpreting Gödel: critical essays. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-00266-4 (hardcover). xi + 279 pp. LCCN QA9.2 .I58

2014. URL http://assets.cambridge.org/97811070/02661/
cover/9781107002661.jpg.

Kilner:2019:CP

[KF19] Steven J. Kilner and David L. Farnsworth. Characterisations of the parabola. The Mathematical Gazette, 103 (558):416–430, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/characterisations-of-the-parabola/43D383EC4B90CB5A7CB67916EF2782F2.

Koshy:2011:SDP

[KG11] Thomas Koshy and Zhenguang Gao. 95.13 Some divisibility properties of Catalan numbers. The Mathematical Gazette, 95(532): 96–102, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Khan:2010:RNE

[Kha10] Rasul A. Khan. 94.32 The recursive nature of Euler’s formula for harmonic series. The Mathematical Gazette, 94(531):488–492, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Khan:2012:PNW

[Kha12] Rasul A. Khan. The probabilistic nature of Wallis’s formula. The Mathematical Gazette, 96(536):193–200, July 2012. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Khan:2013:CRT

[Kha13] Rasul A. Khan. 97.26 The cosine rule in three dimensions and de Gua’s theorem. The Mathematical Gazette, 97(539):281–284, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Kilmister:2010:BRB

[Kil10] C. W. Kilmister. Book review: Exotic infinity and a long decline of confidence in mathematics, Christopher Ormell. pp. 55. 2008. ISBN 0-907669-46-8 (Ingleside-Ashby for the PER Group). The Mathematical Gazette, 94(530):381–382, July 2010. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

King:2010:RPI

[Kin10] Jeremy D. King. 94.34 Regular polygons with integer coordinates. The Mathematical Gazette, 94(531):495–498, November 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

King:2015:WDF

[Kin15] Jeremy D. King. 99.25 The work done by friction. The Mathematical Gazette, 99(545):352–356, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Kirfel:2013:GAM

[Kir13] Christoph Kirfel. A generalisation of Archimedes’ method. The Mathematical Gazette, 97(538):43–52, March 2013. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Kominers:2010:IBS

[KK10] Paul Myer Kominers and Scott Duke Kominers. 94.05 Improved bounds on the sizes of S·P numbers. The Mathematical Gazette, 94 (529):127–129, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Kaplan:2011:HHL

[KK11] Robert Kaplan and Ellen Kaplan. Hidden harmonies: the lives and times of the Pythagorean theorem. Bloomsbury Press, New York, NY, USA, 2011. ISBN 1-59691-522-6 (hardcover). xii + 290 pp. LCCN QA460.P8 K37 2011. URL http://catdir.loc.gov/

catdir/enhancements/fy1010/2010019959-b.html;   http:/
/catdir.loc.gov/catdir/enhancements/fy1010/2010019959-

d.html. Illustrations by Ellen Kaplan.

Kirfel:2017:IS

[KK17] Christoph Kirfel and Rainer Kaenders. 101.05 Integration by symmetry. The Mathematical Gazette, 101(550):99–103, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
D4AB623192CF8F668D5449738F602FD9.

Kirousis:2018:PMR

[KK18] Lefteris Kirousis and Georgios Kontogeorgiou. 102.18 the problème des ménages revisited. The Mathematical Gazette, 102(553):147–149, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10218-the-probleme-des-menages-revisited/C14D9ECE07E7FEEB04F4BDE1CC3C1BA2.

King:2011:MSC

[KL11] Jeremy D. King and Nick Lord. 95.68 The mechanics of sliding down curves. The Mathematical Gazette, 95(534):549–553, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

King:2014:RRE

[KL14] Jeremy D. King and Nick Lord. 98.17 The return of the returning explorer. The Mathematical Gazette, 98(542):339–342, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Kantrowitz:2019:LPC

[KN19]  Robert Kantrowitz and Michael M. Neumann.  Launch-

ing a projectile to cover maximal area. The Mathematical Gazette, 103(557):277–284, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/launching-a-projectile-to-cover-maximal-area/4A3BCE0935E78F3A91F6A13F1A0DCEA7.

Knudsen:2014:SJE

[Knu14a] Toke Lindegaard Knudsen. The Siddhantāsundara of Jñānarāja: an English translation with commentary. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1442-2 (hardcover), 1-4214-1443-0 (e-book). xx + 351

pp. LCCN BF1714.H5. URL http://muse.jhu.edu/books/
9781421414430/.

Jananaraja:2014:SJE

[Knu14b] Toke Lindegaard Knudsen, editor. The Siddhāntasundara of Jänānarāja: an English translation with commentary. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1442-2, 1-4214-1443-0 (e-book). xvii + 351 pp. LCCN BF1714.H5; QB18 J6313 2014. URL http://muse.jhu.edu/

books/9781421414430/.

Kominers:2010:IRR

[Kom10] Scorr Duke Kominers. 94.16 Irrational roots revisited. The Mathematical Gazette, 94(530):303–304, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Konig:2012:MIP

[Kön12] Heinz König. Measure and integration: publications 1997–2011. Birkhäuser, New York, NY, USA, 2012. ISBN 3-0348-0381-8 (hard cover). xi + 508 + 4 pp. LCCN QA312 .K5865 2012.

Koo:2011:GER

[Koo11] Reginald Koo. 95.03 Geometric enumeration of the rationals between any two rational numbers. The Mathematical Gazette, 95 (532):63–66, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Koo:2017:DLF

[Koo17] Reginald Koo. 101.14 the direction of limiting friction. The Mathematical Gazette, 101(550):130–135, March 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
3003B78DCB7257C4C58F84551B9CB0E9.

Korner:2014:CA

[Kör14] T. W. (Thomas William) Körner. Calculus for the ambitious. Cambridge University Press, Cambridge, UK, 2014. ISBN 1-107-06392-2, 1-107-68674-1. xii + 165 pp. LCCN QA303.2

.K665 2014. URL http://www.loc.gov/catdir/enhancements/
fy1501/2014451373-b.html;   http://www.loc.gov/catdir/
enhancements/fy1501/2014451373-d.html;   http://www.loc.
gov/catdir/enhancements/fy1501/2014451373-t.html.

Koshy:2011:CBC

[Kos11] Thomas Koshy. 95.63 Central binomial coefficients and Catalan congruences. The Mathematical Gazette, 95(534):533–538, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Koshy:2012:SCI

[Kos12a] Thomas Koshy. 96.46 Some Catalan identities with interesting consequence. The Mathematical Gazette, 96(536):323–328, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Koshy:2012:LGC

[Kos12b] Thomas Koshy. Lobb’s generalisation of Catalan’s parenthesisation problem revisited. The Mathematical Gazette, 96(535):56–65, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Koshy:2013:PW

[Kos13] Thomas Koshy. Pell walks. The Mathematical Gazette, 97(538): 27–35, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Koshy:2014:GTM

[Kos14] Thomas Koshy. Graph-theoretic models for the Fibonacci family. The Mathematical Gazette, 98(542):256–265, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Koshkin:2016:MAT

[Kos16] Sergiy Koshkin. 100.28 mixing angle trisection with Pythagorean triples. The Mathematical Gazette, 100(549):492–495, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
06D350E00A2DC42EDE979AF23EB8E8F8.

Katz:2014:TUH

[KP14] Victor J. Katz and Karen Hunger Parshall. Taming the unknown: a history of algebra from antiquity to the early twentieth century. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-14905-4 (hardcover). xiii + 485 pp. LCCN QA151 .K38 2014.

Kuhapatanakul:2017:SSR

[KR17] Kantaphon Kuhapatanakul and Pongpol Ruankong. 101.29 sums of square roots. The Mathematical Gazette, 101(552):476– 480, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/D3D9A9AA633966D72D2D71DDC5408C4E.

Kranat:2009:SGA

[Kra09] Jayne Kranat. Statistics GCSE: for AQA. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2009. ISBN 0-19-913430-8. LCCN ????

Krantz:2012:EAM

[Kra12] Steven G. (Steven George) Krantz. Elements of advanced mathematics. Chapman and Hall/CRC, Boca Raton, FL, USA, third edition, 2012. ISBN 1-4398-9834-0 (hardcover). xvi + 351 pp.

LCCN QA37.3 .K73 2012. URL http://marc.crcnetbase.com/
isbn/9781439898345.

Krasopoulos:2015:IIA

[Kra15] Panagiotis T. Krasopoulos. 99.33 an integral inequality with applications to probability theory. The Mathematical Gazette, 99(546): 528–530, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Kress:2014:LIE

[Kre14] Rainer Kress. Linear Integral Equations, volume 82 of Applied mathematical sciences. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2014. ISBN 1-4614-9592-X (hardcover), 1-4614-9593-8 (e-book). ISSN 0066-5452. xvi + 412 pp. LCCN QA431 .K776 2014; QA1 .A647

vol. 82. URL http://www.loc.gov/catdir/enhancements/
fy1404/2013954393-d.html;   http://www.loc.gov/catdir/
enhancements/fy1404/2013954393-t.html.

Krishna:2018:ETA

[Kri18] Dasari Naga Vijay Krishna. 102.47 an equilateral triangle associated with a hexagon. The Mathematical Gazette, 102 (555):515–516, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10247-an-equilateral-triangle-associated-with-
a-hexagon/EC89D2EA243E4024AFCA15161FE518EC.

Kroopnick:2018:LNL

[Kro18] Allan J. Kroopnick. Linear and non-linear models for national health expenditures in the USA. The Mathematical Gazette, 102 (554):193–197, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/linear-
and-nonlinear-models-for-national-health-expenditures-
in-the-usa/02E530C03C2299914FB1DFB44A987ABD.

Kallianpur:2014:SAD

[KS14] G. Kallianpur and P. (Padmanabhan) Sundar. Stochastic analysis and diffusion processes, volume 24 of Oxford graduate texts in mathematics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2014. ISBN 0-19-965706-8 (hardcover), 0-19-965707-6 (paperback). xi + 352 pp. LCCN QA274 .K338 2014.

Kuchel:2012:CYB

[Kuc12] Philip W. Kuchel. 96.45 Can you ‘bend’ a truncated truncated tetrahedron? The Mathematical Gazette, 96(536):317–323, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Kuchel:2015:KC

[Kuc15] Philip W. Kuchel. 99.23 Kicking a cube. The Mathematical Gazette, 99(545):346–347, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Kulkarni:2013:IRC

[Kul13a] Ragha Vendra G. Kulkarni. 97.03 Insert a root to crack the cubic. The Mathematical Gazette, 97(538):115–120, March 2013. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Kulkarni:2013:SQE

[Kul13b] Ragha Vendra G. Kulkarni. 97.04 The solution of quintic equations. The Mathematical Gazette, 97(538):120–122, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Kulkarni:2013:SOS

[Kul13c] Raghavendra G. Kulkarni. 97.19 Shifting the origin to solve quartic equations. The Mathematical Gazette, 97(539):268–270, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Kulkarni:2016:PGZ

[Kul16] Raghavendra G. Kulkarni. Picking genuine zeros of cubics in the Tschirnhaus method. The Mathematical Gazette, 100(547):48–53, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Kumano:2010:GLC

[Kum10] Mitsuhiro Kumano. 94.07 A generalisation of the limits of certain sequences. The Mathematical Gazette, 94(529):132–135, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Kumano:2011:ECI

[Kum11] Mitsuhiro Kumano. 95.38 An extension of a certain inequality in the 24th IMO. The Mathematical Gazette, 95(533):320–321, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

L:2010:PCa

[L.10a] N. J. L. Problem corner. The Mathematical Gazette, 94(530): 342–349, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

L:2010:PCb

[L.10b] N. J. L. Problem corner. The Mathematical Gazette, 94(531):544– 550, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

L:2011:PC

[L.11] N. J. L. Problem corner. The Mathematical Gazette, 95(533): 362–368, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

L:2015:PCa

[L.15a] N. J. L. Problem corner. The Mathematical Gazette, 99(544):165– 170, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

L:2015:PCb

[L.15b] N. J. L. Problem corner. The Mathematical Gazette, 99(545): 368–375, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

L:2015:PCc

[L.15c] N. J. L. Problem corner. The Mathematical Gazette, 99(546):547– 554, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

L:2016:PC

[L.16] N. J. L. Problem corner. The Mathematical Gazette, 100(547):164– 171, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

L:2017:PCa

[L.17a] N. J. L. Problem corner. The Mathematical Gazette, 101(550): 150–156, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/768841C170BBB0A65D5FAEF87215B3C5.

L:2017:PCb

[L.17b] N. J. L. Problem corner. The Mathematical Gazette, 101 (551):339–343, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/BD19437BB9AAE903C018BD3FE4FF2EE0.

L:2017:PCc

[L.17c] N. J. L. Problem corner. The Mathematical Gazette, 101 (552):547–551, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/04A0376ECBE19F9EBFBB06EF26D972BA.

L:2018:PCa

[L.18a] N. J. L. Problem corner. The Mathematical Gazette, 102 (553):174–181, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/7949EB6F19A284E6D7DFF7A8678AB68B.

L:2018:PCb

[L.18b] N. J. L. Problem corner. The Mathematical Gazette, 102(554):355–360, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/BB15AABE04A0A9BB81138D2F0B729ABC.

L:2018:PCc

[L.18c] N. J. L. Problem corner. The Mathematical Gazette, 102 (555):543–550, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/C83D8EADDDD253A66EEF6460C6D239E8.

L:2019:PCa

[L.19a] N. J. L. Problem corner. The Mathematical Gazette, 103 (556):172–178, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/E9EEEDC8BC9BDB115BE4B9B5A0DF99E6.

L:2019:PCb

[L.19b] N. J. L. Problem corner. The Mathematical Gazette, 103 (558):559–566, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/62DB165AAB89C3933BE4ABA338BB4252.

Lamphere:2019:HPT

[Lam19] Robert L. Lamphere. A hyperbolic proportionality theorem and hyperbolic surveying problems. The Mathematical Gazette, 103(557):212–221, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/hyperbolic-proportionality-theorem-and-hyperbolic-
surveying-problems/03F3AAFE8C8937E8D592ACDAE32DFC7E.

Lang:2012:ODS

[Lan12] Robert J. (Robert James) Lang. Origami design secrets: mathematical methods for an ancient art. A K Peters/CRC Press, Boca Raton, second edition, 2012. ISBN 1-56881-436-4 (paperback). xi + 758 pp. LCCN TT870 .L2614 2012. URL http://www.

crcnetbase.com/isbn/9781439867747.

Leader:2010:BRB

[Lea10] Imre Leader. Book review: Problems from the Book, by Titu Andreescu and Gabriel Dospinescu. Pp. 554. $69.95. 2008. ISBN 978-0-9799269-0-7 (XYZ Press). The Mathematical Gazette, 94(530): 380–381, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leek:2017:BRL

[Lee17] Robert Leek. Book review: The logic of infinity, by Sheppard, Barnaby, pp 473, £27.99 (paper), ISBN 978-1-107-67866-8, Cambridge University Press (2014). The Mathematical Gazette, 101 (550):184–185, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/423B5D5E000B0011D9850D3FCA2FD048.

Lehmann:2011:FNC

[Leh11] E. L. (Erich Leo) Lehmann. Fisher, Neyman, and the creation of classical statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 1-4419-9499-

8. viii + 115 pp. LCCN QA276.15 .L44 2011. URL http:

//www.loc.gov/catdir/enhancements/fy1402/2011930669-b.
html;   http://www.loc.gov/catdir/enhancements/fy1402/
2011930669-d.html;   http://www.loc.gov/catdir/enhancements/
fy1402/2011930669-t.html.

Leung:2012:DRA

[Leu12a] King-Shun Leung. 96.42 Dividing a right-angled trapezium into two similar quadrilaterals. The Mathematical Gazette, 96(536): 304–309, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leung:2012:STC

[Leu12b] King-Shun Leung. Similar triangles and the cosine rule. The Mathematical Gazette, 96(535):169–172, March 2012. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leung:2017:AQE

[Leu17] King-Shun Leung. Approaching quadratic equations from a right angle. The Mathematical Gazette, 101(552):424–438, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
6BBB78B4166F8553534351F2E99AFC4E.

Leung:2018:PFC

[Leu18] King-Shun Leung. Paper-folding and cutting activities to demonstrate five-fold symmetry. The Mathematical Gazette, 102 (555):413–421, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/paperfolding-and-cutting-activities-to-demonstrate-
fivefold-symmetry/0A4161EF5D1BA722969BA79581CA78DD.

Leversha:2008:CB

[Lev08] Garry Leversha. Crossing the bridge, volume 1 of Pathways. The United Kingdom Mathematics Trust, Leeds, UK, 2008. ISBN 1-906001-06-5 (paperback). xviii + 180 pp. LCCN ????

Levi:2009:MMU

[Lev09] Mark Levi. The mathematical mechanic: using physical reasoning to solve problems. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-14020-0 (hardcover). viii + 186 pp. LCCN QA63 .L48 2009.

Leversha:2010:BRBd

[Lev10a] Gerry Leversha. Book review: 50 mathematical ideas you really need to know, by Tony Crilly. Pp. 208. £8.99. 2008. ISBN 977-1-84724-008-8 (Quercus). The Mathematical Gazette, 94(530):367– 368, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2010:BRBe

[Lev10b] Gerry Leversha. Book review: How to think like a mathematician, by Kevin Houston. Pp. 265. £18.99. 2009. ISBN 978-0-521-71978-0 (CUP). The Mathematical Gazette, 94(530):369–371, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2010:BRBf

[Lev10c] Gerry Leversha. Book review: Music, a mathematical offering, by David J. Benson. Pp. 411. £25.99. 2007. ISBN 0-521-61999-8 (Cambridge University Press). The Mathematical Gazette, 94 (531):571–572, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2010:BRBa

[Lev10d] Gerry Leversha. Book review: The annotated Flatland, by Edwin A. Abbott and Ian Stewart. Pp. 242. £10.99. 2008. ISBN 978-0-465-01123-0 (Basic Books). The Mathematical Gazette, 94 (529):190–191, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2010:BRBc

[Lev10e] Gerry Leversha. Book review: The shape of content: creative writing in mathematics and science, Chandler Davis, Marjorie Wikler Senechal and Jan Zwicky (eds.). Pp. 194. $39. 2008. ISBN 978-1-56881-444-5 (AK Peters). Strange attractors: poems of love and mathematics, Sarah Glaz and JoAnne Growney (eds.). Pp 255. $39. 2008. ISBN 978-1-56881-341-7 (AK Peters). The Mathematical Gazette, 94(530):363–364, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2010:BRBb

[Lev10f] Gerry Leversha. Book review: The unfinished game, by Keith Devlin. pp. 170. $24.95. 2008. ISBN 978-0-465-00910-7 (Basic Books). The Mathematical Gazette, 94(530):361–362, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2011:BRBc

[Lev11a] Gerry Leversha. Book review: Gems of geometry, by John Barnes. Pp. 311. £35.99 (hbk). 2009. ISBN 978-3-642-05091-6 (Springer). The Mathematical Gazette, 95(532):162–163, March 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2011:BRBa

[Lev11b] Gerry Leversha. Book review: Mathematics and music, by David Wright. Pp. 161. £29.50. 2009. ISBN 978-0-8218-4873-9 (American Mathematical Society). The Mathematical Gazette, 95(532):154– 155, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2011:BRBe

[Lev11c] Gerry Leversha. Book review: Professor Stewart’s hoard of mathematical treasures, by Ian Stewart. Pp. 339. £12.99. 2009. ISBN: 978-1-84668-292-6 (Basic Books). The Mathematical Gazette, 95 (532):172–173, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2011:BRBb

[Lev11d] Gerry Leversha. Book review: Sacred mathematics, by Hidetoshi Fukagawa and Tony Rothman. Pp. 348. £26.95. 2008. ISBN 978-0-691-12745-3 (Princeton University Press). Traditional Japanese mathematics problems, by Hidetoshi Fukagawa and John Rigby. Pp. 191. $50. 2002. ISBN 981-04-2759-X (SCT Publishing, Singapore). The Mathematical Gazette, 95(532):159–160, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2011:BRBd

[Lev11e] Gerry Leversha. Book review: Superbrain: the first twenty-five years, by Early Diarmuid and MacHale. Des Pp. 230. £13.00. 2009. ISBN: 978-1-906338-01-5 (HighPerception). The Mathematical Gazette, 95(532):171–172, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2011:BRBf

[Lev11f] Gerry Leversha. Book review: The big questions: mathematics, by Tony Crilly. Pp. 208. £9.99. 2011. ISBN 978-1-84916-240-1 (Quercus). The Mathematical Gazette, 95(533):379–380, July 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2011:E

[Lev11g] Gerry Leversha. Editorial. The Mathematical Gazette, 95(534): 394–396, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2012:BRBa

[Lev12a] Gerry Leversha. Book review: A transition to advanced mathematics, by William Johnston and Alex M. McAllister. Pp. 745. $79.95 (Hardback). 2009. ISBN 978-0-19-531076-4 (Oxford University Press). The Mathematical Gazette, 96(535):186–187, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2012:BRBc

[Lev12b] Gerry Leversha. Book review: Geometry revealed, by Marcel Berger. Pp. 831. £53.99. 2010. ISBN 978-3-540-70996-1 (Springer). The Mathematical Gazette, 96(536):376–377, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2012:BRBd

[Lev12c] Gerry Leversha. Book review: Six gems of geometry, by Thomas Reale. Pp. 128. $14.99 (hbk). 2010. ISBN: 978-1-935638-02-5 (Perfectly Scientific Press). The Mathematical Gazette, 96(537):570– 572, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2012:BRBb

[Lev12d] Gerry Leversha. Book review: The best writing on mathematics 2010, edited by Mircea Pitici. Pp. 407. £13.95. 2011. ISBN 978-0-691-14841-0 (Princeton University Press). The Mathematical Gazette, 96(536):370–371, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Levi:2012:WCL

[Lev12e] Mark Levi. Why Cats Land on Their Feet: And 76 Other Physical Paradoxes and Puzzles. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-14854-6 (paperback), 1-4008-4172-0 (e-book). x + 4 + 190 pp. LCCN Q173 .L545 2012. URL http:/

/www.jstor.org/stable/10.2307/j.ctt7rw2b.

Levrie:2012:WPF

[Lev12f] Paul Levrie. 96.14 Wallis’s product formula for π: a symbolic approach. The Mathematical Gazette, 96(535):127–129, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2013:BRBe

[Lev13a] Gerry Leversha. Book review: A wealth of numbers, by Benjamin Wardhaugh. Pp. 370 £30.95. 2012. ISBN: 978-0-691-14775-8 (Princeton University Press). The Mathematical Gazette, 97(540): 555–556, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2013:BRBc

[Lev13b] Gerry Leversha. Book review: Hidden harmonies: the lives and times of the Pythagorean theorem, Robert and Ellen Kaplan. Pp.

280. £25. 2011. ISBN 978-1-59691-522-0 (Bloomsbury Press). The Mathematical Gazette, 97(539):360–361, July 2013. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2013:BRBb

[Lev13c] Gerry Leversha. Book review: Math for the frightened, by Colin Pask. Pp. 380. £16.99. 2011. ISBN: 978-1-61614-421-0 (Prometheus Books). The Mathematical Gazette, 97(539):353–354, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2013:BRBh

[Lev13d] Gerry Leversha. Book review: Math goes to the movies, by Burkhard Polster and Marty Ross Pp. 286. £18.00. 2012. ISBN: 978-1-4214-0484-4 (Johns Hopkins University Press). The Mathematical Gazette, 97(540):562–563, November 2013. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2013:BRBa

[Lev13e] Gerry Leversha. Book review: Mathematics in Victorian Britain, by Raymond Flood, Adrian Rice and Robin Wilson (eds.). Pp.

466. £29.99 (hbk). 2011. ISBN: 978-0-19-960139-4 (Oxford University Press). The Mathematical Gazette, 97(538):188–190, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2013:BRBg

[Lev13f] Gerry Leversha. Book review: Maths on trial: how numbers get used and abused, by Leila Schneps and Coralie Colmez. Pp. 236. £17.99. 2013. ISBN: 978-0-465-03292-1 (Basic Books). The Mathematical Gazette, 97(540):560–562, November 2013. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2013:BRBf

[Lev13g] Gerry Leversha. Book review: The crossing of heaven: memoirs of a mathematician, by Karl Gustafson. Pp. 176. $59.95. 2012. ISBN 978-3-642-22557-4 (Springer). The Mathematical Gazette, 97(540): 557, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2013:BRBd

[Lev13h] Gerry Leversha. Book review: The great mathematicians, by Raymond Flood and Robin Wilson. Pp. 208. £9.99. 2011 ISBN: 987-1-84837-902-2 (Arcturus). The Mathematical Gazette, 97(539):361– 362, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2013:DDC

[Lev13i] Gerry Leversha. A direct derivation of the Catalan formula. The Mathematical Gazette, 97(538):53–60, March 2013. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2013:GT

[Lev13j] Gerry Leversha. The geometry of the triangle, volume 2 of Pathways. The United Kingdom Mathematics Trust, Leeds, UK, 2013. ISBN 1-906001-17-0 (paperback). xxxii + 541 pp. LCCN ????

Leversha:2014:BRBm

[Lev14a] Gerry Leversha. Book review: Beautiful geometry by Eli Maor and Eugen Jost, pp 187, £19.95 (hbk), ISBN 978-0-691-15099-4, Princeton University Press (2014). The Mathematical Gazette, 98 (543):562–563, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRBk

[Lev14b] Gerry Leversha. Book review: Combinatorics: ancient and modern, edited by Robin Wilson and John J. Watkins. Pp. 381. £55. hardcover. 2013, ISBN 978-0-19-965659-2 (Oxford University Press). The Mathematical Gazette, 98(542):381–382, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRBh

[Lev14c] Gerry Leversha. Book review: How to teach relativity to your dog, by Chad Orzel. Pp. 328. £11.99. 2012. ISBN: 978-0-465-02331-8 (Basic Books). The Mathematical Gazette, 98(541):185, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRL

[Lev14d] Gerry Leversha. Book review: Love and Math by Edward Frankel, pp 292, £18.99, ISBN 978-0-465-05074-1, Basic Books (2013). The Mathematical Gazette, 98(543):554–555, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRBa

[Lev14e] Gerry Leversha. Book review: Magnificent mistakes in mathematics, by Alfred S Posamentier and Ingmar Lehmann. Pp. 296. $24 (hbk). 2013.ISBN 978-1-61614-747-1 (Prometheus Books). The Mathematical Gazette, 98(541):148–150, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRN

[Lev14f] Gerry Leversha. Book review: Naked statistics, by Charles Wheelan. Pp. 269. £18.99. 2013. ISBN 978-0-393-07195-5. (W.W.Norton). The Mathematical Gazette, 98(543):564–567, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRO

[Lev14g] Gerry Leversha. Book review: Oxford figures: eight centuries of the mathematical sciences by John Fauvel, Raymond Flood and Robin Wilson (eds), pp 406, £39.99, ISBN 978-0-19-968197-6, Oxford University Press (2013). The Mathematical Gazette, 98(543): 549–550, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRBi

[Lev14h] Gerry Leversha. Book review: Phi, the glorious golden ratio, by Alfred S. Posamentier and Ingmar Lehmann. Pp. 363. $27-00. 2012. ISBN: 978-1- 61614-423-4 (Prometheus Books). The Mathematical Gazette, 98(541):191, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRBb

[Lev14i] Gerry Leversha. Book review: Poor Robin’s prophesies, by Benjamin Wardhaugh. Pp. 248. £14.99 (hbk). 2012, ISBN 978-0-19-960542-2 (Oxford University Press). The Mathematical Gazette, 98 (541):150–151, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRBg

[Lev14j] Gerry Leversha. Book review: Ptolemy through the looking-glass, by William Wynne Willson. Pp. 92. £13.00. 2012. ISBN 978-0-9558042-3-6 (Garland Publications). The Mathematical Gazette, 98(541):177–178, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRBe

[Lev14k] Gerry Leversha. Book review: Sophie’s diary, by Dora Musielak. Pp. 291. $68.95 (hardback). 2012. ISBN: 978-88385-577-5 (Mathematical Association of America). The Mathematical Gazette, 98 (541):156–157, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRBf

[Lev14l] Gerry Leversha. Book review: The best writing on mathematics 2011, by Mircea Pitici (ed.). Pp. 383. £13.95. 2012. ISBN: 978-0-691-15315-5 (Princeton University Press). The Mathematical Gazette, 98(541):161–163, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRBl

[Lev14m] Gerry Leversha. Book review: The best writing on mathematics 2012 edited by Mircea Pitici, pp 288, £13.95, ISBN 978-0-691-15655-2, Princeton University Press (2013). The Mathematical Gazette, 98(543):556–557, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRBc

[Lev14n] Gerry Leversha. Book review: The cult of Pythagoras: math and myths by Alberto A. Martinez. Pp. 264. $27.95 (hbk), 2012. ISBN 978-0-8229-4418-8 (University of Pittsburg Press). The Mathematical Gazette, 98(541):151–152, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRBd

[Lev14o] Gerry Leversha. Book review: The history of mathematics, a very short introduction, by Jacqueline Stedall. Pp. 123. £7.99. 2012. ISBN: 978-0-19-959968-4 (Oxford University Press). The Mathematical Gazette, 98(541):155–156, March 2014. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:BRBj

[Lev14p] Gerry Leversha. Book review: The secrets of triangles, by Alfred S. Posamentier and Ingmar Lehmann, Pp. 385, $26. 2012. ISBN 978-1-61614-587-3. (Prometheus Books). The Mathematical Gazette, 98(542):371–373, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2014:RR

[Lev14q] Gerry Leversha. Request for referees. The Mathematical Gazette, 98(543):575, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2015:BRM

[Lev15a] Gerry Leversha. Book review: Mathematics in 20th century literature and art by Robert Tubbs, pp 162, £19.50, ISBN 978-1-4214-1380-8, Johns Hopkins University Press (2014). The Mathematical Gazette, 99(546):565–566, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2015:BRB

[Lev15b] Gerry Leversha. Book review: The best writing on mathematics 2013 edited by Mircea Pitici, pp 244, £14.95, ISBN 978-0-691-16041-2, Princeton (2014). The Mathematical Gazette, 99(546): 566–567, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2015:DJF

[Lev15c] Gerry Leversha. Dr John Frankland Rigby. The Mathematical Gazette, 99(544):174–175, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2015:E

[Lev15d] Gerry Leversha. Editorial. The Mathematical Gazette, 99(544):1, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2015:RN

[Lev15e] Gerry Leversha. Reviewers needed. The Mathematical Gazette, 99 (544):191–192, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Leversha:2017:BRCb

[Lev17a] Gerry Leversha. Book review: Catalan numbers, by Stanley, Richard P., pp. 216, £29.99, ISBN 978-1-107142774-7, Cambridge University Press (2015). The Mathematical Gazette, 101 (551):377–378, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/A6501F8D01842FAE20ABC93AD0E6B4DB.

Leversha:2017:BRCc

[Lev17b] Gerry Leversha. Book review: Creating symmetry: the artful mathematics of wallpaper patterns by Frank A. Farris, pp. 230, £24.95, ISBN 978-0-691-16173-0, Princeton University Press (2015). The Mathematical Gazette, 101(552):560–561, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/creating-symmetry-
the-artful-mathematics-of-wallpaper-patterns-by-frank-
a-farris-pp-230-2495-isbn-9780691161730-princeton-university-
press-2015/B516E9362F137AAFD691E173FD12E3BC.

Leversha:2017:BRM

[Lev17c] Gerry Leversha. Book review: Mathematics and the real world, by Artstein, Zvi, pp. 426, £25, ISBN 978-1-61614-546-0, Prometheus Books (2014). The Mathematical Gazette, 101(551):356–357, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
B14AB7B63BC51E8B245D98441B286D65.

Leversha:2017:BRPb

[Lev17d] Gerry Leversha. Book review: Plain plane geometry, by Sasane, Amol, pp. 270, £40, ISBN 978-9-481474-043-2, World Scientific (2016). The Mathematical Gazette, 101(551):372–373, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
1B205B52E95D49C0C193EEFEA0642B63.

Leversha:2017:BRPa

[Lev17e] Gerry Leversha. Book review: Problem solving tactics, by di Pasquale, A., Do, N. and Mathews, D., pp. 304, $A115, ISBN 978-1-87642-075-8, Australian Mathematics Trust (2014). The Mathematical Gazette, 101(550):186–187, March 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
087CA7DB50AF7161D3733AD0EF35ABAE.

Leversha:2017:BRS

[Lev17f] Gerry Leversha. Book review: Special secondary schools for the mathematically talented: an international panorama by Vogeli, Bruce R. (ed), pp. 438, £134, ISBN 978-9-81466-746-3, World Scientific (2015). The Mathematical Gazette, 101(550): 161–162, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/D022E2CD7C0B8AD795C0CD085C4B8740.

Leversha:2017:BRB

[Lev17g] Gerry Leversha. Book review: The best writing on mathematics 2015, by Pitici, Mircea (ed), pp 392, £18.95, ISBN 978-0-691-16965-2, Princeton (2016). The Mathematical Gazette, 101 (551):362–363, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/9F925CA25BD002023CB4E91AC8896F53.

Leversha:2017:BRCa

[Lev17h] Gerry Leversha. Book review: The circle; mathematical exploration beyond the line, by Posamentier, Alfred S. and Geretschläger, Robert, pp. 350, £19.28, ISBN 978-1-63388-167-9, Prometheus Books (2016). The Mathematical Gazette, 101 (551):373–375, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/1AC2F32F03816BFC295CEECD620BCB56.

Leversha:2017:BRT

[Lev17i] Gerry Leversha. Book review: Trigonometry: notes, problems and exercises by Roger Delbourgo, pp. 172, £48, ISBN 978-9-81320-311-2, World Scientific Publishing (2017). The Mathematical Gazette, 101(552):563–565, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/trigonometry-notes-problems-and-exercises-by-roger-
delbourgo-pp-172-48-isbn-9789813203112-world-scientific-
publishing-2017/86B3257D95D19D88A6010F157F6B0E6B.

Leversha:2018:AR

[Lev18a] Gerry Leversha. Appeal for referees. The Mathematical Gazette, 102(555):574, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/appeal-for-referees/4FAB93FBFDE0848E089F85E596278D6C.

Leversha:2018:BRS

[Lev18b] Gerry Leversha. Book review: A square peg in a round hole: adventures in the mathematics of area by Chris Pritchard, pp. 208, £17.99, ISBN 978-0-906588-82-6, The Mathematical Association (2016). The Mathematical Gazette, 102(555):554–555, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/square-peg-
in-a-round-hole-adventures-in-the-mathematics-of-area-
by-chris-pritchard-pp-208-1799-isbn-9780906588826-the-
mathematical-association-2016/179F918AB91DB1152A804D0DFBC25B98.

Leversha:2018:BRQ

[Lev18c] Gerry Leversha. Book review: Questions pupils ask by Colin Foster, pp. 199, £15.95 (MA members £13.45), ISBN 978-0-90658-891-8, The Mathematical Association (2017). The Mathematical Gazette, 102(555):555–557, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/questions-pupils-ask-by-colin-foster-pp-199-1595-
ma-members-1345-isbn-9780906588918-the-mathematical-association-
2017/6DEE8662B1E325CE394092B481DB2638.

Leversha:2018:BRB

[Lev18d] Gerry Leversha. Book review: The best writing on Mathematics 2016 by Mircea Pitici (ed.), pp. 377, £27.95, ISBN 978-0-691-17529-4, Princeton University Press (2017). The Mathematical Gazette, 102(554):365–366, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/best-writing-on-mathematics-2016-by-mircea-pitici-
ed-pp-377-2795-isbn-9780691175294-princeton-university-
press-2017/574972020B6B674646C266DBF3806263.

Leversha:2019:A

[Lev19a] Gerry Leversha. An appreciation. The Mathematical Gazette, 103(558):488, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/an-appreciation/35A8C3F56E7AE2AC35AB3B3E8A705821.

Leversha:2019:BRC

[Lev19b] Gerry Leversha. Book review: A common family weakness for statistics by Chris Pritchard, pp. 127, £9.00, ISBN 978-1-911616-03-0, The Mathematical Association (2018). The Mathematical Gazette, 103(556):188–189, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/common-family-weakness-for-statistics-by-chris-
pritchard-pp-127-900-isbn-9781911616030-the-mathematical-
association-2018/DEAF85537CBDD7B26222757145A6CD26.

Levrie:2019:CFE

[Lev19c] Paul Levrie. 103.21 Closed form evaluation of a class of improper integrals. The Mathematical Gazette, 103(557):323–328, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10321-closed-form-evaluation-
of-a-class-of-improper-integrals/E76BC6AEFE9E79323C3ED0EE94D6950D.

Lewand:2010:PC

[Lew10] Robert Edward Lewand. The perfect cipher. The Mathematical Gazette, 94(531):401–411, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lewis:2013:HMD

[Lew13] Barry Lewis. How many dead ends in a derangement? The Mathematical Gazette, 97(538):81–94, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lewis:2014:RSA

[Lew14] Barry Lewis. Runs, strings and alphabets. The Mathematical Gazette, 98(542):204–226, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lesmoir-Gordon:2010:CIB

[LG10] Nigel Lesmoir-Gordon, editor. The colours of infinity: the beauty and power of fractals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-84996-485-8. 207 pp. LCCN QA614.86 .C65 2010. URL http://extras. springer.com. With contributions by Ian Stewart, Sir Arthur C. Clarke, Benoît B. Mandelbrot, Michael and Louisa Barnsley, Will Rood, Gary Flake, and David Pennock.

Loveland:2014:TVL

[LHS14] Donald W. Loveland, Richard E. Hodel, and Susan G. Sterrett. Three views of logic: mathematics, philosophy, and computer science. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-16044-9 (paperback), 1-4008-4875-X. xv + 322 pp. LCCN QA9.54 .L68 2014.

li:2017:SPK

[li17] Jun li. 101.35 some properties of the Kepler triangle. The Mathematical Gazette, 101(552):494–495, November 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
4382FCBBCC963D57A8F95E849F9E4C15.

Lin:2011:RVH

[Lin11] C.-S. Lin. 95.66 The reciprocal volume of a Heron tetrahedron. The Mathematical Gazette, 95(534):542–545, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lipscombe:2013:MMM

[Lip13] Trevor Lipscombe. Mental mathematics for multiples of π. The Mathematical Gazette, 97(538):167–169, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Little:2013:FF

[Lit13] Chris Little. Fortuitous functions. The Mathematical Gazette, 97(538):172, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Levitin:2011:AP

[LL11] Anany Levitin and Maria Levitin. Algorithmic puzzles. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-974044-5 (paperback). xxi + 257 pp. LCCN QA95 .L475 2011.

Levrie:2017:RSS

[LN17] Paul Levrie and Amrik Singh Nimbran. Recycling Stirling’s series acceleration technique. The Mathematical Gazette, 101 (550):69–82, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/65EDE94E608218FB23629E50420A8715.

LoBello:2013:OMW

[Lo 13] Anthony Lo Bello. The origins of mathematical words: a comprehensive dictionary of Latin, Greek, and Arabic roots. The Johns Hopkins University Press, Baltimore, MD, USA, 2013. ISBN 1-4214-1098-2 (paperback), 1-4214-1099-0 (electronic). xv + 350 pp.

LCCN QA41.3 .B45 2013. URL http://muse.jhu.edu/books/
9781421410999/;   http://public.eblib.com/choice/publicfullrecord.
aspx?p=3318723;   http://site.ebrary.com/id/10756847.

Lockhart:2012:M

[Loc12] Paul Lockhart. Measurement. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-05755-4 (hardcover). 407 pp. LCCN QA447 .L625 2012.

Loomis:2011:TRS

[Loo11] Paul Loomis. 95.54 Two results on sums of cubes. The Mathematical Gazette, 95(534):506–510, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2010:ICC

[Lor10a] Nick Lord. 94.20 An interesting constraint on cutting corners. The Mathematical Gazette, 94(530):312–314, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2010:BAW

[Lor10b] Nick Lord. Binomial averages when the mean is an integer. The Mathematical Gazette, 94(530):331–332, July 2010. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2010:BRBc

[Lor10c] Nick Lord. Book review: A course in commutative Banach algebras, by E. Kaniuth. pp. 353. 46.95 Euros. 2009. ISBN 978-0-387-72475-1 (Springer). The Mathematical Gazette, 94(531):566–567, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2010:BRBb

[Lor10d] Nick Lord. Book review: Classical algebra, by Roger Cooke. Pp. 206. £28.95. 2008. ISBN 978 0 470 25952 8 (Wiley). The Mathematical Gazette, 94(530):375–376, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2010:BRBa

[Lor10e] Nick Lord. Book review: Penguin dictionary of mathematics (4th edn), David Nelson (ed.). Pp. 480. £10.99. 2008. ISBN 978-0-14-103023-4 (Penguin). The Mathematical Gazette, 94(530):355, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2010:C

[Lor10f] Nick Lord. Correspondence. The Mathematical Gazette, 94(530): 335–336, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2010:PSG

[Lor10g] Nick Lord. A power series graphing challenge. The Mathematical Gazette, 94(531):527–528, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2010:PC

[Lor10h] Nick Lord. Problem corner. The Mathematical Gazette, 94(529): 159–165, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2010:VTS

[Lor10i] Nick Lord. Variations on a theme — squares in a circular sector. The Mathematical Gazette, 94(529):147–149, March 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:SAR

[Lor11a] Nick Lord. 95.36 Sequences of averages revisited. The Mathematical Gazette, 95(533):314–317, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:IAI

[Lor11b] Nick Lord. 95.37 An instructive application of the inequality of means. The Mathematical Gazette, 95(533):317–319, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:USM

[Lor11c] Nick Lord. 95.39 A UKMT senior mathematical challenge problem: frustration leads to joy! The Mathematical Gazette, 95(533): 322, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:NTO

[Lor11d] Nick Lord. 95.44 Newton tackles an Olympiad problem. The Mathematical Gazette, 95(533):334–341, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:BM

[Lor11e] Nick Lord. Binomial modes. The Mathematical Gazette, 95(533): 356–357, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:BRBc

[Lor11f] Nick Lord. Book review: Making mathematics come to life: a guide for teachers and students, by O. A. Ivanov. pp. 337. £55.95. 2009. ISBN 978-0-8218-4808-1 (American Mathematical Society). The Mathematical Gazette, 95(533):372–373, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:BRBa

[Lor11g] Nick Lord. Book review: Riot at the calc. exam and other Mathematically Bent stories, by Colin Adams. Pp. 271. $32. 2009. ISBN 978-0-8218-4817-3 (American Mathematical Society). The Mathematical Gazette, 95(532):155–156, March 2011. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:BRBb

[Lor11h] Nick Lord. Book review: The mathematical colouring book: mathematics of colouring and the colourful life of its creators, by Alexander Soifer. Pp. 607. 42.95 euros. 2009. ISBN 978-0-387-74640-1 (Springer). The Mathematical Gazette, 95(532):163–165, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:BRBd

[Lor11i] Nick Lord. Book review: The mathematics of A-level statistics, by Owen Toller. Pp. 56. £13.99. 2010. ISBN 978-0-906588-68-0 (The Mathematical Association). The Mathematical Gazette, 95 (533):383, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:CBI

[Lor11j] Nick Lord. The confusion between inverse and reciprocal functions. The Mathematical Gazette, 95(532):124–125, March 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:PCa

[Lor11k] Nick Lord. Problem corner. The Mathematical Gazette, 95(532): 133–141, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:PCb

[Lor11l] Nick Lord. Problem corner. The Mathematical Gazette, 95(534): 562–567, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:PW

[Lor11m] Nick Lord. Proof without words that 2 < e < 3. The Mathematical Gazette, 95(532):125–126, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:RPR

[Lor11n] Nick Lord. Reflection paths with restitution. The Mathematical Gazette, 95(533):354–356, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2012:HWP

[Lor12a] Nick Lord. 96.13 How Wallis’s product approaches π. The Mathematical Gazette, 96(535):124–127, March 2012. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2012:EIU

[Lor12b] Nick Lord. 96.38 Evaluating integrals using polar areas. The Mathematical Gazette, 96(536):289–296, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2012:ST

[Lor12c] Nick Lord. 96.39 Squares in triangles. The Mathematical Gazette, 96(536):296–299, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2012:PCa

[Lor12d] Nick Lord. Problem corner. The Mathematical Gazette, 96(535): 177–182, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2012:PCb

[Lor12e] Nick Lord. Problem corner. The Mathematical Gazette, 96(536): 351–356, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2012:PCc

[Lor12f] Nick Lord. Problem corner. The Mathematical Gazette, 96(537): 560–566, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2012:QAF

[Lor12g] Nick Lord. Quadrilaterals with area formula \sqrt{abcd}. The Mathematical Gazette, 96(536):345–347, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2012:SPH

[Lor12h] Nick Lord. Shot putt with a height constraint. The Mathematical Gazette, 96(536):348–350, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2012:SCR

[Lor12i] Nick Lord. Simultaneous construction of regular 5, 15 and 60-gons. The Mathematical Gazette, 96(536):347–348, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2012:VTE

[Lor12j] Nick Lord. Variations on a theme — Euler and the logsine integral. The Mathematical Gazette, 96(537):451–458, November 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2013:FHT

[Lor13a] Nick Lord. 97.29 A family of Heron-type formulae for the triangle. The Mathematical Gazette, 97(539):294–297, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2013:AIS

[Lor13b] Nick Lord. 97.34 An area inequality for subtriangles of a quadrilateral. The Mathematical Gazette, 97(539):308–311, July 2013. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2013:BAE

[Lor13c] Nick Lord. Binomial averages: an erratum. The Mathematical Gazette, 97(539):336, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2013:BRB

[Lor13d] Nick Lord. Book review: A course in commutative algebra, by Gregor Kemper. Pp. 246. £44.99. 2011. ISBN: 978-3-642-03544-9 (Springer). The Mathematical Gazette, 97(539):369, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2013:PCa

[Lor13e] Nick Lord. Problem corner. The Mathematical Gazette, 97(538): 178–183, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2013:PCb

[Lor13f] Nick Lord. Problem corner. The Mathematical Gazette, 97(539): 341–349, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2013:PCc

[Lor13g] Nick Lord. Problem corner. The Mathematical Gazette, 97(540): 544–550, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2013:TSM

[Lor13h] Nick Lord. Two surprising maximisation problems. The Mathematical Gazette, 97(540):535–537, November 2013. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2013:WPQ

[Lor13i] Nick Lord. Why the path of quickest descent is not a straight line. The Mathematical Gazette, 97(538):169–171, March 2013. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:MIE

[Lor14a] Nick Lord. 98.08 The moment of inertia of an elliptical wire. The Mathematical Gazette, 98(541):121–125, March 2014. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:ENF

[Lor14b] Nick Lord. 98.21 Euclid numbers are free from powers. The Mathematical Gazette, 98(543):482–483, November 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:BBP

[Lor14c] Nick Lord. Beyond the Basel problem: Euler’s derivation of the general formula for ζ(2n). The Mathematical Gazette, 98(543): 459–474, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:BRM

[Lor14d] Nick Lord. Book review: More problem solving — the creative side of mathematics by Derek Holton, pp 256, £14.99, £11.99 (MA members), ISBN 978-0-906588-78-9, The Mathematical Association (2013). The Mathematical Gazette, 98(543):573–574, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:BRBb

[Lor14e] Nick Lord. Book review: Real analysis: a historical approach (2nd edn.), by S. Stahl. Pp. 293. £80.50. 2011. ISBN 978-0-470-87890-3 (Wiley). The Mathematical Gazette, 98(541):168–169, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:BRBa

[Lor14f] Nick Lord. Book review: What’s happening in the mathematical sciences, Volume 8, by Dana Mackenzie. Pp. 129. £17.95. 2011. ISBN 978-0-8218-4999-6 (American Mathematical Society). The Mathematical Gazette, 98(541):163, March 2014. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:CWV

[Lor14g] Nick Lord. Cubics whose vertical translates factorise. The Mathematical Gazette, 98(542):349–350, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:DRV

[Lor14h] Nick Lord. Discrete random variables having equal mean and variance. The Mathematical Gazette, 98(542):347–349, July 2014. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:EUM

[Lor14i] Nick Lord. Evaluating ∑_{1}^{n} r^{2} and ∑_{1}^{n} r(r + 1) using moments. The Mathematical Gazette, 98(542):346–347, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:FIS

[Lor14j] Nick Lord. A fallacy involving similar quadrilaterals. The Mathematical Gazette, 98(543):525–527, November 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:NCC

[Lor14k] Nick Lord. A non-calculator challenge: show that ln 2 < 1\sqrt{2}. The Mathematical Gazette, 98(542):346, July 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:PNS

[Lor14l] Nick Lord. Prime numbers and the sequences 6n ± 1. The Mathematical Gazette, 98(541):126–128, March 2014. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:PCa

[Lor14m] Nick Lord. Problem corner. The Mathematical Gazette, 98(541): 136–142, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:PCb

[Lor14n] Nick Lord. Problem corner. The Mathematical Gazette, 98(542): 358–366, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2014:PCc

[Lor14o] Nick Lord. Problem corner. The Mathematical Gazette, 98(543): 536–542, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:QPC

[Lor15a] Nick Lord. 99.03 Quick proofs that certain sums of fractions are not integers. The Mathematical Gazette, 99(544):128–130, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:TWB

[Lor15b] Nick Lord. 99.15 Triangles in which a^{n} = b^{n} + c^{n}. The Mathematical Gazette, 99(545):326–328, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:JSF

[Lor15c] Nick Lord. 99.18 Justifying the \sqrt{2π} factor in Stirling’s formula. The Mathematical Gazette, 99(545):336–337, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:WDI

[Lor15d] Nick Lord. 99.19 When do the inradius and exradii form progressions? The Mathematical Gazette, 99(545):337–339, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:SP

[Lor15e] Nick Lord. 99.34 a special parallelogram. The Mathematical Gazette, 99(546):530–532, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:AIS

[Lor15f] Nick Lord. Areas of images of singular 3 × 3 matrix transformations. The Mathematical Gazette, 99(544):155–156, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:BRBa

[Lor15g] Nick Lord. Book review: Enlightening symbols: a short history of mathematical notation and its hidden powers by J. Mazur, pp. 285, £19.95, ISBN 978-0-691-15463-3, Princeton University Press (2014). The Mathematical Gazette, 99(544):177–178, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:BRL

[Lor15h] Nick Lord. Book review: Locally convex spaces by M. Scott Osborne, pp 213, £44.99 (hard), ISBN 978-3-319-02044-0, Springer Verlag (2014). The Mathematical Gazette, 99(546):568–569, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:BRBb

[Lor15i] Nick Lord. Book review: Quand les mathématiques posent problème(s) by Michel Bataille, pp. 216, EUR19, ISBN 978-2-7298-86738, Ellipses www.editions-ellipses.fr (2014). The Mathematical Gazette, 99(544):190–191, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:CSC

[Lor15j] Nick Lord. Cauchy–Schwarz via collisions. The Mathematical Gazette, 99(546):541–542, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:CAF

[Lor15k] Nick Lord. The constant acceleration formulae imply constant acceleration! The Mathematical Gazette, 99(546):542–543, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:EPP

[Lor15l] Nick Lord. Extending the parity proof that \sqrt{2} is irrational. The Mathematical Gazette, 99(544):155, March 2015. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:RCP

[Lor15m] Nick Lord. A rich calculus problem about a frustum. The Mathematical Gazette, 99(545):359–362, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2015:RPD

[Lor15n] Nick Lord. A rich problem on digital roots. The Mathematical Gazette, 99(546):543–545, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2016:QE

[Lor16a] Nick Lord. 100.19 a quick evaluation of lim_{m→∞} ∑_{k=0}^{m} (\frac{k}{m})^{m}. The Mathematical Gazette, 100(548):321–323, July 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
476DB809786A9AA8D176EA53E234A768.

Lord:2016:BRR

[Lor16b] Nick Lord. Book review: A readable introduction to real mathematics by Daniel Rosenthal, David Rosenthal and Peter Rosenthal, pp. 161, £35.99 (hard), ISBN 978-3-319-05653-1, Springer (2014). The Mathematical Gazette, 100(547):175–176, March 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2016:BRP

[Lor16c] Nick Lord. Book review: Problems and proofs in numbers and algebra by R. S. Millman, P. J. Shiue and E. B. Kahn, pp. 223, £35.99 (hard), ISBN 978-3-319-14426-9, Springer (2015). The Mathematical Gazette, 100(547):186–187, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2016:CPT

[Lor16d] Nick Lord. Constructing the 15th pentagon that tiles the plane. The Mathematical Gazette, 100(547):154–158, March 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2016:EC

[Lor16e] Nick Lord. Euler, the clothoid and ∫_{0}^{∞} \frac{sin x}{/}x dx. The Mathematical Gazette, 100(548):266–273, July 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
2E73DA0C1B3207B7CC3C5DD8A0A6E43B.

Lord:2016:MEP

[Lor16f] Nick Lord. The most elementary proof that ∑_{k=1}^{∞} 1/k^{2} = π^{2}/6. The Mathematical Gazette, 100(549):429–434, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
513F59660D22469C55B1B3DD49D81E0C.

Lord:2016:WST

[Lor16g] Nick Lord. When is the sum of two triangular numbers triangular? The Mathematical Gazette, 100(547):152–154, March 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2017:UAG

[Lor17a] Nick Lord. 101.04 An unexpected appearance of the golden ratio. The Mathematical Gazette, 101(550):98–99, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
61256DE694E36B49A0D038E72D23419C.

Lord:2017:WDD

[Lor17b]  Nick Lord. 101.16 when does n! have n digits?  The

Mathematical Gazette, 101(551):293–296, July 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
75401401C645DEE26B27457AAAE4B46B.

Lord:2017:VEI

[Lor17c] Nick Lord. 101.28 the versatile exponential inequality e^{x} ≥ 1 + x. The Mathematical Gazette, 101(552):470–475, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
F527D4AB73C55003AAECCB0EC4ED019A.

Lord:2017:AC

[Lor17d] Nick Lord. An area conundrum. The Mathematical Gazette, 101 (551):325–327, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/2E56BA422E47DA86728713AD79FD3A1B.

Lord:2017:BRE

[Lor17e] Nick Lord. Book review: Elements of mathematics: from Euclid to Gödel, by Stillwell, John, pp. 422, $39.95, ISBN 978-0-691-17168-5, Princeton University Press (2016). The Mathematical Gazette, 101(551):353–355, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/CC4A7E48BD7517E60FF1826827DA264B.

Lord:2017:BRP

[Lor17f] Nick Lord. Book review: Problem-solving strategies in mathematics, by Posamentier, A. S. & Krulik, S., pp. 163, £20, ISBN 9-789814-651639, World Scientific (2015). The Mathematical Gazette, 101(550):187–188, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.

cambridge.org/core/product/DFED0F97A5C6C100775B878AE7E4EB89.

Lord:2017:BRC

[Lor17g] Nick Lord. Book review: The call of the primes, by O’Shea, Owen, pp. 330, $19, ISBN 978-1-63388-148-8, Prometheus Books (2016). The Mathematical Gazette, 101(550):173–174, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
44DC6BFE08AC5793486213387E112EEE.

Lord:2017:UAS

[Lor17h] Nick Lord. Using A4-sized paper to illustrate that \sqrt{2} is irrational. The Mathematical Gazette, 101(550):142–145, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
128350D2F13D4F0430BA576603A525B9.

Lord:2017:VCP

[Lor17i] Nick Lord. The volume of a cone for pre-calculus students. The Mathematical Gazette, 101(552):534–537, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/volume-of-a-cone-for-precalculus-
students/6F44E90FBE421CA04C9B3BC05C62FCA9.

Lord:2018:VPI

[Lor18a] Nick Lord. 102.23 a visual proof that \sqrt{N} is irrational. The Mathematical Gazette, 102(554):311–312, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10223-a-visual-proof-that-is-irrational/6A28B90702AE39278A85D06CDA536892.

Lord:2018:AFA

[Lor18b] Nick Lord. 102.25 on an alternative formula for the area of a spherical cap. The Mathematical Gazette, 102(554):314–316, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10225-on-an-alternative-formula-
for-the-area-of-a-spherical-cap/75E88788830E92923B65EE31208A619F.

Lord:2018:API

[Lor18c] Nick Lord. 102.44 another proof of the irrationality of square roots. The Mathematical Gazette, 102(555):511–512, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10244-another-proof-of-the-
irrationality-of-square-roots/1EFB65C04D4A6E26CA4B1E013C8D04E1.

Lord:2018:BRI

[Lor18d] Nick Lord. Book review: An introduction to analysis by Jan Mikusiński & Piotr Mikusiński, pp. 303, £73.00 (hard), ISBN 9-789813-2026-10, World Scientific (2017). The Mathematical Gazette, 102(555):562–563, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/an-introduction-to-analysis-by-jan-mikusinski-
piotr-mikusinski-pp-303-7300-hard-isbn-9789813202610-world-
scientific-2017/230735100B5C0E0811FADEF0519E4414.

Lord:2018:BRP

[Lor18e] Nick Lord. Book review: Proofs without words III by Roger B. Nelsen, pp. 187, $50, ISBN 978-0-88585-790-8, MAA Press (2015). The Mathematical Gazette, 102(555):569–570, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/proofs-without-words-iii-
by-roger-b-nelsen-pp-187-50-isbn-9780885857908-maa-press-
2015/D221EEEE54E7C88461E9E162920D24AE.

Lord:2018:BRE

[Lor18f] Nick Lord. Book reviews: Euler’s pioneering equation by Robin Wilson, pp. 162, £14.99, ISBN 978-0-19-879492-9, Oxford University Press (2018). — A most elegant equation by David Stipp, pp. 221, $27, ISBN 978-0-465-09377-9, Basic Books (2017). The Mathematical Gazette, 102 (555):564–565, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/eulers-pioneering-equation-by-robin-wilson-pp-
162-1499-isbn-9780198794929-oxford-university-press-2018-
a-most-elegant-equation-by-david-stipp-pp-221-27-isbn-
9780465093779-basic-books-2017/A41A3A0E21BC36039B88BCC90F25D28D.

Lord:2018:ERC

[Lor18g] Nick Lord. Extending runs of composite numbers. The Mathematical Gazette, 102(554):351–352, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/extending-runs-of-composite-numbers/2DC4F191F4C08D98F1EF53A7FF537CBE.

Lord:2018:CSQ

[Lor18h] Nick Lord. On the converse of a STEP question about centroids. The Mathematical Gazette, 102(555):538–539, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/on-the-converse-
of-a-step-question-about-centroids/E29B99FA872141BE804B22F977D80C16.

Lord:2018:TCI

[Lor18i] Nick Lord. Three cameos involving e. The Mathematical Gazette, 102(553):159–164, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/three-cameos-involving-e/5922B2A87AAA4FDF98D68966E57D4984.

Lord:2018:TQO

[Lor18j] Nick Lord. Two quick observations about rough inclined planes. The Mathematical Gazette, 102(554):352–353, July 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/two-quick-observations-about-
rough-inclined-planes/B249594ADE3CDAC074D394887DBFBDCE.

Lord:2018:UIC

[Lor18k] Nick Lord. An unusual inequality challenge. The Mathematical Gazette, 102(553):155–159, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/an-unusual-inequality-challenge/EBC33A2BDA1BD374E607260F71DC52AB.

Lord:2019:TFC

[Lor19a] Nick Lord. 103.36 Three footnotes to Cartwright’s proof that π is irrational. The Mathematical Gazette, 103(558):514–517, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10336-three-footnotes-
to-cartwrights-proof-that-is-irrational/363A5B83C7DD47CE0756716635816FA7.

Lord:2019:MEV

[Lor19b] Nick Lord. More equal volumes of revolution. The Mathematical Gazette, 103(556):168–169, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/more-equal-volumes-of-revolution/44E8755207D668AA7EEC57F982294AAA.

Lord:2019:PC

[Lor19c] Nick Lord. Problem corner. The Mathematical Gazette, 103(557):359–364, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/problem-corner/CF33C67E1A8A3916425C6FDF76AC3585.

Lord:2019:SEA

[Lor19d] Nick Lord. Separating the effects of air resistance and restitution. The Mathematical Gazette, 103(556):163–165, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/separating-
the-effects-of-air-resistance-and-restitution/AC497031BAC2CBBE670E6E0077889B61.

Lorentziadi:2019:CN

[Lor19e] Athina Lorentziadi. Calculation of π with a needle. The Mathematical Gazette, 103(556):111–116, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/calculation-of-with-a-needle/6232D75E33C50F20341D607FCB5BEF10.

Lovett:2010:DGM

[Lov10] Stephen (Stephen T.) Lovett. Differential geometry of manifolds. A. K. Peters, Ltd., Wellesley, MA, USA, 2010. ISBN 1-56881-457-7. xiii + 421 pp. LCCN QA649 .L68 2010.

Lam:2010:BRB

[LP10] Tung Ken Lam and Sue Pope. Book review: Marvelous modular origami, by Meenakshi Mukerji. Pp. 92. $14.95. 2007. ISBN: 978-1-56881-316-5 (A. K. Peters). The Mathematical Gazette, 94(531): 573, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Levrie:2016:EEA

[LP16] Paul Levrie and Rudi Penne. The extended Euclidean algorithm made easy. The Mathematical Gazette, 100(547):147–149, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lery:2012:ESS

[LPE^{+}12] Thibaut Lery, Mario Primicerio, Maria J. Esteban, Magnus Fontes, Yvon Maday, Volker Mehrmann, Gonçalo Quadros, Wil Schilders, Andreas Schuppert, and Heather Tewkesbury. European success stories in industrial mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-23847-5. 135 pp. LCCN ????

URL http://deposit.d-nb.de/cgi-bin/dokserv?id=3866353\
%26prov=M\%26dok\_var=1\%26dok\_ext=htm;   http://www.
loc.gov/catdir/enhancements/fy1208/2011938138-b.html;
http://www.loc.gov/catdir/enhancements/fy1208/2011938138-
d.html;   http://www.loc.gov/catdir/enhancements/fy1208/
2011938138-t.html.

Lord:2010:BC

[LRQ10] Nick Lord, John Rigby, and Douglas Quadling. (the Bride’s chair)^{2}. The Mathematical Gazette, 94(530):239–246, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lucas:2010:ODT

[Luc10] T. Nigel Lucas. Obtaining derivatives of trigonometric functions using mechanics. The Mathematical Gazette, 94(531):522–526, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lucas:2013:FDT

[Luc13] T. Nigel Lucas. 97.50 Further derivatives of trigonometric functions using mechanics. The Mathematical Gazette, 97(540):520– 523, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lukarevski:2017:ETE

[Luk17] Martin Lukarevski. 101.10 Exradii of the triangle and Euler’s inequality. The Mathematical Gazette, 101(550):123, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
7567F94FE60C2925DA9580A52D1CB18B.

Lukarevski:2018:ETC

[Luk18] Martin Lukarevski. 102.53 the excentral triangle and a curious application to inequalities. The Mathematical Gazette, 102(555):531–533, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10253-the-excentral-triangle-and-a-curious-application-
to-inequalities/73057F3C3FA9E75E0A304D10A962D41F.

Lukarevski:2019:IAI

[Luk19] Martin Lukarevski. 103.43 An inequality arising from the inarc centres of a triangle. The Mathematical Gazette, 103 (558):538–541, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10343-an-inequality-arising-from-the-inarc-centres-
of-a-triangle/7FDABBF7600BEA8C48EDC439C0FBE039.

Luzia:2016:PTS

[Luz16] Nuno Luzia. 100.08 Pythagoras theorem by sliding a triangle. The Mathematical Gazette, 100(547):126–128, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Lord:2011:DBM

[LY11] Nick Lord and Alex Yan. A difference in the behaviour of moments of inertia and variance. The Mathematical Gazette, 95(534):556– 557, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Menghini:2008:FCI

[M^{+}08] Marta Menghini et al., editors. The first century of the International Commission on Mathematical Instruction, 1908-2008: reflecting and shaping the world of mathematics education, Biblioteca dell’Enciclopedia. Scienze e filosofia. Istituto della enciclopedia italiana, Roma, Italy, 2008. ISBN 88-12-00015-0. LCCN

QA11.A1 F615 2008. URL http://www.loc.gov/catdir/toc/
casalini08/08917164.pdf.

MacCluer:2009:EFA

[Mac09] Barbara D. MacCluer. Elementary functional analysis, volume 253 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-85528-9 (hardcover), 0-387-85529-7 (electronic). x + 207 pp. LCCN QA320 .M25 2009.

MacGregor:2010:BRBc

[Mac10a] P. MacGregor. Book review: Mathematical modeling and simulation, introduction for scientists and engineers, by Kai Velten. Pp.

348. £70.00. 2009. ISBN: 978-3-527-40758-8 (Wiley-VCH). The Mathematical Gazette, 94(530):379–380, July 2010. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2010:BRBa

[Mac10b] Peter MacGregor. Book review: Feedback systems: an introduction for scientists and engineers, by K. J. Åstrom and R. M. Murray. Pp. 396. £26.95. 2008. ISBN: 0-691-13576-2 (Princeton). The Mathematical Gazette, 94(529):187, March 2010. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2010:BRBb

[Mac10c] Peter MacGregor. Book review: Mathematical techniques: an introduction for the engineering, physical and mathematical sciences (4th edn), by D. W. Jordan and P. Smith. Pp. 976. £31.99. 2008. ISBN 978-0-19-928201-2 (Oxford University Press). The Mathematical Gazette, 94(530):357–358, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacHale:2010:A

[Mac10d] Des MacHale. 94.31 Approximations for π and e. The Mathematical Gazette, 94(531):486–488, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Macfarlane:2011:MTS

[Mac11a] A. J. Macfarlane. The Morley tricorn a solid figure constructed from the diagram for Morley’s theorem. The Mathematical Gazette, 95(532):49–51, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2011:BRB

[Mac11b] P. MacGregor. Book review: Stability and stabilization: an introduction, by William J. Terrell. pp. 457. £40.95 (Hardback). 2009. ISBN: 978-0-691-13444-4 (Princeton University Press). The Mathematical Gazette, 95(532):166–167, March 2011. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Machale:2011:GV

[Mac11c] Des Machale. 95.25 Goldbach variations. The Mathematical Gazette, 95(533):271–273, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mackenzie:2011:WHM

[Mac11d] Dana Mackenzie. What’s happening in the mathematical sciences, volume 8 of What’s happening in the mathematical sciences. American Mathematical Society, Providence, RI, USA, 2011. ISBN 0-8218-4999-9 (paperback). vi + 129 pp. LCCN QA36 .M223 2011.

MacCormick:2012:NAC

[Mac12a] John MacCormick. Nine algorithms that changed the future: the ingenious ideas that drive today’s computers. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-14714-0 (hardcover), 0-691-15819-3 (paperback). x + 2 + 219 pp.

LCCN QA76 .M21453 2012. URL http://press.princeton.
edu/chapters/s9528.pdf;   http://www.jstor.org/stable/

10.2307/j.ctt7t71s. With a foreword by Christopher M. Bishop.

MacGregor:2012:BRB

[Mac12b] P. MacGregor. Book review: Introduction to inequalities, by C. J. Bradley. Pp. 210. £11.00. 2010. ISBN: 978-1-906001-11-7 (United Kingdom Mathematics Trust). The Mathematical Gazette, 96 (536):374, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacHale:2012:MH

[Mac12c] Des MacHale. Mathematics and humour. The Mathematical Gazette, 96(537):542–543, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2013:BRBb

[Mac13a] P. MacGregor. Book review: Fundamental maths for engineering and science, by Mark Breach. Pp. 380. £25.99. 2011. ISBN: 978-0-230-25208-0 (Palgrave Macmillan). The Mathematical Gazette, 97 (539):355, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2013:BRBd

[Mac13b] P. G. MacGregor. Book review: Orieami 5, by Patsy Wang-Iverson, Robert J. Lang and Mark Yim (eds) Pp. 646. £43.99. 2011. ISBN: 978-1-56881-714-9 (CRC Press). The Mathematical Gazette, 97(539):381–382, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2013:BRBc

[Mac13c] Peter MacGregor. Book review: Differential geometry of manifolds, by Stephen Lovett. Pp. 421. £53.99 (hbk). 2010. ISBN 978-1-56881-457-5 (A. K. Peters). The Mathematical Gazette, 97(539): 372–373, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2013:BRBe

[Mac13d] Peter MacGregor. Book review: Linear algebra and probability for computer science applications, by Ernest Davis. Pp. 413. 2012. ISBN 978-1-4665-0155-3 (CRC press). The Mathematical Gazette, 97(540):572, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2013:BRBa

[Mac13e] Peter MacGregor. Book review: Origami design secrets: mathematical methods for an ancient art, by Robert Lang. Pp. 758. £38.99. 2011. ISBN: 978-1-56881-436-0 (CRC Press). The Mathematical Gazette, 97(538):191, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacHale:2013:IMP

[Mac13f] Des MacHale. 97.40 Infinitely many proofs that there are infinitely many primes. The Mathematical Gazette, 97(540):495–498, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2014:BRBc

[Mac14a] P. G. MacGregor. Book review: Magnificent Principia — exploring Isaac Newton’s masterpiece, by Colin Pask. Pp. 528. £16.05. 2013. ISBN 978-1-61614-745-7 (Prometheus Books). The Mathematical Gazette, 98(542):378–379, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2014:BRA

[Mac14b] Peter MacGregor. Book review: Advanced mechanics, by S.G. Rajeev. Pp 163. 2013. ISBN 978-0-19-967085-7 £45.00 (hbk), £21.99 (pbk). ISBN: 978-0-19-967086-4 (Oxford University Press). The Mathematical Gazette, 98(543):563–564, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2014:BRBa

[Mac14c] Peter MacGregor. Book review: Introduction to differential calculus. Systematic studies with engineering applications. U. L. Rohde, G. C. Jain, A. K. Poddar, A. K. Ghosh. Pp. 747. £93.50 (hbk) 2012. ISBN 978-1-118-11775-0 (John Wiley and Sons). Introduction to integral calculus. Systematic studies with engineering applications. U. L. Rohde, G. C. Jain, A. K. Poddar, A. K. Ghosh. Pp. 402. £76.95 (hbk) 2012. ISBN 978-1-118-11776-7 (John Wiley and Sons). The Mathematical Gazette, 98(541):169–170, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2014:BRBb

[Mac14d] Peter MacGregor. Book review: The shape of inner space, by Shing-Tung Yau and Steve Nadis. Pp. 377. $18.99. 2010. ISBN 978-0-465-02837-5 (Basic Books). The Mathematical Gazette, 98 (542):370, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacHale:2014:ASS

[Mac14e] Des MacHale. 98.10 Associativity of symmetric set difference. The Mathematical Gazette, 98(542):319–320, July 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2015:BRG

[Mac15a] Peter MacGregor. Book review: Games and mathematics by David Wells, pp 246, £15.99 (paper), ISBN 978-1-107-69091-2, Cambridge University Press (2012). The Mathematical Gazette, 99 (546):574, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2015:BRJ

[Mac15b] Peter MacGregor. Book review: James Clerk Maxwell, edited by [Raymond Flood,] Marc McCartney and Andrew Whitaker, pp 364, £39.99 (hard), ISBN 978-0-19-966437-5, Oxford University Press (2014). The Mathematical Gazette, 99(546):564, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacGregor:2015:WAS

[Mac15c] Peter MacGregor. Book review: Wizards, aliens and starships by Charles L. Adler, pp. 378, £29.95 (hard), ISBN 978-0-691-14715-4, Princeton University Press (2014). The Mathematical Gazette, 99 (545):381–382, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacKinnon:2015:PF

[Mac15d] Nick MacKinnon. 99.07 On products of factorials. The Mathematical Gazette, 99(544):144–149, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Macgregor:2016:BRL

[Mac16a] Peter Macgregor. Book review: Linear algebra, step by step by Kuldeep Singh, pp. 608, £29.99 (paper), ISBN 978-0-19-965444-4, Oxford University Press (2014). The Mathematical Gazette, 100 (547):183–184, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Macgregor:2016:BRP

[Mac16b] Peter Macgregor. Book review: Partial differential equations, by Michael Shearer and Rachel Levy, pp. 274, £59.00 (hard), ISBN 978-0-691-16129-7, Princeton University Press (2015). The Mathematical Gazette, 100(547):182–183, March 2016. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Macgregor:2016:BRT

[Mac16c] Peter Macgregor. Book review: Tensor calculus for physics, by Dwight E. Neuenschwander, pp. 227, $45.00 (paper), ISBN 978-1-421-41565-9, Johns Hopkins University Press (2015). The Mathematical Gazette, 100(547):184, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Macgregor:2017:BRP

[Mac17] Peter Macgregor. Book review: Prime numbers, friends who give problems, a trialogue with Papa Paulo by Paulo Ribenboim, pp. 320, £32 (paper), ISBN 978-9-814-72581-1, also available as e-book, World Scientific, 2016. The Mathematical Gazette, 101(552):559–560, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/prime-numbers-friends-who-give-problems-a-trialogue-
with-papa-paulo-by-paulo-ribenboim-pp-320-32-paper-isbn-
9789814725811-also-available-as-ebook-world-scientific-
2016/64A1D64A9778DF675015AA8CBFBADDF9.

Macgregor:2018:BRQ

[Mac18] Peter Macgregor. Book review: The quantum labyrinth by Paul Halpern, pp. 311, $30.00 (hard), ISBN 978-0-46509-758-6, Basic Books (2017). The Mathematical Gazette, 102(555):558, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/quantum-labyrinth-
by-paul-halpern-pp-311-3000-hard-isbn-9780465097586-basic-
books-2017/503AA07B8BD8E31A8FC2BEE77AC33F16.

MacGregor:2019:AMV

[Mac19] Peter MacGregor. Book review: Applied Mathematics, a very short introduction by Alain Goriely, pp. 141, £7.99 (paper), ISBN 978-0-19875-404-6, Oxford University Press (2018). The Mathematical Gazette, 103(557):376–377, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/applied-mathematics-a-very-short-introduction-
by-alain-goriely-pp-141-799-paper-isbn-9780198754046-oxford-
university-press-2018/C5972624545ABFD37497D92EC044FA4A.

Mahajan:2010:SFM

[Mah10a] Sanjoy Mahajan. Street-fighting mathematics: the art of educated guessing and opportunistic problem solving. MIT Press, Cambridge, MA, USA, 2010. ISBN 0-262-51429-X (paperback), 0-262-13507-8 (paperback). xv + 134 pp. LCCN QA63 .M34 2010. Foreword by Carver A. Mead.

Maher:2010:BRB

[Mah10b] Philip Maher. Book review: Practical linear algebra: a geometry toolbox, by Gerald Farin and Dianne Hansford. Pp. 384. £56-00. 2005. ISBN: 1-56881-234-5 (A. K. Peters). The Mathematical Gazette, 94(529):185–187, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mahony:2010:PEA

[Mah10c] John Mahony. The polarisation ellipse and associated properties. The Mathematical Gazette, 94(529):94–102, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mahony:2011:C

[Mah11] John Mahony. Correspondence. The Mathematical Gazette, 95 (534):561, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mahony:2013:GE

[Mah13a] John D. Mahony. Gauging the earth. The Mathematical Gazette, 97(540):413–420, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mahwood:2013:PWC

[Mah13b] Munir Mahwood. Proof without words: a characterisation of the difference of two squares. The Mathematical Gazette, 97(539):333– 334, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mahony:2014:WT

[Mah14] John D. Mahony. Wind over tide. The Mathematical Gazette, 98 (542):273–280, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mahony:2015:C

[Mah15a] John D. Mahony. Correspondence. The Mathematical Gazette, 99 (545):363, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mahony:2015:PIT

[Mah15b] John D. Mahony. Piecing it together. The Mathematical Gazette, 99(544):40–44, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mahmoudi:2018:GCI

[Mah18a] M. G. Mahmoudi. 102.33 a geometrical construction involving two intersecting circles. The Mathematical Gazette, 102(554):335–339, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10233-a-geometrical-construction-involving-two-
intersecting-circles/8CBDB401EEDE2AFA63E659119247AC0A.

Mahony:2018:DD

[Mah18b] John D. Mahony. From dawn to dusk. The Mathematical Gazette, 102(554):246–256, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/from-dawn-to-dusk/0E83122A0AB4B221757AA2C31911CDEE.

Mahony:2018:TPN

[Mah18c] John D. Mahony. Tethering in pastures new. The Mathematical Gazette, 102(554):233–245, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/tethering-in-pastures-new/3AB4ED4ED6BE5A76EBD9205728F00D8E.

Mahony:2019:DVT

[Mah19a] John D. Mahony. “DIYMathematics” v. “ThePackage”. The Mathematical Gazette, 103(556):128–129, March 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/diymathematics-v-thepackage/
402A49A1A9B1E7FF1D9BBDAB1F08E078.

Mahony:2019:LPI

[Mah19b] John D. Mahony. Locating parameters of interest in a conic section. The Mathematical Gazette, 103(557):196–203, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/locating-parameters-of-interest-
in-a-conic-section/48BE161F9D5F79E039D6B41960201F11.

Marrero:2011:OEZ

[Mar11] Osvaldo Marrero. One equals zero: a specious argument. The Mathematical Gazette, 95(534):554–555, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Martinez:2012:CPM

[Mar12] Alberto A. Martinez. The cult of Pythagoras: math and myths. University of Pittsburgh Press, Pittsburgh, PA, USA, 2012. ISBN 0-8229-4418-9 (hardcover). 263 pp. LCCN QA21 .M335 2012.

URL http://muse.jhu.edu/books/9780822978534/;   http:/
/www.jstor.org/stable/10.2307/j.ctt7zw974.

Marrero:2016:WBS

[Mar16]  Osvaldo Marrero. What is Bayesian statistics?  The

Mathematical Gazette, 100(548):247–256, July 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
F004C0EE0AB4E6FD86739DC06A2D1125.

Mason:2014:UPG

[Mas14] John Mason. Uniqueness of patterns generated by repetition. The Mathematical Gazette, 98(541):1–7, March 2014. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Matthews:2016:SWP

[Mat16] J. N. S. Matthews. Should we play on your lawn or mine? The Mathematical Gazette, 100(547):24–33, March 2016. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Matsko:2018:LDP

[Mat18a] Vincent J. Matsko. Lattice dissections: plainer and less fancy. The Mathematical Gazette, 102(555):454–459, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/

mathematical-gazette/article/lattice-dissections-plainer-
and-less-fancy/40A1F5BB22DAAB00A8432598263F9FD9.

Matsuda:2018:CIE

[Mat18b] Yasuo Matsuda. 102.12 on circles inscribed in an ellipse. The Mathematical Gazette, 102(553):131–132, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10212-on-circles-inscribed-
in-an-ellipse/5BBA1A3770D017A83E87A009B56E5909.

May:2016:DPG

[May16] Russell J. May. 100.10 dice problems with generating function solutions. The Mathematical Gazette, 100(547):131–137, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mazur:2014:ESS

[Maz14] Joseph Mazur. Enlightening symbols: a short history of mathematical notation and its hidden powers. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-15463-5 (hardcover), 1-4008-5011-8 (e-book), 1-306-51798-2. xxiii + 285 pp. LCCN QA41 .E39 2014; QA41 .M39 2014. URL http://www.jstor.

org/stable/10.2307/j.ctt5hhpnp.

Mason:2010:TM

[MBS10] John Mason, Leone Burton, and Kaye Stacey. Thinking mathematically. Pearson, Harlow, second edition, 2010. ISBN 0-273-72891-1. xvi + 248 pp. LCCN QA39.2 .M373 2010.

Mapp:2009:KMC

[MC09] Fiona C. Mapp and Steven Croft. KS3 maths coursebook: covers all three years. Essentials. Lonsdale, London, UK, 2009. ISBN 1-84419-239-3 (paperback). 288 pp. LCCN QA139 .M37 2009.

McAndrew:2010:ELF

[McA10] Alasdair McAndrew. An elementary, limit-free calculus for polynomials. The Mathematical Gazette, 94(529):67–83, March 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

McBride:2018:BRC

[McB18] Adam McBride. Book review: The calculus story: a mathematical adventure by David Acheson, pp. 192, £11.99 (hard), ISBN 978-0-19-880454-3, Oxford University Press (2017). The Mathematical Gazette, 102(555):566–567, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/calculus-story-a-mathematical-adventure-by-david-
acheson-pp-192-1199-hard-isbn-9780198804543-oxford-university-
press-2017/5969DE11F6EB9FC1307F8DEE78E2ADCD.

McClure:2015:LBL

[McC15] Lynne McClure. Looking backward, looking forward. The Mathematical Gazette, 99(546):386–390, November 2015. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

McGovern:2013:MDS

[McG13] Iggy McGovern. A mystic dream of 4: a sonnet sequence based on the life of William Rowan Hamilton. Quaternia Press, Dublin, Ireland, 2013. ISBN 0-9926294-0-3 (hardcover), 0-9926294-1-1 (paperback). 92 pp. LCCN PR6113.C476 M97 2013.

McLean:2011:C

[McL11] K. Robin McLean. Correspondence. The Mathematical Gazette, 95 (533):358, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

McLean:2014:CC

[McL14] K. Robin McLean. Conics and convexity. The Mathematical Gazette, 98(542):266–272, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

McLean:2018:BDD

[McL18] K. Robin McLean. Blocks of decimal digits. The Mathematical Gazette, 102(554):198–202, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/blocks-of-decimal-digits/B9BEE0BD932A61C47C2921B8CB8BCA2E.

McLean:2019:DEC

[McL19a] K. Robin McLean. 103.04 does Euclid’s construction miss infinitely many primes? The Mathematical Gazette, 103(556):135– 137, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10304-does-
euclids-construction-miss-infinitely-many-primes/E3D10B9645644275ECDCEFDAEE3FDCC6.

McLean:2019:LSM

[McL19b] K. Robin McLean. Latin square matrices and their inverses. The Mathematical Gazette, 103(557):265–276, July 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/latin-square-matrices-and-
their-inverses/C61E15060466407E8E634124A8FF1361.

McNamee:2014:OCN

[McN14] John Michael McNamee. 98.07 The order of convergence of Newton’s Method in special cases. The Mathematical Gazette, 98 (541):119–120, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mcnamee:2016:OCI

[Mcn16] John M. Mcnamee. 100.14 optimum choice of interpolation points in Newton’s divided difference formula. The Mathematical Gazette, 100(547):143–145, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mumford:2010:PTS

[MD10] David Mumford and Agnès Desolneux. Pattern theory: the stochastic analysis of real-world signals. Applying mathematics. A. K. Peters, Ltd., Wellesley, MA, USA, 2010. ISBN 1-56881-579-4. xi + 407 + 8 pp. LCCN Q327 .M85 2010. URL http://marc.

crcnetbase.com/isbn/9781439865569.

Miet:2016:C

[MD16] James Arathoon Miet and Kevin J. Davis. Correspondence. The Mathematical Gazette, 100(547):159–160, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Megyesi:2012:BRB

[Meg12] Gábor Megyesi. Book review: Deformation theory, by Robin Hartshorne. Pp. 234. £58.80 (hbk). 2010. ISBN: 978-1-4419-1595-5 (Springer). The Mathematical Gazette, 96(537):570, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Melville:2013:SSR

[Mel13] John Melville. A simple series representation for Apéry’s constant. The Mathematical Gazette, 97(540):455–460, November 2013. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mercer:2010:SPJ

[Mer10] A. McD. Mercer. 94.33 Short proofs of Jensen’s and Levinson’s inequalities. The Mathematical Gazette, 94(531):492–495, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mercer:2013:NMV

[Mer13] A. McD. Mercer. 97.46 A new mean value theorem for integrals. The Mathematical Gazette, 97(540):510–512, November 2013. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Merttens:2015:TSS

[Mer15] Ruth Merttens. Textbooks from Shanghai and Singapore: A national debate. The Mathematical Gazette, 99(546):391–401, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mercer:2017:CMV

[Mer17] Peter R. Mercer. 101.07 Cauchy’s mean value theorem meets the logarithmic mean. The Mathematical Gazette, 101(550): 108–115, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/34A986EDB75D49760D24D3AE41D47743.

Mercer:2019:AA

[Mer19] Peter R. Mercer. 103.06 another approach to ∫_{0}^{π} log(2 sin u)du. The Mathematical Gazette, 103(556):138–140, March 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10306-another-approach-to/
542AAF48ACD5A3F077CC73457614045A.

Mezo:2014:CDS

[Mez14] István Mező. 98.15 A class of dense subsets on the real line. The Mathematical Gazette, 98(542):332–334, July 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Moreno:2013:MTJ

[MGC13a] Samuel G. Moreno and Esther M. García-Caballero. Γ^{2}(1/2) is more than just π. The Mathematical Gazette, 97(540):430–434, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Moreno:2013:ERN

[MGC13b] Samuel G. Moreno and Esther M. García-Caballero. Entry 1.414 … in Ramanujan’s notebooks: \sqrt{2} is irrational. The Mathematical Gazette, 97(539):329, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Moreno:2019:SI

[MGC19] Samuel G. Moreno and Esther M. García-Caballero. 103.05 sums of integers. The Mathematical Gazette, 103(556):137–138, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10305-sums-
of-integers/2F5ABBA4D109EB86A6D9503118EF7348.

Michael:2009:HGA

[Mic09] T. S. Michael. How to guard an art gallery: and other discrete mathematical adventures. The Johns Hopkins University Press, Baltimore, MD, USA, 2009. ISBN 0-8018-9298-8 (hardcover), 0-8018-9299-6 (paperback), 0-8018-9704-1 (e-book). xi + 3 + 257 pp.

LCCN QA164 .M53 2009. URL http://muse.jhu.edu/books/
9780801897047/.

Milatovic:2011:POT

[Mil11] Ognjen Milatovic. 95.21 A property of odd terms of a generalised Fibonacci sequence. The Mathematical Gazette, 95(533):262–264, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Milner:2012:VAS

[Mil12] Walter Milner. A visual approach to Schnell and Mendoza. The Mathematical Gazette, 96(535):172–173, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Minton:2012:GNH

[Min12] Roland B. Minton. Golf by the numbers: How stats, math, and physics affect your game. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0315-3. xvi + 291 pp. LCCN GV965 .M478 2012. URL http:

//www.loc.gov/catdir/enhancements/fy1408/2011929971-b.
html;   http://www.loc.gov/catdir/enhancements/fy1408/
2011929971-d.html.

Mitchell:2012:PHT

[Mit12] Douglas W. Mitchell. 96.40 A property of hyperbolas and their asymptotes. The Mathematical Gazette, 96(536):299–301, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Maor:2014:BG

[MJ14] Eli Maor and Eugen Jost. Beautiful geometry. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-15099-0 (hardcover). x + 187 pp. LCCN QA447 .M37 2014.

Mungan:2013:PAR

[ML13] Carl Mungan and Trevor Lipscombe. 97.16 Probability analysis for rolls of a square cuboidal die. The Mathematical Gazette, 97 (538):163–166, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacHale:2012:MS

[MM12] Des MacHale and Peter MacHale. 96.23 Minimal Sudoku. The Mathematical Gazette, 96(535):153–156, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacHale:2013:RS

[MM13a] Des MacHale and Peter MacHale. Relations on sets. The Mathematical Gazette, 97(539):224–233, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

MacHale:2013:VTP

[MM13b] Des MacHale and Joseph Manning. 97.18 Variations on twin primes. The Mathematical Gazette, 97(539):265–268, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mahmood:2013:TPW

[MM13c] Munir Mahmood and Ibtihal Mahmood. 97.05 Two proofs without words. The Mathematical Gazette, 97(538):123–124, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Machale:2015:MRS

[MM15] Des Machale and Joseph Manning. Maximal runs of strictly composite integers. The Mathematical Gazette, 99(545):213–219, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Moore:2009:IPS

[MMC09] David S. Moore, George P. McCabe, and Bruce A. Craig. Introduction to the practice of statistics. W. H. Freeman, New York, NY, USA, sixth edition, 2009. ISBN 1-4292-1623-9 (extended version, casebound), 1-4292-1622-0 (casebound), 1-4292-1621-2 (paperback), 1-4292-1612-3 (CDROM). xxxiii + 709 + 76 pp. LCCN QA276.12 .M65 2009.

Matsui:2010:PLT

[MMYA10] H. Matsui, D. Minematsu, T. Yamauchi, and R. Miy Adera. Pascal-like triangles and Fibonacci-like sequences. The Mathematical Gazette, 94(529):27–41, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

McGregor:2010:FUM

[MNS10] Colin M. McGregor, Jonathan J. C. Nimmo, and Wilson W. Stothers. Fundamentals of University Mathematics. Woodhead Publishing in Mathematics. Woodhead Publishing, Oxford, UK, third edition, 2010. ISBN 0-85709-223-5 (paperback), 0-85709-224-3 (e-book), 1-904275-45-1. xii + 552 pp. LCCN QA37.3

.M3474 2010. URL http://www.sciencedirect.com/science/
book/9780857092236.

Molitierno:2012:ACM

[Mol12] Jason J. Molitierno. Applications of combinatorial matrix theory to Laplacian matrices of graphs. Discrete mathematics and its applications. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-6337-7 (hardcover), 1-4398-6339-3 (e-book). 405 pp. LCCN QA166.243 .M65

2012. URL http://www.crcnetbase.com/isbn/9781439863374;
 http://www.crcnetbase.com/isbn/9781439863398.

Molokach:2017:NOG

[Mol17] John Molokach. 101.20 a natural occurrence of the Golden Ratio. The Mathematical Gazette, 101(551):303, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
8A294A44AFE00AA05FC3A20EDC5D314D.

Morland:2019:IPA

[Mor19] Tim Morland. Book review: Intermediate problems by Andrew Jobbings, pp. 279, £16 (paper), ISBN 978-1-90600-129-2, UK Mathematics Trust (2018). The Mathematical Gazette, 103(557):382, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/intermediate-problems-by-andrew-jobbings-pp-279-
16-paper-isbn-9781906001292-uk-mathematics-trust-2018/
5D66A6BB43A70BF6E8DBC2790123F2CD.

Mathew:2011:FTC

[MP11] Panakkal Mathew and Imre Patyi. 95.10 The fundamental theorem of calculus implies that of algebra. The Mathematical Gazette, 95 (532):88–89, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mercer:2013:ETJ

[MS13] Peter R. Mercer and Allan A. Sesay. 97.25 Error terms for Jensen’s and Levinson’s inequalities. The Mathematical Gazette, 97(539): 277–280, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Mukerji:2007:MMO

[Muk07] Meenakshi Mukerji. Marvelous modular origami. A. K. Peters, Ltd., Wellesley, MA, USA, 2007. ISBN 1-56881-316-3. xvi + 75 pp. LCCN TT870 .M82 2007.

Mukerji:2010:OI

[Muk10] Meenakshi Mukerji. Origami inspirations. A. K. Peters, Ltd., Wellesley, MA, USA, 2010. ISBN 1-56881-584-0. xi + 120 pp. LCCN TT870 .M823 2010.

Munro:2012:BRB

[Mun12] Chris Munro. Book review: Advanced engineering mathematics with MATLAB (3rd edn.), by D. Duffy. Pp. 1079. £70-99. 2010. ISBN: 978-1-4398-1624-0 (CRC Press). The Mathematical Gazette, 96(536):364–365, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Murphy:2011:ACS

[Mur11a] Robin Murphy. 95.12 Accelerating the convergence of some alternating series. The Mathematical Gazette, 95(532):91–96, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Murphy:2011:NIS

[Mur11b] Robin Murphy. 95.62 Numerical integration in a spreadsheet. The Mathematical Gazette, 95(534):528–532, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Murphy:2011:I

[Mur11c] Robin V. W. Murphy. 95.33 Integrating ∫_{0}^{∞} | sin^{a}(x) + cos^{b}(x)|/x^{p} dx. The Mathematical Gazette, 95(533):307–311, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Murphy:2012:EUC

[Mur12] R. V. W. Murphy. 96.37 Evaluating I(α) = ∫_{0}^{∞} 1/(1 + x^{α})^{2} dx using complex numbers. The Mathematical Gazette, 96(536):287– 289, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Murphy:2017:ISP

[Mur17] R. V. W. Murphy. 101.46 improved simple pendulum results. The Mathematical Gazette, 101(552):525–528, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10146-improved-simple-pendulum-
results/A09F9D3B25EE946C623B76F739FA8E73.

Murphy:2019:IEN

[Mur19] Robin Murphy. 103.45 Improving elementary numerical integration using numerical differentiation. The Mathematical Gazette, 103(558):548–556, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10345-improving-elementary-numerical-integration-
using-numerical-differentiation/4AF6B0B19C79ACE00EE9E8C7A15051DC.

MacHale:2012:GRA

[MvdB12] Des MacHale and Christian van den Bosch. 96.01 Generalising a result about Pythagorean triples. The Mathematical Gazette, 96 (535):91–96, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Matkowski:2018:FVS

[MW18] Janusz Matkowski and Alfred Witkowski. 102.48 the function f(x)/x is v-shaped for convex f. The Mathematical Gazette, 102(555):517–520, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10248-the-function-fx-x-is-vshaped-for-convex-
f/23CE2324D473E5BAD210A9077C98A895.

Naidoo:2012:BRB

[Nai12] Jayaluxmi Naidoo. Book review: Origami inspirations, by Meenakshi Mukerji. Pp. 120. £15.99. 2010. ISBN: 978-1-56881-584-8 (A. K. Peters). The Mathematical Gazette, 96(536):383, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nelson:2008:PDM

[Nel08] R. D. (R. David) Nelson, editor. The Penguin dictionary of mathematics. Penguin reference library. Penguin, London, UK and New York, NY, USA, fourth edition, 2008. ISBN 0-14-103023-2 (paperback). 480 pp. LCCN QA5 .P4425 2008.

Neumann:2011:MWE

[Neu11] Peter M. Neumann. The mathematical writings of Évariste Galois. European Mathematical Society, Zürich, Switzerland, 2011. ISBN 3-03719-104-X. x + 410 pp. LCCN QA3 .G18 2011.

Neumann:2016:IT

[Neu16] Peter Neumann. Inspiring teachers. The Mathematical Gazette, 100(549):385–395, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.

cambridge.org/core/product/008EB977AD31BE1B99B830B20E685D1D.

Neal:2015:CCC

[NG15] David K. Neal and Dustin Gentile. Crapoulette: card craps with one deck. The Mathematical Gazette, 99(545):196–204, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nicholson:2008:LME

[Nic08a] James Nicholson. A-level mathematics for Edexcel: statistics S1. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-911782-9 (paperback). ???? pp. LCCN ????

Nicholson:2008:S

[Nic08b] James (James R.) Nicholson. Statistics. A Level mathematics for Edexcel. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2008. ISBN 0-19-911782-9 (S1: paperback), 0-19-911787-X (S2: paperback). ???? pp. LCCN ????

Nickalls:2011:NBP

[Nic11] R. W. D. Nickalls. 95.60 A new bound for polynomials when all the roots are real. The Mathematical Gazette, 95(534):520–526, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nicholson:2012:IAA

[Nic12a] W. Keith Nicholson. Introduction to abstract algebra. John Wiley, New York, NY, USA, fourth edition, 2012. ISBN 1-118-13535-0 (hardcover). xxiv + 535 pp. LCCN QA162 .N53 2012.

Nickalls:2012:QEA

[Nic12b] R. W. D. Nickalls. The quartic equation: alignment with an equivalent tetrahedron. The Mathematical Gazette, 96(535):49– 55, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nicollier:2014:RSS

[Nic14a] Grégoire Nicollier. 98.06 The relative size of the squares inscribed in a triangle. The Mathematical Gazette, 98(541):117–119, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nicollier:2014:TTA

[Nic14b] Grégoire Nicollier. 98.32 triangles with two angles in the ratio 1: 2. The Mathematical Gazette, 98(543):508–509, November 2014. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nicollier:2015:GPI

[Nic15] Grégoire Nicollier. 99.31 a generalisation of Proizvolov’s identity. The Mathematical Gazette, 99(546):525–526, November 2015. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nicollier:2016:EDR

[Nic16] Grégoire Nicollier. 100.09 extremal distance ratios. The Mathematical Gazette, 100(547):129–130, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nicollier:2017:HPO

[Nic17] Grégoire Nicollier. 101.21 a homothetic property of orthic triangles. The Mathematical Gazette, 101(551):303–306, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
570A29A9AFB8B12C5FDBF7AE7D14E8ED.

Nievergelt:2010:TTL

[Nie10] Yves Nievergelt. 94.38 The truth table of the logical implication. The Mathematical Gazette, 94(531):509–513, November 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nillsen:2013:PSC

[Nil13] Rodney Nillsen. 97.22 A proof that the series ∑_{n=1}^{∞} 1/n^{p} is convergent for p > 1. The Mathematical Gazette, 97(539):273–274, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nimbran:2016:SOS

[NL16] Amrik Nimbran and Paul Levrie. 100.29 some odd series for π. The Mathematical Gazette, 100(549):496–500, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
FDFC8970DA45F3C47352E48CA387495C.

Nelsen:1987:PSR

[NS87] Roger B. Nelsen and James E. Schultz. The probability that the “sum of the rounds” equals the “round of the sum”. College Mathematics Journal, 18(5):390–396, November 1987. CO-DEN ???? ISSN 0746-8342 (print), 1931-1346 (electronic).

URL http://www.jstor.org/stable/2686963;   http://www.
tandfonline.com/doi/abs/10.1080/07468342.1987.11973061.

Nunes:2013:ICS

[Nun13] Katharine Nunes. 97.28 Inscribed circles and spheres. The Mathematical Gazette, 97(539):290–294, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nurcombe:2013:ASA

[Nur13] J. R. Nurcombe. Alteration in the sum of alternating series by simple rearrangement. The Mathematical Gazette, 97(539):193– 197, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nurcombe:2014:RSA

[Nur14a] J. R. Nurcombe. 98.11 Rearranging the signs of the alternating harmonic series. The Mathematical Gazette, 98(542):321–324, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nurcombe:2014:NCC

[Nur14b] J. R. Nurcombe. 98.27 necessary conditions for the convergence of series of positive terms. The Mathematical Gazette, 98(543): 493–496, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Nurcombe:2016:CCT

[Nur16] J. R. Nurcombe. 100.04 converses of a ‘curious’ theorem of Cesàro on conditionally convergent series. The Mathematical Gazette, 100 (547):113–118, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Oai:2015:FNT

[Oai15] Dao Thanh Oai. 99.09 A family of Napoleon triangles associated with the Kiepert configuration. The Mathematical Gazette, 99 (544):151–153, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Osler:2014:RFR

[OC14] Thomas J. Osler and Tirupathi R. Chandrupatla. 98.23 recursive formulas related to the Arithmetic–Geometric Mean. The Mathematical Gazette, 98(543):484–486, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Osler:2013:DSD

[OD13] Thomas J. Osler and Steven Donahue. 97.09 The devil’s series — did it fool Euler? The Mathematical Gazette, 97(538):133– 137, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Osler:2016:IPN

[OJN16] Thomas J. Osler, Walter Jacob, and Ryo Nishimura. An infinite product of nested radicals for log x from the Archimedean algorithm. The Mathematical Gazette, 100(548):274–278, July 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
B4FFD1BF546538EDD0EB6D5F4B94F7CD.

Oktem:2015:PS

[Ökt15] I. Ferit Öktem. On the product ∏_{n=2}^{∞} ζ(n). The Mathematical Gazette, 99(546):422–426, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Okumura:2013:AGA

[Oku13a] Hiroshi Okumura. 97.14 The area of a generalised arbelos. The Mathematical Gazette, 97(538):157–160, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Okumura:2013:ATC

[Oku13b] Hiroshi Okumura. 97.47 Archimedean twin circles in the arbelos. The Mathematical Gazette, 97(540):512, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

OLoughlin:2010:HAI

[O’L10] Michael O’Loughlin. 94.10 Half angles and the inradius of a Pythagorean triangle. The Mathematical Gazette, 94(529):144– 146, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ormell:2008:EIL

[Orm08] C. P. (Christopher Peter) Ormell. Exotic infinity and a long decline of confidence in mathematics, volume 4 of PER review. Ingleside-Ashby, London, UK, 2008. ISBN 0-907669-46-8 (paperback). iii + 15 pp. LCCN ????

Ormell:2013:MTW

[Orm13] Christopher Ormell. Mathematics, a third way. The Mathematical Gazette, 97(538):101–106, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ormell:2014:A

[Orm14] Christopher Ormell. 98.26 algorithms for |x| and [x]. The Mathematical Gazette, 98(543):490–492, November 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Orzel:2012:HTR

[Orz12] Chad Orzel. How to teach relativity to your dog. Basic Books, New York, NY, USA, 2012. ISBN 0-465-02331-2 (paperback), 0-465-02937-X (e-book). vi + 327 pp. LCCN QC173.57 .O79 2012.

Oxman:2015:ESC

[OS15] Victor Oxman and Moshe Stupel. Elegant special cases of Van Aubel’s theorem. The Mathematical Gazette, 99(545):256–262, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Osborne:2013:LCS

[Osb13] M. Scott Osborne. Locally convex spaces. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 3-319-02044-7 (hardcover), 3-319-02045-5 (ebook). viii + 213 pp. LCCN QA322 .O73 2014. URL http:

//www.loc.gov/catdir/enhancements/fy1404/2013951652-b.
html;   http://www.loc.gov/catdir/enhancements/fy1404/
2013951652-d.html;   http://www.loc.gov/catdir/enhancements/
fy1404/2013951652-t.html.

Osler:2010:TJW

[Osl10a] Thomas J. Osler. The tables of John Wallis and the discovery of his product π. The Mathematical Gazette, 94(531):430–437, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Osler:2010:VLP

[Osl10b] Thomas J. Osler. Vieta-like products of nested radicals. The Mathematical Gazette, 94(529):62–66, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Osler:2011:EPO

[Osl11a] Thomas J. Osler. Euler’s parallel oblique-angled diameters. The Mathematical Gazette, 95(532):40–48, March 2011. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Osler:2011:MLB

[Osl11b] Thomas J. Osler. Morphing Lord Brouncker’s continued fraction for π into the product of Wallis. The Mathematical Gazette, 95 (532):17–22, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Osler:2012:PSS

[Osl12a] Thomas J. Osler. 96.53 Partial sums of series that cannot be an integer. The Mathematical Gazette, 96(537):515–519, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Osler:2012:MFB

[Osl12b] Thomas J. Osler. The missing fractions in Brouncker’s sequence of continued fractions for π. The Mathematical Gazette, 96(536): 221–225, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Osler:2015:SFD

[Osl15] Thomas J. Osler. Short and fuzzy derivations of five remarkable formulas for primes. The Mathematical Gazette, 99(545):220–225, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Osler:2016:BTA

[Osl16] Thomas J. Osler. Bisecting and trisecting the arc of the lemniscate. The Mathematical Gazette, 100(549):471–481, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
70DF18422939FA7D9856CB1C3C770C7D.

Osler:2014:VPP

[OW14] Thomas J. Osler and Sky Waterpeace. Vieta’s product for pi from the Archimedean algorithm. The Mathematical Gazette, 98(543): 429–431, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Pasles:2008:BFN

[Pas08] Paul C. Pasles. Benjamin Franklin’s numbers: an unsung mathematical odyssey. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-12956-8 (hardcover). 254 + 8 pp. LCCN QA24

.P37 2008. URL http://www.loc.gov/catdir/enhancements/
fy0726/2006102508-d.html;   http://www.loc.gov/catdir/
enhancements/fy0734/2006102508-b.html;   http://www.loc.
gov/catdir/toc/ecip077/2006102508.html.

Paseau:2011:FP

[Pas11a] A. C. Paseau. Family planning. The Mathematical Gazette, 95 (533):213–217, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Pask:2011:MFF

[Pas11b] Colin Pask. Math for the frightened: facing scary symbols and everything else that freaks you out about mathematics: [recognize and overcome your math anxieties; why does math need those crazy symbols?; find the patterns and beauty in math; visualize the mathematical world]. Prometheus Books, Amherst, NY, USA, 2011. ISBN 1-61614-421-1 (paperback), 1-61614-422-X (e-book). 380 pp. LCCN QA41 .P37 2011.

Pask:2013:MPE

[Pas13] Colin Pask. Magnificent Principia: exploring Isaac Newton’s masterpiece. Prometheus Books, Amherst, NY, USA, 2013. ISBN 1-61614-745-8 (hardcover). 528 pp. LCCN QA803 .P37 2013.

Paseau:2014:BRJ

[Pas14] A. C. Paseau. Book review: L. E. J. Brouwer: Topologist, intuitionist, philosopher by Dirk van Dalen, pp 875, £18.99, ISBN 978-1-4471-4615-5, Springer-Verlag (2013). The Mathematical Gazette, 98(543):552–554, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Paseau:2016:LGM

[Pas16] A. C. Paseau. Letter games: a metamathematical taster. The Mathematical Gazette, 100(549):442–449, November 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
EBB109D4AFF3FFCA093EB536C3561E5B.

Paseau:2018:BRP

[Pas18] A. C. Paseau. Book review: Philosophy of mathematics by Øystein Linnebo, pp. 203, £24.95 (hard), ISBN 978-0-691-16140-2, Princeton University Press (2017). The Mathematical Gazette, 102(554): 379–381, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/philosophy-
of-mathematics-by-oystein-linnebo-pp-203-2495-hard-isbn-
9780691161402-princeton-university-press-2017/BFE9DF446991402A210CCEAF1602D7CC.

Patterson:2013:TRI

[Pat13a] T. N. L. Patterson. 97.27 Two rectangles and an intersection. The Mathematical Gazette, 97(539):284–290, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Patyi:2013:EBP

[Pat13b] Imre Patyi. 97.45 On Euler’s Basel problem. The Mathematical Gazette, 97(540):508–510, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Patyi:2015:SEF

[Pat15] Imre Patyi. On some elementary functions. The Mathematical Gazette, 99(545):263–275, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Patterson:2017:BRM

[Pat17] Roger Patterson. Book reviews: Mathematics HL: Sets, relations & groups, by Quinn, Catherine; Blythe, Peter; Sangwin, Chris; Haese, Robert and Haese, Michael pp. 176, AU$41.00 (paper), ISBN 978-1-92197-232-4, Haese Mathematics (2013). Mathematics HL: Calculus, by Quinn, Catherine; Blythe, Peter; Sangwin, Chris; Haese, Robert and Haese Michael, pp. 216, AU$41.00 (paper), ISBN 978-1-92197-233-1, Haese Mathematics (2013). Mathematics HL: Linear algebra & geometry, by Quinn, Catherine; Blythe, Peter; Sangwin, Chris; Haese Robert and Haese Michael, pp. 344, AU$54.50 (paper), ISBN 978-1-92197-235-5, Haese Mathematics (2013). Mathematics HL: Discrete mathematics, by Quinn, Catherine; Blythe, Peter; Sangwin, Chris; Haese Robert and Haese Michael pp. 240, AU$41.00 (paper), ISBN 978-1-92197-234-8, Haese Mathematics (2013). Mathematics HL: Statistics & probability, by Quinn, Catherine; Blythe, Peter; Sangwin, Chris; Haese, Robert and Haese Michael, pp 208, AU$41.00 (paper), ISBN 978-1-92197-231-7, Haese Mathematics (2013). The Mathematical Gazette, 101(551):351–353, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
574010DC256776D34F6A4CD190232171.

Paseau:2014:SAG

[PdVF14] A. C. Paseau, Michael de Villiers, and Michael Fox. The ‘stop after k girls or N children’ policy. The Mathematical Gazette, 98(543): 402–423, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Pearce:2011:NMF

[Pea11] Chris Pearce. New maths frameworking. Level 6. Booster workbook. Collins Educational, London, UK, 2011. ISBN 0-00-743808-7 (paperback). ???? pp. LCCN ????

Percy:2012:GHC

[Per12] Andrew Percy. 96.25 Going halfway with circular boundaries. The Mathematical Gazette, 96(535):159–160, March 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Petkovic:2009:FPG

[Pet09] Miodrag Petković. Famous puzzles of great mathematicians. American Mathematical Society, Providence, RI, USA, 2009. ISBN 0-8218-4814-3. xviii + 325 pp. LCCN QA95 .P4358 2009.

Proctor:2008:EGM

[PF08] John Proctor and Tina Foster. Edexcel GCSE maths foundation tier: revision and classroom companion. Letts and Lonsdale, ????, 2008. ISBN 1-84419-191-5 (paperback). 120 pp. LCCN ????

Plaza:2010:PWT

[PF10] Ángel Plaza and Sergio Falcón. 94.18 Proof without words: Two inequalities proved by convexity. The Mathematical Gazette, 94 (530):306–308, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Pal:2017:AWA

[PG17] Suman Pal and Baptiste Goujaud. 101.33 the application of the Weierstrass approximation theorem in the Riemann–Lebesgue lemma. The Mathematical Gazette, 101(552):489–492, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
843F13A492FDAC31EAB7973C5FBA2E94.

Perkins:2017:C

[PH17] Martin Perkins and David Hopkins. Correspondence. The Mathematical Gazette, 101(551):335–336, July 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
633ED15CCBD7DE8D73B57D05027116A0.

Piskorz:2016:SPN

[Pis16] Pawel Jan Piskorz. 100.02 square pyramidal numbers and the multiplication table. The Mathematical Gazette, 100(547):108– 111, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Pitici:2011:BWM

[Pit11] Mircea Pitici, editor. The best writing on mathematics 2010. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14841-4. xxviii + 440 pp. LCCN QA8.6 .B476 2011. Foreword by William P. Thurston.

Pitici:2012:BWM

[Pit12] Mircea Pitici, editor. The best writing on mathematics 2011. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-15315-9 (paperback). xxx + 383 pp. LCCN QA8.6.

Pitici:2014:BWM

[Pit14] Mircea Pitici, editor. The best writing on mathematics 2013. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-16041-4 (paperback), 1-306-15416-2 (e-book), 1-4008-4799-0 (e-book). xxvi + 244 pp. LCCN QA8.6 .B47 2014; QA36.B4 2014;

QA93 .B47 2013. URL http://proquest.safaribooksonline.
com/?fpi=9780691160412;   http://www.jstor.org/stable/
10.2307/j.ctt4cgb74.

Posamentier:2012:STM

[PL12] Alfred S. Posamentier and Ingmar Lehmann. The secrets of triangles: a mathematical journey. Prometheus Books, Amherst, NY, USA, 2012. ISBN 1-61614-587-0 (hardcover). 387 pp. LCCN QA531 .P67 2012.

Posamentier:2013:MMM

[PL13] Alfred S. Posamentier and Ingmar Lehmann. Magnificent mistakes in mathematics. Prometheus Books, Amherst, NY, USA, 2013. ISBN 1-61614-747-4 (hardcover). 296 pp. LCCN QA99 .P664 2013.

Pla:2010:C

[Pla10a] Juan Pla. Correspondence. The Mathematical Gazette, 94(531): 536, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Pla:2010:FTD

[Pla10b] Juan Pla. A footnote to the theory of double integrals. The Mathematical Gazette, 94(530):262–269, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Pla:2010:KSD

[Pla10c] Juan Pla. A kaleidoscope of solutions for a Diophantine system. The Mathematical Gazette, 94(529):42–50, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Pla:2011:SSO

[Pla11a] Juan Pla. 95.07 Series for π ln k and some other mathematical constants. The Mathematical Gazette, 95(532):78–82, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Pla:2011:MC

[Pla11b] Juan Pla. 95.51 On a multigrade congruence. The Mathematical Gazette, 95(534):494–496, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Pla:2011:SRA

[Pla11c] Juan Pla. Some remarks about a formula of Charles Dodgson. The Mathematical Gazette, 95(533):235–239, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Pla:2012:SAF

[Pla12] Juan Pla. On a sequence of algebraic formulae of Ramanujan. The Mathematical Gazette, 96(536):201–206, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Pla:2014:SSR

[Pla14] Juan Pla. On some subsets of the rational solutions of the equations aX^{2} + bXY + cY^{2} = dZ^{2}. The Mathematical Gazette, 98 (543):424–428, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Plaza:2016:VPL

[Pla16] Ángel Plaza. 100.12 visual proof of the limit of f-mean recurrence sequences. The Mathematical Gazette, 100(547):139–141, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Plaza:2018:PWA

[Pla18] Ángel Plaza. 102.42 proof without words: An alternating geometrical series. The Mathematical Gazette, 102(555):504–505, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10242-proof-without-
words-an-alternating-geometrical-series/268E082309DBA3856112A47EF5F1E027.

Plofker:2009:MI

[Plo09] Kim Plofker. Mathematics in India. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-12067-6 (hardcover). xi + 357 pp. LCCN QA27.I4 P56 2009. URL http://www.loc.gov/

catdir/toc/ecip0821/2008028186.html.

Pitici:2013:BWM

[PM13] Mircea Pitici and David Mumford, editors. The best writing on mathematics, 2012. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 1-4008-4467-3 (e-book), 1-283-65678-7, 0-691-15655-7. xxix + 288 pp. LCCN QA8.6 .B476 2012; QA36; QA93

.B476 2012. URL http://press.princeton.edu/TOCs/c9821.
html.

Popko:2012:DSG

[Pop12] Edward Popko. Divided spheres: geodesics and the orderly subdivision of the sphere. CRC Press/Taylor and Francis Group, Boca Raton, 2012. ISBN 1-4665-0429-3 (hardcover). xx + 509 pp. LCCN TA660.P73 P67 2012.

Polster:2010:MR

[PR10] Burkard Polster and Marty Ross. Mathematical rugby. The Mathematical Gazette, 94(531):450–463, November 2010. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Polster:2012:MGM

[PR12] Burkard Polster and Marty Ross. Math goes to the movies. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0483-4, 1-4214-0484-2 (paperback). xiv + 286 pp. LCCN PN1995.9.M395 P65 2012.

Parramore:2014:TGV

[PS14] Keith Parramore and Joan Stephens. Two girls — the value of information. The Mathematical Gazette, 98(542):243–249, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Patel:2011:CQ

[PU11] Pratik Patel and Robert Underwood. 95.43 On convex quadrilaterals. The Mathematical Gazette, 95(533):330–333, July 2011. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Puritz:2011:MYO

[Pur11] C. W. Puritz. 95.04 Make your own log table! The Mathematical Gazette, 95(532):66–72, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Puritz:2018:COS

[Pur18] Christian Puritz. 102.55 cable to outer space? The Mathematical Gazette, 102(555):535–537, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10255-cable-to-outer-space/B9E47664725F276AF706652449617589.

Purwanto:2019:CTA

[Pur19] Purwanto. 103.33 Conditions for the terms of an arithmetic sequence to form a multiplicative modular group. The Mathematical Gazette, 103(558):504–508, November 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10333-conditions-for-the-
terms-of-an-arithmetic-sequence-to-form-a-multiplicative-
modular-group/A6CAE0F07BAAEB077ACA408164CDCB4E.

Price:2013:RGM

[PW13] Mike Price and Mary Walmsley. R. L. Goodstein and the Mathematical Association. The Mathematical Gazette, 97(540):398–408, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Quadling:2012:GA

[Qua12a] Douglas Quadling. 96.24 Genealogical algebra. The Mathematical Gazette, 96(535):156–159, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Quadling:2012:CME

[Qua12b] Douglas Quadling. A curious misattribution: the early history of ‘Simson’s line’. The Mathematical Gazette, 96(537):420–427, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Quadling:2012:EK

[Qua12c] Douglas Quadling. Edwin Kerr (1926–2012). The Mathematical Gazette, 96(537):551–552, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Quadling:2013:IR

[Qua13] Douglas Quadling. 97.36 The isles to the road. The Mathematical Gazette, 97(539):315–317, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Rohde:2012:IIC

[R^{+}12] Ulrich L. Rohde et al. Introduction to integral calculus: systematic studies with engineering applications for beginners. John Wiley, New York, NY, USA, 2012. ISBN 1-118-11776-X (hardcover). xxv + 402 pp. LCCN QA308 .I58 2012.

Rajeev:2013:AME

[Raj13] S. G. (Sarada G.) Rajeev. Advanced mechanics: from Euler’s determinism to Arnold’s chaos. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-967085-4 (hardcover), 0-19-967086-2 (paperback). xiii + 163 pp. LCCN QA805 .R18 2013.

Rajan:2018:AEP

[Raj18] Ashvin Rajan. 102.21 an addendum to Estermann’s proof of the irrationality of \sqrt{2}. The Mathematical Gazette, 102(554):308–309, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10221-an-addendum-
to-estermanns-proof-of-the-irrationality-of/1458288BB61A0E9D324C89A6A89AD867.

Ramsden:2009:WTS

[Ram09] Jenny Ramsden. Whoever thought of that?: short biographies of some contributors to the history of mathematics. Mathematical Association, Leicester, UK, 2009. ISBN 0-906588-68-5 (Paper). 86 pp. LCCN QA28 .R36 2009. With supplementary notes by Nick Lord, Martin Perkins, and Chris Weeks.

Ransom:2014:TTT

[Ran14] Peter Ransom. Triumphs and tribulations in teaching. The Mathematical Gazette, 98(543):386–395, November 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Range:2018:UHS

[Ran18] R. Michael Range. Using high school algebra for a natural approach to derivatives and continuity. The Mathematical Gazette, 102(555):435–446, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/using-high-school-algebra-for-a-natural-approach-
to-derivatives-and-continuity/2FE7BBB725BFD77E00B368549997E132.

Razminia:2019:HIR

[Raz19] K. Razminia. 103.35 Hölder’s inequality revisited. The Mathematical Gazette, 103(558):512–514, November 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10335-holders-inequality-
revisited/CBE9EF5CD4E46405D527395937134D1A.

Read:2010:DT

[Rea10a] Emrys Read. 94.01 On digit transfers. The Mathematical Gazette, 94(529):114–117, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Reale:2010:SGG

[Rea10b] Thomas Reale. Six gems of geometry: with imaginary dialogue between Newton, Euclid, and the old gods. PS Press, Portland, OR, USA, 2010. ISBN 1-935638-02-5. vi + 126 pp. LCCN QA451 .R42 2010.

Read:2014:IT

[Rea14] Emrys Read. 98.02 On integer triangles. The Mathematical Gazette, 98(541):107–112, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Read:2015:PIT

[Rea15] Emrys Read. 99.12 Primitive integer triangles. The Mathematical Gazette, 99(545):317–319, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Read:2016:HT

[Rea16] Emrys Read. 100.01 Heronian triangles. The Mathematical Gazette, 100(547):103–108, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Read:2019:E

[Rea19] Emrys Read. On the equation x^{2} + ry^{2} = z^{2}. The Mathematical Gazette, 103(556):87–93, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/on-the-equation-x2-ry2-z2/F10C6540ED2EBBCFCF58515BE2B92B15.

Rebaza:2012:FCA

[Reb12] Jorge Rebaza. A first course in applied mathematics. John Wiley, New York, NY, USA, 2012. ISBN 1-118-22962-2. xvi + 439 pp. LCCN TA342 .R43 2012. URL http:

//www.loc.gov/catdir/enhancements/fy1201/2011043340-d.
html;   http://www.loc.gov/catdir/enhancements/fy1201/
2011043340-t.html;   http://www.loc.gov/catdir/enhancements/
fy1210/2011043340-b.html.

Reddy:2015:GIR

[Red15] Varun Amar Reddy. 99.28 a generalisation of an intriguing ratio. The Mathematical Gazette, 99(546):516–522, November 2015. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Reimer:2014:CLE

[Rei14] David Reimer. Count like an Egyptian: a hands-on introduction to ancient mathematics. Princeton University Press, Princeton, NJ, USA, 2014. ISBN 0-691-16012-0 (hardcover), 1-4008-5141-6 (e-book). xiii + 3 + 237 pp. LCCN QA22 .R38 2014. URL http:

//www.jstor.org/stable/10.2307/j.ctt5vjvdr.

Retkes:2017:GSP

[Ret17] Zoltan Retkes. 101.24 generalisation and sharpening of Problem 100.C. The Mathematical Gazette, 101(551):310–313, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
B8E70BA9274F6425C7EC7A96652DDF89.

Retkes:2018:GTI

[Ret18] Zoltan Retkes. Generalising a triangle inequality. The Mathematical Gazette, 102(555):422–427, November 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/generalising-a-triangle-inequality/
7B82A24E699FDC181C32BDAB14F84856.

Rajagopal:2014:FNF

[RG14] Surajit Rajagopal and Martin Griffiths. 98.01 On Fibonacci numbers that are factorials. The Mathematical Gazette, 98(541):104– 107, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ribeiro:2018:APF

[Rib18] Pedro Ribeiro. 102.29 another proof of the Fundamental Theorem of Algebra. The Mathematical Gazette, 102(554):326–328, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10229-another-proof-
of-the-fundamental-theorem-of-algebra/3429DB58512040772AE5448C754C94B8.

Ribeiro:2019:ACF

[Rib19] Pedro Ribeiro. 103.08 another classic Fourier analysis proof of the Basel Problem. The Mathematical Gazette, 103(556):142–145, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10308-another-classic-
fourier-analysis-proof-of-the-basel-problem/E1E1970740FA77029A3CD798694BB954.

Richardson:2008:ACI

[Ric08a] Leonard F. Richardson. Advanced calculus: an introduction to linear analysis. Wiley-Interscience, New York, NY, USA, 2008.

ISBN 0-470-23288-9 (hardcover). xxv + 387 pp. LCCN QA303.2

.R53 2008. URL http://www.loc.gov/catdir/enhancements/
fy0810/2008007377-d.html;   http://www.loc.gov/catdir/
enhancements/fy0810/2008007377-t.html;   http://www.loc.
gov/catdir/enhancements/fy0827/2008007377-b.html.

Richeson:2008:EGP

[Ric08b] David S. (David Scott) Richeson. Euler’s gem: the polyhedron formula and the birth of topology. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-12677-1. xii + 317 pp. LCCN QA611.A3 R53 2008. URL http:

//www.loc.gov/catdir/enhancements/fy0834/2008062108-b.
html;   http://www.loc.gov/catdir/enhancements/fy0834/
2008062108-d.html.

Richardson:2010:DF

[Ric10] Bill Richardson. Doug French. The Mathematical Gazette, 94(529): 1–4, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Rohde:2012:IDC

[RJPG12] Ulrich L. Rohde, G. C. Jain, Ajay K. Poddar, and A. K. Ghosh. Introduction to differential calculus: systematic studies with engineering applications for beginners. John Wiley, New York, NY, USA, 2012. ISBN 1-118-11775-1 (hardcover). xxix + 747 pp. LCCN QA304 .I59 2012.

Robin:2010:BWB

[Rob10] Anthony C. Robin. 94.25 Binomial within binomial yields binomial. The Mathematical Gazette, 94(530):327–330, July 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Robin:2012:SFL

[Rob12] Anthony C. Robin. 96.57 Stability of flotation, a look at the metacentre. The Mathematical Gazette, 96(537):529–536, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Robin:2013:EMD

[Rob13] Anthony C. Robin. 97.37 End moves in a dice game. The Mathematical Gazette, 97(539):318–324, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Rocchi:2014:JFP

[Roc14] Paolo Rocchi. Janus-Faced Probability. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014. ISBN 3-319-04860-0, 3-319-04861-9 (e-book). xiii + 147 pp. LCCN QA273.4 .R63 2014.

Ronan:2007:SMO

[Ron07] Mark Andrew Ronan. Symmetry and the monster: one of the greatest quests of mathematics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-280723-4 (paperback), 0-19-280722-6 (hardcover), 1-4294-8724-0 (e-book). vii + 255 pp. LCCN QA174.7.S96 R66 2006. URL http:/

/catdir.loc.gov/catdir/enhancements/fy0636/2006005354-
d.html;   http://catdir.loc.gov/catdir/enhancements/fy0723/
2006005354-b.html;   http://catdir.loc.gov/catdir/toc/
ecip069/2006005354.html;   http://catdir.loc.gov/catdir/
toc/fy0614/2006005354.html.

Roper:2018:ASS

[Rop18] Tom Roper. Adventures in shape and space — and time. The Mathematical Gazette, 102(555):385–400, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/adventures-in-shape-and-space-
and-time/450BA3657B04098432868E1BB0F1CA58.

Rosenhouse:2009:MHP

[Ros09] Jason Rosenhouse. The Monty Hall problem: the remarkable story of math’s most contentious brainteaser. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-536789-8 (hardcover), 0-19-970990-4 (e-book). xii + 194 pp. LCCN QA95 .R67 2009.

Roth:2010:BRBd

[Rot10a] Bruce Roth. Book review: Crosspacks & cross hypersolids, by Patrick Taylor. Pp. 80. £16.95. 2007. ISBN 0-9516701-8-2 (Nattygrafix, Ipswich). The Mathematical Gazette, 94(531):567, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Roth:2010:BRBc

[Rot10b] Bruce Roth. Book review: Decision D1: A Level mathematics for Edexcel, by Brian Jefferson. Pp. 192. £10. 2008. ISBN: 978-0-19-911780-2 (Oxford University Press). The Mathematical Gazette, 94 (530):356, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Roth:2010:BRBa

[Rot10c] Bruce Roth. Book review: Euler’s gem, by David S. Richeson. Pp. 336. £16.95. 2008. ISBN 978-0-691-12677-7 (Princeton University Press). The Mathematical Gazette, 94(529):176–177, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Roth:2010:BRBb

[Rot10d] Bruce Roth. Book review: Flips for 3-folds and 4-folds, Alessio Corti (ed.). Pp. 200. £41.00. 2007. ISBN 978-0-19-857061-5 (Oxford University Press). The Mathematical Gazette, 94(529):188– 189, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Routledge:2011:GFP

[Rou11] Norman Routledge. 95.49 Generalised Fibonacci properties. The Mathematical Gazette, 95(534):483–486, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Rout:2013:BRB

[Rou13a] Stephen Rout. Book review: Fuelling innovation and discovery. The mathematical sciences in the 21st century, by The National Research Council of the National Academies. Pp. 57. £13.99. 2012. ISBN 978-0-309-25473-1 (The National Academies Press.). The Mathematical Gazette, 97(540):558–559, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Routledge:2013:IFS

[Rou13b] Norman Routledge. 97.15 Iterating to find sines and cosines. The Mathematical Gazette, 97(538):160–162, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Routledge:2013:PSD

[Rou13c] Norman Routledge. 97.38 The 6666 problem and state diagrams. The Mathematical Gazette, 97(539):324–328, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Rout:2014:BRP

[Rou14] Stephen Rout. Book review: A problem solver’s handbook. A guide to intermediate mathematical olympiads, Andrew Jobbings. Pp.

264. £14 (pbk) ISBN: 978-1-90600-1193 (UKMT). The Mathematical Gazette, 98(543):570–572, November 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Rout:2016:BRT

[Rou16] Stephen Rout. Book review: Topics in algebra and analysis by Radmila Bulajich Manfrino, José Antonio Gómez Ortega and Rogelio Valdez Delgado, pp. 303, £40.99 (paper), ISBN 978-3-319-11945-8, Birkhauser (2015). The Mathematical Gazette, 100 (547):187–189, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Rowland:2019:CVS

[Row19] Dominic Rowland. Book review: Combinatorics: a very short introduction by Robin Wilson, pp. 150, £7.99 (paper), ISBN 978-0-19872-349-3, Oxford University Press (2016). The Mathematical Gazette, 103(557):375–376, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/combinatorics-a-very-short-introduction-by-robin-
wilson-pp-150-799-paper-isbn-9780198723493-oxford-university-
press-2016/BA9357277B5801C8CF67862E993B7154.

Rockett:2011:SGF

[RR11] Andrew M. Rockett and Joseph P. Ruggerio. 95.30 On Stråhle’s guitar frets. The Mathematical Gazette, 95(533):300–303, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Rosenhouse:2011:TSS

[RT11] Jason Rosenhouse and Laura Taalman. Taking sudoku seriously: the math behind the world’s most popular pencil puzzle. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-975656-2 (hardcover). xii + 214 pp. LCCN GV1507.S83 R67 2011.

Ruane:2010:BRB

[Rua10] P. N. Ruane. Book review: Lewis Carroll in Numberland. His fantastical mathematical logical life, by Robin Wilson. pp. 237.

£16.99. 2008. ISBN 978-0-713-99757-6 (Allen Lane). The Mathematical Gazette, 94(530):359–360, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ruane:2011:BRB

[Rua11] P. N. Ruane. Book review: The mathematics of the heavens and the earth: The early history of trigonometry, by Glen van Brummelen. Pp. 329. £23.95 (hardback). 2009. ISBN 978-0-691-12973-0 (Princeton University Press). The Mathematical Gazette, 95 (532):150–151, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ruane:2012:BRB

[Rua12] P. N. Ruane. Book review: Newton’s philosophy of nature: selections from his writings, by Isaac Newton (edited by H. S. Thayer). Pp. 207. $12-95. 2008. ISBN 978-0-486-44593-9 (Dover). The Mathematical Gazette, 96(536):367–368, July 2012. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ruane:2013:BRB

[Rua13] P. N. Ruane. Book review: Ancient Indian leaps into mathematics, B. S. Yadav and Man Mohan (eds). Pp. 217. £66.99 (Hardback). 2011 ISBN: 978-0-8176-4694-3 (Birkhäuser). The Mathematical Gazette, 97(539):362–363, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ruane:2014:BRBb

[Rua14a] P. N. Ruane. Book review: Applications of combinatorial matrix theory to Laplacian matrices of graphs, by Jason J. Molitierno. Pp. 405. £57.99 (Hardback). 2012. ISBN: 978-1-4398-6337-4 (CRC Press). The Mathematical Gazette, 98(541):179–180, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ruane:2014:BRH

[Rua14b] P. N. Ruane. Book review: Heavenly mathematics: the forgotten art of spherical trigonometry by Glen Van Brummelen, pp 192, £24.95, ISBN 978-0-691-14892-2, Princeton University Press (2013). The Mathematical Gazette, 98(543):561–562, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ruane:2014:BRBa

[Rua14c] P. N. Ruane. Book review: Motifs of mathematics: history and application of the mediant and the Farey sequence, by Scott B. Guthery. Pp. 243. £11.32. 2011. ISBN 978-1-4538-1057-6. (Docent Press). The Mathematical Gazette, 98(541):163–164, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ruane:2017:BRB

[Rua17] P. N. Ruane. Book review: A brief history of numbers, by Corry, Leo, pp. 309, £24.99 (hard), ISBN 978-0-19-870259-7, Oxford University Press (2015). The Mathematical Gazette, 101 (551):357–358, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/73F6EAB4EF99EB85E509B66619844080.

Rzadkowski:2010:CBA

[Rza10] Grzegorz Rzadkowski. 94.19 A calculus-based approach to the von Staudt–Clausen theorem. The Mathematical Gazette, 94(530): 308–312, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Stubbs:2016:UPC

[SA16] John Stubbs and Jacob Adetunji. UK pension changes in 2015: some mathematical considerations. The Mathematical Gazette, 100 (548):193–202, July 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/52BFC37E606E7BA27EC0219FDA36C356.

Sangalli:2009:PRM

[San09] Arturo Sangalli. Pythagoras’ revenge: a mathematical mystery. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-04955-6. xv + 183 pp. LCCN PS3619.A566 P97 2009.

Sangwin:2011:LFD

[San11] Christopher J. Sangwin. Limit-free derivatives. The Mathematical Gazette, 95(534):469–482, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sangwin:2015:AEA

[San15] C. J. Sangwin. An audited elementary algebra. The Mathematical Gazette, 99(545):298–316, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Saouter:2017:DEF

[Sao17] Yannick Saouter. 101.02 A (doubly) elementary formula for prime numbers. The Mathematical Gazette, 101(550):93–95, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
9BD0069492B42678D9E95A2B0C5E7554.

Seshadri:2012:CPC

[SB^{+}12] C. S. Seshadri, Vikraman Balaji, et al., editors. Collected papers of C. S. [Conjeevaram Srirangachari] Seshadri. Hindustan Book Agency, New Delhi, India, 2012. ISBN 93-80250-17-7. 1008 + 633 pp. LCCN QA29.S415 A2 2012.

Surendran:2019:ERA

[SB19] Ken Surendran and Desarazu Krishna Babu. Expressions for rational approximations to square roots of integers using Pell’s equation. The Mathematical Gazette, 103(556):101–110, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/expressions-for-rational-
approximations-to-square-roots-of-integers-using-pells-
equation/E31A09448088EB6FC48753175841F5A7.

Schneps:2013:MTH

[SC13] Leila Schneps and Coralie Colmez. Math on trial: how numbers get used and abused in the courtroom. Basic Books, New York, NY, USA, 2013. ISBN 0-465-03292-3 (hardcover), 0-465-03794-1 (e-book). xi + 256 pp. LCCN K2290.S73 S36 2013.

Schilling:2010:BRB

[Sch10] René L. Schilling. Book review: A journey through statistical physics. From foundations to the quantum Hall effect: Selecta of Jürg Fröhlich. Giovanni Felder, Gian Michele Graf, Klaus Hepp (eds). Pp. 857. £126.00. 2009. ISBN: 13-978-1-4020-8886-5 (Springer). The Mathematical Gazette, 94(531):570, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Schneider:2011:PAL

[Sch11] P. (Peter) Schneider. p-Adic Lie groups, volume 344 of Grundlehren der mathematischen Wissenschaften. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. ISBN 3-642-21146-1, 3-642-21147-X (e). ISSN 0072-7830. xi + 254 pp. LCCN QA387 .S36 2011.

Schilling:2013:BRB

[Sch13] René L. Schilling. Book review: The Oxford handbook of random matrix theory, Gernot Akemann, Jinho Baik, Philippe Di Francesco (eds) Pp. 919. £110. 2011. ISBN: 978-0-19-957400-1 (Oxford University Press). The Mathematical Gazette, 97(539): 375–376, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Schilling:2014:BRM

[Sch14] René L. Schilling. Book review: Measure and integration: Publications 1997-2011 by Heinz König, pp 512, £85.50, $129.00, EUR101.60, ISBN 978-3-0348-0381-6, Birkhäuser (2012). The Mathematical Gazette, 98(543):561, November 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Schenk:2016:TST

[Sch16a] Robert Schenk. 100.34 tilings from a square template with asymmetric edges. The Mathematical Gazette, 100(549):511– 516, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/3C87DC76B6E895B8AF79CA5E51FA1D1D.

Schilling:2016:BRC

[Sch16b] René L. Schilling. Book review: A course on integration theory by Nicolas Lerner, pp. 492, £62.99 (paper), ISBN 978-3-0348-0693-0, also available as e-book, Birkhäuser (2014). The Mathematical Gazette, 100(547):181, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Schilling:2016:BRU

[Sch16c] René L. Schilling. Book review: Upper and lower bounds for stochastic processes, modern methods and classical problems by Michel Talagrand, pp. 626, £99.00 (hard), ISBN 978-3-642-54074-5, also available as e-book, Springer Verlag (2014). The Mathematical Gazette, 100(547):180–181, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scimone:2015:EWE

[Sci15] Aldo Scimone. Ellipse: what else? The Mathematical Gazette, 99 (546):481–490, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scimone:2019:IGE

[Sci19] Aldo Scimone. 103.40 Identifying golden and equilateral triangles that arise from the golden Arbelos. The Mathematical Gazette, 103(558):525–531, November 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10340-identifying-golden-
and-equilateral-triangles-that-arise-from-the-golden-arbelos/
E3BF1C6D4881071458EF3FAA9D0A588D.

Scott:2010:EEC

[Sco10a] J. A. Scott. 94.17 Evaluating the Euler constant. The Mathematical Gazette, 94(530):304–306, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2010:MTP

[Sco10b] J. A. Scott. Miscellaneous triangle properties. The Mathematical Gazette, 94(529):103–111, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2011:HSR

[Sco11a] J. A. Scott. 95.06 The harmonic series revisited. The Mathematical Gazette, 95(532):77–78, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2011:NEL

[Sco11b] J. A. Scott. 95.18 Normals to the Euler line. The Mathematical Gazette, 95(532):115–117, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2011:ISP

[Sco11c] J. A. Scott. 95.58 On infinite series over the primes. The Mathematical Gazette, 95(534):517–518, November 2011. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2011:ASI

[Sco11d] J. A. Scott. 95.59 Another series for the inverse tangent. The Mathematical Gazette, 95(534):518–520, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2012:EPR

[Sco12] J. A. Scott. 96.26 The exeter point revisited. The Mathematical Gazette, 96(535):160–161, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2013:AVM

[Sco13a] J. A. Scott. 97.32 An areal view of the Morley triangle. The Mathematical Gazette, 97(539):304–306, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2013:KHR

[Sco13b] J. A. Scott. 97.48 The Kiepert hyperbola revisited. The Mathematical Gazette, 97(540):513–515, November 2013. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2014:NRE

[Sco14a] J. A. Scott. 98.04 A nested radical for the Euler constant. The Mathematical Gazette, 98(541):114–116, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2014:AMT

[Sco14b] J. A. Scott. 98.05 Another minimum for the triangle. The Mathematical Gazette, 98(541):116–117, March 2014. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2014:CSZ

[Sco14c] J. A. Scott. 98.12 A convergent series with zero sum. The Mathematical Gazette, 98(542):325–327, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2014:UCF

[Sco14d] J. A. Scott. 98.14 An unexpected closed form. The Mathematical Gazette, 98(542):331–332, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2014:ALO

[Sco14e] J. A. Scott. 98.18 Another look at the orthocentroidal disc. The Mathematical Gazette, 98(542):343–345, July 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2014:II

[Sco14f] J. A. Scott. 98.24 an intriguing inequality. The Mathematical Gazette, 98(543):487–488, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2015:SWS

[Sco15a] J. A. Scott. 99.32 on the square wave series. The Mathematical Gazette, 99(546):526–528, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2015:AKH

[Sco15b] J. A. Scott. 99.35 on the asymptotes of the Kiepert hyperbola. The Mathematical Gazette, 99(546):532–533, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Scott:2016:LQK

[Sco16] J. A. Scott. 100.32 linear quadrisection and the Kiepert hyperbola. The Mathematical Gazette, 100(549):508–510, November 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
D4E34F1FA8777D498D79AFBB514A98A3.

Scott:2017:LCO

[Sco17a] J. A. Scott. 101.13 the Lester circle once again. The Mathematical Gazette, 101(550):129–130, March 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
D87B42298536C86DBD6C55AB535E2FBC.

Scott:2017:SEC

[Sco17b] J. A. Scott. 101.31 on series for the Euler constant. The Mathematical Gazette, 101(552):486–488, November 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
790476AB490AA66B4F92174D4BA32BBD.

Scott:2017:YAC

[Sco17c] J. A. Scott. 101.32 yet another Cantor series. The Mathematical Gazette, 101(552):488–489, November 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
29464C1F43C534C5AD7DFCF192B117E4.

Scott:2017:SPR

[Sco17d] J. A. Scott. 101.42 the Steiner point revisited. The Mathematical Gazette, 101(552):515–516, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10142-the-steiner-point-revisited/E7EFC81DCF49834A213133B22F3E8E48.

Scott:2017:LDP

[Sco17e] J. A. Scott. 101.43 on the least distance from a point to a line. The Mathematical Gazette, 101(552):517–519, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10143-on-the-least-distance-
from-a-point-to-a-line/2A91C3BEF5FED33F612B1EE64A3DB084.

Scott:2018:SFR

[Sco18a] J. A. Scott. 102.04 on the series of Fibonacci reciprocals. The Mathematical Gazette, 102(553):100–102, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10204-on-the-series-of-fibonacci-
reciprocals/874B3D016DC21BAC874EE884E26D4D96.

Scott:2018:AVC

[Sco18b] J. A. Scott. 102.16 an areal view of curvature. The Mathematical Gazette, 102(553):140–143, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10216-an-areal-view-of-curvature/F98C84D178E5547635202CE228892C66.

Scott:2018:AHL

[Sco18c] J. A. Scott. 102.26 an ad hoc limiting process. The Mathematical Gazette, 102(554):317, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10226-an-ad-hoc-limiting-process/A569EBF229B6FB670A6BCF0BBB2ED1B7.

Scott:2018:TMT

[Sco18d] J. A. Scott. 102.27 on the theorem of means for terms in AP and GP. The Mathematical Gazette, 102(554):317–320, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10227-on-the-theorem-of-means-
for-terms-in-ap-and-gp/3F27D6A2A16E80AF8B49EB170E232FC4.

Scott:2018:VKH

[Sco18e] J. A. Scott. 102.34 on the vertices of the Kiepert hyperbola. The Mathematical Gazette, 102(554):339–343, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10234-on-the-vertices-of-
the-kiepert-hyperbola/40F0B0B08CF3DE58F3401737AFFF6C53.

Scott:2018:SSL

[Sco18f] J. A. Scott. 102.52 a sum of squares for the Lester minimum. The Mathematical Gazette, 102(555):529–531, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10252-a-sum-of-squares-for-
the-lester-minimum/F61C51D377973EB54C519FED803EE7E6.

Scott:2019:ANP

[Sco19a] J. A. Scott. 103.10 around the nine-point circle. The Mathematical Gazette, 103(556):150–152, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10310-around-the-ninepoint-circle/C823BC1E9B1D6365543FD565848411E4.

Scott:2019:SCS

[Sco19b] J. A. Scott. 103.11 on the Spieker centre of a scalene triangle. The Mathematical Gazette, 103(556):153–154, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10311-on-the-spieker-centre-
of-a-scalene-triangle/31650AB8B8CB0443BF17272C4ED53383.

Scott:2019:UCF

[Sco19c] J. A. Scott. 103.17 An unexpected closed form revisited. The Mathematical Gazette, 103(557):312–313, July 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10317-an-unexpected-closed-
form-revisited/31D545290DE669445E6FA019E1EA776A.

Scott:2019:FHR

[Sco19d] J. A. Scott. 103.42 The Feuerbach hyperbola revisited. The Mathematical Gazette, 103(558):536–538, November 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10342-the-feuerbach-hyperbola-
revisited/21DF2BABBF455B91DE3418CB512DCEB1.

Sherran:2010:EGMb

[SD10] Peter Sherran and Peter Derych. Edexcel GCSE mathematics for A* Revision and classroom companion. Revision plus. Lonsdale, London, UK, 2010. ISBN 1-84419-269-5 (paperback). 112 pp. LCCN ????

Sadek:2013:DPF

[SE13] Jawad Sadek and Russell Euler. Divisibility properties for Fibonacci and related numbers. The Mathematical Gazette, 97(540): 461–464, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Segel:2009:RCM

[Seg09] Joel Segel, editor. Recountings: conversations with MIT mathematicians. A. K. Peters, Ltd., Wellesley, MA, USA, 2009. ISBN 1-56881-449-6. xxi + 441 pp. LCCN QA28 .R43 2009.

Senechal:2012:DBD

[Sen12] Marjorie Senechal. I died for beauty: Dorothy Wrinch and the cultures of science. Oxford University Press, USA, Oxford, UK, 2012. ISBN 0-19-973259-0 (hardcover). ix + 300 pp. LCCN QA29.W75 S56 2012.

Serre:2008:TGT

[Ser08] Jean-Pierre Serre. Topics in Galois theory, volume 1 of Research notes in mathematics. A. K. Peters, Ltd., Wellesley, MA, USA, second edition, 2008. ISBN 1-56881-412-7. xvi + 120 pp.

LCCN QA214 .S47 2008. URL http://www.loc.gov/catdir/
toc/ecip0723/2007030849.html. Notes written by Henri Dar-

mon.

Sewell:2014:AL

[Sew14] Michael Sewell. Areas of lunes. The Mathematical Gazette, 98 (541):129–131, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sewell:2015:OC

[Sew15a] Michael Sewell. An octet of circles. The Mathematical Gazette, 99 (546):468–473, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sewell:2015:PE

[Sew15b] Michael Sewell. Parametric envelopes. The Mathematical Gazette, 99(545):226–229, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Simoson:2012:SF

[SG12] Andrew J. Simoson and C. W. Groetsch. 96.28 Spurtacular fountains. The Mathematical Gazette, 96(535):166–168, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sharp:2015:APT

[Sha15] John Sharp. Artzt parabolas of a triangle. The Mathematical Gazette, 99(546):444–463, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sherran:2010:EGMc

[She10a] Peter Sherran. Edexcel GCSE mathematics. Higher. Revision and classroom companion. Revision plus. Lonsdale, London, UK, 2010. ISBN 1-84419-261-X (paperback). 112 pp. LCCN ????

Sherran:2010:EGMa

[She10b] Peter Sherran. Edexcel GCSE mathematics higher. Workbook. Revision plus. Lonsdale, London, UK, 2010. ISBN 1-84419-262-8 (paperback). LCCN ????

Shephard:2013:RPC

[She13] G. C. Shephard. Regular polyhedral clones. The Mathematical Gazette, 97(540):421–429, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shephard:2016:CPC

[She16] G. C. Shephard. Convex polygons as carriers. The Mathematical Gazette, 100(547):93–102, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shimura:2008:MML

[Shi08] Gorāo Shimura. The map of my life. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 0-387-79714-9, 0-387-79715-7 (e-book). vi + 211 pp. LCCN QA29.S428 A3 2008.

Shimura:2010:AQF

[Shi10a] Goråo Shimura. Arithmetic of quadratic forms. Springer monographs in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 1-4419-1731-4 (hardcover), 1-4419-1732-2 (e-book). ISSN 1439-7382. xi + 237 pp. LCCN QA243 .S492 2010.

Shiu:2010:FBA

[Shi10b] Peter Shiu. 94.06 A footnote to Besicovitch’s arc. The Mathematical Gazette, 94(529):130–132, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2010:BRBc

[Shi10c] Peter Shiu. Book review: A first course in functional analysis, by S. David Promislow. Pp. 307. £52.95 (hardback). 2008. ISBN: 978-0-470-14619-7 (Wiley). The Mathematical Gazette, 94(530): 373, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2010:BRBf

[Shi10d] Peter Shiu. Book review: Elementary functional analysis, by Barbara D. MacCluer. Pp. 207. £29.99. 2009. ISBN 978-0-387-85528-8 (Springer). The Mathematical Gazette, 94(531):564, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2010:BRBa

[Shi10e] Peter Shiu. Book review: Experimental number theory, by Fernando Rodriguez Villegas. Pp. 214. £31.00 (pbk). 2007. ISBN 978-0-19-922730-3 (Oxford University Press). The Mathematical Gazette, 94(529):183, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2010:BRBb

[Shi10f] Peter Shiu. Book review: Experimental number theory, by Fernando Rodriguez Villegas. Pp. 214. £31.00 (pbk). 2007. ISBN 978-0-19-922730-3 (Oxford University Press). The Mathematical Gazette, 94(529):183, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2010:BRBe

[Shi10g] Peter Shiu. Book review: Introduction to metric & topological spaces (2nd edn.), by Wilson A. Sutherland. Pp. 206 (hbk). £45.00. 2009. ISBN: 978-0-19-956307-4 (Oxford University Press). The Mathematical Gazette, 94(531):563–564, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2010:BRBg

[Shi10h] Peter Shiu. Book review: Pythagoras’ revenge: a mathematical mystery, by Arturo Sangalli. Pp. 183. £14.95. 2009. ISBN 978-0-691-04955-7 (Princeton University Press). The Mathematical Gazette, 94(531):574, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2010:BRBd

[Shi10i] Peter Shiu. Book review: Topics in Galois theory (2nd edn), by Jean-Pierre Serre. Pp. 120. $39.00. 2008. ISBN 978-1-56881-412-4 (A. K. Peters). The Mathematical Gazette, 94(530):374, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shirali:2011:BPG

[Shi11a] Shailesh A. Shirali. 95.09 A bootstrapping proof of the AM–GM inequality for three variables. The Mathematical Gazette, 95(532): 86–87, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shirali:2011:GAT

[Shi11b] Shailesh A. Shirali. A generalisation of the Arbelos Theorem of Archimedes. The Mathematical Gazette, 95(533):197–205, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shirali:2011:TOA

[Shi11c] Shailesh A. Shirali. Triangles with one angle equal to 60 degrees. The Mathematical Gazette, 95(533):227–234, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2011:BRBb

[Shi11d] Peter Shiu. Book review: Functional equations and inequalities with applications, by PI. Kannappan. Pp. 810, £117-00 (Hardback), 2009. ISBN: 978-0-387-89491-1. (Springer). The Mathematical Gazette, 95(532):160–162, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2011:BRBa

[Shi11e] Peter Shiu. Book review: The finite simple groups, by Robert A. Wilson. Pp. 298. £53.45 (Hardback). 2009. ISBN: 978-1-84800-987-5. (Springer). The Mathematical Gazette, 95(532):157–159, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2011:FMD

[Shi11f] Peter Shiu. Finding a missing digit. The Mathematical Gazette, 95(533):206–212, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiral:2012:PTF

[Shi12a] Shailesh A. Shiral. 96.29 A party trick featuring the triangular numbers. The Mathematical Gazette, 96(536):261–263, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shirali:2012:DRU

[Shi12b] Shailesh A. Shirali. Digging for roots using the Pascal spade. The Mathematical Gazette, 96(536):251–260, July 2012. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2012:BRB

[Shi12c] Peter Shiu. Book review: Arithmetic of quadratic forms, by Goro Shimura. Pp. 237, £64.99 (hbk), 2010. ISBN: 978-1-4419-1731-7 (Springer). The Mathematical Gazette, 96(536):371–372, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2012:PT

[Shi12d] Peter Shiu. Partitioning triangles. The Mathematical Gazette, 96 (537):546–550, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shirali:2013:CL

[Shi13a] Shailesh A. Shirali. 97.13 A circular locus. The Mathematical Gazette, 97(538):153–157, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2013:S

[Shi13b] Peter Shiu. 97.02 Solutions to ϕ(m) = k! and σ(n) = k!. The Mathematical Gazette, 97(538):110–115, March 2013. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2013:BRB

[Shi13c] Peter Shiu. Book review: Mathematical inequalities: a perspective, by Pietro Cerone and Sever S. Dragomir. Pp. 391. £76.99 (hbk). 2011. ISBN: 978-1-4398-4896-8 (CRC Press). The Mathematical Gazette, 97(539):374–375, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2013:GBS

[Shi13d] Peter Shiu. The gaps between sums of two squares. The Mathematical Gazette, 97(539):256–262, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2014:ITI

[Shi14a] Peter Shiu. 98.28 if e < 2 then π is irrational. The Mathematical Gazette, 98(543):497–501, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2014:BRBb

[Shi14b] Peter Shiu. Book review: Alan Turing’s system of logic: the Princeton thesis, Andrew W. Appel (Ed.) Pp. 142. £16.95 (hbk). 2012. ISBN: 978-0-691-15574-6 (Princeton University Press). The Mathematical Gazette, 98(541):186–187, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2014:BRBc

[Shi14c] Peter Shiu. Book review: Combinatorial set theory: with a gentle introduction to forcing, by Lorenz J Halbeisen. Pp. 453. £81.00 (Hardback). 2012. ISBN: 978-1-4471-2172-5 (Springer). The Mathematical Gazette, 98(541):188–190, March 2014. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2014:BRBa

[Shi14d] Peter Shiu. Book review: Inequalities: Theorems, techniques and selected problems, by Zdravko Cvetkovski. Pp. 444, £35.99 (pbk), 2012. ISBN: 978-3-642-23791-1 (Springer). The Mathematical Gazette, 98(541):171–172, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2014:BRN

[Shi14e] Peter Shiu. Book review: Number, shape, and symmetry: an introduction to number theory, geometry and group theory by Diane L. Herrmann and Paul J. Sally Jr, pp 426, £44.99, ISBN 978-1-4665-5464-1, AK Peters, CRC Press (2012). The Mathematical Gazette, 98(543):559–560, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2015:BRBb

[Shi15a] Peter Shiu. Book review: Basic modern algebra with applications by Mahima Ranjan Adhikari and Avishek Adhikari, pp 637, £42.99 (hbk), ISBN 978-81-322-1598-1, also available as e-book, Springer (2014). The Mathematical Gazette, 99(544):183–185, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2015:BRBc

[Shi15b] Peter Shiu. Book review: Introduction to abstract algebra by Benjamin Fine, Anthony M. Gaglione & Gerhard Rosenberger, pp. 566, £64.60 (hard), ISBN 978-1-4214-1176-7, Johns Hopkins University Press (2014). The Mathematical Gazette, 99(544):187–189, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2015:BRBa

[Shi15c] Peter Shiu. Book review: Linear integral equations (3rd edn.) by Rainer Kress, pp 412, £53.99 (hbk), ISBN 978-1-4614-9592-5, also available as e-book, Springer (2014). The Mathematical Gazette, 99(544):180–181, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2015:F

[Shi15d] Peter Shiu. The factorisation of 2^{67}−1. The Mathematical Gazette, 99(544):104–108, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2015:FMF

[Shi15e] Peter Shiu. Fermat’s method of factorisation. The Mathematical Gazette, 99(544):97–103, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2016:AE

[Shi16a] Peter Shiu. 100.05 another evaluation of ζ(2). The Mathematical Gazette, 100(547):118–119, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2016:BRM

[Shi16b] Peter Shiu. Book review: Multivariate calculus and geometry (3rd edn.), by Seán Dineen, pp. 257, $36.39 (paper), ISBN 978-1-4471-6418-0, also available as e-book, Springer Verlag (2014). The Mathematical Gazette, 100(547):181–182, March 2016. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2016:CH

[Shi16c] Peter Shiu. Cyclotomy and the heptadecagon. The Mathematical Gazette, 100(548):288–297, July 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.

cambridge.org/core/product/8E14D43530C969C10CE4191E5B30623A.

Shiu:2016:MSV

[Shi16d] Peter Shiu. The mean-square value of the divisor function. The Mathematical Gazette, 100(548):203–212, July 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
6691207B631979E7788761A3A8231C84.

Shiu:2017:BRL

[Shi17a] Peter Shiu. Book review: Lecture notes on calculus of variations by Kung Ching Chang; translated by Tan Zhang, pp. 324, $56.00 (paper), ISBN 978-9-81314-623-5, World Scientific (2017). The Mathematical Gazette, 101(552):568–569, November 2017.

CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/lecture-notes-on-calculus-
of-variations-by-kung-ching-chang-translated-by-tan-zhang-
pp-324-5600-paper-isbn-9789813146235-world-scientific-
2017/28AA8D996741C3ED1A1865EAFE0714B5.

Shiu:2017:BRS

[Shi17b] Peter Shiu. Book review: Summing it up: from one plus one to modern number theory, by Ash, Avner and Gross, Robert, pp. 229, US$27.95 (hard), ISBN 978-0-691-17019-0, Princeton University Press (2016). The Mathematical Gazette, 101(551):363–365, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
0C702DD9E9D6B7E013858AC1376A8210.

Shiu:2018:BRS

[Shi18a] Peter Shiu. Book review: A second course in linear algebra by Stephan Ramon Garcia and Roger A. Horn, pp. 426, £49.99 (hard), ISBN 978-1-107-10381-8, Cambridge University Press (2017). The Mathematical Gazette, 102(554): 374–375, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/second-
course-in-linear-algebra-by-stephan-ramon-garcia-and-roger-
a-horn-pp-426-4999-hard-isbn-9781107103818-cambridge-university-
press-2017/F764B6C9A4645704D144478179234BEE.

Shiu:2018:BRC

[Shi18b] Peter Shiu. Book review: Calculus in context: background, basics, and applications by Alexander J. Hahn, pp. 712, $70.50 (hard), £70.50, ISBN 978-1-4214-2230-5, Johns Hopkins University Press (2017). The Mathematical Gazette, 102(554):372–374, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/calculus-in-context-background-basics-and-applications-
by-alexander-j-hahn-pp-712-7050-hard-7050-isbn-9781421422305-
johns-hopkins-university-press-2017/D097E5E09E5EE1448DB5D0B17BCEA524.

Shiu:2018:COC

[Shi18c] Peter Shiu. Can our coinage system be improved? The Mathematical Gazette, 102(554):226–232, July 2018. CODEN MAGAAS.

ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/can-our-coinage-system-be-improved/CA2AE6AA47EA8A88C2ADAAFE000E0339.

Shiu:2019:MSC

[Shi19a] Peter Shiu. 103.07 more on slowly convergent, or divergent, series. The Mathematical Gazette, 103(556):140–142, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10307-more-
on-slowly-convergent-or-divergent-series/B202A9CD91C0CA1FF3D62EC3C3185D6F.

Shiu:2019:FNO

[Shi19b] Peter Shiu. 103.31 Factorising numbers with oracles. The Mathematical Gazette, 103(558):495–499, November 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10331-factorising-numbers-
with-oracles/90DF80483BD34CCDA26C7D8F9D1B16C9.

Shiu:2019:RAA

[Shi19c] Peter Shiu. Book review: Real analysis and applications by Fabio Silva Botelho, pp. 567, £44.99 (hard), ISBN 978-3-31978-630-8, Springer Verlag (2018). The Mathematical Gazette, 103(557):371–372, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/real-analysis-and-applications-by-fabio-silva-
botelho-pp-567-4499-hard-isbn-9783319786308-springer-verlag-
2018/CAF85993BCCFFAAF49BF3BF6F06C18EE.

Shiu:2019:WCV

[Shi19d] Peter Shiu. The whole can be very much less than the sum of its parts. The Mathematical Gazette, 103(556):117–127, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/whole-can-be-very-much-less-
than-the-sum-of-its-parts/BB4AD603F5272DDDF1F666E7DE2998A2.

Siklos:2012:MIR

[Sik12] Stephen Siklos. Maclaurin’s inequalities: reflections on a STEP question. The Mathematical Gazette, 96(537):499–507, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Siklos:2018:ME

[Sik18] Stephen Siklos. 102.08 a method of evaluating ζ(2) and ∫_{0}^{∞} sin(x)/x dx. The Mathematical Gazette, 102(553):114–121, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10208-a-method-of-
evaluating-2-and/DAB61D6D96484DB29D5FBA08A061C7F4.

Silvester:2010:CWK

[Sil10] John R. Silvester. C. W. Kilmister 1924–2010. The Mathematical Gazette, 94(531):529–531, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Silvester:2012:DFC

[Sil12] John Silvester. Droz–Farny constructions. The Mathematical Gazette, 96(535):19–38, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Silvester:2017:EAE

[Sil17] John R. Silvester. Extremal area ellipses of a convex quadrilateral. The Mathematical Gazette, 101(550):11–26, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
5BFD5A045A464E6538A60FF47ABE3B0E.

Silvester:2018:TC

[Sil18a] John R. Silvester. On touching circles. The Mathematical Gazette, 102(555):460–470, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/on-touching-circles/544BE72726D4DFCC28BE622A03902B97.

Silvester:2018:SCT

[Sil18b] John R. Silvester. The seven circles theorem revisited. The Mathematical Gazette, 102(554):280–301, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/seven-circles-theorem-revisited/A4A2727DE243AEFF70F55F849F288673.

Silvester:2018:VJT

[Sil18c] John R. Silvester. Variations on a Japanese Temple theorem. The Mathematical Gazette, 102(553):38–49, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/variations-on-a-japanese-
temple-theorem/C1C288AD3861831785116ECBF4039BFA.

Simons:2010:GC

[Sim10] Stuart Simons. 94.40 Getting to the centre. The Mathematical Gazette, 94(531):517–521, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Simons:2011:BWH

[Sim11a] Stuart Simons. 95.48 Balls and walls — how energy can be lost in elastic collisions. The Mathematical Gazette, 95(533):349–353, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Simons:2011:SRS

[Sim11b] Stuart Simons. 95.67 SHM on a rough surface. The Mathematical Gazette, 95(534):545–549, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Simons:2011:CAR

[Sim11c] Stuart Simons. Constant acceleration revisited. The Mathematical Gazette, 95(534):553–554, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Simons:2012:HSR

[Sim12] Stuart Simons. 96.58 Hauling a sledge and related matters. The Mathematical Gazette, 96(537):536–541, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Simons:2013:EP

[Sim13] Stuart Simons. 97.33 An extremum problem. The Mathematical Gazette, 97(539):306–308, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Simons:2016:OB

[Sim16] Jim Simons. 100.20 outside the box. The Mathematical Gazette, 100(548):323–329, July 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/64919183456B442F999D0F7C9A40226A.

Simoson:2018:BHC

[Sim18] Andrew J. Simoson. 102.19 a black hole conundrum. The Mathematical Gazette, 102(553):150–152, March 2018. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10219-a-black-hole-conundrum/
222C8E1A20F4FDB0F4B8AC032D19DB11.

Simoson:2019:M

[Sim19] Andrew J. Simoson. The morphology of Z\sqrt{[10]}. The Mathematical Gazette, 103(558):442–460, November 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/morphology-of-mathbbzsqrt-
rm10/6F90DD74EEDB7C8DD5192248B24DCCF5.

Sintamarian:2010:RGI

[Sîn10] Alina Sîntamarian. Regarding a generalisation of Ioachimescu’s constant. The Mathematical Gazette, 94(530):270–283, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Siu:2013:SME

[Siu13] Man Keung Siu. 97.21 Some more on Estermann and Pythagoras. The Mathematical Gazette, 97(539):272–273, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shi:2017:GNI

[SL17] Yehao Shi and Ling Lin. 101.09 A generalisation of Nesbitt inequality. The Mathematical Gazette, 101(550):122, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
A86D984010D43F3D782F3D5D1A4A9E44.

Slessenger:2012:SGI

[Sle12] Wolfgang Slessenger. A simple geometrical inequality. The Mathematical Gazette, 96(536):343–344, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Spahn:2010:UUP

[SLMRR10] Mark Spahn, Ron Lancaster, Deborah Moore-Russo, and Gerald Rising. An unexpected use of primes: solving sudokus by calculator. The Mathematical Gazette, 94(530):224–232, July 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Small:2012:DBS

[Sma12] Michael Small, (Professor). Dynamics of biological systems. Chapman and Hall/CRC mathematical and computational biology series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-4398-5336-3 (hardcover). xv + 262 pp. LCCN QH323.5 .S566 2012.

Smith:2008:MOP

[Smi08] Geoff Smith. A mathematical olympiad primer, volume 3 of Handbooks. The United Kingdom Mathematics Trust, Leeds, UK, 2008. ISBN 1-906001-03-0 (paperback), 1-906001-05-7. x + 178 pp. LCCN QA43 .S65 2008.

Smith:2015:PSL

[Smi15] Geoff Smith. 99.20 A projective Simson line. The Mathematical Gazette, 99(545):339–341, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Smith:2016:CA

[Smi16] Geoff C. Smith. Circles in areals. The Mathematical Gazette, 100 (548):239–246, July 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/83BEF0FA3702B3C9CF13D33E3FC54BD7.

Smyth:2010:ISB

[Smy10] R. W. Smyth. An interesting strategic blunder. The Mathematical Gazette, 94(529):84–93, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sharp:2011:AAL

[SNH11] John Sharp, B. G. Nickel, and J. L. Hunt. Anamorphoscopes another look at circle inverting mirrors. The Mathematical Gazette, 95(532):1–16, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sofo:2010:GGD

[Sof10] Anthony Sofo. 94.21 Grandma, grandaunts and distant cousins. The Mathematical Gazette, 94(530):314–316, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sofair:2013:DR

[Sof13] Isaac Sofair. Derangement revisited. The Mathematical Gazette, 97(540):435–440, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Soifer:2009:MCB

[Soi09a] Alexander Soifer. The mathematical coloring book: mathematics of coloring and the colorful life of its creators. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 0-387-74640-4 (hardcover). xxx + 607 pp. LCCN QA166 .S65 2009.

Soifer:2009:MPS

[Soi09b] Alexander Soifer. Mathematics as problem solving. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2009. ISBN 0-387-74646-3 (paperback), 0-387-74647-1 (e-book). xviii + 106 pp. LCCN

QA63 .S68 2009. URL http://www.gbv.de/dms/bowker/toc/
9780387746463.pdf;   http://www.zentralblatt-math.org/
zmath/en/search/?an=1171.00005.

Spivey:2010:VCF

[Spi10] Michael Z. Spivey. 94.11 Visualising continued fractions. The Mathematical Gazette, 94(530):284–290, July 2010. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sporn:2018:ACF

[Spo18] Howard Sporn. 102.24 area of a circle via the Fundamental Theorem of Calculus. The Mathematical Gazette, 102(554):312–314, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10224-area-of-a-
circle-via-the-fundamental-theorem-of-calculus/CDB8CAFE5ED9DD92D2764BAB3E7C3888.

Sprent:2009:MF

[Spr09] Peter Sprent. Mathematics without fears. www.lulu.com, ????, 2009. ISBN 1-4092-5670-7 (paperback). xii + 453 pp. LCCN QA40 .S67 2009.

Shiu:2011:LBP

[SS11] Daniel Shiu and Peter Shiu. A lower bound for the prime counting function. The Mathematical Gazette, 95(534):433–436, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Shiu:2012:PPA

[SS12] Daniel Shiu and Peter Shiu. A poor person’s approximation to π. The Mathematical Gazette, 96(537):408–414, November 2012. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sigler:2016:PW

[SS16] Avi Sigler and Moshe Stupel. 100.15 proof without words. The Mathematical Gazette, 100(547):146, March 2016. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Stupel:2019:TPW

[SSJ19] Moshe Stupel, Avi Sigler, and Jay Jahangiri. Teaching proofs without words using dynamic geometry. The Mathematical Gazette, 103(557):204–211, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/teaching-proofs-without-words-using-dynamic-geometry/
5065F46C0B07FD6E1B852D0775756FAD.

Stahl:2011:RAH

[Sta11] Saul Stahl. Real analysis: a historical approach. Pure and applied mathematics. John Wiley, New York, NY, USA, second edition, 2011. ISBN 0-470-87890-8 (hardcover). xv + 293 pp. LCCN QA300 .S882 2011.

Stanley:2018:NSE

[Sta18] P. Stanley. 102.40 a nomogram for the solution of the equation p/ sin θ + q/ cos θ = 1. The Mathematical Gazette, 102 (555):501–503, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10240-a-nomogram-for-the-solution-of-the-equation/
62044A55171C411661C463E17DBA03E1.

Stanley:2019:FHY

[Sta19a] P. Stanley. 103.46 A four-hundred year calendar. The Mathematical Gazette, 103(558):556–558, November 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10346-a-fourhundred-year-
calendar/AE88FFB93D1480C0E0AF1E652894D16A.

Stanley:2019:ALN

[Sta19b] P. Stanley. An appraisal of Lalanne’s nomogram. The Mathematical Gazette, 103(556):47–51, March 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/an-appraisal-of-lalannes-
nomogram/98DC74B52C2AB1999829650CC4E075CE.

Stewart:2008:PSC

[Ste08] Ian Stewart. Professor Stewart’s cabinet of mathematical curiosities. Profile Books, London, UK, 2008. ISBN 1-84668-064-6 (paperback). x + 310 pp. LCCN QA95 .S7257 2008. URL http://

www.gbv.de/dms/goettingen/573354561.pdf.

Stewart:2009:PSC

[Ste09a] Ian Stewart. Professor Stewart’s cabinet of mathematical curiosities. Basic Books, New York, NY, USA, 2009. ISBN 0-465-01302-3, 1-84668-064-6. x + 310 pp. LCCN QA95 .S7257 2009.

Stewart:2009:PSH

[Ste09b] Ian Stewart. Professor Stewart’s hoard of mathematical treasures: another drawer from the cabinet of curiosities. Basic Books, New York, NY, USA, 2009. ISBN 1-84668-292-4 (hardcover), 0-465-01775-4 (paperback), 1-282-74737-1 (e-book). x + 2 + 339 pp. LCCN QA99 .S85 2009.

Stewart:2010:CM

[Ste10] Ian Stewart. Cows in the maze. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-956207-5 (paperback). xi + 306 pp. LCCN QA95 .S724 2010.

Stein:2011:CNN

[Ste11a] James D. Stein. Cosmic numbers: the numbers that define our universe. Basic Books, New York, NY, USA, 2011. ISBN 0-465-02198-0 (hardcover), 0-465-06379-9 (paperback), 0-465-02788-1 (ebook). xii + 228 pp. LCCN QB991.C658 S74 2011.

Stewart:2011:ML

[Ste11b] Ian Stewart. Mathematics of life: unlocking the secrets of existence. Profile Books, London, UK, 2011. ISBN 0-465-02238-3, 1-84668-198-7 (hardcover), 1-84765-350-2 (e-book), 1-84668-550-8. viii + 358 pp. LCCN QH323.5 .S74 2011; QH323.5 .S84 2011.

Stedall:2012:HM

[Ste12a] Jacqueline A. Stedall. The history of mathematics: a very short introduction, volume 305 of Very short introductions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-959968-8 (paperback), 0-19-177809-5 (e-book). xvii + 123 pp. LCCN QA21 .S74 2012; QA21 .S828 2012.

Stephenson:2012:IDA

[Ste12b] Paul Stephenson. 96.51 Iterated dissections? Approximations to \sqrt{2} in the Indian Sulbasutras. The Mathematical Gazette, 96(537): 508–512, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Stewart:2012:SEC

[Ste12c] Ian Stewart. Seventeen equations that changed the world: a unique insight into humanity told through the equations that define our lives. Profile Books, London, UK, 2012. ISBN 1-84668-531-1 (paperback), 1-84765-769-9 (e-book). x + 342 pp. LCCN QA211 .S775 2012.

Stewart:2013:GMP

[Ste13a] Ian Stewart. The great mathematical problems: marvels and mysteries of mathematics. Profile Books, London, UK, 2013. ISBN 1-84668-199-5 (paperback), 1-84765-351-0 (e-book), 1-306-31411-9 (MyiLibrary). x + 340 pp. LCCN QA39.2 G73 2013; QA93 .S74 2013.

Stewart:2013:VIG

[Ste13b] Ian Stewart. Visions of infinity: the great mathematical problems. Basic Books, New York, NY, USA, 2013. ISBN 0-465-02240-5 (hardcover), 0-465-06599-6 (e-book), 1-84668-199-5, 1-84765-351-0. x + 340 pp. LCCN QA93 .S745 2013. URL http:/

/catdir.loc.gov/catdir/enhancements/fy1313/2012924095-
b.html;   http://catdir.loc.gov/catdir/enhancements/fy1313/
2012924095-d.html.

Stewart:2014:PSC

[Ste14] Ian Stewart. Professor Stewart’s casebook of mathematical mysteries. Basic Books, New York, NY, USA, 2014. ISBN 0-465-05497-8, 0-465-05688-1 (e-book). 307 pp. LCCN QA99

.S848 2014. URL http://www.loc.gov/catdir/enhancements/
fy1503/2014940655-b.html;   http://www.loc.gov/catdir/
enhancements/fy1503/2014940655-d.html.

Stephenson:2016:SFC

[Ste16a] Paul Stephenson. 100.13 the Steinhaus ‘four-colour’ tiling. The Mathematical Gazette, 100(547):141–143, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Stewart:2016:SEC

[Ste16b] Ian Stewart. Sir Erik Christopher Zeeman FRS 1925–2016. The Mathematical Gazette, 100(548):307–313, July 2016. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
046B3C2D06E0D6AF439815DDDF90E35A.

Stephenson:2017:PWR

[Ste17a] Paul Stephenson. 101.18 proof without words: relationship between inverse cotangents. The Mathematical Gazette, 101 (551):300, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/EAB7228B190C4FD2D93D7C62E06123AE.

Stephenson:2017:DC

[Ste17b] Paul Stephenson. 101.40 Dandelin circles. The Mathematical Gazette, 101(552):508–511, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10140-dandelin-circles/64A9F28AC6B1E3A4BF64B423B9BD832D.

Stephenson:2017:PTD

[Ste17c] Paul Stephenson. Pappus in two dimensions, Mamikon in three. The Mathematical Gazette, 101(551):322–325, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
3DD1E99EDCCD9FC35E78A09E25AA13D5.

Stephenson:2017:TIR

[Ste17d] Paul Stephenson. Trigonometric identities from regular polygons. The Mathematical Gazette, 101(551):327–334, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
E9742F9AB30179E0A16FA67A757DD28D.

Stewart:2017:FSI

[Ste17e] Seán M. Stewart. 101.06 Finding some integrals using an interesting self-similar substitution. The Mathematical Gazette, 101 (550):103–108, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/84F2FF4136DCE83996306691165EA679.

Stewart:2017:SAD

[Ste17f] Seán M. Stewart. Some alternative derivations of Craig’s formula. The Mathematical Gazette, 101(551):268–279, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
DE7405D526DF9E9D344CBC2FAA745EBA.

Stephenson:2018:DAT

[Ste18a] Paul Stephenson. 102.17 dimensional analogues of two figurate identities. The Mathematical Gazette, 102(553):143–146, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10217-dimensional-
analogues-of-two-figurate-identities/7DBFD3D561A776088BFB8DBFCACF72D6.

Stephenson:2018:CGP

[Ste18b] Paul Stephenson. 102.31 cyclic gridpoint polygons. The Mathematical Gazette, 102(554):331–332, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10231-cyclic-gridpoint-polygons/2CDF33E4D0C4A3699AAE42FF28A2B7C7.

Stephenson:2018:F

[Ste18c] Paul Stephenson. Feedback. The Mathematical Gazette, 102(553):172–173, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/feedback/DDF0A79855B6434CD5F8EA1D2A24CD6B.

Stephenson:2019:WSF

[Ste19a] Paul Stephenson. 103.15 When is the sum of the first n consecutive odd cubes a square? The Mathematical Gazette, 103(557): 305–306, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10315-when-
is-the-sum-of-the-first-n-consecutive-odd-cubes-a-square/
B61CA720E7E51ECF8AE876934D133E19.

Stewart:2019:C

[Ste19b] Ian Stewart. Correspondence. The Mathematical Gazette, 103 (556):170, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/correspondence/
972DBF1E6B243D0A27C9F04B9E3E041E.

Stillwell:2010:RIM

[Sti10] John Stillwell. Roads to infinity: the mathematics of truth and proof. A. K. Peters, Ltd., Wellesley, MA, USA, 2010. ISBN 1-56881-466-6. xi + 203 pp. LCCN QA248 .S778 2010.

Storr:2010:BRBb

[Sto10a] Graham H. Storr. Book review: Crossing the bridge, by Gerry Leversha. Pp. 180. £11.00. 2008. ISBN 978-1-906001-06-3 (United Kingdom Mathematics Trust). The Mathematical Gazette, 94 (531):567–568, November 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Storr:2010:BRBa

[Sto10b] Graham H. Storr. Book review: Even better mathematics: looking back to move forward, by Afzal Ahmed and Honor Williams. Pp.

99. £19.99. 2007. ISBN 978-1-85539-147-5 (Network Continuum Education). The Mathematical Gazette, 94(529):182–183, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Storr:2012:BRBc

[Sto12a] Graham H. Storr. Book review: Edexcel GCSE mathematics higher workbook, by Peter Sherran. Pp. 112. £5.99. 2010. ISBN 978-1-84419-262-5 (Lonsdale). The Mathematical Gazette, 96(536):366, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Storr:2012:BRBb

[Sto12b] Graham H. Storr. Book review: GCSE revision plus Edexcel mathematics for A* revision guide, by Peter Sherran and Peter Derych. Pp. 112. £5.99. 2010. ISBN 978-1-84419-269-4 (Lonsdale). CSE revision plus Edexcel mathematics higher revision guide, by Peter Sherran. Pp. 112. £5.99. 2010. ISBN 978-1-84419-261-8 (Lonsdale). The Mathematical Gazette, 96(536):363–364, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Storr:2012:BRBa

[Sto12c] Graham H. Storr. Book review: Mathematics under the microscope: notes on cognitive aspects of mathematical practice, by Alexandre V. Borovik. Pp. 317. £44.95 (hbk). 2010. ISBN: 978-0-8218-4761-9 (American Mathematical Society). The Mathematical Gazette, 96(536):362–363, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Storr:2014:BRH

[Sto14a] Graham Storr. Book review: How to study for a mathematics degree by Lara Alcock, pp. 288, £12.99, ISBN 978-019-966132-9, Oxford University Press (2013). The Mathematical Gazette, 98 (543):547, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Storr:2014:BRMa

[Sto14b] Graham H. Storr. Book review: Master class in mathematics education: international perspectives on teaching and learning edited by Paul Andrews and Tim Rowland, pp 223, £24.99, ISBN 978-1-4411-7975-3, Bloomsbury (2014). The Mathematical Gazette, 98 (543):550–551, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Storr:2014:BRMb

[Sto14c] Graham H. Storr. Book review: The mathematical sciences in 2025 by National Research Council, pp 205, £32.00, ISBN 978-0-309-28457-8, The National Academies Press (2013). The Mathematical Gazette, 98(543):558–559, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Storr:2017:BRM

[Sto17a] Graham Storr. Book review: Mathematics and the body: material entanglements in the classroom by De Freitas, Elizabeth and Sinclair, Nathalie, pp. 284, £65.00 (hard), ISBN 978-1-107-03948-3, Cambridge University Press (2014). The Mathematical Gazette, 101(550):160, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/4B35676F8809E0071F8214C77174E5FD.

Storr:2017:BRF

[Sto17b] Graham Storr. Book review: The foundations of mathematics (second edition), by Stewart, Ian and Tall, David, pp. 416, £14.99 (paper), ISBN 978-0-19870643-4, Oxford University Press (2015). The Mathematical Gazette, 101(550):185–186, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
FDECF381791FC4932B3BAFB1AF07DCC1.

Stubhaug:2010:GML

[Stu10] Arild Stubhaug. Gösta Mittag-Leffler: a man of conviction. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. ISBN 3-642-11671-X (hardcover), 3-642-11672-8 (e-book). x + 733 pp. LCCN QA29.M5233 S8913 2010. Translated by Tiina Nunnally.

Subramaniam:2014:NPR

[Sub14a] K. B. Subramaniam. 98.19 A new proof that the rationals are countable. The Mathematical Gazette, 98(542):345, July 2014. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Subramaniam:2014:SSC

[Sub14b] K. B. Subramaniam. ‘summing squares’ by counting cubes. The Mathematical Gazette, 98(541):131–132, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Subramaniam:2015:PW

[Sub15a] K. B. Subramaniam. 99.10 Proof without words: T_{1} + T_{2} + ⋯ +

T_{n} = ( \begin{matrix}n + 2 \\ 3 \end{matrix} ). The Mathematical Gazette, 99(544):153–154,

March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Subramaniam:2015:PWC

[Sub15b] K. B. Subramaniam. 99.38 proof without words: Cotangent expressed as a series of cosecants of multiple angles in powers of two.

The Mathematical Gazette, 99(546):538, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Subramaniam:2018:PTP

[Sub18a] K. B. Subramaniam. 102.10 Pythagoras’ theorem: a proof by contradiction. The Mathematical Gazette, 102(553):128, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10210-pythagoras-theorem-
a-proof-by-contradiction/0C030079BB2802F1FAC06B9B448D475C.

Subramaniam:2018:PWW

[Sub18b] K. B. Subramaniam. 102.41 proof without words: sin^{2} A+sin^{2} B+ sin^{2} C = 4 sin A sin B sin C (where A+B+C = π). The Mathematical Gazette, 102(555):504, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10241-proof-without-words-sin2a-sin2b-sin2c-4-
sinasinbsinc-where-a-b-c/D0DD96EB237FC79F7689C8B4BBD4C49F.

Subramaniam:2019:LBT

[Sub19] K. B. Subramaniam. 103.28 On a link between Triangular and Fibonacci numbers. The Mathematical Gazette, 103(558):489, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/10328-on-a-
link-between-triangular-and-fibonacci-numbers/BAE56357F78DCA5BBE2CCBF593B6610A.

Such:2011:OE

[Suc11] Ondrej Such. 95.29 The order of elements in SL(2, p). The Mathematical Gazette, 95(533):292–300, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sudbury:2011:VNP

[Sud11] Aidan Sudbury. 95.65 Vectors and the nine-point circle. The Mathematical Gazette, 95(534):540–542, November 2011. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sudbery:2013:DMT

[Sud13] Anthony Sudbery. Dürer’s magic tesseract. The Mathematical Gazette, 97(538):8–12, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sury:2010:GTL

[Sur10] B. Sury. 94.13 Group theory lends a hand to number theory. The Mathematical Gazette, 94(530):294–296, July 2010. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sury:2013:CP

[Sur13] B. Sury. Composition of polynomials. The Mathematical Gazette, 97(538):36–42, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sury:2014:ES

[Sur14] B. Sury. 98.03 Explicit solutions of ϕ(m) = k!. The Mathematical Gazette, 98(541):112–114, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sury:2018:VCN

[Sur18] B. Sury. 102.39 variants of Carmichael numbers and Cunningham chains. The Mathematical Gazette, 102(555):498–501, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10239-variants-of-
carmichael-numbers-and-cunningham-chains/4FD957653AF2F7AC5B87057198F4471C.

Sutherland:2009:IMT

[Sut09] W. A. (Wilson Alexander) Sutherland. Introduction to metric and topological spaces. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, second edition, 2009. ISBN 0-19-956307-1 (hardcover), 0-19-956308-X (paperback). 206 pp. LCCN QA611.3 .S93 2009.

Sutcliffe:2010:BRB

[Sut10] Alan Sutcliffe. Book review: Legacy of the Luoshu: the 4,000 year search for the meaning of the magic square of order three (2nd edn), by Frank J. Swetz. Pp. 214. £21.50 (hardback). 2008. ISBN 978 1 56881 427 8 (A. K. Peters). The Mathematical Gazette, 94 (530):365–367, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sutcliffe:2011:BRB

[Sut11] Alan Sutcliffe. Book review: A lifetime of puzzles: A collection of puzzles in honor of Martin Gardner’s 90th birthday, edited by Erik D. Demaine, Martin L. Demaine, and Tom Rodgers. Pp.

349. £38.95 (hbk). 2008. ISBN: 978-1-56881-245-8 (A. K. Peters). Origami, Eleusis, and the Soma cube: Martin Gardner’s mathematical diversions, (Volume 2 of The New Martin Gardner Mathematical Library), by Martin Gardner. Pp. 234. £8.99. 2008. ISBN: 978-0-521-73524-7 (Cambridge University Press). The Mathematical Gazette, 95(532):174–175, March 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Sieminski:2015:PWD

[SV15] E. M. Sieminski and R. Viglione. 99.39 proof without words: A delightful water pouring problem. The Mathematical Gazette, 99 (546):538–539, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Swetz:2012:MEE

[Swe12] Frank J. Swetz. Mathematical expeditions: exploring word problems across the ages. The Johns Hopkins University Press, Baltimore, MD, USA, 2012. ISBN 1-4214-0437-0 (hardcover), 1-4214-0438-9 (paperback). ix + 192 pp. LCCN QA43 .S96 2012.

Szpiro:2010:MMF

[Szp10a] George Szpiro. A mathematical medley: fifty easy pieces on mathematics. American Mathematical Society, Providence, RI, USA, 2010. ISBN 0-8218-4928-X (paperback). x + 236 pp. LCCN QA93 .S973 2010. Translated by Eva Burke.

Szpiro:2010:NRV

[Szp10b] George G. Szpiro. Numbers rule: the vexing mathematics of democracy, from Plato to the present. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-13994-6, 1-282-53153-0, 1-4008-3444-9 (e-book). xi + 3 + 226 pp. LCCN QA99 .S97 2010.

URL http://www.jstor.org/stable/10.2307/j.ctt7s773.

Szyszkowicz:2017:AQC

[Szy17] Mieczyslaw Szyszkowicz. 101.19 approximate quadrature of the circle using a set square. The Mathematical Gazette, 101(551): 301–302, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/product/96CF1E24AB79236FAFACC82F5797CE19.

Szyszkowicz:2019:SCU

[Szy19] Mieczyslaw Szyszkowicz. 103.13 squaring the circle by using proportion. The Mathematical Gazette, 103(556):156–158, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/10313-squaring-the-
circle-by-using-proportion/468B80859736DD196BEAAD57D5B9597E.

Talvila:2013:TGB

[Tal13] Erik Talvila. 97.06 Trigonometry of the gold-bug. The Mathematical Gazette, 97(538):124–127, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Taubes:2011:DGB

[Tau11] Clifford Taubes. Differential geometry: bundles, connections, metrics and curvature, volume 23 of Oxford mathematics; Oxford graduate texts in mathematics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-960588-2 (hardcover), 0-19-960587-4 (paperback). xiii + 298 pp. LCCN QA641 .T238 2011.

Taylor:2007:CHB

[Tay07] Patrick Taylor. Crosspacks and hypersolids: being the seventh part of several comprising the complete polyhedra. Complete polyhedra. Nattygrafix, Ipswich, UK, 2007. ISBN 0-9516701-8-2. 79 pp. LCCN ????

Tent:2008:ENM

[Ten08] M. B. W. (Margaret B. W.) Tent. Emmy Noether: the mother of modern algebra. A. K. Peters, Ltd., Wellesley, MA, USA, 2008. ISBN 1-56881-430-5. xvii + 177 pp. LCCN QA29.N6

T46 2008. URL http://www.loc.gov/catdir/toc/fy0906/
2008021713.html.

Tennant:2012:BRB

[Ten12] Geoff Tennant. Book review: Thinking mathematically (Second Edition), by John Mason, Leone Burton and Kaye Stacey. Pp.

248. £24.99. 2010. ISBN: 978-0-273-72891-7 (Pearson Education). The Mathematical Gazette, 96(536):362, July 2012. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Terrell:2009:SSI

[Ter09] William J. Terrell. Stability and stabilization: an introduction. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-13444-8 (hardcover). xv + 457 pp.

LCCN QA871 .T44 2009; QA871 .T46 2009. URL http:

//www.gbv.de/dms/bowker/toc/9780691134444.pdf;   http:
//www.loc.gov/catdir/enhancements/fy1204/2008926510-b.
html;   http://www.loc.gov/catdir/enhancements/fy1204/
2008926510-d.html;   http://www.loc.gov/catdir/enhancements/
fy1204/2008926510-t.html.

Thayer:2005:NPN

[Tha05] H. S. (Horace Standish) Thayer, editor. Newton’s philosophy of nature: selections from his writings. Dover Publications, Inc., New York, NY, USA, 2005. ISBN 0-486-44593-3 (paperback). xvi + 207 pp. LCCN B1299.N31 T4 2005. URL http://catdir.loc.

gov/catdir/enhancements/fy0624/2005048374-d.html.

Ting:2017:BRM

[Tin17] Vincent Ting. Book review: Math bytes by Tim Chartier, pp, 136, £16.95 (hard), ISBN 978-0-691-16060-3, Princeton University Press (2014). The Mathematical Gazette, 101(552):558–559, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/math-bytes-
by-tim-chartier-pp-136-1695-hard-isbn-9780691160603-princeton-
university-press-2014/6E98B2A1D957BBA2E9BDDFCEEBA1FF9C.

Tehrani:2016:SAS

[TL16] Fleur T. Tehrani and Gerry Leversha. A simple approach to solving cubic equations. The Mathematical Gazette, 100(548):225–232, July 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
B2D1D3ECF0CB9985DC6C7FB9A005A622.

Tobies:2012:IRL

[Tob12] Renate Tobies. Iris Runge: a life at the crossroads of mathematics, science and industry, volume 43 of Science networks historical studies. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2012. ISBN 3-0348-0251-X, 3-0348-0229-3 (print). xxii + 442, 16of plates. pp. LCCN ???? With a foreword by Helmut Neunzert. Revised by the author and translated by Valentine A. Pakis.

Todd:2015:MPL

[Tod15] Philip Todd. 99.08 A mechanical ‘proof’ of the largest outscribed equilateral triangle. The Mathematical Gazette, 99(544):149–151, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2010:BRB

[Tol10] Owen Toller. Book review: Professor Stewart’s cabinet of mathematical curiosities, by Ian Stewart. Pp. 256. £10.99. 2008. ISBN 978-1-84668-064-9 (Basic Books (NY), Profile Books Ltd (UK)). The Mathematical Gazette, 94(530):381, July 2010. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2013:BRB

[Tol13] Owen Toller. Book review: Progress to advanced mathematics, by Mary Teresa Fyfe, Andrew Jobbings, Kitty Kilday and Mary Read Pp. 168 £9.00 plus delivery 2011 ISBN: 978-0-9555477-3-7 (Arbelos). The Mathematical Gazette, 97(539):356–357, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2014:BRL

[Tol14a] Owen Toller. Book review: Looking at numbers, by Tom Johnson and Franck Jedrzejewski. Pp 120. £76.50. 2014. ISBN 978-3-0348-0553-7 (Birkhäuser). The Mathematical Gazette, 98(543): 564, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2014:BRB

[Tol14b] Owen Toller. Book review: Seduced by logic — Émilie du Châtelet, Mary Somerville and the Newtonian revolution, by Robyn Arianrhod. Pp. 338. £18.99, (hbk). 2012. ISBN 978-0-19-993161-3 (Oxford University Press). The Mathematical Gazette, 98(542):375, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2014:BRT

[Tol14c] Owen Toller. Book review: Transcending tradition: Jewish mathematicians in German-speaking academic culture edited by Birgit Bergmann, Moritz Epple and Rudi Ungar, pp 290, £35.99 (hard), £29.99 (e-book), ISBN 978-3-642-22463-8 (hard), 978-3-642-22464-5 (ebook), Springer Verlag (2013). The Mathematical Gazette, 98 (543):551–552, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2014:MLS

[Tol14d] Owen Toller. Misunderstandings in A-level statistics. The Mathematical Gazette, 98(541):40–50, March 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2014:SEF

[Tol14e] Owen Toller. Statistics education: fit for purpose? The Mathematical Gazette, 98(543):475–478, November 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2015:CAW

[Tol15a] Owen Toller. Book review: Calculus for the ambitious by T. W. Körner, pp 165, £17.99 (hard), ISBN 978-1-107-06392-1, Cambridge University Press (2014). The Mathematical Gazette, 99 (545):382–383, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2015:BRBa

[Tol15b] Owen Toller. Book review: Count like an Egyptian by David Reimer, pp 237 (hbk), £19.95, ISBN 978-0-691-16012-2, Princeton University Press (2014). The Mathematical Gazette, 99(544):176– 177, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2015:BRE

[Tol15c] Owen Toller. Book review: Elliptic tales by Avner Ash and Robert Gross, pp 253, $16.95 (paper), ISBN 978-0-691-16350-5, Princeton University Press (2012). The Mathematical Gazette, 99(546):564– 565, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2015:OMW

[Tol15d] Owen Toller. Book review: Origins of mathematical words by Anthony Lo Bello, pp 350, $32.00 (paper), ISBN 978-1-4214-1098-2, Johns Hopkins University Press (2013). The Mathematical Gazette, 99(545):380–381, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2015:BRBb

[Tol15e] Owen Toller. Book review: Padlock challenges, using number and algebra by Rachel Horsman, pp 80 (paper), £7.59 (members), £9.99 (non-members), ISBN 978-0-906588-79-6, The Mathematical Association (2013). Padlock challenges, statistics and probability by Rachel Horsman, pp 72 (paper), £7.59 (members), £9.99 (non-members), ISBN 978-0-906588-80-2, The Mathematical Association (2014). The Mathematical Gazette, 99(544):189–190, March 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2015:PSC

[Tol15f] Owen Toller. Book review: Professor Stewart’s casebook of mathematical mysteries by Ian Stewart, pp. 307. $16.99 (paper), ISBN 978-0-465-05497-8, Basic Books (2014). The Mathematical Gazette, 99(545):383, July 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2015:BRM

[Tol15g] Owen Toller. Book review: The Mathematical Association Calendar 2016, £6.99 (reductions for multiple purchases), The Mathematical Association. The Mathematical Gazette, 99(546):557, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2016:BRMa

[Tol16a] Owen Toller. Book review: Mathematics without apologies by Michael Harris, pp. 421, £19.95 (hard), ISBN 978-0-691-15423-7, Princeton University Press (2015). The Mathematical Gazette, 100 (547):178–179, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2016:BRNb

[Tol16b] Owen Toller. Book review: Nautical mathematics (paperback edition) by S. A. Watling and J. C. Hill, pp. 143, £24.99 (paper), ISBN 978-1-107-64252-2, Cambridge University Press (2014). The Mathematical Gazette, 100(547):184–185, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2016:BRNa

[Tol16c] Owen Toller. Book review: Numbers and functions, steps into analysis (3rd edn.) by R.P. Burn, pp. 347, £34.99 (paper), ISBN 978-1-107-44453-9, Cambridge University Press (2015). The Mathematical Gazette, 100(547):183, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2016:BRMb

[Tol16d] Owen Toller. Book review: The mathematics of map projections (paperback edition) by R. K. Melluish, pp. 145, £24.99 (paper), ISBN 978-1-107-65848-6, Cambridge University Press (2014). The Mathematical Gazette, 100(547):184–185, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2016:BRW

[Tol16e] Owen Toller. Book review: The world of numbers by Herbert McKay, pp. 198, £26.99 (paper), ISBN 978-1-107-49499-2, Cambridge University Press (2015). The Mathematical Gazette, 100 (547):179–180, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Toller:2017:BRBa

[Tol17a] Owen Toller. Book review: Basic discrete mathematics, by Kohar, Richard, pp. 706, £56 (paper), ISBN 978-9-81314-754-6, World Scientific (2016). The Mathematical Gazette, 101(551):375–376, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
903F24B6E090AC56DE90F27C1FE33F63.

Toller:2017:BRBb

[Tol17b] Owen Toller. Book review: Burn math class by Jason Wilkes, pp. 416, $29.99 (paper), ISBN 978-0-46505-373-5, Basic Books (2016). The Mathematical Gazette, 101(552):555–556, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/burn-math-class-by-jason-
wilkes-pp-416-2999-paper-isbn-9780465053735-basic-books-
2016/5A26D2E1893BB1E039F7ADBF6F97F789.

Toller:2017:BRC

[Tol17c] Owen Toller. Book review: Chancing it, by Matthews, Robert, pp. 290, £14.99 (hard), ISBN 978-1-781-25030-3, Profile Books (2015). The Mathematical Gazette, 101(550):181–182, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
2B08B8BCE9FBD5EFFE1A27F3B204E5AE.

Toller:2017:BRFa

[Tol17d] Owen Toller. Book review: Finding Fibonacci by Keith Devlin, pp. 241, £24.95, ISBN 978-0-691-17486-0, Princeton University Press (2017). The Mathematical Gazette, 101 (552):569–570, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/finding-fibonacci-by-keith-devlin-pp-241-2495-
isbn-9780691174860-princeton-university-press-2017/263E75E701900510EA12D5BE5458A741.

Toller:2017:BRFb

[Tol17e] Owen Toller. Book review: Fluke, the math and myth of coincidence by Joseph Mazur, pp. 288, $26.99 (paper), ISBN 978-0-46506-095-5, Basic Books (2016). The Mathematical Gazette, 101(552):570–571, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/fluke-the-math-and-myth-of-coincidence-by-joseph-
mazur-pp-288-2699-paper-isbn-9780465060955-basic-books-
2016/D72705DDFF7C209AAFEEC214547D7052.

Toller:2017:BRMb

[Tol17f] Owen Toller. Book review: From music to mathematics, by Roberts, Gareth E., pp. 320, $49.95 (hard), ISBN 978-1-42141-918-3, Johns Hopkins University Press (2016). The Mathematical Gazette, 101(550):170–171, March 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
7B12F90C2DFD32709441ACB4FCE1F57B.

Toller:2017:BRMa

[Tol17g] Owen Toller. Book review: L. A. Math by Stein, James D., pp. 240, $24.95 (hard), ISBN 978-0-69116-828-9, Princeton University Press (2015). The Mathematical Gazette, 101(550):168–169, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
F78F886EA177307737FE30D536065FF1.

Toller:2017:BRMc

[Tol17h] Owen Toller. Book review: My search for Ramanujan, by Ono, Ken and Aczel, Amir D., pp. 238, £19.99 (hard), ISBN 978-3-319-25566-8, also available as e-book, Springer Verlag (2016). The Mathematical Gazette, 101(551):355–356, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
5FF64DCA3E29D0122C8857EF4CFCBCB6.

Toller:2017:BRSa

[Tol17i] Owen Toller. Book review: Summing it up: from one plus one to modern number theory, by Ash, Avner and Gross, Robert, pp. 229, $27.95 (hard), ISBN 978-0-691-17019-0, Princeton University Press (2016). The Mathematical Gazette, 101(550):174–175, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
8219F88550DEBCD3D3302AE4CFCB0EFB.

Toller:2017:BRT

[Tol17j] Owen Toller. Book review: Teaching probability, by Gage, Jenny and Speigelhalter, David, pp. 197, £17.95 (paper), ISBN 978-1-316-60589-9, Cambridge University Press (2016). The Mathematical Gazette, 101(550):183–184, March 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
DEAC8D0B7B0B76EDC32B5887DC21DF61.

Toller:2017:BRSb

[Tol17k] Owen Toller. Book review: The seven pillars of statistical wisdom, by Stigler, Stephen M., pp. 230, £16.95 (paper), ISBN 978-0-674-08891-7, Harvard University Press (2016). The Mathematical Gazette, 101(550):182–183, March 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
0EB39D615E03888BBA24B47F9DA69992.

Toller:2018:BRC

[Tol18a] Owen Toller. Book review: Closing the gap by Vicky Neale, pp. 164, £19.99 (hard), ISBN 978-0-19878-828-7, Oxford University Press (2017). The Mathematical Gazette, 102(555):561, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/closing-the-gap-
by-vicky-neale-pp-164-1999-hard-isbn-9780198788287-oxford-
university-press-2017/8AE5E573478D1E807BF3BDE2404EE68E.

Toller:2018:BRI

[Tol18b] Owen Toller. Book review: Introductory statistics by Sheldon M. Ross, pp. 828, £76 (hard), ISBN 978-0-12804-317-2, also available as e-book, Academic Press (Elsevir) (2017). The Mathematical Gazette, 102(554):377–378, July 2018. CODEN MAGAAS.

ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/introductory-statistics-by-sheldon-m-ross-pp-828-
76-hard-isbn-9780128043172-also-available-as-ebook-academic-
press-elsevir-2017/5E73DA529201FF3EC89D17A9AB0EF5EB.

Toller:2018:BRP

[Tol18c] Owen Toller. Book review: Pushing limits by Ted Hill, pp. 224, £46.50 (hard), ISBN 978-1-47043-584-4, AMS/Mathematical Association of America (2017). The Mathematical Gazette, 102(553): 185, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print),

2056-6328 (electronic). URL https://www.cambridge.org/
core/journals/mathematical-gazette/article/pushing-limits-
by-ted-hill-pp-224-4650-hard-isbn-9781470435844-amsmathematical-
association-of-america-2017/7119CFF871AA45D99B0D0A63D255226C.

Toller:2018:BRT

[Tol18d] Owen Toller. Book review: Ten great ideas about chance by Persi Diaconis and Brian Skyrms, pp. 272, £22.95 (hard), ISBN 978-0-69117-416-7, Princeton University Press (2017). The Mathematical Gazette, 102(555):567–568, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/ten-great-ideas-about-chance-by-persi-diaconis-
and-brian-skyrms-pp-272-2295-hard-isbn-9780691174167-princeton-
university-press-2017/E765081A5A754CAF1F7243405CE08E2F.

Toller:2018:BRJ

[Tol18e] Owen Toller. Book review: The joy of mathematics by Alfred S. Posamentier, Christian Li, Christian Spreitzer and Robert Geretschläger, pp. 302, £15.99 (paper), ISBN 978-1-63388-297-3, also available as e-book, Prometheus Books (2017). The Mathematical Gazette, 102(554):364–365, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/joy-of-mathematics-by-alfred-s-posamentier-christian-
li-christian-spreitzer-and-robert-geretschlager-pp-302-
1599-paper-isbn-9781633882973-also-available-as-ebook-
prometheus-books-2017/49B307EE80643EF022986F85FFA251D4.

Toller:2018:BRU

[Tol18f] Owen Toller. Book review: Unsolved! by Craig P. Bauer, pp. 620, £27.95 (hard), ISBN 978-0-691-16767-1, Princeton University Press (2017). The Mathematical Gazette, 102(554):375–376, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/unsolved-by-craig-
p-bauer-pp-620-2795-hard-isbn-9780691167671-princeton-
university-press-2017/E8BF574892CFC985F678AFED28B1CF71.

Toller:2019:BRA

[Tol19a] Owen Toller. Book review: All sides to an oval by Angelo Alessandro Mazzotti, pp. 160, £27.99 (hard), ISBN 978-3-319-39374-2, also available as e-book, Springer Verlag (2017). The Mathematical Gazette, 103(558):572, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/all-sides-to-an-
oval-by-angelo-alessandro-mazzotti-pp-160-2799-hard-isbn-
9783319393742-also-available-as-ebook-springer-verlag-
2017/89290F690347F9D5980B94AFA4A36FA8.

Toller:2019:DGT

[Tol19b] Owen Toller. Book review: Discovering group theory by Tony Barnard and Hugh Neill, pp. 219, £31.99 (paper), ISBN 978-1-13803-016-9, CRC Press (2017). The Mathematical Gazette, 103(557):368–369, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/discovering-group-theory-by-tony-barnard-and-hugh-
neill-pp-219-3199-paper-isbn-9781138030169-crc-press-2017/
8B6306C51C0FFFC241815E627854268D.

Toller:2019:IAL

[Tol19c] Owen Toller. Book review: Introduction to applied linear algebra by Stephen Boyd and Lieven Vandenberghe, pp. 463, £35.99 (hard), ISBN 978-1-31651-896-0, Cambridge University Press (2019). The Mathematical Gazette, 103(558):573–574, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/introduction-to-
applied-linear-algebra-by-stephen-boyd-and-lieven-vandenberghe-
pp-463-3599-hard-isbn-9781316518960-cambridge-university-
press-2019/2A419040D388D0B14085AA199C421827.

Toller:2019:BRM

[Tol19d] Owen Toller. Book review: Music by the numbers by Eli Maor, pp. 185, £19.95 (hard), ISBN 978-0-69117-690-1, Princeton University Press (2018). The Mathematical Gazette, 103 (556):186–187, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/music-by-the-numbers-by-eli-maor-pp-185-1995-hard-
isbn-9780691176901-princeton-university-press-2018/110B47F203F511D94EC6850E3CC726F1.

Toller:2019:TSD

[Tol19e] Owen Toller. Book review: Teaching statistics by Darren Macey and Will Hornby, pp. 190, £19.95 (paper), ISBN 978-1-10840-630-7, Cambridge University Press (2018). The Mathematical Gazette, 103(557):377–378, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/teaching-statistics-by-darren-macey-and-will-hornby-
pp-190-1995-paper-isbn-9781108406307-cambridge-university-
press-2018/A1BE532B2340B056B712B11FD81DC841.

Toller:2019:MTS

[Tol19f] Owen Toller. Book review: The model thinker by Scott E. Page, pp. 410, $19.99 (paper), ISBN 978-0-465-09463-9, also available as e-book, Basic Books (2018). The Mathematical Gazette, 103(557):376, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/model-thinker-by-scott-e-page-pp-410-1999-paper-
isbn-9780465094639-also-available-as-ebook-basic-books-
2018/3136660E3FDA77C2FCA1C08172E334CC.

Toller:2019:BRT

[Tol19g] Owen Toller. Book review: Towards higher mathematics, a companion by Richard Earl, pp. 527, £29.99 (paper), ISBN 978-1-31661-483-9, Cambridge University Press (2018). The Mathematical Gazette, 103(556):182–183, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/towards-higher-mathematics-a-companion-by-richard-
earl-pp-527-2999-paper-isbn-9781316614839-cambridge-university-
press-2018/41E8D46AE5AC8AA6B61D7D5EDD61CC7F.

Toller:2019:PFT

[Tol19h] Owen Toller. Partial fractions and telescoping series. The Mathematical Gazette, 103(556):165–167, March 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/partial-fractions-and-telescoping-
series/CA001F746BA5FF52A3C804B7BD18C72D.

Tonien:2016:SPE

[Ton16] Joseph Tonien. A simple proof of Euler’s continued fraction of e^{1/M}. The Mathematical Gazette, 100(548):279–287, July 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
13F06F0985DC789AB8CC3A785D2685F1.

Tonien:2017:CFI

[Ton17] Joseph Tonien. 101.08 A continued fraction inspired by an identity of Euler. The Mathematical Gazette, 101(550):115–121, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
05633BCAB00FA4979EF8B6502687CAF7.

Tonien:2018:NEP

[Ton18]  Joseph Tonien.  102.06 a new elementary proof of Eu-

ler’s continued fractions. The Mathematical Gazette, 102 (553):105–111, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10206-a-new-elementary-proof-of-eulers-continued-
fractions/B38A60A4EE600B83EA98F3D7AEEE9303.

Tonien:2019:NYI

[Ton19] Joseph Tonien. 103.01 New Year identities for 2019. The Mathematical Gazette, 103(556):130, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10301-new-year-identities-for-2019/F2E1EE1C40C4B57196AA533E57E9533B.

Torsey:2017:NBE

[Tor17] Bridget M. Torsey. 101.45 a note on bounds for the expected value of a random variable. The Mathematical Gazette, 101 (552):522–525, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10145-a-note-on-bounds-for-the-expected-value-
of-a-random-variable/58248BBE2395C1E03B226E8476CCB057.

Trainin:2010:IEF

[Tra10] J. Trainin. Integrating expressions of the form sin^{n} x/x^{m} and others. The Mathematical Gazette, 94(530):216–223, July 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Trainin:2013:BAS

[Tra13] J. E. T. Trainin. 97.42 The Bombelli approach to solving cubic equations. The Mathematical Gazette, 97(540):501–504, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Treeby:2017:FSP

[Tre17] David Treeby. Finding sums of powers using physical arguments. The Mathematical Gazette, 101(551):227–235, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
449642B1DC3EDD01B5B3ADB2EC7ADC0C.

Tsao:2011:SPE

[Tsa11] Hung-Ping Tsao. 95.47 Sums of powers and Eulerian numbers. The Mathematical Gazette, 95(533):347–349, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Tsaknakis:2018:BRJ

[Tsa18] George Tsaknakis. Book review: Junior problems by Andrew Jobbings, pp. 250, £15. (paper), ISBN 978-1-90600-130-8, UK Mathematics Trust (2017). The Mathematical Gazette, 102(554):382, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/junior-problems-
by-andrew-jobbings-pp-250-15-paper-isbn-9781906001308-
uk-mathematics-trust-2017/43BB0DD9FAC446A0CD03957B3353C8D5.

Tubbs:2009:WNM

[Tub09] Robert Tubbs. What is a number?: Mathematical Concepts and their Origins. The Johns Hopkins University Press, Baltimore, MD, USA, 2009. ISBN 0-8018-9018-7 (paperback), 0-8018-9017-9 (hardcover). x + 2 + 305 pp.

LCCN QA8.4 .T83 2009. URL http://catdir.loc.gov/
catdir/enhancements/fy0829/2008013745-b.html;   http:/
/catdir.loc.gov/catdir/enhancements/fy0829/2008013745-
d.html;   http://catdir.loc.gov/catdir/toc/fy0904/2008013745.
html.

Tubbs:2014:MTC

[Tub14] Robert Tubbs. Mathematics in twentieth-century literature and art: content, form, meaning. The Johns Hopkins University Press, Baltimore, MD, USA, 2014. ISBN 1-4214-1379-5 (hardcover), 1-4214-1380-9 (paperback), 1-4214-1402-3 (electronic). xvi + 162 pp. LCCN PN56.M36 T83 2014.

VanBrummelen:2009:MHE

[Van09] Glen Van Brummelen. The mathematics of the heavens and the earth: the early history of trigonometry. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-12973-8 (hardcover). xvii + 329 pp. LCCN QA24 .V36 2009. URL http://www.loc.

gov/catdir/toc/ecip0824/2008032521.html.

VanBrummelen:2013:HMF

[Van13] Glen Van Brummelen. Heavenly mathematics: the forgotten art of spherical trigonometry. Princeton University Press, Princeton, NJ, USA, 2013. ISBN 0-691-14892-9 (hardcover). xvi + 192, [8] col. pp. LCCN QA535 .V36 2013.

vanDalen:2013:JBT

[vDD13] D. van (Dirk) Dalen. L. E. J. Brouwer: topologist, intuitionist, philosopher: how mathematics is rooted in life. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4471-4615-8 (hardcover). xii + 875 pp. LCCN QA29.B697 D343 2013.

vanderVen:2012:HDG

[vdV12] A. H. G. S. van der Ven. The Hardy distribution for golf hole scores. The Mathematical Gazette, 96(537):428–438, November 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Vedral:2010:DRU

[Ved10] Vlatko Vedral. Decoding reality: the universe as quantum information. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2010. ISBN 0-19-923769-7 (hardcover), 0-19-969574-1 (paperback), 0-19-157267-5 (e-book). x + 229 pp. LCCN Q360 .V43 2010; QC6 .V42 2010.

Velten:2009:MMS

[Vel09] Kai Velten. Mathematical modeling and simulation: introduction for scientists and engineers. Wiley-VCH, Weinheim, Germany, 2009. ISBN 3-527-40758-8 (paperback). xiv + 348 pp. LCCN QA401 .V38 2009.

Viglione:2019:ETS

[Vig19] Raymond Viglione. 103.25 An extension of Thébault’s second problem. The Mathematical Gazette, 103(557):343–346, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10325-an-extension-of-thebaults-
second-problem/67FD4D9176B9F14F2736D6BC166B8A40.

Vijiitu:2019:BIN

[Vîj19] Horia Vîjîitu. 103.19 Bernoulli’s inequality for negative exponents. The Mathematical Gazette, 103(557):316–317, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10319-bernoullis-inequality-
for-negative-exponents/61CF64752EC503F651F4C94D0A8DB0B5.

Villegas:2007:ENT

[Vil07] Fernando Rodriguez Villegas. Experimental number theory, volume 13 of Oxford graduate texts in mathematics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2007. ISBN 0-19-852822-1 (hardcover), 0-19-922730-6 (paperback). xvi + 214 pp.

LCCN QA241 .V43 2007. URL http://www.loc.gov/catdir/
toc/fy0801/2007299209.html.

Villarino:2011:BMO

[Vil11] Mark B. Villarino. 95.56 Burning mirrors on opening day. The Mathematical Gazette, 95(534):512–514, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Villarino:2014:CSR

[Vil14a] Mark B. Villarino. A cubic surface of revolution. The Mathematical Gazette, 98(542):281–290, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Villarino:2014:QPP

[Vil14b] Mark B. Villarino. Quotient polynomials with positive coefficients. The Mathematical Gazette, 98(542):250–255, July 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Villarino:2015:TS

[Vil15] Mark B. Villarino. A Tripos surd. The Mathematical Gazette, 99 (546):474–480, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Vieli:2019:SCP

[VM19] Francisco Javier González Vieli and Marion Maillard. 103.41 A spatial characterisation of Pascal limaçons. The Mathematical Gazette, 103(558):531–536, November 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/10341-a-spatial-characterisation-of-pascal-limacons/
718635E630E286EF6A208EA685EB26CA.

Volpert:2011:EPD

[Vol11] Vitaly A. Volpert. Elliptic Partial Differential Equations: Volume 1: Fredholm Theory of Elliptic Problems in Unbounded Domains, volume 101, 104 of Monographs in Mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2011. ISBN 3-0346-0536-6 (vol. 1: hardcover), 3-0346-0537-4 (vol. 1: ebook), 3-0348-0812-7 (vol. 2: hardcover), 3-0348-0813-5 (vol. 2: e-book). ISSN 1017-0480. xvii + 1 + 639 pp. LCCN QA377 .V638

2011; QA377 .V65 2011. URL http://www.gbv.de/dms/bowker/
toc/9783034605366.pdf.

VonStengel:2017:BRJ

[Von17] Bernhard Von Stengel. Book review: Jane Austen, game theorist (updated edition), by Chwe, Michael Suk-Young, pp 296, £15.95 (paper), ISBN: 978-0-691-16244-7, Princeton University Press (2014). The Mathematical Gazette, 101(550):172, March 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
46A1B01F6D76B1AEBE59326FB7E77135.

Villarino:2013:MCT

[VSR13] Mark B. Villarino, Daniel Campos Salas, and Javier Carvajal Rojas. 91.23 On the monotonocity of the correction term in Ramanujan’s factorial approximation. The Mathematical Gazette, 97(539):

274–275, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Volchkov:2009:HAM

[VV09] V. V. (Valeri V.) Volchkov and Vitaly V. Volchkov. Harmonic analysis of mean periodic functions on symmetric spaces and the Heisenberg group. Springer monographs in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-84882-532-3. ISSN 1439-7382. xi + 671 pp. LCCN QA403 .V65 2009. URL http:

//www.loc.gov/catdir/enhancements/fy1106/2009925998-t.
html;   http://www.loc.gov/catdir/enhancements/fy1409/
2009925998-d.html.

Vandebril:2008:MCS

[VvBM08] Raf Vandebril, Marc van Barel, and Nicola Mastronardi. Matrix Computations and Semiseparable Matrices. The Johns Hopkins University Press, Baltimore, MD, USA, 2008. ISBN 0-8018-8714-3 (vol. 1, hardcover), 0-8018-9052-7 (vol. 2). xviii + 575 (vol. 1), xvi + 498 (vol. 2) pp. LCCN QA188 .V36 2008. URL http:

//www.loc.gov/catdir/enhancements/fy0740/2007030657-b.
html;   http://www.loc.gov/catdir/enhancements/fy0740/
2007030657-d.html;   http://www.loc.gov/catdir/toc/ecip0723/
2007030657.html.

vanWyk:2019:RVT

[vW19] Leonard van Wyk. Rounding versus truncation estimates in difference calculations. The Mathematical Gazette, 103(557):285–292, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-

6328 (electronic). URL https://www.cambridge.org/core/
journals/mathematical-gazette/article/rounding-versus-
truncation-estimates-in-difference-calculations/57BD92568C03F0FCDE63BBBDB5E6FF6D.

Wainer:2009:PUW

[Wai09] Howard Wainer. Picturing the uncertain world: how to understand, communicate, and control uncertainty through graphical display. Princeton University Press, Princeton, NJ, USA, 2009. ISBN 0-691-13759-5 (hardcover). xviii + 244 + 16 pp. LCCN Q375 .W35 2009.

Wapner:2012:UEC

[Wap12] Leonard M. Wapner. Unexpected expectations: the curiosities of a mathematical crystal ball. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2012. ISBN 1-56881-721-5 (hardback). xvi + 204 pp. LCCN QA95 .W34 2012.

Wapner:2019:PPI

[Wap19] Leonard M. Wapner. The prescient power of indifference. The Mathematical Gazette, 103(556):41–46, March 2019. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/prescient-power-of-indifference/
7C187369FF8F9E48063D26DA01824EFE.

Ward:2010:BRBa

[War10a] Michael Ward. Book review: 50 mathematical ideas you really need to know, by Tony Crilly. pp. 208. £8.99. 2008. ISBN 978-1-84724-008-8 (Quercus Publishing plc). The Mathematical Gazette, 94 (530):365, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ward:2010:BRBc

[War10b] Michael Ward. Book review: Before sudoku: the world of magic squares, by Seymour S. Block and Santiago A. Tavares. Pp. 239. £7.99. 2009. ISBN: 978-0-19-536790-4 (Oxford University Press). The Mathematical Gazette, 94(531):572–573, November 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ward:2010:BRBb

[War10c] Michael Ward. Book review: Invariant descriptive set theory, by Su Gao. Pp. 383. £49.99. 2008. ISBN 978-1-58488-793-5 (Chapman and Hall). The Mathematical Gazette, 94(530):383, July 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Wardhaugh:2010:HRH

[War10d] Benjamin Wardhaugh. How to read historical mathematics. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-14014-6 (hardcover). xii + 116 pp. LCCN QA21 .W324 2010.

Ward:2011:BRBa

[War11a] Michael Ward. Book review: Decoding reality: the universe as quantum information, by Vlatko Vedral. Pp. 229. £16.99. 2010. ISBN 978-0-19-923769-2 (Oxford University Press). The Mathematical Gazette, 95(532):151–152, March 2011. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ward:2011:BRBb

[War11b] Michael Ward. Book review: Flatland: an edition with notes and commentary, by Edwin A. Abbott, William F. Lindgren and Thomas F. Banchof. pp. 294. £9.99. 2010. ISBN 978-0-521-75994-6. (Cambridge University Press). The Mathematical Gazette, 95 (533):381–382, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ward:2012:BRB

[War12a] Michael Ward. Book review: Figuring it out: entertaining encounters with everyday math, by Nuno Crato. Pp. 227. £24.99. 2010. ISBN: 978-3-642-04832-6 (Springer). The Mathematical Gazette, 96(536):382–383, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Wardhaugh:2012:PRP

[War12b] Benjamin Wardhaugh. Poor Robin’s prophecies: a curious Almanac, and the everyday mathematics of Georgian Britain. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2012. ISBN 0-19-960542-4 (hardcover), 0-19-164456-0 (e-book), 1-283-60967-3 (MyiLibrary). viii + 248 pp. LCCN QA27.G7 W37 2012.

Wardhaugh:2012:WNA

[War12c] Benjamin Wardhaugh, editor. A wealth of numbers: an anthology of 500 years of popular mathematics writing. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-14775-2 (hardcover). xv + 370 pp. LCCN QA7 .W43 2012.

Ward:2013:BRBa

[War13a] Michael Ward. Book review: Mathematics is precise, period, vs. mathematics is precise, strings attached, by William J. Adams. Pp. 492. £23.00. 2010 ISBN: 978-1-4363-3419-8 (Xlibris Corporation). The Mathematical Gazette, 97(539):367, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ward:2013:BRBb

[War13b] Michael Ward. Book review: Train your brain: a year’s worth of puzzles, by George Grätzer. Pp. 235. £18.99. 2011. ISBN: 978-1-56881-710-1 (CRC Press). The Mathematical Gazette, 97(539):380, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Ward:2013:BRBc

[War13c] Michael Ward. Book review: Train your brain: a year’s worth of puzzles, by George Grätzer. Pp. 235. £18.99. 2011. ISBN: 978-1-56881-710-1 (CRC Press). The Mathematical Gazette, 97(539):380, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Weinstein:2012:RZP

[Wei12] Jonathan Weinstein. 96.48 A result on zigzag permutations: a combinatorial proof. The Mathematical Gazette, 96(536):331–333, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Wells:2008:WPM

[Wel08] D. G. (David G.) Wells. What’s the point?: Motivation and the mathematics crisis. Rain, London, UK, 2008. ISBN 0-907944-11-6 (paperback). 309 pp. LCCN ????

Wells:2012:GMS

[Wel12] D. G. (David G.) Wells. Games and mathematics: subtle connections. Cambridge University Press, Cambridge, UK, 2012. ISBN 1-107-69091-9 (paperback), 1-107-02460-9 (hardcover). x + 246

pp. LCCN QA95 .W438 2012. URL http://assets.cambridge.
org/97811070/24601/cover/9781107024601.jpg.

Werthamer:2009:RRS

[Wer09] N. Richard Werthamer. Risk and Reward: The Science of Casino Blackjack. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-4419-0252-X (hardcover). xviii + 131 pp. LCCN GV1295.B55 W47 2009.

Wheelan:2013:NSS

[Whe13] Charles J. Wheelan. Naked statistics: stripping the dread from the data. W. W. Norton and Company, New York, NY, USA, 2013. ISBN 0-393-07195-2 (hardcover). xviii + 282 pp. LCCN QA276 .W458 2013.

Wilson:2008:LCN

[Wil08] Robin James Wilson. Lewis Carroll in Numberland: his fantastical mathematical logical life: an agony in eight fits. Penguin Group, London, UK, 2008. ISBN 0-7139-9757-5 (hardcover). xi + 237 pp. LCCN PR4612 .W55 2008; QA29.C353 W55 2008.

Wilson:2009:FSG

[Wil09] Robert Wilson. The finite simple groups, volume 251 of Graduate texts in mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 1-84800-987-9, 1-84800-988-7 (e). ISSN 0072-5285. xv + 298 pp. LCCN QA177 .W55 2009.

Williams:2010:BRB

[Wil10] Hugh Williams. Book review: Symmetry and the monster: one of the greatest quests of mathematics, by Mark Ronan. Pp. 272. £8-99. 2006. ISBN: 978-0-19-280722-9 (Oxford University Press). The Mathematical Gazette, 94(529):173–174, March 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Willson:2012:PTL

[Wil12] William Wynne Willson. Ptolemy through the looking-glass: a proof of his theorem using transformations. Garland Publications, ????, 2012. ISBN 0-9558042-3-X. LCCN QA455 .W55 2012.

Williams:2014:LVU

[Wil14] Paul G. Williams. Leaving the variance unchanged. The Mathematical Gazette, 98(543):520–525, November 2014. CODEN MA-GAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Wilson:2017:EPA

[Wil17] Robert A. Wilson. 101.15 an elementary proof that not all principal ideal domains are Euclidean domains. The Mathematical Gazette, 101(551):289–293, July 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
C862415E89978A0796EE1FC2738112C7.

Williams:2018:MPC

[Wil18] Gareth Williams. 102.09 measurement of polygons and circles. The Mathematical Gazette, 102(553):121–127, March 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/10209-measurement-of-polygons-
and-circles/A60BA36519271F956EB47D198E39A19B.

Wang-Iverson:2011:OFI

[WILY11] Patsy Wang-Iverson, Robert J. (Robert James) Lang, and Mark Yim, editors. Origami 5: Fifth International Meeting of Origami Science, Mathematics, and Education. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-56881-714-2 (paperback). xiv + 646 pp. LCCN QA491 .I55 2011.

Wiorkowski:2019:FPR

[Wio19] John Wiorkowski. Finding the probability of a rare real world event. The Mathematical Gazette, 103(557):240–247, July 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/journals/
mathematical-gazette/article/finding-the-probability-of-
a-rare-real-world-event/4AA2C76A72BF3E31F370F0246574BED6.

Williams:2010:MWL

[WMW10] Kim Williams, Lionel March, and Stephen R. Wassell. The mathematical works of Leon Battista Alberti. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2010. ISBN 3-0346-0473-4 (hardcover). x + 221 pp. LCCN QA32 .M38 2010. Translations and commentary by Kim Williams, Lionel March, and Stephen R. Wassell. Foreword by Robert Tavernor. With contributions by Richard Schofield and Angela Pintore.

Wright:2011:IGH

[WO11] Marcus Wright and Thomas J. Osler. 95.41 An interesting generalisation of the half-angle formulas to complex numbers. The Mathematical Gazette, 95(533):327, July 2011. CODEN MAG-AAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Wright:2013:IBS

[WO13] Cory Wright and Thomas J. Osler. 97.08 Interesting bilateral series generalising a result of Euler. The Mathematical Gazette, 97 (538):131–133, March 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Wright:2009:MM

[Wri09] David Wright. Mathematics and music, volume 28 of Mathematical world. American Mathematical Society, Providence, RI, USA, 2009. ISBN 0-8218-4873-9. xiv + 161 pp. LCCN ML3809 .W85

2009. URL http://www.ams.org/bookpages/mawrld-28.

Wilson:2013:CAM

[WW13] Robin J. Wilson and John J. Watkins, editors. Combinatorics: ancient and modern. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2013. ISBN 0-19-965659-2 (hardcover). x + 381 pp. LCCN QA164 .C66335 2013. URL http://librarytitles.

ebrary.com/id/10722743.

Yang:2010:CQF

[Yan10] Kung-Wei Yang. 94.26 Computing quadratic fundamental units. The Mathematical Gazette, 94(531):464–467, November 2010. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Yang:2014:SCF

[Yan14] Kung-Wei Yang. 98.29 symmetry of continued fraction for \sqrt{d}. The Mathematical Gazette, 98(543):501–503, November 2014. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Yang:2015:PEP

[Yan15] Kung-Wei Yang. 99.30 the positive eigenvector of a positive matrix. The Mathematical Gazette, 99(546):523–525, November 2015. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Yeo:2016:BRB

[Yeo16] Dominic Yeo. Book review: Balkan Mathematical Olympiads by Mircea Becheanu & Bogdan Enescupp pp. 273, $59.95 (hard), ISBN 978-0-98856-225-7, XYZ Press (2014). The Mathematical Gazette, 100(547):185–186, March 2016. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Yeo:2017:BRC

[Yeo17a] Dominic Yeo. Book review: 1000 challenges mathématiques, algèbre, by Aassila, Mohammed, pp. 643, EUR42.00 (paper), ISBN 978-2-34001-108-3, Editions Ellipses (2016). 1000 challenges mathématiques, analyse by Aassila Mohammed, pp. 644, EUR42.00 (paper), ISBN 978-2-34001-109-0, Editions Ellipses (2016). The Mathematical Gazette, 101(551):381–382, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
21C6F761BE274F8F1F877964D73E916F.

Yeo:2017:BRM

[Yeo17b] Dominic Yeo. Book review: Methods and techniques for proving inequalities, by Su, Yong and Xiong, Bin, pp. 222, £21.00 (paper), ISBN 978-9-81469-645-6, World Scientific (2016).

The Mathematical Gazette, 101(551):366–367, July 2017. CO-DEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (elec-

tronic). URL https://www.cambridge.org/core/product/
6F568D6EE542A69061F7F68CC090DD39.

Yeo:2017:BRP

[Yeo17c] Dominic Yeo. Book review: Problem-solving challenges for secondary school students and above, by Linker, David and Sultan, Alan, pp. 181, £18.00 (paper), ISBN 978-9-81473-003-7, World Scientific (2016). The Mathematical Gazette, 101(551):379–380, July 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328

(electronic). URL https://www.cambridge.org/core/product/
A6E9E8C94A02A13FC857C4A50C7E227D.

Yeo:2017:BRR

[Yeo17d] Dominic Yeo. Book review: The real analysis lifesaver by Raffi Grinberg, pp. 193, £22.95 (paper), ISBN 978-0-69117-293-4, Princeton University Press (2017). The Mathematical Gazette, 101(552):561–562, November 2017. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/the-real-analysis-lifesaver-by-raffi-grinberg-
pp-193-2295-paper-isbn-9780691172934-princeton-university-
press-2017/CF9D941F49E768BAC6D62ED0B0C54459.

Yeo:2018:BRC

[Yeo18a] Dominic Yeo. Book review: Coming to know: the introduction of new concepts in undergraduate mathematics by R. P. Burn, pp. 114, £8.59, ISBN 978-0-955-59142-6, MSOR Network (available via lulu.com) (2009). The Mathematical Gazette, 102(555): 557–558, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https://www.cambridge.

org/core/journals/mathematical-gazette/article/coming-
to-know-the-introduction-of-new-concepts-in-undergraduate-
mathematics-by-r-p-burn-pp-114-859-isbn-9780955591426-
msor-network-available-via-lulucom-2009/7C6B27C390AA4EF008EB97119042BD48.

Yeo:2018:BRM

[Yeo18b] Dominic Yeo. Book review: Making and breaking mathematical sense by Roi Wanger, pp. 236, £37.95 (hard), ISBN 978-0-69117-171-5, Princeton University Press (2017). The Mathematical Gazette, 102(554):381–382, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/

article/making-and-breaking-mathematical-sense-by-roi-
wanger-pp-236-3795-hard-isbn-9780691171715-princeton-university-
press-2017/37419C1BA8840329E129AECD22591E68.

Yeo:2018:BRU

[Yeo18c] Dominic Yeo. Book review: Undergraduate convexity, problems and solutions by Mikkel Slot Nielsena and Victor Ulrich Rohde, pp. 184, £32.00 (paper), ISBN 978-9-81314-364-7, World Scientific (2017). The Mathematical Gazette, 102 (555):561–562, November 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/undergraduate-convexity-problems-and-solutions-
by-mikkel-slot-nielsena-and-victor-ulrich-rohde-pp-184-
3200-paper-isbn-9789813143647-world-scientific-2017/98FB3C731C2F46209B0F7B28860EFC41.

Yacoub:2012:SQE

[YF12] Michel Daoud Yacoub and Gustavo Fraidenraich. 96.33 A solution to the quartic equation. The Mathematical Gazette, 96(536):271– 275, July 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Yadav:2011:AIL

[YM11] B. S. (Bhuri Singh) Yadav and Man Mohan, editors. Ancient Indian leaps into mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 2011. ISBN 0-8176-4694-9. xx + 217 pp. LCCN QA27.I4 A53 2011. URL http:

//www.loc.gov/catdir/enhancements/fy1114/2010937771-d.
html;   http://www.loc.gov/catdir/enhancements/fy1114/
2010937771-t.html.

Yau:2010:SIS

[YN10] Shing-Tung Yau and Steven J. Nadis. The shape of inner space: string theory and the geometry of the Universe’s hidden dimensions. Basic Books, New York, NY, USA, 2010. ISBN 0-465-02023-2, 0-465-02837-3 (paperback). xvii + 377 pp. LCCN QA691 .Y38 2010.

Yau:2012:SIS

[YN12] Shing-Tung Yau and Steven J. Nadis. The shape of inner space: string theory and the geometry of the Universe’s hidden dimensions. Basic Books, New York, NY, USA, 2012. ISBN 0-465-02837-3 (paperback), 0-465-02023-2, 0-465-02266-9 (e-book). xvii + 1 + 377 pp. LCCN QA691 .Y38 2012.

Young:2011:PT

[You11a] Robert M. Young. 95.31 ∫_{0}^{∞} 1/(1 + x^{2}) dx = π/2: A proof without trigonometry. The Mathematical Gazette, 95(533):304, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Young:2011:EP

[You11b] Robert M. Young. 95.32 int_{0}^{1}1/(1 + x^{2}) dx = π/4: an elementary proof. The Mathematical Gazette, 95(533):304–307, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Young:2011:EPI

[You11c] Robert M. Young. 95.34 On evaluating the probability integral. The Mathematical Gazette, 95(533):311–313, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Young:2012:EPL

[You12] Robert M. Young. 96.10 On an elementary proof of the Leibniz formula for π. The Mathematical Gazette, 96(535):116–117, March 2012. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Young:2013:AQP

[You13] Robert M. Young. Another quick proof that e < 3. The Mathematical Gazette, 97(539):333–334, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Zarzycki:2010:BRBb

[Zar10a] Piotr Zarzycki. Book review: A Mathematical Olympiad Primer, by Geoff Smith. Pp. 178. £10.00. 2008. ISBN 978-1-906001-03-2 (The United Kingdom Mathematics Trust). The Mathematical Gazette, 94(529):187–188, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Zarzycki:2010:BRBa

[Zar10b] Piotr Zarzycki. Book review: The teaching assistant’s guide to numeracy, by Sara Fielder. Pp. 117. £12.99. 2007. ISBN 978-0-8264-9606-5 (Continuum). The Mathematical Gazette, 94(529):181–182, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Zarzycki:2011:BRB

[Zar11] Piotr Zarzycki. Book review: Introduction to the practice of statistics, by David S. Moore, George P. McCabe, Bruce A. Craig. pp. 710. £49.99. 2009. ISBN 978-1-4292-1623-4 (W. H. Freeman and Company). The Mathematical Gazette, 95(533):380–381, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Zetie:2013:BRB

[Zet13] Ken Zetie. Book review: Cosmic numbers: the numbers that define our universe, by James D. Stein. Pp. 272. £17.99. 2011. ISBN 978-0-465-02198-7 (Basic Books). The Mathematical Gazette, 97(539): 376–378, July 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Zetie:2018:BRE

[Zet18] Ken Zetie. Book review: Emmy Noether’s wonderful theorem (revised and updated edition) by Dwight E. Neuenschwander, pp. 344, £22 (paper), ISBN 978-1-42142-267-1, Johns Hopkins University Press (2017). The Mathematical Gazette, 102(554):378–379, July 2018. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/emmy-noethers-wonderful-theorem-revised-and-updated-
edition-by-dwight-e-neuenschwander-pp-344-22-paper-isbn-
9781421422671-johns-hopkins-university-press-2017/988A71D612D290D8747F7CCA1B9908CC.

Zetie:2019:BRQ

[Zet19] Ken Zetie. Book review: Q is for quantum by Terry Rudolph, pp. 164, £9.95 (paper), ISBN 978-0-99906-350-7, Terrence Rudolph (2017). The Mathematical Gazette, 103 (556):185–186, March 2019. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic). URL https:/

/www.cambridge.org/core/journals/mathematical-gazette/
article/q-is-for-quantum-by-terry-rudolph-pp-164-995-paper-
isbn-9780999063507-terrence-rudolph-2017/F7BF2F8D83DBBA507062D6DEC722E0A1.

Zelator:2010:NEI

[ZF10] Konstantine Zelator and Ovidiu Furdui. 94.12 The non-existence of integer solutions to py^{2} = x^{4} + x. The Mathematical Gazette, 94(530):290–294, July 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Zawaira:2009:PMC

[ZH09] Alexander Zawaira and Gavin Hitchcock. A primer for mathematics competitions. Oxford mathematics. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2009. ISBN 0-19-953987-1 (hardcover), 0-19-953988-X (paperback). xvii + 344 pp.

LCCN QA43 .Z345 2009. URL http://www.loc.gov/catdir/
toc/fy0903/2008301216.html.

Zhou:2011:IPH

[Zho11] Li Zhou. Irrationality proofs à la Hermite. The Mathematical Gazette, 95(534):407–413, November 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Zhou:2013:FS

[Zho13] Li Zhou. Fun with sign. The Mathematical Gazette, 97(540):465– 473, November 2013. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Zorzi:2010:LPN

[Zor10] Alberto Zorzi. 94.09 Lattice polygons and the number 12: an elementary proof. The Mathematical Gazette, 94(529):137–143, March 2010. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).

Zorzi:2011:BL

[Zor11] Alberto Zorzi. 95.22 Benford’s law and π. The Mathematical Gazette, 95(533):264–266, July 2011. CODEN MAGAAS. ISSN 0025-5572 (print), 2056-6328 (electronic).